### v1

Abstract
In this paper we study a class of weighted estimands, which we define as parameters that can be expressed as weighted averages of the underlying heterogeneous treatment effects. The popular ordinary least squares (OLS), two-stage least squares (2SLS), and two-way fixed effects (TWFE) estimands are all special cases within our framework. Our focus is on answering two questions concerning weighted estimands. First, under what conditions can they be interpreted as the average treatment effect for some (possibly latent) subpopulation? Second, when these conditions are satisfied, what is the upper bound on the size of that subpopulation, either in absolute terms or relative to a target population of interest? We argue that this upper bound provides a valuable diagnostic for empirical research. When a given weighted estimand corresponds to the average treatment effect for a small subset of the population of interest, we say its internal validity is low. Our paper develops practical tools to quantify the internal validity of weighted estimands.
Keywords: internal validity, ordinary least squares, representativeness, treatment effects, two-stage least squares, two-way fixed effects, weakly causal estimands, weighted estimands
JEL classification: C20, C21, C23, C26
1 Introduction
Estimating average treatment effects is an important objective of empirical research in many areas of economics. Applied researchers usually believe that treatment effects are heterogeneous, which means that they vary between units with different observable and unobservable characteristics. Yet, many researchers also favor using well-established estimation methods that were not originally designed with treatment effect heterogeneity in mind. These methods may be chosen because of their computational simplicity, comparability across studies, effectiveness at incorporating high-dimensional covariates, and other reasons. In turn, these methods often lead to estimands that can be represented as weighted averages of the underlying treatment effects of interest.
For example, consider a scenario where unconfoundedness holds given covariates $X$. Let treatment $D$ be binary, $(Y(1),Y(0))$ be potential outcomes, and let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$ be the conditional average treatment effect, or CATE, for covariate value $X$. Following Angrist (1998), if we additionally assume that $\mathbb{E}[D\mid X]$ is linear in $X$, the regression of $Y$ on a constant, treatment $D$, and covariates $X$ yields a coefficient on $D$ that can be written as
$\beta_{\text{OLS}}=\frac{\mathbb{E}[\operatorname{var}(D\mid X)\tau_{0}(X)]}{\mathbb{E}[\operatorname{var}(D\mid X)]},$
a weighted average of CATEs with nonnegative weights that integrate to 1.
In this paper we are concerned with a general class of weighted estimands that can be expressed as follows:
$\mu(a,\tau_{0})$ $\equiv\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$ (1.1)
where $W_{0}\in\{0,1\}$ is an indicator for a subpopulation, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ are the CATEs given covariates $X$ in subpopulation $W_{0}$, $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$ is the probability of being in subpopulation $W_{0}$ given $X$, and $a(X)$ is an identified weight function satisfying $\mathbb{E}[a(X)w_{0}(X)]>0$. The regression estimand above belongs to this class, which can be seen by letting $W_{0}=1$ with probability 1, and letting the weight function $a(X)$ be the conditional variance of treatment given covariates. Under some assumptions, this class also includes the two-stage least squares (2SLS) and two-way fixed effects (TWFE) estimands in instrumental variables and difference-in-differences settings, as well as many other parameters.
There are two main questions that this paper seeks to answer. The first is whether, and under what circumstances, the estimand in (1.1) corresponds to an average treatment effect of the form $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$, where $W^{*}\in\{0,1\}$ is an indicator for a (possibly latent) subpopulation of $W_{0}$. An affirmative answer to this question would endow a specific weighted estimand with some degree of validity as a causal parameter, given that it would then measure the average effect of treatment for a subset of all units.
The second and primary aim of this paper is to quantify the degree of validity of $\mu(a,\tau_{0})$ as a causal parameter. To do this, we characterize the size, and the size relative to $W_{0}$, of subpopulations $W^{*}$ associated with the estimand in (1.1). More plainly, we ask how large $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ can be in the representation $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$. If these probabilities can be large, the estimand corresponds to the average treatment effect for a (relatively) large subpopulation, and when they are small, it corresponds to the average effect for a (relatively) small number of units. If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the parameter of interest, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target. If $\mathbb{P}(W^{*}=1)$, the corresponding marginal probability, is large, we say that $\mu(a,\tau_{0})$ is highly representative of the underlying population.
The answer to our questions about subpopulation existence and size depends on the information we have about the CATE function, $\tau_{0}$. Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any knowledge of this function. If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form. We can also answer the second question about the maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ without needing to estimate or know the structure of the CATEs. In a second case, we may want to know how representative $\mu(a,\tau_{0})$ is given knowledge of the CATE function. While the resulting maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are less useful as measures of robustness than in the first case—after all, if the researcher knows or estimates the entire CATE function, they can as well report any average of $\tau_{0}(X)$ that may be of interest—we consider this problem to be of independent theoretical interest. Additionally, if the researcher estimates and compares the maximum values of $\mathbb{P}(W^{*}=1)$ or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ in both cases, they can evaluate the impact of the nonuniform weights in $\mu(a,\tau_{0})$ in a given application.
In the first case, when the CATE function is unrestricted, we formally show that $\mu(a,\tau_{0})$ can be written as the average treatment effect for a subpopulation of $W_{0}$ if and only if $a(X)\geq 0$ with probability 1 given $W_{0}=1$. The contrapositive of this statement is that the incidence of “negative weights,” i.e. $\mathbb{P}(a(X)<0)>0$, implies that $\mu(a,\tau_{0})$ cannot be represented as an average treatment effect for some subpopulation uniformly in $\tau_{0}$. In a related contribution, Blandhol, Bonney, Mogstad, and Torgovitsky (2022) have shown that, for estimands that do not depend on potential outcome levels, the lack of negative weights is a sufficient and necessary condition for the weighted estimand to be “weakly causal,” that is, to guarantee that the sign of $\tau_{0}$ will be preserved whenever it is uniform across all units. We also provide simple expressions for the maxima of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$. We propose an analog estimator for these objects and briefly discuss associated estimation and inferential issues.
In the second case, when the CATE function is assumed to be known, we show that $\mu(a,\tau_{0})$ can be written as an average treatment effect whenever it lies in the convex hull of CATE values, a weaker criterion than having nonnegative weights. The maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ now depend on $\tau_{0}$, and can be obtained via linear programming when $X$ is discrete. We show the solution to this linear program also admits a closed-form expression even when the support of $X$ includes discrete, continuous, and mixed components. This expression can be used to derive plug-in estimators.
Literature Review
This paper is related to a large literature studying weighted average representations of common estimands, including ordinary least squares (OLS), 2SLS, and TWFE in additive linear models. Some of the contributions to this literature include Angrist (1998), Humphreys (2009), Aronow and Samii (2016), Blandhol, Bonney, Mogstad, and Torgovitsky (2022), Słoczyński (2022), and Goldsmith-Pinkham, Hull, and Kolesár (2024) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
A common view in much of this literature, traceable to Imbens and Angrist (1994), is that causal interpretability of weighted estimands requires all weights to be positive. For example, Sun and Abraham (2021) explicitly associate “reasonable weights” with weights that “sum to one and are non-negative.” Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that the lack of negative weights and level dependence is necessary and sufficient for an estimand to be “weakly causal,” that is, to guarantee sign preservation when all treatment effects have the same sign. In this paper we focus on the related problem of whether a weighted estimand can be written as the average treatment effect for some (possibly latent) subpopulation. While the lack of negative weights is essential in our framework when the CATE function is unrestricted, negative weights and nonuniform weights play a similar role when the CATE function is assumed to be known, at least as long as the weighted estimand lies in the convex hull of CATE values. This point is related to the negative view of both negative and nonuniform weights in Callaway, Goodman-Bacon, and Sant’Anna (2024).
Some papers focus on weighted averages of heterogeneous treatment effects as legitimate targets in their own right rather than as probability limits of existing estimators. Hirano, Imbens, and Ridder (2003) introduce the class of weighted average treatment effects, which are a subclass of the more general class of estimands in (1.1). Li, Morgan, and Zaslavsky (2018) discuss the connection between weighted average treatment effects and implicit target subpopulations. However, the internal validity and representativeness of weighted estimands have received very little attention to date.
One exception is Aronow and Samii (2016), who explicitly acknowledge that the OLS estimand, like the local average treatment effect of Imbens and Angrist (1994), corresponds to the average effect for a “highly specific subpopulation” rather than the entire population, and consequently is not necessarily representative of that population. Then, Aronow and Samii (2016) focus on whether mean covariate values are similar in the entire sample and in the “effective sample” used by OLS. We focus on the size of the implicit subpopulation, which is different and complementary.
Another exception is Miller, Shenhav, and Grosz (2023), who focus on (one-way) fixed effects estimands and argue that it is problematic if “switchers,” that is, fixed-effect groups with nonzero variation in treatment, are a small subset of the sample. They also recommend that applied researchers report the sample size when limited to “switcher groups.” In this paper we build a general framework to study the internal validity and representativeness of weighted estimands, with the fixed effects estimand (equivalent to OLS) as a special case. We argue, similar to Miller, Shenhav, and Grosz (2023), that if a given weighted estimand corresponds to the average treatment effect for a small subpopulation, then it may not be an appropriate target parameter, unless that subpopulation is interesting in its own right.
Plan of the Paper
We organize the paper as follows. In Section 2, we discuss several examples of commonly used estimands that satisfy the representation in (1.1). In Section 3, we develop our theoretical framework and examine the conditions under which the estimand in (1.1) has a causal representation as an average treatment effect over a population. In Section 4, we establish our main results on the absolute and relative size of subpopulations associated with the estimand in (1.1), which we propose as measures of representativeness and internal validity of weighted estimands. We also generalize our results to subpopulations and revisit the examples from Section 2. In Section 5, we briefly discuss estimation and inference for the proposed measures. In Section 6, we provide an empirical application to the effects of unilateral divorce laws on female suicide, as in Stevenson and Wolfers (2006) and Goodman-Bacon (2021). In Section 7, we conclude. Appendix contains our proofs as well as several additional results and derivations.
2 Applications
Here we consider three identification strategies where commonly used estimands follow the structure of equation (1.1).
2.1 Unconfoundedness
In this setup, we have a binary treatment $D\in\{0,1\}$, potential outcomes $(Y(1),Y(0))$, covariate vector $X$, and realized outcome $Y=Y(D)$. We make the following two assumptions.
Assumption 2.1 (Unconfoundedness).
Let
  1. 1.
    Conditional independence: $(Y(1),Y(0))\perp\!\!\!\perp D\mid X$;
  2. 2.
    Overlap: $p(X)\equiv\mathbb{P}(D=1\mid X)\in(0,1)$ almost surely.
Following Angrist (1998), we can establish that $\beta_{\text{OLS}}$, the coefficient on $D$ in the linear projection of $Y$ on $(1,D,X)$, satisfies the representation in (1.1). The following proposition summarizes Angrist’s (1998) result.
Proposition 2.1.
Let Assumption 2.1 hold and let $p(X)$ be linear in $X$. Then
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[p(X)(1-p(X))\cdot\mathbb{E}[Y(1)-Y(0)\mid X]]}{\mathbb{E}[p(X)(1-p(X))]}.$
The linearity assumption can be removed if we instead regress $Y$ on $(1,D,h(X))$ where $h(X)$ is a vector of functions of $X$ such that $p(X)$ is in their linear span. The overlap assumption can also be weakened since it is not required for $\beta_{\text{OLS}}$ to be defined.
Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)]}{\mathbb{E}[a(X)]},$
where $a(X)=p(X)(1-p(X))$ and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$. Here we implicitly set $W_{0}=1$ with probability 1. Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are $p(X)(1-p(X))$. Note that $\beta_{\text{OLS}}=\text{ATE}\equiv\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case when $p(X)$ or $\tau_{0}(X)$ is constant.
An alternative representation of this estimand can be obtained by focusing on the subpopulation of treated units, $D=1$. Let $W_{0}=D$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,X]=\mathbb{E}[Y(1)-Y(0)\mid X]$, which follows from conditional independence, and let $\tilde{a}(X)=1-p(X)$. Then, we can write
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]}=\frac{\mathbb{E}[\tilde{a}(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[\tilde{a}(X)w_{0}(X)]},$
where $w_{0}(X)=\mathbb{P}(D=1\mid X)=p(X)$. Yet another representation can be obtained when focusing on the subpopulation of untreated units by letting $W_{0}=1-D$. We omit details for brevity.
2.2 Instrumental Variables
Consider an endogenous binary treatment $D\in\{0,1\}$ and a binary instrument $Z\in\{0,1\}$. Potential treatments, denoted by $(D(1),D(0))$, are linked to the realized treatment through $Z$, that is, $D=D(Z)$. Potential outcomes, $Y(d,z)$ for $d,z\in\{0,1\}$, may depend on both $D$ and $Z$ in the absence of an exclusion restriction. Let $Y=Y(D,Z)$ be the realized outcome. As before, let $X$ denote covariates. We make the following assumptions.
Assumption 2.2 (Instrument validity).
We have
  1. 1.
    Exogeneity: $(Y(0,0),Y(1,0),Y(0,1),Y(1,1),D(1),D(0))\perp\!\!\!\perp Z\mid X$;
  2. 2.
    Exclusion: $\mathbb{P}(Y(d,0)=Y(d,1))=1$ for $d\in\{0,1\}$;
  3. 3.
    First stage: $e(X)\equiv\mathbb{P}(Z=1\mid X)\in(0,1)$ and $\mathbb{P}(D(1)=1\mid X)\neq\mathbb{P}(D(0)=1\mid X)$ almost surely.
We also make one of the following two nested monotonicity assumptions.
Assumption 2.3 (Strong monotonicity).
$\mathbb{P}(D(1)\geq D(0)\mid X)=1$ almost surely.
Assumption 2.4 (Weak monotonicity).
There exists a subset of the support of $X$ such that $\mathbb{P}(D(1)\geq D(0)\mid X)=1$ on it and $\mathbb{P}(D(1)\leq D(0)\mid X)=1$ on its complement.
The first instrumental variables estimand we consider was originally studied by Angrist and Imbens (1995). In addition to the assumptions on the instrument, suppose that the model for $X$ is saturated, with $K$ possible combinations of covariate values, i.e. let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, where $\text{supp}(\cdot)$ is used to denote the support. Let $X_{S}=\left(1,\mathbbm{1}(X=x_{1}),\ldots,\mathbbm{1}(X=x_{K-1})\right)$ and $Z_{S}=\left(Z,Z\mathbbm{1}(X=x_{1}),\ldots,Z\mathbbm{1}(X=x_{K-1})\right)=ZX_{S}$, where $Z_{S}$ is the constructed instrument vector. The estimand in Angrist and Imbens (1995) is the following 2SLS estimand:
$\beta_{\text{2SLS}}\equiv\left[\left(\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}W_{S}\right]\right)^{-1}\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}Y\right]\right]_{1},$
where $W_{S}=\left(D,X_{S}\right)$, $Q_{S}=\left(Z_{S},X_{S}\right)$, and $\left[\cdot\right]_{k}$ denotes the $k$th element of the corresponding vector. This estimand has been studied by Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022), and the specific representation in Proposition 2.2 follows Słoczyński (2020).
Proposition 2.2.
Let Assumptions 2.2 and 2.4 hold, and let $X$ be discrete with finite support. Then
$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}\left[e(X)(1-e(X))\cdot\mathbb{P}\left(D(1)\neq D(0)\mid X\right)^{2}\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]\right]}{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)^{2}]}.$
This means that we can write $\beta_{\text{2SLS}}$ as
$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]},$
where $W_{0}=\mathbbm{1}(D(1)\neq D(0))$ is the population of compliers/defiers, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]$, and $a(X)=e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)$, a nonnegative weight function. Note that $\beta_{\text{2SLS}}=\text{LATE}\equiv\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0)]$ if and only if $a(X)$ is uncorrelated with $\tau_{0}(X)$ given $D(1)\neq D(0)$.
The practical limitation of focusing on $\beta_{\text{2SLS}}$ is that applied researchers rarely create additional instruments by interacting the original instrument with covariates (cf. Blandhol, Bonney, Mogstad, and Torgovitsky, 2022), which is how $Z_{S}$ is constructed to obtain $\beta_{\text{2SLS}}$ above. A more practically relevant estimand is the “just-identified” IV estimand,
$\beta_{\text{IV}}\equiv\left[\left(\mathbb{E}\left[Q^{\prime}W\right]\right)^{-1}\mathbb{E}\left[Q^{\prime}Y\right]\right]_{1},$
where $Q=\left(Z,X\right)$ and $W=\left(D,X\right)$. To introduce one of the representations of $\beta_{\text{IV}}$ below, define
$c(X)\equiv\operatorname{sign}\Big{(}\mathbb{P}\left[D(1)\geq D(0)\mid X\right]-\mathbb{P}\left[D(1)\leq D(0)\mid X\right]\Big{)},$
where $\operatorname{sign}(\cdot)$ is the sign function. We also make the following “rich covariates” assumption on the instrument propensity score, which is implied by the saturated specification in Proposition 2.2.
Assumption 2.5 (Rich covariates).
$e(X)$ is linear in $X$.
Under the instrument validity assumption, the rich covariates assumption, and either monotonicity assumption, Słoczyński (2020) obtains the following representations for the “just-identified” IV estimand.
Proposition 2.3.
Let Assumptions 2.2, 2.4, and 2.5 hold. Then
$\beta_{\text{IV}}=\frac{\mathbb{E}[c(X)\cdot e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]]}{\mathbb{E}[c(X)\cdot e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)]}.$
If Assumptions 2.2, 2.3, and 2.5 hold instead, then
$\beta_{\text{IV}}=\frac{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)>D(0)\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]]}{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)>D(0)\mid X)]}.$
It follows that we can write $\beta_{\text{IV}}$ as
$\beta_{\text{IV}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}$
under either monotonicity assumption. Under weak monotonicity, $W_{0}=\mathbbm{1}(D(1)\neq D(0))$, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]$, and possibly negative weights $a(X)=c(X)\cdot e(X)(1-e(X))$. Under strong monotonicity, $W_{0}=\mathbbm{1}(D(1)>D(0))$, $w_{0}(X)=\mathbb{P}(D(1)>D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]$, and nonnegative weights $a(X)=e(X)(1-e(X))\geq 0$.
2.3 Difference-in-Differences
Now suppose units are observed for $T$ periods and, for $t\in\{1,\ldots,T\}$, denote binary treatment $D_{t}\in\{0,1\}$, potential outcomes $(Y_{t}(1),Y_{t}(0))$, and realized outcome $Y_{t}=Y_{t}(D_{t})$. We assume units are untreated prior to period $G\in\mathcal{G}=\{2,3,\ldots,T\}\cup\{+\infty\}$, receive the treatment in period $G$, and remain treated thereafter. We assume no units are treated in the first time period. This may include a group that remains untreated throughout, for which $G=+\infty$. Thus, $D_{t}=\mathbbm{1}(G\leq t)$. The panel is balanced, that is, no group appears or disappears over time.
The two-way fixed effects estimand is often used in this setting, and consists of regressing the outcome on the treatment indicator, group indicators, and period indicators. Using partitioned regression results, it is defined as
$\beta_{\text{TWFE}}\equiv\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}Y_{t}\right]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}^{2}\right]},$
where $\ddot{D}_{t}=D_{t}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]$.
We assume a version of parallel trends most similar to the one in de Chaisemartin and D’Haultfœuille (2020).
Assumption 2.6 (Difference-in-differences).
We have
  1. 1.
    $\mathcal{G}\equiv\text{supp}(G)=\{2,3,\ldots,T\}\cup\{+\infty\}$;
  2. 2.
    For all $t\in\{2,\ldots,T\}$ and $g,g^{\prime}\in\mathcal{G}$, we have that $\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g]=\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g^{\prime}]$.
We use a proposition that is essentially a special case of Theorem 1 in de Chaisemartin and D’Haultfœuille (2020) to obtain a representation of the two-way fixed effects estimand as a weighted average.
Proposition 2.4.
Let Assumption 2.6 hold. Then
$\beta_{\text{TWFE}}=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}\big{(}1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]\big{)}\cdot\mathbb{P}(D_{t}=1\mid G)\cdot\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid G,D_{t}=1]\Big{]}}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}\big{(}1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]\big{)}\cdot\mathbb{P}(D_{t}=1\mid G)\Big{]}}.$
We show the above representation satisfies equation (1.1) by introducing an auxiliary variable $P$ that is uniformly distributed on $\{1,\ldots,T\}$ independently from $\{(Y_{t}(0),Y_{t}(1),G)\}_{t=1}^{T}$. This period variable denotes the time period and we use it to define $(Y(1),Y(0))\equiv(Y_{P}(1),Y_{P}(0))$, $Y=Y_{P}$, and $D=D_{P}$, which are potential outcomes, the realized outcome, and treatment at a random time period, respectively.
Letting $X=(G,P)$, this means we can write $\beta_{\text{TWFE}}$ as
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]},$
where $W_{0}=D$, $w_{0}(X)=\mathbb{P}(D=1\mid G,P)\in\{0,1\}$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G,P]$, and the weight function $a(X)=1-\mathbb{P}(D=1\mid G)-\mathbb{P}(D=1\mid P)+\mathbb{P}(D=1)$ is not generally nonnegative. [^2] Here, $\tau_{0}(X)$ is what Callaway and Sant’Anna (2021) call “the group-time average treatment effect.” A nonnegative weight function can be obtained under the assumption that group-level average treatment effects are constant over time. This property was described in Goodman-Bacon (2021, Section 3.1.1) and a representation of the two-way fixed effects estimand is given in his equation (16). The following proposition yields a simpler (but equivalent) expression for the weights in our setting.
Proposition 2.5.
Let Assumption 2.6 hold and let $\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid D=1,G]=\mathbb{E}[Y_{s}(1)-Y_{s}(0)\mid D=1,G]$ for any $s,t\in\{1,\ldots,T\}$. Then
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\mathbb{E}[Y(1)-Y(0)\mid D=1,G]]}{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)]},$
where $a_{B}(g)=\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\geq 0$ for $g\in\{2,\ldots,T\}$.
As is the case of the representation in Proposition 2.4, the two-way fixed effects estimand in Proposition 2.5 satisfies the representation in (1.1), with $X=G$, $W_{0}=D$, $w_{0}(X)=\mathbb{P}(D=1\mid G)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G]$, and the weight function $a(X)=a_{B}(G)\geq 0$. This weight function is derived in the proof of the proposition, and we show it is equivalent to equation (16) in Goodman-Bacon (2021) in Appendix D.
3 Causal Representation of Weighted Estimands
We now consider a class of weighted estimands that subsumes all our previous examples. In this section, we show necessary and sufficient conditions for an estimand of this class to have a causal representation as an average treatment effect over a population. Then, we compute the largest size of a population that is compatible with the weighted estimand. The size of this population relative to $\mathbb{P}(W_{0}=1)$ will be a measure of the internal validity of our estimand relative to the target estimand $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. The absolute size of this population will be a measure of the representativeness of the estimand with respect to the population distribution the data is drawn from.
Recall the earlier setting where we let $D\in\{0,1\}$ denote a binary treatment variable, and let $(Y(1),Y(0))$ denote the corresponding potential outcomes under treatment and control, respectively. Let $X\in\text{supp}(X)\subseteq\mathbb{R}^{d_{X}}$ denote a $d_{X}$-vector of covariates. We suppose that $(Y(1),Y(0),D,X)$ are drawn from a common population distribution $F_{Y(1),Y(0),D,X}$.
Let $W_{0}\in\{0,1\}$ be an indicator variable used to denote a subpopulation $\{W_{0}=1\}$. Throughout this paper, we assume that $\mathbb{P}(W_{0}=1)>0$, so that this subpopulation has a positive mass, which avoids technical issues associated with conditioning on zero-probability events. For example, this subpopulation can be the entire population by setting $W_{0}=1$ almost surely, the treated units by setting $W_{0}=D$, or units complying with binary instrument $Z$ by setting $W_{0}=\mathbbm{1}(D(1)>D(0))$. Let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ denote the conditional average treatment effect given $X$ in subpopulation $W_{0}$. This is defined for all values of $X$ such that $w_{0}(X)>0$.
Also recall the weighted estimands of equation (1.1):
$\mu(a,\tau_{0})$ $\equiv\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$
where $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$. [^3] While $\tau_{0}(X)$ is only defined for $w_{0}(X)>0$, we set $\tau_{0}(X)w_{0}(X)=0$ when $w_{0}(X)=0$.
The estimands we consider have the above representation and satisfy the following regularity conditions.
Assumption 3.1 (Regularity).
Let $\mathbb{E}[\tau_{0}(X)^{2}]<\infty$, $\mathbb{E}[a(X)^{2}]<\infty$, and $\mathbb{E}[a(X)\mid W_{0}=1]>0$.
The first two restrictions are weak regularity assumptions that ensure the finiteness of the numerator of $\mu(a,\tau_{0})$. We rule out $\mathbb{E}[a(X)\mid W_{0}=1]=0$ since it implies the estimand does not exist. The estimand in (1.1) is unchanged if the sign of $a(X)$ is reversed, so $\mathbb{E}[a(X)\mid W_{0}=1]>0$ is a sign normalization.
3.1 Alternative Representations of Weighted Estimands
The weighted estimands of equation (1.1) can also be written as a weighted sum when $X$ is discrete, or an integral when $X$ is continuous. In the discrete case, let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, $W_{0}=1$ almost surely, and let $p_{k}\equiv\mathbb{P}(X=x_{k})>0$ for $k=1,\ldots,K$. Then,
$\mu(a,\tau_{0})$ $=\sum_{k=1}^{K}\omega_{k}\tau_{0}(x_{k})\qquad\text{ where }\qquad\omega_{k}=\frac{a(x_{k})p_{k}}{\sum_{l=1}^{K}a(x_{l})p_{l}},$ (3.1)
which are weights that sum to one. The representations in (1.1) and (3.1) are equivalent as we can obtain $a(x_{k})$ (up to scale) as the ratio $\omega_{k}/p_{k}$, and $\omega_{k}$ is defined as a function of $(a(\cdot),p_{\cdot})$ in equation (3.1) above.
From (3.1), we can see that $a(x_{k})$ being constant ensures $\omega_{k}=p_{k}$, or that the estimand is the ATE. Moreover,
$\frac{a(x_{k})}{a(x_{k^{\prime}})}=\frac{\omega_{k}}{\omega_{k^{\prime}}}\big{/}\frac{p_{k}}{p_{k^{\prime}}},$
which is the ratio of the relative weights of covariate cells $\{X=x_{k}\}$ and $\{X=x_{k^{\prime}}\}$ in the estimand $(\omega_{k}/\omega_{k^{\prime}})$ and in the population $(p_{k}/p_{k^{\prime}})$. The inequality $a(x_{k})>a(x_{k^{\prime}})$ indicates that covariate cell $\{X=x_{k}\}$ is overweighted relative to $\{X=x_{k^{\prime}}\}$, when compared to their relative weights in the population.
Alternatively, consider the case where $X$ is continuously distributed with density [^4] With respect to the Lebesgue measure. $f_{X}$. Still assuming $W_{0}=1$ almost surely, we can write
[…]
is a weight function that integrates to 1. We focus on the representation in equation (1.1) since it seamlessly accommodates discrete, continuous, and mixed covariates.
3.2 Regular Subpopulations
The first question we address is whether an estimand defined by (1.1) can be represented as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$, where $W^{*}\in\{0,1\}$ is binary and $\{W^{*}=1\}$ characterizes a subpopulation of $\{W_{0}=1\}$. Formally, $\{W^{*}=1\}$ forms a subpopulation if $\{W^{*}=1\}\subseteq\{W_{0}=1\}$ or, equivalently, if $W^{*}\leq W_{0}$ almost surely.
We impose some structure on this problem by restricting how these subpopulations may be formed. We will consider what we call “regular subpopulations,” which we define here.
Definition 3.1.
Let $W^{*}\in\{0,1\}$ such that $\mathbb{P}(W^{*}=1)>0$. Say $\{W^{*}=1\}$ is a regular subpopulation of $\{W_{0}=1\}$ if
  1. 1.
    (Inclusion) $\mathbb{P}(W^{*}=1\mid W_{0}=0)=0$;
  2. 2.
    (Conditional Independence) $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$.
For convenience, we will abbreviate this as “$W^{*}$ is a regular subpopulation of $W_{0}$”. We denote the set of regular subpopulations of $W_{0}$ as
$\text{SP}(W_{0})$ =$=\{W^{*}\in\{0,1\}:W^{*}\text{ is a regular subpopulation of }W_{0}\}.$
We consider subpopulations with positive masses that are subsets of $\{W_{0}=1\}$. The second and main requirement is that regular subpopulations do not depend on potential outcomes when conditioning on $X$ and the original population $W_{0}=1$. While this may seem restrictive, it allows for rich and natural classes of subpopulations. For example, consider the unconfoundedness restriction of Section 2.1 and let $W_{0}$ be the entire population, i.e. $\mathbb{P}(W_{0}=1)=1$. In this case, regular subpopulations must satisfy $\;\;W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X$, or be unconfounded. Regular subpopulations include the population of all treated (or untreated) individuals, i.e. $W^{*}=D$ (or $W^{*}=1-D$), and any subpopulation characterized by a subset of $\text{supp}(X)$. More generally, it includes any subpopulation that can be described through a combination of $(D,X,U)$ where $U$ is independent from $(Y(1),Y(0),X)$. For example, a subpopulation characterized by “fraction $a(x)$ of units with covariate $X=x$ for all $x\in\text{supp}(X)$” can be constructed as $W^{*}=\mathbbm{1}(U\leq a(X))$ where $U\sim\text{Unif}(0,1)$ is independent from $(Y(1),Y(0),X)$.
The conditional independence requirement rules out subpopulations that directly depend on the potential outcomes such as $W^{*}=\mathbbm{1}(Y(1)\geq Y(0))$, the subpopulation of those who benefit from treatment. Note that $\mathbb{P}(W^{*}=1\mid X)=\mathbb{P}(Y(1)\geq Y(0)\mid X)$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(Y(1)\geq Y(0))$ are not point-identified under unconfoundedness.
These particular subpopulations enjoy a number of useful properties. Two of them are characterized in the following proposition.
Proposition 3.1 (Properties of regular subpopulations).
Let $\mathbb{P}(W_{0}=1)>0$ and $W^{*}\in\text{SP}(W_{0})$.
  1. 1.
    Let $\mathbb{P}(W^{*}=1\mid W_{0}=1,X)>0$. Then,
    $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X]$ $=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X].$ (3.3)
  2. 2.
    Let $\mathbb{E}[\tau_{0}(X)^{2}]<\infty$. Then,
    $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}=\mu(\underline{w}^{*},\tau_{0}),$ (3.4)
    where $\underline{w}^{*}(x)=\mathbb{P}(W^{*}=1\mid X=x,W_{0}=1)$.
The first part of this proposition shows that average effects within the original population $W_{0}$ and regular subpopulation $W^{*}$ are the same when conditioning on $X$. For example, this holds under unconfoundedness for the subpopulation of treated individuals, $W^{*}=D$. The second property allows us to write the average effect for $W^{*}=1$ using the same functional $\mu(\cdot,\cdot)$ that was used to characterize the class of estimands we analyze. This property will be used when studying the mapping between weighted estimands and average effects for regular subpopulations of $W_{0}$.
We conclude this subsection by showing that regular subpopulations are transitive, or that regular subpopulations of a regular subpopulation of $W_{0}$ are also regular subpopulations of $W_{0}$.
Lemma 3.1 (Transitivity of regular subpopulations).
Let $W^{*}$ be a regular subpopulation of $W^{\prime}$ and let $W^{\prime}$ be a regular subpopulation of $W_{0}$. Then, $W^{*}$ is a regular subpopulation of $W_{0}$.
3.3 Existence of a Causal Representation for Weighted Estimands
We now consider necessary and sufficient conditions for the weighted estimand $\mu(a,\tau_{0})$ to be written as the average treatment effect within a regular subpopulation of $W_{0}$. As we will show, these conditions depend on what is assumed about the function $\tau_{0}=\mathbb{E}[Y(1)-Y(0)\mid X=\cdot,W_{0}=1]$.
For example, if $\tau_{0}$ is constant in $X$, then any weighted estimand satisfying (1.1) equals $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the average treatment effect within population $\{W_{0}=1\}$. This is the case even when the sign of weight function $a(X)$ varies with $X$. However, if $\tau_{0}$ is nonconstant, the existence of causal representations will depend on the weight function $a(X)$. Among other cases, we will consider the case where no restrictions are placed on function $\tau_{0}$, and in this case, the existence of a causal representation of $\mu(a,\tau_{0})$ will require the sign of $a(X)$ to be constant.
To formalize this, let $\mathcal{T}$ denote a class of functions such that $\tau_{0}\in\mathcal{T}$ and define
$\mathcal{W}(a;W_{0},\mathcal{T})$ =$=\{W^{*}\in\text{SP}(W_{0}):\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\text{ for all }\tau_{0}\in\mathcal{T}\}.$
This is the set of regular subpopulations of $W_{0}$ such that the estimand $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for all $\tau_{0}$ functions in the set $\mathcal{T}$. If it is empty, then estimand $\mu(a,\tau_{0})$ cannot be written as an average treatment effect over a regular subpopulation of $W_{0}$ uniformly in $\tau_{0}\in\mathcal{T}$. We use this set to formally define a notion of uniform causal representation.
Definition 3.2.
A weighted estimand $\mu(a,\tau_{0})$ has a causal representation uniformly in $\tau_{0}\in\mathcal{T}$ if
$\mathcal{W}(a;W_{0},\mathcal{T})\neq\emptyset.$
Recall that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu(\underline{w}^{*},\tau_{0})$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$, so $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})$ if $\mu(a,\tau_{0})=\mu(\underline{w}^{*},\tau_{0})$ for all $\tau_{0}\in\mathcal{T}$. We examine further two main cases for the set $\mathcal{T}$.
3.3.1 Existence Uniformly in $\tau_{0}$
We begin by considering the largest class of functions in which $\tau_{0}$ lies: the class of all functions, subject to the moment condition in Assumption 3.1 that ensures the existence of $\mu(a,\tau_{0})$. We denote this class by
$\mathcal{T}_{\text{all}}$$\equiv\{\tau_{0}:\mathbb{E}[\tau_{0}(X)^{2}]<\infty\}.$
In this function class, the existence of a causal representation is equivalent to the weights being nonnegative.
Theorem 3.1.
Let Assumption 3.1 hold and let $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$ if and only if $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$.
A uniform (in $\mathcal{T}_{\text{all}}$) causal representation exists if and only if $a(X)$ is nonnegative on the support of $X\mid W_{0}=1$. To give some intuition on why the sign of $a(X)$ must be nonnegative with probability 1, partition $\text{supp}(X\mid W_{0}=1)$ in $\mathcal{X}^{+}\equiv\{x\in\text{supp}(X\mid W_{0}=1):a(x)\geq 0\}$ and $\mathcal{X}^{-}\equiv\{x\in\text{supp}(X\mid W_{0}=1):a(x)<0\}$. If $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$, then $\mathbb{P}(X\in\mathcal{X}^{+}\mid W_{0}=1)>0$ and $\mathbb{P}(X\in\mathcal{X}^{-}\mid W_{0}=1)>0$ under Assumption 3.1. Then, setting $\tau^{-}(X)=\mathbbm{1}(X\in\mathcal{X}^{-})$ and $\tau^{+}(X)=\mathbbm{1}(X\in\mathcal{X}^{+})$ yields estimands of different signs, i.e. $\mu(a,\tau^{-})<0<\mu(a,\tau^{+})$, despite $\tau^{-}$ and $\tau^{+}$ being of the same sign. In other words, as is well known, nonnegative CATEs can be combined into a negative estimand when weights are negative. Hence, such an estimand is not weakly causal (Blandhol, Bonney, Mogstad, and Torgovitsky, 2022).
Conversely, if $a(X)\geq 0$, our proof constructively defines a regular subpopulation $W^{*}$ for which the average effect is equal to the weighted estimand $\mu(a,\tau_{0})$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$. Let
$W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0},$
where $U\sim\text{Unif}(0,1)\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$. This is a regular subpopulation of $W_{0}$ for which the probability of inclusion, conditional on $X$ and $W_{0}=1$, is proportional to $a(X)$. The condition $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$ restricts our attention to subpopulations with positive mass. From this construction, we can see that $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ is proportional to $a(X)$, and therefore $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu(\underline{w}^{*},\tau_{0})=\mu(a,\tau_{0})$ uniformly in $\tau_{0}$.
3.3.2 Existence for a Given $\tau_{0}$
We now provide an existence result that requires the causal representation to exist only for the given $\tau_{0}$, rather than uniformly for $\tau_{0}$ in the larger set $\mathcal{T}_{\text{all}}$. This result’s assumptions depend on the nature of the CATE function $\tau_{0}$ in the population, whereas Theorem 3.1’s condition depended only on the weight function $a(X)$, and the nature of the covariates’ support. Thus, the distribution of the potential outcomes will have an impact on the existence of a causal representation given $\tau_{0}$. Using the notation from Definition 3.2, a causal representation exists if and only if $\mathcal{W}(a;W_{0},\{\tau_{0}\})\neq\emptyset$. The following theorem characterizes this existence.
Theorem 3.2.
Let Assumption 3.1 hold. Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ if and only if
$\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W_{0})\equiv\{t\in\mathbb{R}:\mathbb{P}(\tau_{0}(X)\leq t\mid W_{0}=1)>0\text{ and }\mathbb{P}(\tau_{0}(X)\geq t\mid W_{0}=1)>0\}.$ (3.5)
The existence condition in Theorem 3.2 is weaker than the one in Theorem 3.1 since we no longer require this representation to be valid for all $\tau_{0}\in\mathcal{T}_{\text{all}}$, but rather for just one its element. The necessary and sufficient condition in this theorem is rather weak, since it only requires that the estimand is in the convex hull of the support of the CATEs. This means $\mu(a,\tau_{0})$ has a causal representation even with negative weights, as long as there are CATEs smaller and greater than $\mu(a,\tau_{0})$. We can see this support condition holds for all $\tau_{0}\in\mathcal{T}_{\text{all}}$ if and only if $\mu(a,\tau_{0})$ is in the support of $\tau_{0}(X)$ for all $\tau_{0}$. This is precisely the case when the weights $a(X)$ are nonnegative since it guarantees $\inf(\text{supp}(\tau_{0}(X)\mid W_{0}=1))\leq\mu(a,\tau_{0})\leq\sup(\text{supp}(\tau_{0}(X)\mid W_{0}=1))$ for any function $\tau_{0}$.
3.3.3 Intermediate Cases
Analyzing the representativeness of an estimand under no restrictions on $\tau_{0}$ could be viewed as unnecessarily conservative in some settings. At the other extreme, assuming knowledge of $\tau_{0}$ may be unrealistic, especially in scenarios where $X$ has many components which makes the estimation of $\tau_{0}$ more challenging. For example, some shape constraints may be known to hold for $\tau_{0}$. In some economic applications one may posit that $\tau_{0}$ is monotonic or convex in some components of $X$, or positive/negative over a subset of $\text{supp}(X\mid W_{0}=1)$. In these cases, the existence of a causal representation may occur under weaker conditions than Theorem 3.1, but stronger than Theorem 3.2. In particular, one may be able to relax the requirement that $a(X)\geq 0$ without requiring that $\tau_{0}$ be completely known to the researcher. The following proposition shows this is the case when $\tau_{0}(X)$ is assumed to be linear in $X$.
Proposition 3.2.
Let Assumption 3.1 hold and define
$\mathcal{T}_{\text{lin}}$ =$=\{\tau_{0}(X)=c+d^{\prime}X:(c,d)\in\mathbb{R}^{1+d_{X}}\}.$
Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$ if and only if
$\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}\in\text{conv}(\text{supp}(X\mid W_{0}=1)),$
where $\text{conv}(\cdot)$ denotes the convex hull.
The above proposition shows that once the class of CATE functions is restricted, the requirement that $a(X)$ be nonnegative is no longer needed for the existence of a uniform causal representation for an estimand. In particular, the requirement here is that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$ lies in the convex hull of the support of $X$ given $W_{0}=1$. When $X$ is scalar, this consists of an interval. This condition does not require $a(X)$ be nonnegative: for example, if $\text{supp}(X)=\{0,1,2\}$, $\mathbb{P}(X=x)=1/3$ for $x\in\{0,1,2\}$, and $W_{0}=1$ almost surely, then any combination of $(a(0),a(1),a(2))$ values such that $\frac{a(1)+2a(2)}{a(0)+a(1)+a(2)}\in[0,2]$ implies a causal representation. For example, $(a(0),a(1),a(2))=(1,-1,1)$ contains a negative element, and thus $\mathbb{P}(a(X)\geq 0)=2/3<1$, but $\frac{a(1)+2a(2)}{a(0)+a(1)+a(2)}=1\in[0,2]$, implying that the corresponding weighted estimand has a causal representation.
We consider another class of CATE functions that restricts their heterogeneity. For $K\geq 0$, let
$\mathcal{T}_{\text{BV}}(K)=\left\{\tau\in\mathcal{T}_{\text{all}}:\sup_{x,x^{\prime}\in\text{supp}(X\mid W_{0}=1)}|\tau(x)-\tau(x^{\prime})|\leq K\right\}.$
This function class bounds the variation of the CATE function. When $K=0$, the CATE function is constant, and thus equal to $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. When $K>0$, CATEs may differ in value, but the maximum discrepancy between two CATEs is bounded above by $K$. We show that restricting the CATEs to satisfy this bounded variation assumption does not remove the requirement that $a(X)$ be nonnegative, unless $K=0$, in which case all $a$ functions yield a causal representation uniformly in $\mathcal{T}_{\text{BV}}(0)$. We formalize this in the next proposition.
Proposition 3.3.
Let Assumption 3.1 hold. If $K>0$, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ if and only if $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$. If $K=0$, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(0))$ for all $a$.
The last two propositions show that the impact of restrictions on $\tau_{0}$ on the requirement that $a(X)$ be nonnegative depends on the nature of these restrictions. Generalizations to additional or empirically motivated function classes is left for future work.
4 Quantifying the Internal Validity of Weighted Estimands
Many estimands will admit causal representations, but their associated subpopulations $\{W^{*}=1\}$ will generally differ. Also, a weighted estimand may not always correspond to the target estimand a researcher is interested in, such as $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ if the average effect in $\{W_{0}=1\}$ is the parameter of interest. However, the set of $W^{*}$ corresponding to a weighted estimand can be used to understand how representative the weighted estimand is of the target. For example, we may seek estimands for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ attains values closest to 1, since they have a higher degree of internal validity with respect to the target $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. At one extreme, an estimand for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)=1$ is perfectly representative of $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ and would be deemed to have the highest degree of internal validity for this target parameter.
Formally, we ask what is the sharp upper bound on $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ for any regular subpopulation $W^{*}$ of $W_{0}$ such that the weighted estimand $\mu(a,\tau_{0})$ has a causal representation as the average treatment effect over subpopulation $W^{*}$. We denote this object using the following notation:
$\overline{P}(a,W_{0};\mathcal{T})=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})}\mathbb{P}(W^{*}=1\mid W_{0}=1).$ (4.1)
We also set $\overline{P}(a,W_{0};\mathcal{T})=0$ when $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty. This object depends on the chosen function class $\mathcal{T}$, as did Theorems 3.1 and 3.2 in the previous section. Given the above terminology, we call $\overline{P}(a,W_{0};\mathcal{T})$ a measure of the internal validity of estimand $\mu(a,\tau_{0})$, assuming that $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target.
We can also compute the maximum value of $\mathbb{P}(W^{*}=1)$ across $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})$, which equals $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W_{0}=1)\cdot\mathbb{P}(W_{0}=1)$ since $W^{*}$ is a subpopulation of $W_{0}$. This maximum value of $\mathbb{P}(W^{*}=1)$ gives the internal validity of the weighted estimand with respect to target estimand $\mathbb{E}[Y(1)-Y(0)]$, the average treatment effect in the population from which the sample is drawn. It can also be understood as a measure of representativeness of the weighted estimand with respect to that population.
As earlier, we break down our results in two cases, the first being when $\tau_{0}$ is unrestricted.
4.1 Quantifying Internal Validity Uniformly in $\tau_{0}$
Without imposing any restrictions on the CATE function, except for the existence of second moments, the maximum value that $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ can achieve is given by the following theorem.
Theorem 4.1.
Let Assumption 3.1 hold and let $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. Then
$\sup_{W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})}\mathbb{P}(W^{*}=1\mid W_{0}=1)=\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\begin{cases}\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\sup(\text{supp}(a(X)\mid W_{0}=1))}&\text{ if }\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1\\ 0&\text{ if }\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1.\end{cases}$ (4.2)
Here we see that the maximum size of a subpopulation characterizing the estimand $\mu(a,\tau_{0})$ depends on $a(X)$ through two terms: the denominator of the estimand, $\mathbb{E}[a(X)\mid W_{0}=1]$, and the maximum value of $a(x)$ over the support of $X\mid W_{0}=1$. Note that $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ is invariant to the scale of $a(X)$, as is the estimand $\mu(a,\tau_{0})$. To understand the supremum’s role in this expression, let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$ and note that $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ is equivalent to writing
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}=\mu(\underline{w}^{*},\tau_{0})$ (4.3)
for all $\tau_{0}$. Equation (4.3) holding for all $\tau_{0}$ requires $\underline{w}^{*}(X)$ be exactly proportional to $a(X)$. While $a(X)$’s range is unconstrained, $\underline{w}^{*}(X)$ must lie in $[0,1]$ to be a valid conditional probability. Since we seek to maximize $\mathbb{P}(W^{*}=1\mid W_{0}=1)=\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$, we let $\underline{w}^{*}(X)$ be the largest multiple of $a(X)$ that lies in $[0,1]$ with probability 1, which is defined below:
$W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0}\qquad\text{ and }\qquad\underline{w}^{*}(X)=\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}.$
Here, $U\sim\text{Unif}(0,1)$ and $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$. This population places relatively more weight on units with larger values of $a(X)$. Specifically, the population $\{W^{*}=1\}$ contains a random subset of $\{W_{0}=1\}$ where the probability of inclusion is proportional to $a(X)$. Thus, units with larger values of $a(X)$ are more likely to be included in $W^{*}$. All units in $\{W_{0}=1\}$ with $X$ such that $a(X)=\sup(\text{supp}(a(X)\mid W_{0}=1))$ are included in $W^{*}$, whereas no units where $a(X)=0$ are included.
Given a data-generating process, this bound can be computed without any knowledge of the outcome’s conditional distribution and depends solely on weight function $a(\cdot)$ and the distribution of $X\mid W_{0}=1$.
Remark 4.1 (Defining the target estimand from the weighted estimand).
Suppose we consider $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ to be the target estimand, where $W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0}$ is the subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ uniformly in $\tau_{0}$. For example, under Example 2.1, this consists of a subpopulation where the probability of inclusion is proportional to $\operatorname{var}(D\mid X)$, the conditional variance of treatment. If this subpopulation is the target, it would be reasonable to infer that the measure of internal validity of the estimand $\mu(a,\tau_{0})$ is the maximum value of 1. Indeed, this is the case because the estimand can be written as
$\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
and $\{W^{*}=1\}$ is trivially the largest regular subpopulation of $\{W^{*}=1\}$. This illustrates how the internal validity of an estimand is entirely dependent on the target estimand. On the other hand, the representativeness of the weighted estimand, as measured by the largest value of $\mathbb{P}(W^{*}=1)$, is less than one unless $W^{*}=W_{0}$ almost surely, or that the weighted estimand actually equals the ATE. Theorem 4.4 below can also be applied to obtain this intuition.
We now consider a simple example to give further intuition for Theorem 4.1.
4.1.1 Illustrative Example: A Single Binary Covariate
Consider an estimand $\mu(a,\tau_{0})$ where $W_{0}=1$ almost surely, $a(X)\geq 0$, and where $X$ is binary with support $\text{supp}(X)=\{0,1\}$. Let $p_{x}=\mathbb{P}(X=x)>0$ for $x\in\{0,1\}$. As in Section 3.1, the weighted estimand can be written as a linear combination of the two CATEs:
$\mu(a,\tau_{0})$ $=\frac{a(0)p_{0}}{\mathbb{E}[a(X)]}\tau_{0}(0)+\frac{a(1)p_{1}}{\mathbb{E}[a(X)]}\tau_{0}(1)\equiv\omega_{0}\tau_{0}(0)+\omega_{1}\tau_{0}(1).$
Let $\mathbb{E}[Y(1)-Y(0)]=\text{ATE}$ be the target estimand, which can be written as
ATE $=p_{0}\tau_{0}(0)+p_{1}\tau_{0}(1).$
If $a(0)=a(1)$, the relative weights placed on $\{X=0\}$ and $\{X=1\}$ by the estimand are equal to $p_{0}/p_{1}$, the ratio of the weights placed by the ATE. Therefore, the estimand equals the ATE and thus clearly has the maximum degree of internal validity with respect to the ATE. Applying Theorem 4.1, we can directly see that
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)]}{\sup_{x\in\{0,1\}}a(x)}=\frac{a(0)p_{0}+a(1)p_{1}}{\max\{a(0),a(1)\}}=\frac{a(1)(p_{0}+p_{1})}{a(1)}=1$
when $a(0)=a(1)$.
However, when $a(0)\neq a(1)$, the estimand’s weights differ from $(p_{0},p_{1})$, the population weights for the two covariate cells. For concreteness, let $a(1)>a(0)$. In this case, the estimand overrepresents the population with $X=1$ and underrepresents the population with $X=0$, relative to the ATE. The largest subpopulation $\{W^{*}=1\}$ that places weights $(a(0),a(1))$ can be constructed by combining subsets of the subpopulations defined by $\{X=0\}$ and $\{X=1\}$. Specifically, let
$W^{*}=\mathbbm{1}(X=1)+\mathbbm{1}\left(U\leq\frac{a(0)}{a(1)},X=0\right),$
where $U\sim\text{Unif}(0,1)$ is independent from $(Y(1),Y(0),X)$. This is a subpopulation that contains all units with $X=1$ and fraction $a(0)/a(1)<1$ of units with $X=0$, selected uniformly at random. Its size is
$\mathbb{P}(W^{*}=1)=\mathbb{P}(X=1)+\mathbb{P}\left(U\leq\frac{a(0)}{a(1)},X=0\right)=p_{1}+\frac{a(0)}{a(1)}p_{0}<1,$
which is the same as $\frac{\mathbb{E}[a(X)]}{\sup_{x\in\{0,1\}}a(x)}=\frac{a(1)p_{1}+a(0)p_{0}}{a(1)}$ obtainable from Theorem 4.1. The average treatment effect in this subpopulation is given by
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mathbb{P}(X=0\mid W^{*}=1)\tau_{0}(0)+\mathbb{P}(X=1\mid W^{*}=1)\tau_{0}(1)$
$=\frac{\frac{a(0)}{a(1)}p_{0}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}\tau_{0}(0)+\frac{p_{1}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}\tau_{0}(1)$
$=\frac{a(0)p_{0}}{a(0)p_{0}+a(1)p_{1}}\tau_{0}(0)+\frac{a(1)p_{1}}{a(0)p_{0}+a(1)p_{1}}\tau_{0}(1)$
$=\mu(a,\tau_{0}).$
The relative weights placed on $\{X=0\}$ and $\{X=1\}$ in subpopulation $\{W^{*}=1\}$ are given by
$\frac{\mathbb{P}(X=0\mid W^{*}=1)}{\mathbb{P}(X=1\mid W^{*}=1)}$ $=\frac{\frac{a(0)}{a(1)}p_{0}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}/\frac{p_{1}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}=\frac{\omega_{0}}{\omega_{1}},$
matching the ratio of the weights on $\{X=0\}$ and $\{X=1\}$ assigned by the estimand. The subpopulation $\{W^{*}=1\}$ cannot expand while preserving this ratio since it already encompasses all units with $X=1$. Therefore, $W^{*}$ is the largest subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$.
4.1.2 Partial Identification of Average Effects
This subpopulation size can be used to bound target estimand $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ using only the weighted estimand and bounds on $Y$. Given knowledge of the weighted estimand $\mu(a,\tau_{0})$ and its representativeness, given by $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$, we can decompose the target estimand as
$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\cdot\mathbb{P}(W^{*}=1\mid W_{0}=1)$
$\qquad+\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]\cdot(1-\mathbb{P}(W^{*}=1\mid W_{0}=1))$
$=\mu(a,\tau_{0})\cdot\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]\cdot(1-\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$
for $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$. Therefore, $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is point-identified given knowledge of $(\mu(a,\tau_{0}),\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$ if $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=1$, and partially identified if $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})<1$. If we have knowledge of bounds for the treatment effect $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]$, e.g. from the support of the potential outcomes, we can obtain bounds on $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$. For example, if $\text{supp}(Y(1)-Y(0))=[B_{\ell},B_{u}]$, bounds for the target estimand are given by
$\left[(\mu(a,\tau_{0})-B_{\ell})\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+B_{\ell},(\mu(a,\tau_{0})-B_{u})\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+B_{u}\right].$
The width of these bounds is $(B_{u}-B_{\ell})(1-\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$. Hence for fixed $(B_{\ell},B_{u})$, this width decreases linearly with $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ and values of $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ close to 1, or high degrees of internal validity, lead to narrow bounds.
4.2 Quantifying Internal Validity Given $\tau_{0}$
As in the previous subsection, we can also ask how internally valid a weighted estimand can be, given knowledge of the CATE function. In this case, the object of interest is
$\overline{P}(a,W_{0};\{\tau_{0}\})=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})}\mathbb{P}(W^{*}=1\mid W_{0}=1).$ (4.4)
where $\tau_{0}$ is a given CATE function. Since $\tau_{0}$ is known, the condition $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ can be written as
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
or, equivalently,
$\mathbb{E}[(\tau_{0}(X)-\mu(a,\tau_{0}))\underline{w}^{*}(X)\mid W_{0}=1]=0,$ (4.5)
where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$. Equation (4.5) is a linear constraint on the conditional probability of being in subpopulation $W^{*}$. Additionally, the objective function $\mathbb{P}(W^{*}=1\mid W_{0}=1)=\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ is linear in $\underline{w}^{*}$. Thus, the optimization in (4.4) can be cast as a linear program. To see this, consider as an example the case where $W_{0}=1$ almost surely and where $X$ is discrete with finite support, i.e. $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$. Let $f_{k}=\mathbb{P}(W^{*}=1,X=x_{k})$ denote the probability of being in subpopulation $W^{*}$ and having covariate value $x_{k}$. Note that $f_{k}\in[0,p_{k}]$ where $p_{k}\equiv\mathbb{P}(X=x_{k})$.
We can write the above optimization problem as
$\max_{(f_{1},\ldots,f_{K})\geq\textbf{0}}\sum_{k=1}^{K}f_{k}\qquad\text{ such that }\qquad$ $f_{k}\leq p_{k}\text{ for }k\in\{1,\ldots,K\},$
$\sum_{k=1}^{K}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))f_{k}=0,$
a finite-dimensional linear program. This program has a feasible solution if $\tau_{0}(x_{k})-\mu(a,\tau_{0})$ is not strictly positive or strictly negative for all $k$, which is precisely stated in the condition for Theorem 3.2. While there exist many methods for solving linear programs, the value function can be obtained through an algorithm that is simple to describe analytically.
Let $\textbf{f}=(f_{1},\ldots,f_{K})$, $\textbf{p}=(p_{1},\ldots,p_{K})$, and $\textbf{t}_{\mu}=(\tau_{0}(x_{1})-\mu(a,\tau_{0}),\ldots,\tau_{0}(x_{K})-\mu(a,\tau_{0}))$. Without loss of generality, assume that $\tau_{0}(x_{1})-\mu(a,\tau_{0})\leq\tau_{0}(x_{2})-\mu(a,\tau_{0})\leq\cdots\leq\tau_{0}(x_{K})-\mu(a,\tau_{0})$.
  1. 1.
[…]
  2. 2.
    If $\textbf{t}_{\mu}^{\prime}\textbf{f}=0$, end the algorithm and report $\sum_{k=1}^{K}f_{k}$.
  3. 3.
    If $\textbf{t}_{\mu}^{\prime}\textbf{f}\neq 0$:
    1. (a)
      If $\textbf{t}_{\mu}^{\prime}\textbf{f}>0$, let $k^{*}=\max\{k\in\{1,\ldots,K\}:f_{k}=p_{k}\}$ and set $f_{k^{*}}=\max\left\{0,\frac{-\sum_{k=1}^{k^{*}-1}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))p_{k}}{\tau_{0}(x_{k^{*}})-\mu(a,\tau_{0})}\right\}$.
    2. (b)
      If $\textbf{t}_{\mu}^{\prime}\textbf{f}<0$, let $k^{*}=\min\{k\in\{1,\ldots,K\}:f_{k}=p_{k}\}$ and set $f_{k^{*}}=\max\left\{0,\frac{-\sum_{k=k^{*}+1}^{K}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))p_{k}}{\tau_{0}(x_{k^{*}})-\mu(a,\tau_{0})}\right\}$.
  4. 4.
    Go to step 2.
When $\mu(a,\tau_{0})$ exceeds $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, this algorithm reduces the weights associated with largest CATEs until $\mu(a,\tau_{0})$ equals $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for some subpopulation. When $\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the same procedure is instead applied to the smallest CATEs. The support assumption of Theorem 3.2 guarantees that this algorithm ends.
When $X$ is not discretely supported, the problem can still be cast as a linear program, but its dimension may be infinite, which generates difficulties in implementation. However, we show this program has an analytical solution for the size of the subpopulation of interest, $\mathbb{P}(W^{*}=1\mid W_{0}=1)$. The following theorem gives its expression and is valid when vector $X$ has an arbitrary kind of distribution, with continuous, discrete, and mixed components, as is often the case in empirical applications.
Theorem 4.2.
Let Assumption 3.1 hold. Let $T_{\mu}=\tau_{0}(X)-\mu(a,\tau_{0})$. If $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W_{0})$,
$\overline{P}(a,W_{0};\{\tau_{0}\})$
=$=\begin{cases}\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}\\ \qquad\text{ where }\alpha^{+}=\inf\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]\geq 0\}\text{ if }\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]\\ \mathbb{P}(T_{\mu}\geq\alpha^{-}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}\\ \qquad\text{ where }\alpha^{-}=\sup\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]\leq 0\}\text{ if }\mu(a,\tau_{0})>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]\\ 1\qquad\text{ if }\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1].\end{cases}$ (4.6)
If $\mu(a,\tau_{0})\notin\mathcal{S}(\tau_{0};W_{0})$, then $\overline{P}(a,W_{0};\{\tau_{0}\})=0$.
The computation of these bounds can be done using a linear programming algorithm when $X$ is discrete, or through plug-in estimators of the terms in equation (4.2) regardless of the nature of the support of $X$.
In this setting, the value $\overline{P}(a,W_{0};\{\tau_{0}\})$ is larger when the truncated subpopulations are smaller. In particular, this is the case when there are a few units with extreme values of $\tau_{0}$ whose removal has a large impact on the estimand, but a small impact on the share of the population.
4.2.1 Illustrative Example: A Single Continuous Covariate
To illustrate the previous theorem, let $X$ be a continuously distributed covariate with support $[x_{L},x_{U}]$. Let $W_{0}=1$ almost surely, and let $\mu\equiv\mu(a,\tau_{0})$ denote the estimand. Also, for simplicity assume that $\tau_{0}(x)$ is increasing in $x$ and that $\tau_{0}(x_{L})<\mu<\tau_{0}(x_{U})$. Without loss of generality, let $\mathbb{E}[Y(1)-Y(0)]\geq\mu$. If $\mathbb{E}[Y(1)-Y(0)]=\mu$, then the estimand is perfectly representative of the population since it equals the average treatment effect over it. Now consider the case where $\mathbb{E}[Y(1)-Y(0)]>\mu$. In this case, the estimand is larger than the ATE so it is not representative of the entire population.
We are seeking the largest subpopulation $\{W^{*}=1\}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu$. Equivalently, we can seek the smallest subpopulation $\{W^{-}=1\}$ such that, when it is removed from the original population, the average treatment effect equals $\mu$. The subpopulation $\{W^{-}=1\}$ is related to $\{W^{*}=1\}$ via $W^{-}+W^{*}=1$. We will search for $W^{-}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=0]=\mu$ and $\mathbb{P}(W^{-}=1)$ is minimized.
Since the ATE exceeds $\mu$, we must have that $\mathbb{E}[\tau_{0}(X)\mid W^{-}=1]>\mu$. For a given subpopulation of size $\mathbb{P}(W^{-}=1)$, removing the subpopulation with the largest values of $\tau_{0}(x)$ yields the largest decrease in $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=0]$. Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})$ is obtained by removing a subpopulation of the kind $W^{-}(\delta)=\mathbbm{1}(\tau_{0}(X)\geq\delta)$ for a given threshold $\delta$. This subpopulation consists of units whose CATEs exceed a threshold $\delta$. This threshold is determined by the constraint $\mathbb{E}[\tau_{0}(X)\mid W^{-}(\delta)=0]=\mu$, or
$\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq\delta]=\mu.$ (4.7)
This constraint states that once $\{W^{-}(\delta)=1\}$ is removed, the average treatment effect for the remaining units equals the weighted estimand. Since we assumed $\tau_{0}$ is increasing, $W^{*}$ corresponds to all units with covariate values below threshold $\delta=\tau_{0}^{-1}(\delta)$. Therefore,
$W^{*}=\mathbbm{1}(X\leq\tau_{0}^{-1}(\delta)),$
where $\delta$ is the unique solution to (4.7). How large $\mathbb{P}(W^{*}=1)$ is depends on the size of this fraction of removed units. A smaller subpopulation needs to be removed when $\tau_{0}(X)$ has a long right tail, and a larger subpopulation needs to be removed when the distribution of $\tau_{0}(X)$ in this right tail is shorter.
4.3 Generalizing to Subpopulations
The results from the previous two sections can be generalized to cases where we hold the average effect for a subset of $W_{0}$ as the object of interest. Specifically, let $W^{\prime}$ be a regular subpopulation of $W_{0}$ and $\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]$ be the target parameter. In analogy to the previous subsections, we ask (i) when does there exist a regular subpopulation $W^{*}$ of $W^{\prime}$ such that $\mu(a,\tau_{0})$ can be written as a causal estimand and (ii) how representative of $W^{\prime}$ is this estimand.
For concreteness, in the IV model of Section 2.2, we may be interested in the average treatment effect for compliers/defiers with covariates in given set $\mathcal{X}_{0}$: $\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X\in\mathcal{X}_{0}]$. In this case $W_{0}=\mathbbm{1}(D(1)>D(0))$ and $W^{\prime}=\mathbbm{1}(D(1)>D(0),X\in\mathcal{X}_{0})$, a regular subpopulation of $W_{0}$.
We first present a result showing conditions for the existence of such population $W^{*}$ in the case where $\tau_{0}$ is unrestricted, and when $\tau_{0}$ is fixed.
Theorem 4.3.
Let Assumption 3.1 hold and let $W^{\prime}$ be a regular subpopulation of $W_{0}$.
  1. 1.
    (Uniformly in $\tau_{0}$) Let $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$. [^5] We use the conventions $0/0=1$ and $a/0=\infty$ when $a>0$. Then, there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ if and only if $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$.
  2. 2.
    (Given $\tau_{0}$) There exists $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$ if and only if $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W^{\prime})$.
Note that letting $W^{\prime}=W_{0}$ yields Theorems 3.1 and 3.2 as special cases since $W_{0}$ is a regular subpopulation of $W_{0}$ by definition.
The condition $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ rules out uniform causal representations if there are covariate values for which $a(x)$ is strictly positive but are not represented in $W^{\prime}$, i.e., $\mathbb{P}(W^{\prime}=1\mid X=x,W_{0}=1)=0$ when $a(x)>0$. For example, if $W^{\prime}$ is a subpopulation of $W_{0}$ defined by a subset of covariate values such that $W^{\prime}=\mathbbm{1}(X\in\mathcal{X}_{0})\cdot W_{0}$, then $W^{\prime}$ is a regular subpopulation but will fail the above inequality if $a(X)>0$ for $X\in\text{supp}(X\mid W_{0}=1)\setminus\mathcal{X}_{0}$. In the case where $\tau_{0}$ is known, the existence of a causal subpopulation of $W^{\prime}$ is equivalent to the estimand being in the convex hull of $\text{supp}(\tau_{0}(X)\mid W^{\prime}=1)$.
Regarding the representativeness of a weighted estimand, we generalize Theorems 4.1 and 4.2 and ask how large $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ can be given that $W^{*}$ is a regular subpopulation of $W^{\prime}$, which is itself a regular subpopulation of $W_{0}$. In what follows, we let $\underline{w}^{\prime}(X)=\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$.
Theorem 4.4.
Let Assumption 3.1 hold and let $W^{\prime}$ be a regular subpopulation of $W_{0}$.
  1. 1.
    (Uniformly in $\tau_{0}$) Let $\sup\left(\text{supp}\left(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1\right)\right)<\infty$. If $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$, then
    $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})=\mathbb{E}[a(X)\mid W_{0}=1]\cdot\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}\cdot\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right).$ (4.8)
    If $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)<1$, $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})=0$.
  2. 2.
    (Given $\tau_{0}$) If $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W^{\prime})$, then
    $\overline{P}(a,W^{\prime};\{\tau_{0}\})$
    =$=\begin{cases}\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W^{\prime}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W^{\prime}=1]}{\alpha^{+}}\\ \qquad\text{ where }\alpha^{+}=\inf\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W^{\prime}=1]\geq 0\}\text{ if }\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]\\ \mathbb{P}(T_{\mu}\geq\alpha^{-}\mid W^{\prime}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W^{\prime}=1]}{\alpha^{-}}\\ \qquad\text{ where }\alpha^{-}=\sup\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]\leq 0\}\text{ if }\mu(a,\tau_{0})>\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]\\ 1\qquad\text{if }\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1].\end{cases}$ (4.9)
    If $\mu(a,\tau_{0})\notin\mathcal{S}(\tau_{0};W^{\prime})$, then $\overline{P}(a,W^{\prime};\{\tau_{0}\})=0$.
We can also use this result to obtain bounds on $\mathbb{P}(W^{*}=1)$, or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$. This can be done by noting that $W^{*}$ is a subpopulation of both $W^{\prime}$ and $W_{0}$, so
$\mathbb{P}(W^{*}=1)$ $=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\cdot\mathbb{P}(W^{\prime}=1)\quad\text{ and }\quad\mathbb{P}(W^{*}=1\mid W_{0}=1)=\frac{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\cdot\mathbb{P}(W^{\prime}=1)}{\mathbb{P}(W_{0}=1)}.$
Thus bounds on these probabilities are trivially obtained from the bound on $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ from Theorem 4.4 and knowledge of $(\mathbb{P}(W^{\prime}=1),\mathbb{P}(W_{0}=1))$.
4.4 Examples Revisited
For simplicity, we assume that $\sup(\text{supp}(a(X)\mid W_{0}=1))=\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)$ for the remainder of the paper. This condition is satisfied for example when $a(\cdot)$ is continuous.
4.4.1 Unconfoundedness
Recall the expression for the coefficient on $D$ in a population regression of $Y$ on $(1,D,X)$:
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[p(X)(1-p(X))\tau_{0}(X)]}{\mathbb{E}[p(X)(1-p(X))]}.$
Suppose the target estimand is the average treatment effect, i.e. $W_{0}=1$ almost surely.
By Theorem 3.1, there exists a regular subpopulation $W^{*}$ such that $\beta_{\text{OLS}}$ equals the average treatment effect over $W^{*}$ since the weight function $a(X)\equiv p(X)(1-p(X))$ is non-negative. By Theorem 4.1, the upper bound on the size of subpopulation $W^{*}$ is given by
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[p(X)(1-p(X))]}{\sup_{x\in\text{supp}(X)}p(x)(1-p(x))}.$
The corresponding subpopulation $W^{*}$ satisfies
$\mathbb{P}(W^{*}=1\mid X)$ $=\frac{p(X)(1-p(X))}{\sup_{x\in\text{supp}(X)}p(x)(1-p(x))}.$
This is a subpopulation that places greater weights on units which have a larger variation in treatment given their covariate values. The size of this subpopulation is largest when $\operatorname{var}(D\mid X)=p(X)(1-p(X))$ is constant, in which case $\mathbb{P}(W^{*}=1\mid X)=1$. This is the case if and only if $p(X)$ has support equal to $\{b,1-b\}$ for some $b\in[0,1]$. This is implied by $D\perp\!\!\!\perp X$, or random assignment. It can also be achieved if there exists a partition of $\text{supp}(X)$ where $\mathbb{P}(D=1\mid X)=b$ on one element and $\mathbb{P}(D=1\mid X)=1-b$ on its complement. Whenever $\operatorname{var}(p(X)(1-p(X)))>0$, $\{W^{*}=1\}$ will be a strict subpopulation.
The size of this subpopulation is the expectation of $\operatorname{var}(D\mid X)$ divided by its maximum value. There are a few ways this expression can be further simplified or bounded. Its numerator is bounded above by $\operatorname{var}(D)=\mathbb{P}(D=1)\cdot\mathbb{P}(D=0)$, which is particularly simple to estimate. As for the denominator, it is a non-smooth functional of $p(\cdot)$ which will lead to nonstandard inference. However, if $X$ is continuously distributed, it may be likely that $p(X)$ is continuously distributed as well. In this case, it is likely that $1/2\in\text{supp}(p(X))$. If this is the case, $\sup_{x\in\text{supp}(X)}p(x)(1-p(x))=1/4$. Combining these approximations yields that
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $\leq 4\cdot\mathbb{P}(D=1)\cdot\mathbb{P}(D=0),$
when the support of $p(X)$ includes 1/2. This bound is trivial when $\mathbb{P}(D=1)=1/2$, but is informative when the treatment probability is close to 0 or 1. For example, if $\mathbb{P}(D=1)=0.1$, the OLS estimand cannot causally represent more than 36% of the population.
When $1/2\in\text{supp}(p(X))$, we can also compute bounds on the ATE derived from the OLS estimand, bounds on the support of $(Y(1),Y(0))$, and our measure of representativeness $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$. Following Remark 4.1.2, bounds on the ATE are given by
$\left[(\beta_{\text{OLS}}-B_{\ell})4\mathbb{E}[\operatorname{var}(D\mid X)]+B_{\ell},(\beta_{\text{OLS}}-B_{u})4\mathbb{E}[\operatorname{var}(D\mid X)]+B_{u}\right].$
Estimating these bounds requires the estimation of one additional quantity beyond the OLS estimator, which is the expectation of $\operatorname{var}(D\mid X)$. Of course, these bounds’ width depends crucially on $B_{u}-B_{\ell}$, or the width of the support for unit-level treatment effects.
Alternatively, we can assess the representativeness of $\beta_{\text{OLS}}$ with respect to an alternative estimand such as $\mathbb{E}[Y(1)-Y(0)\mid D=1]$, the average treatment effect on the treated. In this case, we can write
$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]},$
where $w_{0}(X)=\mathbb{P}(D=1\mid X)=p(X)$, the propensity score, and $a(X)=\mathbb{P}(D=0\mid X)=1-p(X)$. Applying Theorem 4.1 (or 4.4) yields that
$\overline{P}(a,D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[1-p(X)\mid D=1]}{\sup_{x\in\text{supp}(X\mid D=1)}(1-p(x))}=\frac{\mathbb{E}[p(X)(1-p(X))]}{\mathbb{P}(D=1)\cdot\sup_{x\in\text{supp}(X\mid D=1)}(1-p(x))}$
is the largest value that $\mathbb{P}(W^{*}=1\mid D=1)$ can take. Once again, this bound depends only on the propensity score and the distribution of $X$. The same bound is obtained by using Theorem 4.4 instead and setting $W_{0}=1$ and $W^{\prime}=D$. This subpopulation satisfies
$\mathbb{P}(W^{*}=1\mid X,D=1)=\frac{1-p(X)}{1-\inf_{x\in\text{supp}(X\mid D=1)}p(X)}$
so units with smaller propensity scores are more likely to be included in $W^{*}$, given that they are treated. This probability is maximized at 1 when $p(X)$ is constant, or if $D\perp\!\!\!\perp X$. In this case, $\mathbb{P}(W^{*}=1\mid D=1)=1$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(D=1)$.
If $p(X)$ takes values close to 0, this bound equals
$\overline{P}(a,D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[p(X)(1-p(X))]}{\mathbb{P}(D=1)}\leq\frac{\mathbb{P}(D=1)\cdot\mathbb{P}(D=0)}{\mathbb{P}(D=1)}=\mathbb{P}(D=0).$
This suggests that the OLS estimand is more representative of the ATT when the fraction of untreated units is larger.
We can also assess the representativeness of $\beta_{\text{OLS}}$ given $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$. For simplicity, assume that $X$ has a continuous distribution and, without loss of generality, assume that $\beta_{\text{OLS}}>\text{ATE}$. Then, using Theorem 4.2, we obtain
$\overline{P}(a,W_{0};\{\tau_{0}\})$ $=\mathbb{P}(\tau_{0}(X)\leq b^{*}),$
where $b^{*}$ satisfies $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq b^{*}]=\beta_{\text{OLS}}$. The quantity $\overline{P}(a,W_{0};\{\tau_{0}\})$ is largest when the least amount of trimming needs to be applied. This is the case when the trimmed values are largest, or when $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\geq b]$ is large when $b$ is near the maximum of $\text{supp}(\tau_{0}(X))$. More concretely, if some covariate values in $\text{supp}(X)$ have large CATEs, then considering a subpopulation that removes only a small subset of $\text{supp}(X)$ — those corresponding to large CATE values — can allow $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq b^{*}]$ and $\beta_{\text{OLS}}$ to be equal.
4.4.2 Instrumental Variables
Consider again the setting of Section 2.2. The estimand $\beta_{\text{2SLS}}$ can be characterized as $\mu(a_{\text{2SLS}},\tau_{0})$ for $a_{\text{2SLS}}(X)=\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)\neq D(0)\mid X)$, $W_{0}=\mathbbm{1}(D(1)\neq D(0))$, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$.
Since $a_{\text{2SLS}}(X)\geq 0$, there exists a subpopulation of $\{W_{0}=1\}$ such that $\beta_{\text{2SLS}}$ is an average treatment effect over that subpopulation. The maximum size of that subpopulation is given by
$\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{2SLS}}(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}$
$=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)\neq D(0)\mid X)^{2}]}{\mathbb{P}(D(1)\neq D(0))\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\operatorname{var}(Z\mid X=x)\cdot\mathbb{P}(D(1)\neq D(0)\mid X=x)}.$
Under Assumptions 2.2 and 2.4, $\mathbb{P}(D(1)\neq D(0)\mid X)$ is identified from the distribution of $(D,X,Z)$ as $w_{0}(X)\equiv|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|$, which is the fraction of compliers or defiers, i.e. units for which $D(1)\neq D(0)$, among those with covariates $X$.
The maximum value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is obtained when $\operatorname{var}(Z\mid X)\cdot|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|=|\operatorname{cov}(Z,D\mid X)|$ does not depend on $X$. This occurs, for example, when the instrument and the fraction of units for which $D(1)\neq D(0)$ are independent of $X$. In this case, we have that $\beta_{\text{2SLS}}=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the average effect of treatment in the complier/defier subpopulation.
The IV estimand has the same $W_{0}$, $w_{0}(X)$, but has $a_{\text{IV}}(X)=\text{sign}(\mathbb{P}(D(1)\geq D(0)\mid X)-\mathbb{P}(D(1)\leq D(0)\mid X))\cdot\operatorname{var}(Z\mid X)\equiv c(X)\operatorname{var}(Z\mid X)$ instead.
First, we notice that if $\mathbb{P}(c(X)=-1)>0$, then weights are negative with positive probability and there does not exist a causal representation for the estimand $\beta_{\text{IV}}$ that is uniform in $\tau_{0}\in\mathcal{T}_{\text{all}}$. However, there will exist a causal representation given $\tau_{0}$ if the support condition of Theorem 3.2 holds, in that $\beta_{\text{IV}}$ lies in the support of $\tau_{0}(X)$.
If we assume strong monotonicity (Assumption 2.3), then $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)\geq 0$ and
$\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{IV}}(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}$
$=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)>D(0)\mid X)]}{\mathbb{P}(D(1)>D(0))\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\operatorname{var}(Z\mid X=x)}.$
Here the representativeness of the IV estimand is maximized when $\operatorname{var}(Z\mid X)$ is constant, which occurs when $Z$ is independent of $X$. In this case, $\beta_{\text{IV}}$ equals LATE. The quantities $\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ and $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ are not ranked uniformly in the distributions of $(D(1),D(0),X,Z)$ as there are data-generating processes that make each of these two quantities larger than the other. For example, if $\operatorname{var}(Z\mid X)$ is constant but $\mathbb{P}(D(1)\neq D(0)\mid X)$ is not, then $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})<\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$. This scenario is plausible if $Z$ is randomly assigned and $X$ is a vector of pre-assignment characteristics. This inequality is reversed if $a_{\text{2SLS}}(X)=|\operatorname{cov}(Z,D\mid X)|$ is constant but $\mathbb{P}(D(1)\neq D(0)\mid X)$ is not. They are equally representative when $\mathbb{P}(D(1)\neq D(0)\mid X)$ is constant. In this case, the estimands are equal, so this is not unexpected.
4.4.3 Differences-in-Differences
We now consider the weights obtained in Proposition 2.5 under its assumptions. These weights are nonnegative and therefore Theorem 3.1 guarantees the existence of a causal representation for $\beta_{\text{TWFE}}$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$. Using Theorem 4.1, the level of internal validity of $\beta_{\text{TWFE}}$ relative to $\mathbb{E}[Y(1)-Y(0)\mid D=1]$ is given by
$\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{B}(G)w_{0}(G)]}{\mathbb{E}[w_{0}(G)]\cdot\sup_{g\in\text{supp}(G\mid D=1)}a_{B}(g)}$
$=\frac{\sum_{g=2}^{T}\operatorname{var}(D\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\cdot\mathbb{P}(G=g)}{\mathbb{P}(D=1)\cdot\sup_{g\in\{2,\ldots,T\}}\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))}.$
Due to the absorbing nature of treatment in our setting, all expressions involving the distribution of $D$ given $P$ or $G$ can be derived as a function of the marginal distribution of $G$. Therefore, $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ depends only on $\{\mathbb{P}(G=g)\}_{g\in\{2,\ldots,T\}}$.
To give some intuition, consider the case where $T=3$ and therefore $G\in\{2,3,+\infty\}$. In this case, $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ will simply be a function of $(\mathbb{P}(G=2),\mathbb{P}(G=3))$.
Calculations yield that
$\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ =$=\begin{cases}\frac{2\mathbb{P}(G=2)+\omega\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)}&\text{ if }\omega<1\\ \frac{\omega^{-1}2\mathbb{P}(G=2)+\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)}&\text{ if }\omega>1\\ 1&\text{ if }\omega=1,\end{cases}$
where
$\omega$ $=\frac{4-2\mathbb{P}(G=2)-4\mathbb{P}(G=3)}{2-2\mathbb{P}(G=2)-\mathbb{P}(G=3)}=\frac{a_{B}(2)}{a_{B}(3)}.$
Therefore, the TWFE estimand is perfectly representative of the ATT if and only if $\omega=1$.
With some algebra, we can see that $\omega<1$ if and only if $\mathbb{P}(G=3)>2/3$, while $\omega>1$ if and only if $\mathbb{P}(G=3)<2/3$. We can also see that $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ is equal to 1 when $\mathbb{P}(G=3)=2/3$, and that the representativeness of the TWFE estimand declines as $|\mathbb{P}(G=3)-2/3|$ increases. This is due to weight function $a_{B}(g)$ being constant in $g$ if and only if $\mathbb{P}(G=3)$ equals 2/3. Whenever the weight function is constant, we have that $\mu(a,\tau_{0})=\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]=\mathbb{E}[\tau_{0}(X)\mid W_{0}=1]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, and therefore the weighted estimand equals the average effect over population $\{W_{0}=1\}$.
5 Discussion of Estimation and Inference
We now briefly discuss procedures to estimate $\overline{P}(a,W_{0};\mathcal{T})$.
First consider the case when $\mathcal{T}=\mathcal{T}_{\text{all}}$. Then, we seek to estimate
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}=\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}.$ (5.1)
Suppose we observe a random sample of size $n$, $\{(W_{i},X_{i})\}_{i=1}^{n}$, where $W_{i}$ are a set of variables that allow the estimation of $a(\cdot)$ and $w_{0}(\cdot)$. For example, under unconfoundedness we can let $W_{i}=(Y_{i},D_{i})$ or $W_{i}=D_{i}$ since only the distribution of $(D,X)$ is needed to identify $a(\cdot)$. In our instrumental variables examples, one can let $W_{i}=(Z_{i},D_{i})$.
Assuming the existence of estimators for $a(\cdot)$ and $w_{0}(\cdot)$, we can propose the following analog estimator:
$\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\widehat{a}(x)}.$
This estimator requires knowledge of the support of $X$ given $W_{0}=1$, but can be implemented by taking the maximum over the empirical distribution of $X\mid W_{0}=1$. Consistency of $\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$ for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ is achieved when
[…]
and $\widehat{a}(x)$ is uniformly consistent for $a(x)$ over $x\in\text{supp}(X\mid W_{0}=1)$. These conditions can be achieved under many standard estimators of $a(\cdot)$ and $w_{0}(\cdot)$. For example, when $a(X)$ and $w_{0}(X)$ consist of continuous functions of conditional means, as is the case when $a(X)=\operatorname{var}(D\mid X)$ or, for the IV estimator, $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ and $w_{0}(X)=|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|$, consistent estimators of these conditional means will result in the consistency of $\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$. Cell-by-cell estimators can be used when $X_{i}$ is discrete, or many nonparametric estimators when $X_{i}$ contains continuous components. See Hansen (2022) for a recent textbook treatment of such methods.
Inference will generally be nonstandard. This is because the asymptotic behavior of the denominator will usually be non-Gaussian due to the lack of Hadamard differentiability of the supremum operator. This lack of differentiability will invalidate standard bootstrap inference: see Fang and Santos (2019). However, the supremum is a Hadamard directionally differentiable functional and nonstandard bootstrap approaches, such as those proposed in Hong and Li (2018) and Fang and Santos (2019), can be used.
We can also view these inferential problems through the lens of intersection or union bounds. For example, we can write
$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\inf_{x\in\text{supp}(X\mid W_{0}=1)}\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot a(x)}\equiv\inf_{x\in\text{supp}(X\mid W_{0}=1)}\overline{P}(x).$
Computing a one-sided confidence interval for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ of the kind $[0,\widehat{\overline{P}}^{+}]$ can be cast as doing inference on intersection bounds. Chernozhukov, Lee, and Rosen (2013) offer methods for such problems. Equivalently, the computation of a one-sided confidence interval $[\widehat{\overline{P}}^{-},1]$ is related to inferential questions in union bounds: see Bei (2024). We leave all details for future work.
Estimation and inference of $\overline{P}(a,W_{0};\{\tau_{0}\})$ is related to the question of estimation and inference in linear programs with estimated constraints. See Fang, Santos, Shaikh, and Torgovitsky (2023) and Cho and Russell (2024) for recent advances on this topic.
6 Empirical Application
In this section, we implement the proposed tools in an application to the effects of unilateral divorce laws in the U.S. on female suicide, as in Stevenson and Wolfers (2006). Between 1969 and 1985, 37 states (including the District of Columbia) reformed their law by enabling each spouse to seek divorce without the other spouse’s consent. Stevenson and Wolfers (2006) argue that these “unilateral” or “no-fault” divorce laws reduced female suicide, domestic violence, and spousal homicide. The results on female suicide are also replicated by Goodman-Bacon (2021), whose analysis we follow here.
Our sample consists of 41 states observed over the 1964–1996 period. The outcome of interest is the state- and year-specific female suicide rate (per million women), as computed by the National Center for Health Statistics (NCHS). The treatment is whether the state allowed unilateral divorce in a given year. Following Goodman-Bacon (2021), our sample omits Alaska and Hawaii. Additionally, we omit Louisiana, Maryland, North Carolina, Oklahoma, Utah, Vermont, Virginia, and West Virginia. These eight states (and Alaska) had unilateral divorce laws preceding 1964 and are therefore always treated within our timeframe.
{adjustwidth}
-1in-1in A. Estimates of the effects of unilateral divorce laws TWFE ATT Callaway and Sant’Anna Wooldridge –0.604 –10.220 –5.530 (2.622) (3.086) (3.650) B. Representativeness of the TWFE estimand based on Proposition 2.4 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216 C. Representativeness of the TWFE estimand based on Proposition 2.5 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
Table 1: Representativeness of the TWFE estimand of the effects of unilateral divorce laws
Panel A of Table 1 reports our baseline estimates of the average effects of unilateral divorce laws on female suicide. After we drop the eight always-treated states, the TWFE estimate, –0.604, becomes much smaller in absolute value than the corresponding estimate in Goodman-Bacon (2021), –3.080. Unlike that estimate, ours is also statistically insignificant, with $p$-value = 0.819.
The conclusion changes, however, when we explicitly target the average treatment effect on the treated (ATT), that is, the average effect for the largest subpopulation for which such an effect is identified. Using the approach of Wooldridge (2021), we obtain an estimate of –5.530 with a $p$-value of 0.138. The approach of Callaway and Sant’Anna (2021) produces an estimate of –10.220 and a $p$-value of 0.001. As is clear, the choice of an estimation approach has fundamental implications for our conclusions in this study.
While the TWFE estimate and both estimates of the ATT are clearly different, this paper focuses on another implication of the nonuniformity of the TWFE weight function. We ask: How representative of the underlying population is the TWFE estimand? What is the internal validity of this estimand if we are interested in the treated subpopulation? Panel B of Table 1 reports our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)=\mathbb{P}(W^{*}=1\mid W_{0}=1)$, based on the representation of the TWFE estimand in de Chaisemartin and D’Haultfœuille (2020), restated in our Proposition 2.4. First, because the weights on some of the group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, ^$\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$. Second, when we estimate the CATE function and use these estimates in constructing the bounds, we conclude that the TWFE estimand corresponds to the average treatment effect for at most 62.16% of the treated units or 38.73% of the entire population.
Panel C of Table 1 revisits these questions on the basis of the representation of the TWFE estimand in Proposition 2.5, based on Goodman-Bacon (2021). Here, we assume that group-level average treatment effects are constant over time, which eliminates the problem of negative weights. Indeed, we now conclude that the TWFE estimand has a causal interpretation uniformly in $\tau_{0}$, even if it is still not particularly representative of the underlying population or the treated subpopulation. Our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$ are equal to 14.00% and 22.46%, respectively. When we use the estimated CATE function in constructing the bounds, these estimates increase to 48.02% and 77.07%. This is obviously much more than our initial estimate of 0, but still substantially less than 1, guaranteed in the case of $\mathbb{P}(W^{*}=1\mid D=1)$ when using the estimation approaches in Callaway and Sant’Anna (2021), Wooldridge (2021), and other recent papers.
7 Conclusion
In this paper, we studied the representativeness and internal validity of a class of weighted estimands, which includes the popular OLS, 2SLS, and TWFE estimands in additive linear models. We examined the conditions under which such estimands can be written as the average treatment effect for some (possibly latent) subpopulation. In our main result, we derived the sharp upper bound on the size of that subpopulation. We consider this bound to be a valuable diagnostic for empirical research. When a given estimand can be shown to correspond to the average treatment effect for a large subset of the population of interest, we say its internal validity is high. In an application to the effects of unilateral divorce laws in the U.S. on female suicide, as in Stevenson and Wolfers (2006) and Goodman-Bacon (2021), we showed that the TWFE estimand has a low degree of internal validity (assuming that the treated subpopulation is of interest), even when we assume away the existence of negative weights. Because this result is then necessarily driven by the nonuniformity of the TWFE weight function, it corroborates the negative view of both negative and nonuniform weights in Callaway, Goodman-Bacon, and Sant’Anna (2024).
Appendix
This appendix is organized as follows: Appendix A contains proofs for Section 2, Appendix B contains proofs for Section 3, Appendix C contains proofs for Section 4, and Appendix D contains additional results and derivations regarding difference-in-differences and associated weighted estimands.
Appendix A Proofs for Section 2
Proof of Proposition 2.4.
We begin by noting that
$\beta_{\text{TWFE}}$ $=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}]}$
$=\frac{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}\mid P=t]\mathbb{P}(P=t)}{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}\mid P=t]\mathbb{P}(P=t)}$
$=\frac{\mathbb{E}[\ddot{D}Y]}{\mathbb{E}[\ddot{D}^{2}]},$
where the second equality follows from the uniform distribution of $P$ which is independent from $(\ddot{D}_{t},Y_{t})$ for all $t\in\{1,\ldots,T\}$. The third equality follows from defining $\ddot{D}\equiv\ddot{D}_{P}$. We also note that
$\ddot{D}_{P}$ $=D_{P}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\sum_{t=1}^{T}\mathbb{E}[D_{t}]\mathbbm{1}(P=t)+\sum_{s=1}^{T}\mathbb{E}[D_{s}]\mathbb{E}[\mathbbm{1}(P=s)]$
$=D-\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}(G\leq s)-\mathbb{E}[D\mid P]+\mathbb{E}[D]$
$=D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D].$
The third equality follows from $\mathbb{E}[D\mid G]=\mathbb{E}[\mathbbm{1}(G\leq P)\mid G]=\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}(G\leq s)=\frac{1}{T}\sum_{s=1}^{T}D_{s}$.
We break down the rest of this proof into four steps.
Step 1: Splitting the Numerator in Two
We have that
$\mathbb{E}[\ddot{D}Y]=\mathbb{E}[\ddot{D}(Y(0)+D(Y(1)-Y(0)))]=\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]+\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid P,G]]$
The first equality follows from $Y=Y(0)+D(Y(1)-Y(0))$ and the second from iterated expectations and $\mathbb{E}[D\mid G,P]=D$.
Step 2: First Numerator Term
We have that
$\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]$ $=\mathbb{E}[\ddot{D}\theta(G,P)]=\mathbb{E}[\ddot{D}\ddot{\theta}(G,P)]$
where $\theta(G,P)=\mathbb{E}[Y(0)\mid G,P]$. The second equality follows by properties of projections and from defining $\ddot{\theta}(G,P)$ as follows:
$\ddot{\theta}(G,P)$ $\equiv\theta(G,P)-\mathbb{E}[\theta(G,P)\mid G]-\mathbb{E}[\theta(G,P)\mid P]+\mathbb{E}[\theta(G,P)]$
$=\mathbb{E}[Y(0)\mid G,P]-\mathbb{E}[Y(0)\mid G]-\mathbb{E}[Y(0)\mid P]+\mathbb{E}[Y(0)].$
Then, we note that
$\theta(g^{\prime},t^{\prime})$ $=\mathbb{E}[Y(0)\mid G=g^{\prime},P=t^{\prime}]-\mathbb{E}[Y(0)\mid G=g^{\prime}]-\mathbb{E}[Y(0)\mid P=t^{\prime}]+\mathbb{E}[Y(0)]$
$=\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g^{\prime}]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g).$
Assumption 2.6.2 implies that for any pair $t,t^{\prime}\in\{1,\ldots,T\}$ and any $g^{\prime}\in\mathcal{G}$
$\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]=\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)].$
This can be shown for $t^{\prime}>t$ by writing $\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]=\sum_{s=t+1}^{t^{\prime}}\mathbb{E}[Y_{s}(0)-Y_{s-1}(0)\mid G=g^{\prime}]=\sum_{s=t+1}^{t^{\prime}}\mathbb{E}[Y_{s}(0)-Y_{s-1}(0)]=\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]$. Similar derivations show this holds for $t^{\prime}<t$. The case where $t^{\prime}=t$ is trivial. Therefore,
$\theta(g^{\prime},t^{\prime})$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]\right)\mathbb{P}(G=g)$
$=0$
for all $(g^{\prime},t^{\prime})\in\mathcal{G}\times\{1,\ldots,T\}$, which implies $\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]=0$.
Step 3: Second Numerator Term
We can write
$\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$ $=\mathbb{E}[(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[D(Y(1)-Y(0))\mid G,P]]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\mathbb{P}(D=1\mid G,P)].$
The first equality is by definition, the second by $D^{2}=D$, the third by $\mathbb{E}[D\mid G,P]=D$, and the fourth by law of total probability.
Step 4: Denominator
In this step, we show that
$\mathbb{E}[\ddot{D}^{2}]$ $=\mathbb{E}[(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]$
$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)].$
The first line is obtained from properties of linear projections, the second from $D^{2}=D$, and the third from $D=\mathbb{P}(D=1\mid G,P)$.
We conclude the proof by noting the equivalence of integrating over the distribution of $P$ and averages over time periods, which shows the equivalence between
$\beta_{\text{TWFE}}=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\mathbb{P}(D=1\mid G,P)]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)]}$
and the expression in Proposition 2.4. ∎
Proof of Proposition 2.5.
Proposition 2.4 and $\mathbb{P}(D=1\mid G,P)=D$ yielded
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]}.$
Since $\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]=\mathbb{E}[Y(1)-Y(0)\mid G,D=1]$ by assumption, we can use the law of iterated expectations to obtain
$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G]\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G]]}.$
We now calculate the conditional expectation $\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G=g]$ for $g\in\mathcal{G}$. If $g=+\infty$, then this conditional expectation is 0 by construction, so we let $g\in\{2,\ldots,T\}$ in what follows:
$\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G=g]$
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[D\mid G=g]-\frac{\mathbb{E}[\mathbb{E}[D\mid P]D\mid G=g]}{\mathbb{E}[D\mid G=g]}+\mathbb{E}[D]\right)$
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(G\leq P)\mid G=g]}{\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]}+\mathbb{E}[\mathbb{E}[\mathbbm{1}(G\leq P)\mid P]]\right)$
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(g\leq P)]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(g\leq P)]}{\mathbb{E}[\mathbbm{1}(g\leq P)]}+\mathbb{E}[F_{G}(P)]\right)$
$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(g\leq P)]-\mathbb{E}[F_{G}(P)\mid g\leq P]\right.$
$\quad\left.+\;\mathbb{E}[F_{G}(P)\mid g\leq P]\mathbb{E}[\mathbbm{1}(g\leq P)]+\mathbb{E}[F_{G}(P)\mid g>P]\mathbb{E}[\mathbbm{1}(g>P)]\right)$
$=\mathbb{E}[D\mid G=g]\mathbb{E}[\mathbbm{1}(g>P)](1+\mathbb{E}[F_{G}(P)\mid g>P]-\mathbb{E}[F_{G}(P)\mid g\leq P])$
$=\mathbb{E}[D\mid G=g](1-\mathbb{E}[D\mid G=g])(1+\mathbb{E}[D\mid g>P]-\mathbb{E}[D\mid g\leq P])$
$=\mathbb{P}(D=1\mid G=g)\mathbb{P}(D=0\mid G=g)(\mathbb{P}(D=1\mid P<g)+\mathbb{P}(D=0\mid P\geq g)).$
The first equality follows from $\mathbb{E}[D\mid G=g]>0$ for $g\in\{2,\ldots,T\}$. The second follows from $D=\mathbbm{1}(G\leq P)$ and the law of iterated expectations, the third from $G\perp\!\!\!\perp P$, the fourth from definitions of conditional expectations and the law of iterated expectations, the fifth from combining terms, the seventh from the law of iterated expectations again, and the last line is obtained by the fact that $D\in\{0,1\}$. The representation in Proposition 2.5 follows. ∎
Appendix B Proofs for Section 3
Proof of Proposition 3.1.
We begin by showing the first claim of the proposition. The equation
$\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]$ $=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ (B.1)
holds since $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, which holds by Definition 3.1. Since $\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ is assumed positive, we can divide both sides of equation (B.1) by it and obtain
$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ $=\frac{\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]}{\mathbb{P}(W^{*}=1\mid W_{0}=1,X)}$
$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1,X]$
$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X],$
where the second equality holds by definition, and the third holds from $W^{*}$ being a subpopulation of $W_{0}$.
We now show the second claim of the proposition. Equation (3.4) can be obtained as follows:
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1]$
$=\frac{\mathbb{E}[W^{*}(Y(1)-Y(0))\mid W_{0}=1]}{\mathbb{E}[W^{*}\mid W_{0}=1]}$
$=\frac{\mathbb{E}[\mathbb{E}[W^{*}(Y(1)-Y(0))\mid X,W_{0}=1]\mid W_{0}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\mid W_{0}=1]}$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid W_{0}=1,X)\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W_{0}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid W_{0}=1,X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\mu(\underline{w}^{*},\tau_{0}).$
The first equality follows from $W^{*}$ being a subpopulation of $W_{0}$, the second from the definition of conditional expectation and $\mathbb{P}(W^{*}=1)>0$, the third from the law of iterated expectations and $W^{*}=1$ implying $W_{0}=1$, the fourth from $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, and the fifth and sixth follow immediately. ∎
Proof of Lemma 3.1.
We verify that $W^{*}$ satisfies the two conditions in Definition 3.1. Condition 1 holds since $\mathbb{P}(W^{*}=1\mid W_{0}=0)=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid W_{0}=1)\leq\mathbb{P}(W^{\prime}=1\mid W_{0}=1)=0$, The first equality follows from $W^{*}\leq W^{\prime}$ almost surely, and the last one from $W^{\prime}\in\text{SP}(W_{0})$.
To verify condition 2, note that
$\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W_{0}=1)$ $=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W^{\prime}=1,W_{0}=1)\mathbb{P}(W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W^{\prime}=1)\mathbb{P}(W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1,W_{0}=1)\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid X,W_{0}=1)$
$=\mathbb{P}(W^{*}=1\mid X,W_{0}=1).$
The first and seventh line follow from $W^{*}\leq W^{\prime}$ almost surely. The second and sixth line follow from factoring conditional probabilities. The third and fifth line follow from $W^{\prime}\leq W_{0}$ almost surely. The fourth line follows from $W^{\prime}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$ and $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W^{\prime}=1$. ∎
Proof of Theorem 3.1.
This theorem follows as a special case of the first part of Theorem 4.3 when $W^{\prime}=W_{0}$. This is because $W_{0}$ trivially regular subpopulation of $W_{0}$, and because the condition
$\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$
is equivalent to $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. They are equivalent since $\underline{w}^{\prime}(X)=\mathbb{P}(W_{0}=1\mid X,W_{0}=1)=1$ almost surely. ∎
Proof of Theorem 3.2.
This is a special case of the second part of Theorem 4.3 where we set $W^{\prime}=W_{0}$. ∎
Proof of Proposition 3.2.
First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. Therefore, by Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{lin}}$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$. Therefore,
00 0 $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
$=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}[a(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=d^{\prime}\left(\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}\right)$
for all $d\in\mathbb{R}^{d_{X}}$, which implies that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}=\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[w^{*}(X)\mid W_{0}=1]}$. Letting $u(x)=\underline{w}^{*}(x)/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$, we have that
[…]
$W^{*}=\mathbbm{1}\left(U\leq\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\right)\cdot W_{0}$
where $U\perp\!\!\!\perp(X,Y(1),Y(0),W_{0})$. Then $W^{*}$ is a regular subpopulation of $W_{0}$ and $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)=\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}$ since $\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\in[0,1]$ with probability 1 given $W_{0}=1$. Therefore, for all $\tau_{0}(x)=c+d^{\prime}x\in\mathcal{T}_{\text{lin}}$ we have that
$\mu(\underline{w}^{*},\tau_{0})$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}\left[\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\tau_{0}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[u(X)\tau_{0}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[u(X)(c+d^{\prime}X)\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
$=c+d^{\prime}\frac{\mathbb{E}\left[u(X)X\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
$=c+d^{\prime}\frac{\mathbb{E}\left[a(X)X\mid W_{0}=1\right]}{\mathbb{E}\left[a(X)\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[a(X)(c+d^{\prime}X)\mid W_{0}=1\right]}{\mathbb{E}\left[a(X)\mid W_{0}=1\right]}$
$=\mu(a,\tau_{0}).$
Therefore, by Proposition 3.1 we have that $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. ∎
Proof of Proposition 3.3.
We consider the $K>0$ case first and the $K=0$ case second.
Case 1: $K>0$.
First, let $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. We will show that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))=\emptyset$ by way of contradiction.
Suppose there is a $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ and let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\in[0,1]$. Let $\tau^{*}(x)=K\cdot\mathbbm{1}(a(x)<0)$. This definition implies $\tau^{*}\in\mathcal{T}_{\text{BV}}(K)$. Since $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ we must have that $\mu(a,\tau_{0})=\mu(\underline{w}^{*},\tau_{0})$ for all $\tau_{0}\in\mathcal{T}_{\text{BV}}(K)$. Since $\tau^{*}\in\mathcal{T}_{\text{BV}}(K)$,
00 0 $=\mu(a,\tau^{*})-\mu(\underline{w}^{*},\tau^{*})$
$=\mathbb{E}\left[\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\tau^{*}(X)\mid W_{0}=1\right]$
$=K\cdot\mathbb{E}\left[\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\mathbbm{1}(a(X)<0)\mid W_{0}=1\right].$ (B.2)
The right-hand side of (B.2) is $K$ times the expectation of a nonpositive function. This follows from $a(X)\mathbbm{1}(a(X)<0)\leq 0$, $\underline{w}^{*}(X)\mathbbm{1}(a(X)<0)\geq 0$ by $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\geq 0$, $\mathbb{E}[a(X)\mid W_{0}=1]>0$ by Assumption 3.1, and $\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]=\mathbb{P}(W^{*}=1\mid W_{0}=1)>0$ by $W^{*}\in\text{SP}(W_{0})$. Since $K>0$, equation (B.2) implies the nonpositive function must equal 0 with probability 1 given $W_{0}=1$:
11 11 $=\mathbb{P}\left(\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\mathbbm{1}(a(X)<0)=0\mid W_{0}=1\right)$
$=\mathbb{P}\left(a(X)=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)]}\cdot\underline{w}^{*}(X)=0\mid a(X)<0,W_{0}=1\right),$
where the second equality follows from $\mathbb{P}(a(X)<0\mid W_{0}=1)>0$. This implies $a(X)$ equals a positive multiple of $\underline{w}^{*}(X)$, a nonnegative quantity, with probability 1 given $\{a(X)<0,W_{0}=1\}$, an event with positive probability that implies $a(X)$ is strictly negative. This is a contradiction and therefore $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))=\emptyset$.
Second, suppose $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$. By Theorem 3.1, $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})\neq\emptyset$. Since $\mathcal{T}_{\text{BV}}(K)\subseteq\mathcal{T}_{\text{all}}$, we have that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})\subseteq\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$. Therefore, $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))\neq\emptyset$ thus $\mu(a,\tau_{0})$ has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{BV}}(K)$.
Case 2: $K=0$.
When $K=0$, the function class $\mathcal{T}_{\text{BV}}(K)$ is the set of all constant functions. In this case, $\tau_{0}(X)=t_{0}$, where $t_{0}\in\mathbb{R}$ denotes a constant. Thus $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(0))\neq\emptyset$ for all weight functions $a(\cdot)$ since $W_{0}\in\text{SP}(W_{0})$ and because $\mu(a,\tau_{0})=\mathbb{E}[a(X)t_{0}\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]=t_{0}=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ for all $a(\cdot)$. ∎
Appendix C Proofs for Section 4
Proof of Theorem 4.1.
This is a special case of the first part of Theorem 4.4 where we set $W^{\prime}=W_{0}$. ∎
We begin with a technical lemma that we use to prove the subsequent theorems.
Lemma C.1.
Let Assumption 3.1 hold. Let $T_{\mu}=\tau_{0}(X)-\mu$. Then, for any $W^{\prime}\in\text{SP}(W_{0})$, we have that
  1. 1.
    The functions $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ and $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]$ are left-continuous.
  2. 2.
    The functions $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}>\alpha)\mid W^{\prime}=1]$ and $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W^{\prime}=1]$ are right-continuous.
Proof of Lemma C.1.
The function $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ is left-continuous if for any strictly increasing sequence $\alpha_{n}\uparrow\alpha$ we have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})\mid W^{\prime}=1]\rightarrow\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$. To see this holds, note that $f_{n}(t)\equiv t\mathbbm{1}(t<\alpha_{n})\rightarrow t\mathbbm{1}(t<\alpha)$ since $t\mathbbm{1}(t<\alpha_{n})=0$ for all $t\geq\alpha$, and $t\mathbbm{1}(t<\alpha_{n})=t$ whenever $t<\alpha$ for sufficiently large $n$. The random variable $|T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})|$ is dominated by $|T_{\mu}|$ and $\mathbb{E}[|T_{\mu}|\mid W^{\prime}=1]<\infty$ by Assumption 3.1 and by $\mathbb{P}(W^{\prime}=1)>0$. Therefore, by dominated convergence, $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})\mid W^{\prime}=1]\rightarrow\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ hence $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ is left-continuous. The function $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]$ is also left-continuous because $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]=\mathbb{E}[T_{\mu}\mid W^{\prime}=1]-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$.
The lemma’s second claim can be similarly shown by considering a sequence $\alpha_{n}\downarrow\alpha$. ∎
Proof of Theorem 4.2.
To simplify the notation in the proof, let $\mu\equiv\mu(a,\tau_{0})$. We break down this proof into four cases.
Case 1: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
We want to maximize $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ over the subpopulations $W^{*}$ in $\mathcal{W}(a;W_{0},\{\tau_{0}\})$. Recall that $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ if $\mu=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$ and $W^{*}\in\text{SP}(W_{0})$ hold, where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$.
Therefore,
$\overline{P}(a,W_{0};\{\tau_{0}\})$ $=\max_{W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})}\mathbb{P}(W^{*}=1\mid W_{0}=1)$
$=\max_{W^{*}\in\{0,1\}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],W^{*}\in\text{SP}(W_{0})}\mathbb{P}(W^{*}=1\mid W_{0}=1)$
$\leq\max_{W^{*}\in\{0,1\}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\mid W_{0}=1]$
$=\max_{\underline{w}^{*}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],\underline{w}^{*}(X)\in[0,1]}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1].$
We will first show a closed-form expression for an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$. Then, we will show that this upper bound can be attained by a corresponding $W^{+}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$, and therefore it equals $\overline{P}(a,W_{0};\{\tau_{0}\})$.
Before proceeding, let $\alpha^{+}=\inf\{\alpha\in\mathbb{R}:R(\alpha)\geq 0\}$ where $R(\alpha)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]$. By construction, $\alpha^{+}\geq 0$. By $\mu<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ we also have that $\alpha^{+}<+\infty$. By Lemma C.1, $R$ is a right-continuous function, and therefore $R(\alpha^{+})=\lim_{\alpha\downarrow\alpha^{+}}R(\alpha)\geq 0$. We now claim that $\alpha^{+}>0$. To show this claim, assume $\alpha^{+}=0$. Then, $0\leq R(\alpha^{+})=R(0)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq 0)\mid W_{0}=1]\leq 0$, which implies $\mathbb{P}(\tau_{0}(X)=\mu\mid W_{0}=1)=1$. This is ruled out by the assumption that $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]=\mathbb{E}[\tau_{0}(X)\mid W_{0}=1]$. Therefore, $\alpha^{+}>0$.
We now show an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$. For all $\underline{w}^{*}$ such that $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ and $\underline{w}^{*}(X)\in[0,1]$, we have that
$\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\alpha^{+}\mid W_{0}=1]}{\alpha^{+}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[\underline{w}^{*}(X)T_{\mu}\mid W_{0}=1]}{\alpha^{+}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mid W_{0}=1]}{\alpha^{+}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}>\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$\leq\frac{\mathbb{E}[1\cdot(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[0\cdot(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}>\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$=\mathbb{E}[\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$\equiv P^{+}.$
The third equality follows from $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$. The inequality follows from $\{0,1\}$ being lower/upper bounds for $\underline{w}^{*}(X)$.
Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})\leq P^{+}$. We now show that this inequality is binding by finding $W^{+}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ such that $\mathbb{P}(W^{+}=1\mid W_{0}=1)=P^{+}$.
Let
$\underline{w}^{+}(X)$ =$=\begin{cases}1&\text{ if }T_{\mu}<\alpha^{+}\\ 1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}&\text{ if }T_{\mu}=\alpha^{+}\\ 0&\text{ if }T_{\mu}>\alpha^{+}.\end{cases}$
This function is bounded above by 1 because $R(\alpha^{+})\geq 0$ and $\alpha^{+}>0$. To show it is bounded below by 0, consider cases where $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$ or $R(\alpha^{+})$ equal and differ from 0. If $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0$ or $R(\alpha^{+})=0$, then $\underline{w}^{+}(X)\in\{0,1\}\subseteq[0,1]$ and it is therefore bounded below by 0. If $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)>0$ and $R(\alpha^{+})>0$, then
$1-\frac{R(\alpha^{+})}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$ $=\frac{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$
$=\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}=\alpha^{+})\mid W_{0}=1]-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$
$=\frac{-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}.$
By the definition of $\alpha^{+}$ as an infimum, we must have that $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, implying that $R(\alpha)$ is discontinuous at $\alpha^{+}$. By Lemma C.1, $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)]$ is left-continuous and satisfies $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)]\leq R(\alpha)$. Therefore, since $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, we must have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+}-\varepsilon)]<0$ for all $\varepsilon>0$. Letting $\varepsilon\downarrow 0$ yields that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})]\leq 0$. Therefore $-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]/(\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1))\geq 0$ and $\underline{w}^{*}(X)\geq 0$ in that case as well.
Next, we compute
$\mathbb{E}[\underline{w}^{+}(X)\mid W_{0}=1]$ $=\mathbb{E}[\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]$
$\qquad+\left(1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\mathbb{E}[\mathbbm{1}(T_{\mu}=\alpha^{+})\mid W_{0}=1]$
$=\mathbb{E}[\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]+\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$=\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
$=P^{+}.$
Finally we verify that $\frac{\mathbb{E}[\underline{w}^{+}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{+}(X)\mid W_{0}=1]}=\mu$. This condition is equivalent to $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]=0$, which we verify here:
$\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ $=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]$
$\qquad+\left(1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\mathbb{E}[T_{\mu}1(T_{\mu}=\alpha^{+})\mid W_{0}=1]$
$=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$
$=R(\alpha^{+})-R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)$
$=R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0).$
Therefore, $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ equals 0 when $R(\alpha^{+})=0$. When $R(\alpha^{+})>0$, we have that $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)>0$ as shown earlier. So $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ is also equal to 0 in this case.
We conclude the proof by showing that $\underline{w}^{+}(X)$ corresponds to $\mathbb{P}(W^{+}=1\mid X,W_{0}=1)$ for some $W^{+}\in\text{SP}(W_{0})$. Let $U\sim\text{Unif}(0,1)$ satisfy $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$ and define
$W^{+}$ $=\left(\mathbbm{1}(T_{\mu}<\alpha^{+})+\mathbbm{1}\left(T_{\mu}=\alpha^{+},U\leq 1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\right)\cdot W_{0}.$
By construction, $W^{+}\in\{0,1\}$, $\mathbb{P}(W^{+}=1\mid X,W_{0}=1)=\underline{w}^{+}(X)$, $\mathbb{P}(W^{+}=1\mid W_{0}=0)=0$, and
$W^{+}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$. Also, since $\mu\in\mathcal{S}(\tau_{0};W_{0})$, $\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu\mid W_{0}=1)>0$. Since $\alpha^{+}>0$ we have that $\mathbb{P}(W^{+}=1\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}<\alpha^{+}\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)>0$. Therefore $W^{+}$ is a regular subpopulation of $W^{+}$ for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)=P^{+}$, hence $P^{+}$ is the maximum.
Case 2: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
As in case 1,
$\overline{P}(a,W_{0};\{\tau_{0}\})$ $\leq\max_{\underline{w}^{*}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],\underline{w}^{*}(X)\in[0,1]}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1].$
Before proceeding, let $\alpha^{-}=\sup\{\alpha\in\mathbb{R}:L(\alpha)\leq 0\}$ where $L(\alpha)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]$. By construction, $\alpha^{-}\leq 0$ and by $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ we have that $\alpha^{-}>-\infty$. By Lemma C.1, $L$ is a left-continuous function, and therefore $L(\alpha^{-})=\lim_{\alpha\uparrow\alpha^{-}}L(\alpha)\leq 0$. Similarly to case 1, we can show that $\alpha^{-}<0$.
We now show an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$. For all $\underline{w}^{*}$ such that $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ and $\underline{w}^{*}(X)\in[0,1]$, we have that
$\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\alpha^{-}\mid W_{0}=1]}{\alpha^{-}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[\underline{w}^{*}(X)T_{\mu}\mid W_{0}=1]}{\alpha^{-}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mid W_{0}=1]}{\alpha^{-}}$
$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}<\alpha^{-})\mid W_{0}=1]}{\alpha^{+}}$
$\leq\frac{\mathbb{E}[1\cdot(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}\geq\alpha^{+})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[0\cdot(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}<\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}$
$=\mathbb{E}[\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}$
$\equiv P^{-},$
which follows a similar argument as above. This implies $\overline{P}(a,W_{0};\{\tau_{0}\})\leq P^{-}$. We now show that this inequality is an equality by finding $W^{-}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ such that $\mathbb{P}(W^{-}=1\mid W_{0}=1)=P^{-}$.
Let
$\underline{w}^{-}(X)$ =$=\begin{cases}1&\text{ if }T_{\mu}>\alpha^{+}\\ 1-\frac{L(\alpha^{-})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{-}\mid W_{0}=1)\neq 0)}{\alpha^{-}\mathbb{P}(T_{\mu}=\alpha^{-}\mid W_{0}=1)}&\text{ if }T_{\mu}=\alpha^{-}\\ 0&\text{ if }T_{\mu}<\alpha^{+}.\end{cases}$
The rest of the proof symmetrically follows the one for the previous case.
Case 3: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
Note that $W^{*}=W_{0}$ is the largest regular subpopulation of $W_{0}$. Since $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, we have that $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is trivially maximized at $1$.
Case 4: $\mu\notin\mathcal{S}(\tau_{0};W_{0})$
By Theorem 3.2, there does not exist a regular subpopulation $W^{*}$ satisfying $\mu=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ and therefore the supremum equals 0 by its definition. ∎
Proof of Theorem 4.3.
Part 1: $\mathcal{T}=\mathcal{T}_{\text{all}}$
()($\Longrightarrow$)( ⟹ ) First, suppose there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$. Using the law of iterated expectations and letting $\underline{w}^{\prime}(X)=\mathbb{P}(W^{\prime}=1\mid W_{0}=1,X)$, we can write
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\underline{w}^{\prime}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\underline{w}^{\prime}(X)\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)W^{\prime}\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\mid W_{0}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1,W_{0}=1\right]\mathbb{P}(W^{\prime}=1\mid W_{0}=1)}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1,W_{0}=1\right]\mathbb{P}(W^{\prime}=1\mid W_{0}=1)}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$\equiv\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right).$
The second equality is valid due to $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$. The third and fourth follow from the law of iterated expectations, and the fifth from $W^{\prime}$ being a subpopulation of $W_{0}$. The second to last line follows from Proposition 3.1 and from $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ implying $\mathbb{P}(\underline{w}^{\prime}(X)>0\mid W_{0}=1)>0$.
Similarly, we can write $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu^{\prime}(\underline{w}^{*},\tau_{0})$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W^{\prime}=1,X)$. Therefore, by Proposition 3.1, we have that $\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right)-\mu^{\prime}(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{all}}$.
Let $\tau^{*}(X)=\frac{a(X)/\underline{w}^{\prime}(X)}{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}$. We have $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ and $\mathbb{E}[\underline{w}^{*}(X)^{2}]\leq 1$ by construction. Hence, $\mathbb{E}[\tau^{*}(X)^{2}]<\infty$ and thus $\tau^{*}\in\mathcal{T}_{\text{all}}$.
Thus, we must have $\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau^{*}\right)-\mu^{\prime}(\underline{w}^{*},\tau^{*})=0$. Expanding this equality yields
00 0 $=\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau^{*}\right)-\mu^{\prime}(\underline{w}^{*},\tau^{*})$
$=\mathbb{E}\left[\tau^{*}(X)\left(\frac{a(X)/\underline{w}^{\prime}(X)}{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}\right)\mid W^{\prime}=1\right]$
$=\mathbb{E}[\tau^{*}(X)^{2}\mid W^{\prime}=1].$
This implies that $\mathbb{P}(\tau^{*}(X)=0\mid W^{\prime}=1)=1$. In turn, this implies that
$\mathbb{P}\left(a(X)=\frac{\underline{w}^{\prime}(X)\underline{w}^{*}(X)\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}\mid W^{\prime}=1\right)=1.$
We have that $\underline{w}^{*}(X)\geq 0$ and $\underline{w}^{\prime}(X)\geq 0$ almost surely, and $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)>0$ by the assumption that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$. Also, $\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]=\mathbb{E}[a(X)\mid W_{0}=1]>0$ by Assumption 3.1. Therefore, $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$.
[…]
$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$\equiv\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right).$
Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0},W^{\prime})$, and define
$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}.$
We verify that $W^{*}$ is a regular subpopulation of $W^{\prime}$. First, we see that $\mathbb{P}(W^{*}=1)>0$ because
$\mathbb{P}(W^{*}=1)$ $=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)+\mathbb{P}(W^{*}=1\mid W^{\prime}=0)\mathbb{P}(W^{\prime}=0)$
$=\mathbb{P}\left(\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}=1\mid W^{\prime}=1\right)\mathbb{P}(W^{\prime}=1)$
$=\mathbb{E}\left[\mathbb{P}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid X,W^{\prime}=1\right)\mid W^{\prime}=1\right]\mathbb{P}(W^{\prime}=1)$
$=\mathbb{E}\left[\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid W^{\prime}=1\right]\mathbb{P}(W^{\prime}=1)$
$=\frac{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]\mathbb{P}(W^{\prime}=1)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}$
$=\frac{\mathbb{E}[a(X)\mid W_{0}=1]\mathbb{P}(W^{\prime}=1)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}$
$>0.$
The second equality follows from $\mathbb{P}(W^{*}=1\mid W^{\prime}=0)=0$, which holds by the construction of $W^{*}$. The fourth equality holds from $U\perp\!\!\!\perp(X,W^{\prime})$ and from $\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\in[0,1]$ almost surely given $W^{\prime}=1$. To establish the final inequality, recall that $\mathbb{P}(W^{\prime}=1)$ and $\mathbb{E}[a(X)\mid W_{0}=1]$ are positive by assumption. Also $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))\leq\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ since $\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1)$ is a subset of $\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1)$. That $W^{*}$ satisfies the two properties of Definition 3.1 holds immediately. Therefore, $W^{*}$ is a regular subpopulation of $W^{\prime}$.
Finally, let
$\underline{w}^{*}(X)$ $=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)$
$=\mathbb{P}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid X,W^{\prime}=1\right)$
$=\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}.$
Letting $\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]=\tau_{0}(X)\in\mathcal{T}_{\text{all}}$ and using Proposition 3.1, we can see that
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[\underline{w}^{\prime}(X)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}\left[\frac{\frac{a(X)}{\underline{w}^{\prime}(X)}}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{\frac{a(X)}{\underline{w}^{\prime}(X)}}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid W^{\prime}=1\right]}$
$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
$=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
$=\mu(a,\tau_{0}).$
Since $\tau_{0}\in\mathcal{T}_{\text{all}}$ was arbitrary, we have that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ which concludes the proof.
Part 2: $\mathcal{T}=\{\tau_{0}\}$
To simplify the notation in the proof, we let $\mu\equiv\mu(a,\tau_{0})$.
()($\Longrightarrow$)( ⟹ ) First, let $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. Since $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$, we can without loss of generality suppose that $\mathbb{P}(\tau_{0}(X)\leq\mu\mid W^{\prime}=1)=0$, which implies $\mathbb{P}(\tau_{0}(X)>\mu\mid W^{\prime}=1)=1$. Since $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$, we can write by Proposition 3.1
$\mu$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}$
$\geq\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mu\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}=\mu.$ (C.1)
The inequality is strict unless $\mathbb{E}[(\underbrace{\tau_{0}(X)-\mu}_{>0\text{ w.p.1}})\underbrace{\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)}_{\in[0,1]}\mid W^{\prime}=1]=0$ holds. This holds if $\mathbb{P}((\tau_{0}(X)-\mu)\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)=0\mid W^{\prime}=1)=1$, which in turns occurs if and only if $\mathbb{P}(\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)=0\mid W^{\prime}=1)=1$. This implies $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)=0$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)=0$, a contradiction of $W^{*}\in\mathcal{W}(a,W^{\prime},\{\tau_{0}\})$. Therefore, the inequality in (C.1) is strict and yields $\mu>\mu$, a contradiction. Therefore, $\mathcal{W}(a;W^{\prime},\{\tau_{0}\})=\emptyset$ when $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$.
[…]
$\mathcal{X}^{-}$ =$=\{x\in\text{supp}(X):\tau_{0}(x)\leq\mu\}\qquad\text{ and }\qquad\mathcal{X}^{+}=\{x\in\text{supp}(X):\tau_{0}(x)\geq\mu\}.$
By $\mu\in\mathcal{S}(\tau_{0};W^{\prime})$, $\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu\mid W^{\prime}=1)>0$. Similarly, $\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)>0$.
Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W^{\prime},W_{0})$. For $u\in[0,1]$ let
$W^{*}(u)$ $=(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+}))\cdot W^{\prime}.$
We can see that $W^{*}(u)\in\{0,1\}$, that $W^{*}(u)\perp\!\!\!\perp(Y(1),Y(0))\mid X,W^{\prime}=1$, and that $\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=0)=0$. To show that $W^{*}(u)$ characterizes a regular subpopulation of $W^{\prime}$, we also show that it is nonzero with positive probability:
$\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=1)$ $=\mathbb{P}(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+})=1\mid W^{\prime}=1)$
$=\mathbb{P}(U>u,X\in\mathcal{X^{-}}\mid W^{\prime}=1)+\mathbb{P}(U\leq u,X\in\mathcal{X^{-}}\mid W^{\prime}=1)$
$=(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)$
$>0$
for all $u\in[0,1]$. Therefore, $\mathbb{P}(W^{*}(u)=1)=\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)>0$ by $\mathbb{P}(W^{\prime}=1)>0$. Hence, $W^{*}(u)\in\text{SP}(W^{\prime})$ for all $u\in[0,1]$.
For a given $u$, we have that $\underline{w}^{*}(X;u)\equiv\mathbb{P}(W^{*}(u)=1\mid X,W^{\prime}=1)=(1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+})$. Therefore, using Proposition 3.1,
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X;u)\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]\mid W^{\prime}=1]}{\mathbb{E}[\underline{w}^{*}(X;u)\mid W^{\prime}=1]}$
$=\frac{\mathbb{E}[((1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+}))\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[((1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+}))\mid W^{\prime}=1]}$
$=\frac{(1-u)\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W^{\prime}=1]+u\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W^{\prime}=1]}{(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}.$
By construction, $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{-})\leq\mu\mathbbm{1}(X\in\mathcal{X}^{-})$ and $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{+})\geq\mu\mathbbm{1}(X\in\mathcal{X}^{+})$ almost surely. Therefore,
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(0)=1]$ $=\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)}\leq\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\mu\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)}=\mu$
and
$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(1)=1]$ $=\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}\geq\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\mu\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}=\mu.$
By the continuity of $\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ in $u$ and the intermediate value theorem, there exists $u^{*}\in[0,1]$ such that $\mu=\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u^{*})=1]$ and $W^{*}(u^{*})\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. ∎
Proof of Theorem 4.4.
Part 1: $\mathcal{T}=\mathcal{T}_{\text{all}}$
First suppose $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$. From Theorem 4.3, there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$. Written differently, we have that
$\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$ $=\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\frac{\mathbb{E}\left[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1\right]}$
for all $\tau_{0}\in\mathcal{T}_{\text{all}}$. From derivations in the proof of Theorem 4.3, we have that $\mathbb{P}(C\cdot\frac{a(X)}{\underline{w}^{\prime}(X)}=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1)=1$ for some positive constant $C>0$.
Since $\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\leq 1$ almost surely given $W^{\prime}=1$, we must have $C\cdot a(X)/\underline{w}^{\prime}(X)\leq 1$ almost surely given $W^{\prime}=1$. This means $C$ is bounded above by $\inf(\text{supp}(\underline{w}^{\prime}(X)/a(X)\mid W^{\prime}=1))$, which is strictly positive by assumption. Therefore,
$\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ $=\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]$
$\leq\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\right]}{\mathbb{P}(W^{\prime}=1)}$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)$
$\qquad\cdot\mathbb{E}\left[\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\mid X,W_{0}=1\right]\mid W_{0}=1\right]\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}$
$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\cdot\mathbb{E}[a(X)\mid W_{0}=1]\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}.$
The fifth line follows from $W^{\prime}$ being a subpopulation of $W_{0}$, and the last line follows from the law of iterated expectations.
This upper bound is sharp because it is attained by setting
$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}$
and noting that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ from Theorem 4.3.
Now suppose $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)<1$. By Theorem 4.3, $\mathcal{W}(a;W^{\prime},\mathcal{T})=\emptyset$ and therefore $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})$ is zero.
Part 2: $\mathcal{T}=\{\tau_{0}\}$
In this case, we seek to maximize $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ subject to $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$. Using Proposition 3.1, we can write $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ as $\mathbb{E}[\tau_{0}(X)\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]/\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]$. The result then follows from a direct application of Theorem 4.2 that replaces $W_{0}$ by $W^{\prime}$ in its statement and $\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$ by $\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)$ in the proofs. ∎
Appendix D Difference-in-Differences
Goodman-Bacon (2021) provides the following representation of the two-way fixed effects estimand under the assumption that group-level average treatment effects are constant over time:
$\beta_{\text{TWFE}}$ $=\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\Bigg{[}\sum_{j=1}^{k-1}\sigma_{jk}^{k}+\sum_{j=k+1}^{K}\sigma_{kj}^{k}\Bigg{]}\cdot\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1],$
where
$\sigma_{jk}^{k}\;=\;\frac{\mathbb{P}(G=j)\cdot\mathbb{P}(G=k)\cdot\mathbb{P}(D=1\mid G=k)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}}{\operatorname{var}(D^{\perp\left(G_{t_{1}},\ldots,G_{t_{K-1}},P_{1},\ldots,P_{T}\right)})}$
and
$\sigma_{kj}^{k}\;=\;\frac{\mathbb{P}(G=j)\cdot\mathbb{P}(G=k)\cdot\big{[}1-\mathbb{P}(D=1\mid G=k)\big{]}\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}}{\operatorname{var}(D^{\perp\left(G_{t_{1}},\ldots,G_{t_{K-1}},P_{1},\ldots,P_{T}\right)})}.$
[…]
$\beta_{\text{TWFE}}$ $=$ $\frac{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)]}$
$=$ $\frac{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)\cdot\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1]}{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)},$
it becomes clear that, for each group $k$ other than the always treated and the never treated,
$a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)$ $=$ $\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\mathbb{P}(D=1\mid G=k)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
$+$ $\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}1-\mathbb{P}(D=1\mid G=k)\big{]}\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]},$
and this, in turn, implies that
$a_{B}(k)$ $=$ $\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$ (D.1)
$+$ $\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}.$
D.1 Equivalence of Weight Functions
We now show that the weights obtained in equation (D.1) are equivalent to those in Proposition 2.5. First, we rewrite the weights in (D.1) as follows:
$a_{B}(k)$ $=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
$+\;\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}$
$=\mathbb{P}(D=1,G<k)-\mathbb{P}(G<k)\mathbb{E}[D\mid G=k]+(1-\mathbb{E}[D\mid G=k])\mathbb{P}(G>k)$
$\quad-\frac{1-\mathbb{E}[D\mid G=k]}{\mathbb{E}[D\mid G=k]}\mathbb{P}(D=1,G>k)$
$=\mathbb{P}(D=1,G<k)-\mathbb{P}(G<k)\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)-\mathbb{E}[D\mid G=k]\mathbb{P}(G>k)$
$\quad-\frac{1}{\mathbb{E}[D\mid G=k]}\mathbb{P}(D=1,G>k)+\mathbb{P}(D=1,G>k)$
$=\mathbb{P}(D=1,G\neq k)-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
$=(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]\mathbb{P}(G=k))-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
$=\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right).$
For $k\in\{2,\ldots,T\}$, the weights in Proposition 2.5 are equal to
$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D].$ (D.2)
because they are the average of the weights in Proposition 2.4 conditional on $G=k$. The proof of Proposition 2.5 explicitly shows that
$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=\mathbb{P}(D=0\mid G=k)\cdot(\mathbb{P}(D=0\mid P\geq k)+\mathbb{P}(D=1\mid P<g)).$
Let us look at the difference between the weights in (D.1) and (D.2). Fix $k\in\{2,\ldots,T\}$ and write
$\left(1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D]\right)-\left(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)\right)$
$=1-\mathbb{E}[D\mid P\geq k]-\mathbb{P}(G>k)+\frac{\mathbb{E}[D\mathbbm{1}(G>k)]}{\mathbb{E}[D\mid G=k]}$
$=\mathbb{E}[\mathbbm{1}(G\leq k)]-\frac{\mathbb{E}[D\mathbbm{1}(P\geq k)]}{\mathbb{E}[\mathbbm{1}(P\geq k)]}+\frac{\mathbb{E}[D\mathbbm{1}(G>k)]}{\mathbb{E}[\mathbbm{1}(k\leq P)]}$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[D\mathbbm{1}(G>k)]-\mathbb{E}[D\mathbbm{1}(P\geq k)]\right)$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[\mathbbm{1}(k<G\leq P)]-\mathbb{E}[\mathbb{E}[D\mid P]\mathbbm{1}(P\geq k)]\right)$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[\mathbb{E}[\mathbbm{1}(k<G\leq P)\mid P]]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[(F_{G}(P)-F_{G}(k))\mathbbm{1}(P\geq k)]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]-F_{G}(k)\mathbb{E}[\mathbbm{1}(P\geq k)]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
$=0.$
Therefore, the weights in Proposition (2.5), (D.1), and (D.2) are all equal to one another.

### v3

@@
-
+Supp_Appendix_v4
 Abstract
@@
 1   Introduction
-Estimating average treatment effects is an important objective of empirical research in many areas of economics.
-Applied researchers usually believe that treatment effects are heterogeneous, which means that they vary between units with different observable and unobservable characteristics.
+Estimating average treatment effects is an important goal in many areas of empirical research.
+Applied researchers usually believe that treatment effects are heterogeneous, which means that they vary across units.
 Yet, many researchers also favor using well-established estimation methods that were not originally designed with treatment effect heterogeneity in mind.
@@
 Let treatment $D$ be binary, $(Y(1),Y(0))$ be potential outcomes, and let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$ be the conditional average treatment effect, or CATE, for covariate value $X$.
-Following Angrist (1998), if we additionally assume that $\mathbb{E}[D\mid X]$ is linear in $X$, the regression of $Y$ on a constant, treatment $D$, and covariates $X$ yields a coefficient on $D$ that can be written as
+Following Angrist (1998), if we additionally assume that $\mathbb{P}(D=1\mid X)$ is linear in $X$, the population regression of $Y$ on a constant, treatment $D$, and covariates $X$ yields a coefficient on $D$ that can be written as
 $\beta_{\text{OLS}}=\frac{\mathbb{E}[\operatorname{var}(D\mid X)\tau_{0}(X)]}{\mathbb{E}[\operatorname{var}(D\mid X)]},$
 a weighted average of CATEs with nonnegative weights that integrate to 1.
+This parameter will be equal to the average treatment effect, $\mathbb{E}[Y(1)-Y(0)]$, if and only if $\operatorname{var}(D\mid X)$ and $\tau_{0}(X)$ are uncorrelated.
 In this paper we are concerned with a general class of weighted estimands that can be expressed as follows:
-$\mu(a,\tau_{0})$ $\equiv\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$ (1.1)
-where $W_{0}\in\{0,1\}$ is an indicator for a subpopulation, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ are the CATEs given covariates $X$ in subpopulation $W_{0}$, $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$ is the probability of being in subpopulation $W_{0}$ given $X$, and $a(X)$ is an identified weight function satisfying $\mathbb{E}[a(X)w_{0}(X)]>0$.
+$\mu(a,\tau_{0})$ $\coloneqq\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$ (1.1)
+where $W_{0}\in\{0,1\}$ is an indicator for a subpopulation, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ are the CATEs given covariates $X$ in the same subpopulation $W_{0}$, $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$ is the probability of being in this subpopulation given $X$, and $a(X)$ is an identified weight function.
 The regression estimand above belongs to this class, which can be seen by letting $W_{0}=1$ with probability 1, and letting the weight function $a(X)$ be the conditional variance of treatment given covariates.
 Under some assumptions, this class also includes the two-stage least squares (2SLS) and two-way fixed effects (TWFE) estimands in instrumental variables and difference-in-differences settings, as well as many other parameters.
+Here, the leading cases of $W_{0}$ are compliers in the case of 2SLS and treated units in the case of TWFE.
 There are two main questions that this paper seeks to answer.
-The first is whether, and under what circumstances, the estimand in (1.1) corresponds to an average treatment effect of the form $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$, where $W^{*}\in\{0,1\}$ is an indicator for a (possibly latent) subpopulation of $W_{0}$.
-An affirmative answer to this question would endow a specific weighted estimand with some degree of validity as a causal parameter, given that it would then measure the average effect of treatment for a subset of all units.
+The first is whether, and under what conditions, the estimand in (1.1) corresponds to an average treatment effect of the form $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$, where $W^{*}\in\{0,1\}$ is an indicator for a (possibly latent) subpopulation of $W_{0}$.
+An affirmative answer to this question would endow a specific estimand with some degree of validity as a causal parameter, given that it would then measure the average effect of treatment for a subset of all units.
 The second and primary aim of this paper is to quantify the degree of validity of $\mu(a,\tau_{0})$ as a causal parameter.
@@
 If these probabilities can be large, the estimand corresponds to the average treatment effect for a (relatively) large subpopulation, and when they are small, it corresponds to the average effect for a (relatively) small number of units.
-If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the parameter of interest, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target.
+If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target parameter, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target.
 If $\mathbb{P}(W^{*}=1)$, the corresponding marginal probability, is large, we say that $\mu(a,\tau_{0})$ is highly representative of the underlying population.
@@
 Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any knowledge of this function.
-If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form.
+If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form, including the most adversarial CATE functions.
 We can also answer the second question about the maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ without needing to estimate or know the structure of the CATEs.
 In a second case, we may want to know how representative $\mu(a,\tau_{0})$ is given knowledge of the CATE function.
-While the resulting maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are less useful as measures of robustness than in the first case—after all, if the researcher knows or estimates the entire CATE function, they can as well report any average of $\tau_{0}(X)$ that may be of interest—we consider this problem to be of independent theoretical interest.
-Additionally, if the researcher estimates and compares the maximum values of $\mathbb{P}(W^{*}=1)$ or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ in both cases, they can evaluate the impact of the nonuniform weights in $\mu(a,\tau_{0})$ in a given application.
+While the resulting maximum values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are less useful as measures of robustness than in the first case—after all, if the researcher knows or estimates the entire CATE function, they can as well report any average of $\tau_{0}(X)$ that may be relevant—we consider this problem to be of independent theoretical interest.
+Additionally, if the researcher estimates and compares the maximum values of $\mathbb{P}(W^{*}=1)$ or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ in both cases, they can evaluate the importance of treatment effect heterogeneity for the interpretation of $\mu(a,\tau_{0})$ in a given application.
 In the first case, when the CATE function is unrestricted, we formally show that $\mu(a,\tau_{0})$ can be written as the average treatment effect for a subpopulation of $W_{0}$ if and only if $a(X)\geq 0$ with probability 1 given $W_{0}=1$.
-The contrapositive of this statement is that the incidence of “negative weights,” i.e.
-$\mathbb{P}(a(X)<0)>0$, implies that $\mu(a,\tau_{0})$ cannot be represented as an average treatment effect for some subpopulation uniformly in $\tau_{0}$.
-In a related contribution, Blandhol, Bonney, Mogstad, and Torgovitsky (2022) have shown that, for estimands that do not depend on potential outcome levels, the lack of negative weights is a sufficient and necessary condition for the weighted estimand to be “weakly causal,” that is, to guarantee that the sign of $\tau_{0}$ will be preserved whenever it is uniform across all units.
+The contrapositive of this statement is that the incidence of “negative weights,” that is, $\mathbb{P}(a(X)<0\mid W_{0}=1)>0$, implies that $\mu(a,\tau_{0})$ cannot be represented as an average treatment effect for some subpopulation uniformly in $\tau_{0}$.
+This result provides a novel justification for the common requirement that the weights underlying a suitable estimand must be nonnegative.
+In a related contribution, Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that the lack of negative weights and level dependence is a sufficient and necessary condition for the weighted estimand to be “weakly causal,” that is, to guarantee that the sign of $\tau_{0}$ will be preserved whenever it is uniform across all units.
 We also provide simple expressions for the maxima of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$.
-We propose an analog estimator for these objects and briefly discuss associated estimation and inferential issues.
+We show how knowledge of the estimand and these expressions can be used to construct simple bounds on the target parameter.
+We also propose an analog estimator for our measure of internal validity.
+We establish the nonstandard limiting distribution of this estimator and describe inference procedures for it in Appendix C.
 In the second case, when the CATE function is assumed to be known, we show that $\mu(a,\tau_{0})$ can be written as an average treatment effect whenever it lies in the convex hull of CATE values, a weaker criterion than having nonnegative weights.
@@
 This expression can be used to derive plug-in estimators.
+Besides theoretical interest, we argue that $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are a practically relevant measure that may be appealing to applied researchers.
+First, as stated above, our initial results demonstrate that two commonly invoked criteria for weighted estimands—that they lack negative weights and that they lie in the convex hull of CATE values—are necessary and sufficient (under different assumptions) for the existence of their causal representation, that is, for $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be strictly positive.
+This suggests that researchers invoking these criteria are (indirectly) interested in whether their estimands can be represented as an average treatment effect for some subpopulation.
+If this is the case, it makes sense to better understand this implicit subpopulation, similar to how it is standard practice in instrumental variables settings to study the subpopulation of compliers.
+Relatedly, even though our main results concern subpopulation size, we also show that the distribution of covariates in the implicit subpopulation is identified.
+Thus, practitioners can examine whether this subpopulation has similar characteristics as the entire population, and report the associated sample statistics.
+Second, we argue that when $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the parameter of interest, it is reassuring for $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be large.
+For one, related claims have been made by other researchers.
+Mogstad and Torgovitsky (2024) argue that “[t]arget parameters that reflect larger subpopulations of the population of interest are more interesting than those that reflect smaller and more specific subpopulations.” In a setting with multiple instrumental variables, van ’t Hoff, Lewbel, and Mellace (2024) suggest that the largest subpopulation of compliers is generally more interesting than other complier subpopulations.
+However, we also formalize this claim and show how to construct bounds on $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ that only depend on $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, $\mu(a,\tau_{0})$, and a support restriction.
+The bounds are easy to compute and collapse to a point as $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ approaches 1.
+Indeed, large values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, our primary measures of interest, guarantee that the weighted estimand is not “too different” from $\mathbb{E}[Y(1)-Y(0)]$ and $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
 Literature Review
 This paper is related to a large literature studying weighted average representations of common estimands, including ordinary least squares (OLS), 2SLS, and TWFE in additive linear models.
-Some of the contributions to this literature include Angrist (1998), Humphreys (2009), Aronow and Samii (2016), Blandhol, Bonney, Mogstad, and Torgovitsky (2022), Słoczyński (2022), and Goldsmith-Pinkham, Hull, and Kolesár (2024) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
-A common view in much of this literature, traceable to Imbens and Angrist (1994), is that causal interpretability of weighted estimands requires all weights to be positive.
-For example, Sun and Abraham (2021) explicitly associate “reasonable weights” with weights that “sum to one and are non-negative.” Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that the lack of negative weights and level dependence is necessary and sufficient for an estimand to be “weakly causal,” that is, to guarantee sign preservation when all treatment effects have the same sign.
-In this paper we focus on the related problem of whether a weighted estimand can be written as the average treatment effect for some (possibly latent) subpopulation.
-While the lack of negative weights is essential in our framework when the CATE function is unrestricted, negative weights and nonuniform weights play a similar role when the CATE function is assumed to be known, at least as long as the weighted estimand lies in the convex hull of CATE values.
+Some of the contributions to this literature include Angrist (1998), Humphreys (2009), Aronow and Samii (2016), Słoczyński (2022), Chen (2024), and Goldsmith-Pinkham, Hull, and Kolesár (2024) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
+A common view in much of this literature, attributable to Imbens and Angrist (1994), is that causal interpretability of weighted estimands requires all weights to be nonnegative.
+Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that the lack of negative weights and level dependence is necessary and sufficient for an estimand to be “weakly causal,” that is, to guarantee sign preservation when all treatment effects have the same sign.
+In this paper we focus on the related problem of whether a weighted estimand can be written as the average treatment effect over a subpopulation.
+While the lack of negative weights is essential for subpopulation existence when the CATE function is unrestricted, nonuniform weights, too, have a detrimental effect on subpopulation size.
 This point is related to the negative view of both negative and nonuniform weights in Callaway, Goodman-Bacon, and Sant’Anna (2024).
@@
 Li, Morgan, and Zaslavsky (2018) discuss the connection between weighted average treatment effects and implicit target subpopulations.
-However, the internal validity and representativeness of weighted estimands have received very little attention to date.
-One exception is Aronow and Samii (2016), who explicitly acknowledge that the OLS estimand, like the local average treatment effect of Imbens and Angrist (1994), corresponds to the average effect for a “highly specific subpopulation” rather than the entire population, and consequently is not necessarily representative of that population.
-Then, Aronow and Samii (2016) focus on whether mean covariate values are similar in the entire sample and in the “effective sample” used by OLS.
+However, the internal validity and representativeness of weighted estimands have received little attention to date.
+One exception is de Chaisemartin (2012, 2017), who focuses on the interpretation of the instrumental variables (IV) estimand.
+First, de Chaisemartin (2012) studies the size of the largest subpopulation whose average treatment effect is equal to that of compliers.
+While this question is similar to ours, the corresponding subpopulation size is not point identified, unlike in our paper.
+Second, when the usual monotonicity assumption is violated, de Chaisemartin (2012, 2017) uses a specific restriction on the CATE function to reinterpret the IV estimand as the average treatment effect for a subset of compliers.
+In our framework, this result can be seen as an existence result in an intermediate case between the setting where $\tau_{0}$ is unrestricted and where it is fixed.
+Another exception is Aronow and Samii (2016), who focus on whether mean covariate values are similar in the entire sample and in the “effective sample” used by OLS.
 We focus on the size of the implicit subpopulation, which is different and complementary.
-Another exception is Miller, Shenhav, and Grosz (2023), who focus on (one-way) fixed effects estimands and argue that it is problematic if “switchers,” that is, fixed-effect groups with nonzero variation in treatment, are a small subset of the sample.
-They also recommend that applied researchers report the sample size when limited to “switcher groups.” In this paper we build a general framework to study the internal validity and representativeness of weighted estimands, with the fixed effects estimand (equivalent to OLS) as a special case.
-We argue, similar to Miller, Shenhav, and Grosz (2023), that if a given weighted estimand corresponds to the average treatment effect for a small subpopulation, then it may not be an appropriate target parameter, unless that subpopulation is interesting in its own right.
+We also extend the results on mean covariate values to the entire distribution of covariates and to other weighted estimands besides OLS.
+Yet another exception is Miller, Shenhav, and Grosz (2023), who focus on fixed effects estimands and argue that it is problematic if “switchers” are a small subset of the sample.
+We similarly argue that if a given weighted estimand corresponds to the average treatment effect for a small subpopulation, then it may not be an appropriate target parameter, unless that subpopulation is interesting in its own right.
 Plan of the Paper
 We organize the paper as follows.
-In Section 2, we discuss several examples of commonly used estimands that satisfy the representation in (1.1).
+In Section 2, we briefly discuss our motivating example of the OLS estimand.
 In Section 3, we develop our theoretical framework and examine the conditions under which the estimand in (1.1) has a causal representation as an average treatment effect over a population.
-In Section 4, we establish our main results on the absolute and relative size of subpopulations associated with the estimand in (1.1), which we propose as measures of representativeness and internal validity of weighted estimands.
-We also generalize our results to subpopulations and revisit the examples from Section 2.
+In Section 4, we establish our main results on the size of subpopulations associated with the estimand in (1.1).
+In Section 5, we revisit our motivating example from Section 2 and apply our theoretical results to additional examples of weighted estimands.
 In Section 6, we briefly discuss estimation and inference for the proposed measures.
@@
 In Section 8, we conclude.
-Appendix contains our proofs as well as several additional results and derivations.
-2 Applications
-Here we consider three identification strategies where commonly used estimands follow the structure of equation (1.1).
-2.1 Unconfoundedness
-In this setup, we have a binary treatment $D\in\{0,1\}$, potential outcomes $(Y(1),Y(0))$, covariate vector $X$, and realized outcome $Y=Y(D)$. We make the following two assumptions.
+The appendices contain our proofs as well as several additional results and derivations.
+2   Motivating Example
+Here we provide further discussion of the OLS estimand.
+We postpone the discussion of the 2SLS and TWFE estimands to Section 5.
+In the initial example, we have a binary treatment $D\in\{0,1\}$, potential outcomes $(Y(1),Y(0))$, covariate vector $X$, and realized outcome $Y=Y(D)$.
+We make the following assumption.
 Assumption 2.1 (Unconfoundedness).
@@
   2. 2.
-    Overlap: $p(X)\equiv\mathbb{P}(D=1\mid X)\in(0,1)$ almost surely.
+    Overlap: $p(X)\coloneqq\mathbb{P}(D=1\mid X)\in(0,1)$ almost surely.
 Following Angrist (1998), we can establish that $\beta_{\text{OLS}}$, the coefficient on $D$ in the linear projection of $Y$ on $(1,D,X)$, satisfies the representation in (1.1).
@@
 Proposition 2.1.
-Let Assumption 2.1 hold and let $p(X)$ be linear in $X$. Then
+Suppose Assumption 2.1 holds. Suppose $p(X)$ is linear in $X$. Then
 $\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[p(X)(1-p(X))\cdot\mathbb{E}[Y(1)-Y(0)\mid X]]}{\mathbb{E}[p(X)(1-p(X))]}.$
@@
 The overlap assumption can also be weakened since it is not required for $\beta_{\text{OLS}}$ to be defined.
-Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as
-$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)]}{\mathbb{E}[a(X)]},$
-where $a(X)=p(X)(1-p(X))$ and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$.
+Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as a weighted estimand satisfying the representation in (1.1) where $a(X)=p(X)(1-p(X))$ and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$.
 Here we implicitly set $W_{0}=1$ with probability 1.
 Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are $p(X)(1-p(X))$.
-Note that $\beta_{\text{OLS}}=\text{ATE}\equiv\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case when $p(X)$ or $\tau_{0}(X)$ is constant.
+Note that $\beta_{\text{OLS}}=\text{ATE}\coloneqq\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case, for example, when $p(X)$ or $\tau_{0}(X)$ is constant.
 An alternative representation of this estimand can be obtained by focusing on the subpopulation of treated units, $D=1$.
-Let $W_{0}=D$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,X]=\mathbb{E}[Y(1)-Y(0)\mid X]$, which follows from conditional independence, and let $\tilde{a}(X)=1-p(X)$.
+Let $W_{0}=D$, $w_{0}(X)=p(X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,X]=\mathbb{E}[Y(1)-Y(0)\mid X]$, which follows from conditional independence, and let $\tilde{a}(X)=1-p(X)$.
 Then, we can write
-$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]}=\frac{\mathbb{E}[\tilde{a}(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[\tilde{a}(X)w_{0}(X)]},$
-where $w_{0}(X)=\mathbb{P}(D=1\mid X)=p(X)$. Yet another representation can be obtained when focusing on the subpopulation of untreated units by letting $W_{0}=1-D$. We omit details for brevity.
-2.2 Instrumental Variables
-Consider an endogenous binary treatment $D\in\{0,1\}$ and a binary instrument $Z\in\{0,1\}$.
-Potential treatments, denoted by $(D(1),D(0))$, are linked to the realized treatment through $Z$, that is, $D=D(Z)$.
-Potential outcomes, $Y(d,z)$ for $d,z\in\{0,1\}$, may depend on both $D$ and $Z$ in the absence of an exclusion restriction.
-Let $Y=Y(D,Z)$ be the realized outcome.
-As before, let $X$ denote covariates.
-We make the following assumptions.
-Assumption 2.2 (Instrument validity).
-We have
-  1. 1.
-    Exogeneity: $(Y(0,0),Y(1,0),Y(0,1),Y(1,1),D(1),D(0))\perp\!\!\!\perp Z\mid X$;
-  2. 2.
-    Exclusion: $\mathbb{P}(Y(d,0)=Y(d,1))=1$ for $d\in\{0,1\}$;
-  3. 3.
-    First stage: $e(X)\equiv\mathbb{P}(Z=1\mid X)\in(0,1)$ and $\mathbb{P}(D(1)=1\mid X)\neq\mathbb{P}(D(0)=1\mid X)$ almost surely.
-We also make one of the following two nested monotonicity assumptions.
-Assumption 2.3 (Strong monotonicity).
-$\mathbb{P}(D(1)\geq D(0)\mid X)=1$ almost surely.
-Assumption 2.4 (Weak monotonicity).
-There exists a subset of the support of $X$ such that $\mathbb{P}(D(1)\geq D(0)\mid X)=1$ on it and $\mathbb{P}(D(1)\leq D(0)\mid X)=1$ on its complement.
-The first instrumental variables estimand we consider was originally studied by Angrist and Imbens (1995).
-In addition to the assumptions on the instrument, suppose that the model for $X$ is saturated, with $K$ possible combinations of covariate values, i.e.
-let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, where $\text{supp}(\cdot)$ is used to denote the support.
-Let $X_{S}=\left(1,\mathbbm{1}(X=x_{1}),\ldots,\mathbbm{1}(X=x_{K-1})\right)$ and $Z_{S}=\left(Z,Z\mathbbm{1}(X=x_{1}),\ldots,Z\mathbbm{1}(X=x_{K-1})\right)=ZX_{S}$, where $Z_{S}$ is the constructed instrument vector.
-The estimand in Angrist and Imbens (1995) is the following 2SLS estimand:
-$\beta_{\text{2SLS}}\equiv\left[\left(\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}W_{S}\right]\right)^{-1}\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}Y\right]\right]_{1},$
-where $W_{S}=\left(D,X_{S}\right)$, $Q_{S}=\left(Z_{S},X_{S}\right)$, and $\left[\cdot\right]_{k}$ denotes the $k$th element of the corresponding vector.
-This estimand has been studied by Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022), and the specific representation in Proposition 2.2 follows Słoczyński (2020).
-Proposition 2.2.
-Let Assumptions 2.2 and 2.4 hold, and let $X$ be discrete with finite support. Then
-$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}\left[e(X)(1-e(X))\cdot\mathbb{P}\left(D(1)\neq D(0)\mid X\right)^{2}\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]\right]}{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)^{2}]}.$
-This means that we can write $\beta_{\text{2SLS}}$ as
-$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]},$
-where $W_{0}=\mathbbm{1}(D(1)\neq D(0))$ is the population of compliers/defiers, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]$, and $a(X)=e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)$, a nonnegative weight function.
-Note that $\beta_{\text{2SLS}}=\text{LATE}\equiv\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0)]$ if and only if $a(X)$ is uncorrelated with $\tau_{0}(X)$ given $D(1)\neq D(0)$.
-The practical limitation of focusing on $\beta_{\text{2SLS}}$ is that applied researchers rarely create additional instruments by interacting the original instrument with covariates (cf.
-Blandhol, Bonney, Mogstad, and Torgovitsky, 2022), which is how $Z_{S}$ is constructed to obtain $\beta_{\text{2SLS}}$ above.
-A more practically relevant estimand is the “just-identified” IV estimand,
-$\beta_{\text{IV}}\equiv\left[\left(\mathbb{E}\left[Q^{\prime}W\right]\right)^{-1}\mathbb{E}\left[Q^{\prime}Y\right]\right]_{1},$
-where $Q=\left(Z,X\right)$ and $W=\left(D,X\right)$. To introduce one of the representations of $\beta_{\text{IV}}$ below, define
-$c(X)\equiv\operatorname{sign}\Big{(}\mathbb{P}\left[D(1)\geq D(0)\mid X\right]-\mathbb{P}\left[D(1)\leq D(0)\mid X\right]\Big{)},$
-where $\operatorname{sign}(\cdot)$ is the sign function.
-We also make the following “rich covariates” assumption on the instrument propensity score, which is implied by the saturated specification in Proposition 2.2.
-Assumption 2.5 (Rich covariates).
-$e(X)$ is linear in $X$.
-Under the instrument validity assumption, the rich covariates assumption, and either monotonicity assumption, Słoczyński (2020) obtains the following representations for the “just-identified” IV estimand.
-Proposition 2.3.
-Let Assumptions 2.2, 2.4, and 2.5 hold. Then
-$\beta_{\text{IV}}=\frac{\mathbb{E}[c(X)\cdot e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]]}{\mathbb{E}[c(X)\cdot e(X)(1-e(X))\cdot\mathbb{P}(D(1)\neq D(0)\mid X)]}.$
-If Assumptions 2.2, 2.3, and 2.5 hold instead, then
-$\beta_{\text{IV}}=\frac{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)>D(0)\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]]}{\mathbb{E}[e(X)(1-e(X))\cdot\mathbb{P}(D(1)>D(0)\mid X)]}.$
-It follows that we can write $\beta_{\text{IV}}$ as
-$\beta_{\text{IV}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}$
-under either monotonicity assumption.
-Under weak monotonicity, $W_{0}=\mathbbm{1}(D(1)\neq D(0))$, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)\neq D(0),X]$, and possibly negative weights $a(X)=c(X)\cdot e(X)(1-e(X))$.
-Under strong monotonicity, $W_{0}=\mathbbm{1}(D(1)>D(0))$, $w_{0}(X)=\mathbb{P}(D(1)>D(0)\mid X)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]$, and nonnegative weights $a(X)=e(X)(1-e(X))\geq 0$.
-2.3 Difference-in-Differences
-Now suppose units are observed for $T$ periods and, for $t\in\{1,\ldots,T\}$, denote binary treatment $D_{t}\in\{0,1\}$, potential outcomes $(Y_{t}(1),Y_{t}(0))$, and realized outcome $Y_{t}=Y_{t}(D_{t})$.
-We assume units are untreated prior to period $G\in\mathcal{G}=\{2,3,\ldots,T\}\cup\{+\infty\}$, receive the treatment in period $G$, and remain treated thereafter.
-We assume no units are treated in the first time period.
-This may include a group that remains untreated throughout, for which $G=+\infty$.
-Thus, $D_{t}=\mathbbm{1}(G\leq t)$.
-The panel is balanced, that is, no group appears or disappears over time.
-The two-way fixed effects estimand is often used in this setting, and consists of regressing the outcome on the treatment indicator, group indicators, and period indicators.
-Using partitioned regression results, it is defined as
-$\beta_{\text{TWFE}}\equiv\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}Y_{t}\right]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}^{2}\right]},$
-where $\ddot{D}_{t}=D_{t}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]$.
-We assume a version of parallel trends most similar to the one in de Chaisemartin and D’Haultfœuille (2020).
-Assumption 2.6 (Difference-in-differences).
-We have
-  1. 1.
-    $\mathcal{G}\equiv\text{supp}(G)=\{2,3,\ldots,T\}\cup\{+\infty\}$;
-  2. 2.
-    For all $t\in\{2,\ldots,T\}$ and $g,g^{\prime}\in\mathcal{G}$, we have that $\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g]=\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g^{\prime}]$.
-We use a proposition that is essentially a special case of Theorem 1 in de Chaisemartin and D’Haultfœuille (2020) to obtain a representation of the two-way fixed effects estimand as a weighted average.
-Proposition 2.4.
-Let Assumption 2.6 hold. Then
-$\beta_{\text{TWFE}}=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}\big{(}1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]\big{)}\cdot\mathbb{P}(D_{t}=1\mid G)\cdot\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid G,D_{t}=1]\Big{]}}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}\big{(}1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]\big{)}\cdot\mathbb{P}(D_{t}=1\mid G)\Big{]}}.$
-We show the above representation satisfies equation (1.1) by introducing an auxiliary variable $P$ that is uniformly distributed on $\{1,\ldots,T\}$ independently from $\{(Y_{t}(0),Y_{t}(1),G)\}_{t=1}^{T}$.
-This period variable denotes the time period and we use it to define $(Y(1),Y(0))\equiv(Y_{P}(1),Y_{P}(0))$, $Y=Y_{P}$, and $D=D_{P}$, which are potential outcomes, the realized outcome, and treatment at a random time period, respectively.
-Letting $X=(G,P)$, this means we can write $\beta_{\text{TWFE}}$ as
-$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]},$
-where $W_{0}=D$, $w_{0}(X)=\mathbb{P}(D=1\mid G,P)\in\{0,1\}$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G,P]$, and the weight function $a(X)=1-\mathbb{P}(D=1\mid G)-\mathbb{P}(D=1\mid P)+\mathbb{P}(D=1)$ is not generally nonnegative.
-[^2] Here, $\tau_{0}(X)$ is what Callaway and Sant’Anna (2021) call “the group-time average treatment effect.” A nonnegative weight function can be obtained under the assumption that group-level average treatment effects are constant over time.
-This property was described in Goodman-Bacon (2021, Section 3.1.1) and a representation of the two-way fixed effects estimand is given in his equation (16).
-The following proposition yields a simpler (but equivalent) expression for the weights in our setting.
-Proposition 2.5.
-Let Assumption 2.6 hold and let $\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid D=1,G]=\mathbb{E}[Y_{s}(1)-Y_{s}(0)\mid D=1,G]$ for any $s,t\in\{1,\ldots,T\}$. Then
-$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\mathbb{E}[Y(1)-Y(0)\mid D=1,G]]}{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)]},$
-where $a_{B}(g)=\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\geq 0$ for $g\in\{2,\ldots,T\}$.
-As is the case of the representation in Proposition 2.4, the two-way fixed effects estimand in Proposition 2.5 satisfies the representation in (1.1), with $X=G$, $W_{0}=D$, $w_{0}(X)=\mathbb{P}(D=1\mid G)$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G]$, and the weight function $a(X)=a_{B}(G)\geq 0$.
-This weight function is derived in the proof of the proposition, and we show it is equivalent to equation (16) in Goodman-Bacon (2021) in Appendix D.
+$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]}=\frac{\mathbb{E}[\tilde{a}(X)\tau_{0}(X)\mid D=1]}{\mathbb{E}[\tilde{a}(X)\mid D=1]}.$
+Yet another representation can be obtained when focusing on the subpopulation of untreated units by letting $W_{0}=1-D$. We omit details for brevity.
+We will return to this example in Section 5 after establishing conditions under which weighted estimands have a causal representation (Section 3) and identifying the size of subpopulations that are represented by these estimands (Section 4).
 3   Causal Representation of Weighted Estimands
-We now consider a class of weighted estimands that subsumes all our previous examples.
-In this section, we show necessary and sufficient conditions for an estimand of this class to have a causal representation as an average treatment effect over a population.
-Then, we compute the largest size of a population that is compatible with the weighted estimand.
-The size of this population relative to $\mathbb{P}(W_{0}=1)$ will be a measure of the internal validity of our estimand relative to the target estimand $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
-The absolute size of this population will be a measure of the representativeness of the estimand with respect to the population distribution the data is drawn from.
-Recall the earlier setting where we let $D\in\{0,1\}$ denote a binary treatment variable, and let $(Y(1),Y(0))$ denote the corresponding potential outcomes under treatment and control, respectively.
-Let $X\in\text{supp}(X)\subseteq\mathbb{R}^{d_{X}}$ denote a $d_{X}$-vector of covariates.
+In this section, we consider a general class of weighted estimands.
+We show necessary and sufficient conditions for an estimand in this class to have a causal representation as an average treatment effect over a subpopulation.
+We provide these conditions under various assumptions—including no assumptions—on treatment effect heterogeneity.
+3.1   Preliminaries
+Recall the earlier setting where we let $D\in\{0,1\}$ denote a treatment variable, and let $(Y(0),Y(1))$ denote the corresponding potential outcomes.
+Let $X\in\text{supp}(X)\subseteq\mathbb{R}^{d_{X}}$ denote a $d_{X}$-vector of covariates, where $\text{supp}(\cdot)$ denotes the support.
 We suppose that $(Y(1),Y(0),D,X)$ are drawn from a common population distribution $F_{Y(1),Y(0),D,X}$.
-Let $W_{0}\in\{0,1\}$ be an indicator variable used to denote a subpopulation $\{W_{0}=1\}$.
-Throughout this paper, we assume that $\mathbb{P}(W_{0}=1)>0$, so that this subpopulation has a positive mass, which avoids technical issues associated with conditioning on zero-probability events.
-For example, this subpopulation can be the entire population by setting $W_{0}=1$ almost surely, the treated units by setting $W_{0}=D$, or units complying with binary instrument $Z$ by setting $W_{0}=\mathbbm{1}(D(1)>D(0))$.
-Let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ denote the conditional average treatment effect given $X$ in subpopulation $W_{0}$.
-This is defined for all values of $X$ such that $w_{0}(X)>0$.
+Let $W_{0}\in\{0,1\}$ be an indicator variable used to denote a subpopulation $\{W_{0}=1\}$ and let $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ denote the conditional average treatment effect given $X$ in that subpopulation.
+For example, this subpopulation can be the entire population by setting $W_{0}=1$ almost surely, in which case $\tau_{0}$ denotes the usual CATE function.
+It can also denote the subpopulation of treated units by setting $W_{0}=D$.
+In the presence of a binary instrument $Z$, the complier subpopulation is defined by setting $W_{0}=\mathbbm{1}(D(1)>D(0))$, where $D(1)$ and $D(0)$ are potential treatments.
+In this case, $\tau_{0}$ denotes the conditional local average treatment effect.
+[…]
 Also recall the weighted estimands of equation (1.1):
-$\mu(a,\tau_{0})$ $\equiv\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]},$
-where $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)$. [^3] While $\tau_{0}(X)$ is only defined for $w_{0}(X)>0$, we set $\tau_{0}(X)w_{0}(X)=0$ when $w_{0}(X)=0$.
-The estimands we consider have the above representation and satisfy the following regularity conditions.
-Assumption 3.1 (Regularity).
+$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[a(X)w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[a(X)w_{0}(X)]}=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}.$
+The estimands we consider have the above representation and satisfy the following regularity condition.
+Assumption 3.1.
 Let $\mathbb{E}[\tau_{0}(X)^{2}]<\infty$, $\mathbb{E}[a(X)^{2}]<\infty$, and $\mathbb{E}[a(X)\mid W_{0}=1]>0$.
-The first two restrictions are weak regularity assumptions that ensure the finiteness of the numerator of $\mu(a,\tau_{0})$.
+The first two restrictions ensure the existence of the numerator of $\mu(a,\tau_{0})$.
 We rule out $\mathbb{E}[a(X)\mid W_{0}=1]=0$ since it implies the estimand does not exist.
 The estimand in (1.1) is unchanged if the sign of $a(X)$ is reversed, so $\mathbb{E}[a(X)\mid W_{0}=1]>0$ is a sign normalization.
-3.1 Alternative Representations of Weighted Estimands
 The weighted estimands of equation (1.1) can also be written as a weighted sum when $X$ is discrete, or an integral when $X$ is continuous.
-In the discrete case, let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, $W_{0}=1$ almost surely, and let $p_{k}\equiv\mathbb{P}(X=x_{k})>0$ for $k=1,\ldots,K$.
+In the discrete case, let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, let $p_{k}\coloneqq\mathbb{P}(X=x_{k})>0$ for $k=1,\ldots,K$, and assume $W_{0}=1$ almost surely for simplicity.
 Then,
-$\mu(a,\tau_{0})$ $=\sum_{k=1}^{K}\omega_{k}\tau_{0}(x_{k})\qquad\text{ where }\qquad\omega_{k}=\frac{a(x_{k})p_{k}}{\sum_{l=1}^{K}a(x_{l})p_{l}},$ (3.1)
+$\mu(a,\tau_{0})$ $=\sum_{k=1}^{K}\omega_{k}\tau_{0}(x_{k})\qquad\text{ where }\qquad\omega_{k}=a(x_{k})p_{k}\bigg{/}\sum_{l=1}^{K}a(x_{l})p_{l},$ (3.1)
 which are weights that sum to one.
-The representations in (1.1) and (3.1) are equivalent as we can obtain $a(x_{k})$ (up to scale) as the ratio $\omega_{k}/p_{k}$, and $\omega_{k}$ is defined as a function of $(a(\cdot),p_{\cdot})$ in equation (3.1) above.
-From (3.1), we can see that $a(x_{k})$ being constant ensures $\omega_{k}=p_{k}$, or that the estimand is the ATE. Moreover,
-$\frac{a(x_{k})}{a(x_{k^{\prime}})}=\frac{\omega_{k}}{\omega_{k^{\prime}}}\big{/}\frac{p_{k}}{p_{k^{\prime}}},$
-which is the ratio of the relative weights of covariate cells $\{X=x_{k}\}$ and $\{X=x_{k^{\prime}}\}$ in the estimand $(\omega_{k}/\omega_{k^{\prime}})$ and in the population $(p_{k}/p_{k^{\prime}})$.
-The inequality $a(x_{k})>a(x_{k^{\prime}})$ indicates that covariate cell $\{X=x_{k}\}$ is overweighted relative to $\{X=x_{k^{\prime}}\}$, when compared to their relative weights in the population.
-Alternatively, consider the case where $X$ is continuously distributed with density [^4] With respect to the Lebesgue measure. $f_{X}$. Still assuming $W_{0}=1$ almost surely, we can write
-[…]
-is a weight function that integrates to 1. We focus on the representation in equation (1.1) since it seamlessly accommodates discrete, continuous, and mixed covariates.
+The representations in (1.1) and (3.1) are equivalent since we can obtain $a(x_{k})$ (up to scale) as the ratio $\omega_{k}/p_{k}$, and $\omega_{k}$ is defined as a function of $\{(a(x_{k}),p_{k})\}_{k=1}^{K}$ in equation (3.1).
+From equation (3.1), we can see that $a(x_{k})$ being constant ensures $\omega_{k}=p_{k}$, or that the estimand is the ATE.
+Moreover, $\frac{a(x_{k})}{a(x_{k^{\prime}})}=\frac{\omega_{k}}{\omega_{k^{\prime}}}\big{/}\frac{p_{k}}{p_{k^{\prime}}}$, which is the ratio of the relative weights of covariate cells $\{X=x_{k}\}$ and $\{X=x_{k^{\prime}}\}$ in the estimand $(\omega_{k}/\omega_{k^{\prime}})$ and in the population $(p_{k}/p_{k^{\prime}})$.
+The inequality $a(x_{k})>a(x_{k^{\prime}})$ indicates that covariate cell $\{X=x_{k}\}$ is overweighted by the estimand relative to $\{X=x_{k^{\prime}}\}$, when compared to their relative weights in the population.
+Similar algebra can be used to write the estimand as an integral when $X$ is continuously distributed.
+We instead focus on the representation in equation (1.1) since it seamlessly accommodates discrete, continuous, and mixed covariates.
 3.2   Regular Subpopulations
@@
 Formally, $\{W^{*}=1\}$ forms a subpopulation if $\{W^{*}=1\}\subseteq\{W_{0}=1\}$ or, equivalently, if $W^{*}\leq W_{0}$ almost surely.
-We impose some structure on this problem by restricting how these subpopulations may be formed. We will consider what we call “regular subpopulations,” which we define here.
+We impose some structure on this problem by restricting how these subpopulations may be formed. We only consider what we call “regular subpopulations,” defined here.
 Definition 3.1.
@@
   1. 1.
-    (Inclusion) $\mathbb{P}(W^{*}=1\mid W_{0}=0)=0$;
+    (Inclusion) $\mathbb{P}(W_{0}=1\mid W^{*}=1)=1$;
   2. 2.
@@
 $\text{SP}(W_{0})$ =$=\{W^{*}\in\{0,1\}:W^{*}\text{ is a regular subpopulation of }W_{0}\}.$
-We consider subpopulations with positive masses that are subsets of $\{W_{0}=1\}$.
-The second and main requirement is that regular subpopulations do not depend on potential outcomes when conditioning on $X$ and the original population $W_{0}=1$.
+These subpopulations have positive masses and are subsets of $\{W_{0}=1\}$.
+The other substantive requirement is that they do not depend on potential outcomes when conditioning on $X$ and the original population $W_{0}=1$.
 While this may seem restrictive, it allows for rich and natural classes of subpopulations.
-For example, consider the unconfoundedness restriction of Section 2.1 and let $W_{0}$ be the entire population, i.e.
-$\mathbb{P}(W_{0}=1)=1$.
+For example, consider the unconfoundedness restriction of Section 2 and let $W_{0}$ be the entire population, i.e., $\mathbb{P}(W_{0}=1)=1$.
 In this case, regular subpopulations must satisfy $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X$, or be unconfounded.
-Regular subpopulations include the population of all treated (or untreated) individuals, i.e.
-$W^{*}=D$ (or $W^{*}=1-D$), and any subpopulation characterized by a subset of $\text{supp}(X)$.
-More generally, it includes any subpopulation that can be described through a combination of $(D,X,U)$ where $U$ is independent from $(Y(1),Y(0),X)$.
+Regular subpopulations include the population of all treated (or untreated) individuals, i.e., $W^{*}=D$ (or $W^{*}=1-D$), and any subpopulation characterized by a subset of $\text{supp}(X)$.
+More generally, they include any subpopulation that can be described through a combination of $(D,X,U)$ where $U$ is independent from $(Y(1),Y(0),X)$.
 For example, a subpopulation characterized by “fraction $a(x)$ of units with covariate $X=x$ for all $x\in\text{supp}(X)$” can be constructed as $W^{*}=\mathbbm{1}(U\leq a(X))$ where $U\sim\text{Unif}(0,1)$ is independent from $(Y(1),Y(0),X)$.
-The conditional independence requirement rules out subpopulations that directly depend on the potential outcomes such as $W^{*}=\mathbbm{1}(Y(1)\geq Y(0))$, the subpopulation of those who benefit from treatment.
+The conditional independence requirement rules out subpopulations that directly depend on the potential outcomes such as $W^{*}=\mathbbm{1}(Y(1)\geq Y(0))$, i.e., the subpopulation of those who benefit from treatment.
 Note that $\mathbb{P}(W^{*}=1\mid X)=\mathbb{P}(Y(1)\geq Y(0)\mid X)$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(Y(1)\geq Y(0))$ are not point-identified under unconfoundedness.
+Another way to view this requirement is that regular subpopulations are policy relevant in the sense that we could design a policy that targets a regular subpopulation.
+Indeed, a policy maker may observe $X$ and can use $U$ to randomly target a fraction of units with specific values of $X$, but cannot observe potential outcomes.
 These particular subpopulations enjoy a number of useful properties. Two of them are characterized in the following proposition.
 Proposition 3.1 (Properties of regular subpopulations).
-Let $\mathbb{P}(W_{0}=1)>0$ and $W^{*}\in\text{SP}(W_{0})$.
+Suppose that $\mathbb{P}(W_{0}=1)>0$ and $W^{*}\in\text{SP}(W_{0})$. Suppose Assumption 3.1 holds. Let $\underline{w}^{*}(x)\coloneqq\mathbb{P}(W^{*}=1\mid X=x,W_{0}=1)$. Then,
   1. 1.
-    Let $\mathbb{P}(W^{*}=1\mid W_{0}=1,X)>0$. Then,
-    $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X]$ $=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X].$ (3.3)
+    $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ when $\underline{w}^{*}(X)>0$;
   2. 2.
-    Let $\mathbb{E}[\tau_{0}(X)^{2}]<\infty$. Then,
-    $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}=\mu(\underline{w}^{*},\tau_{0}),$ (3.4)
-    where $\underline{w}^{*}(x)=\mathbb{P}(W^{*}=1\mid X=x,W_{0}=1)$.
+    $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}=\mu(\underline{w}^{*},\tau_{0})$.
 The first part of this proposition shows that average effects within the original population $W_{0}$ and regular subpopulation $W^{*}$ are the same when conditioning on $X$.
-For example, this holds under unconfoundedness for the subpopulation of treated individuals, $W^{*}=D$.
-The second property allows us to write the average effect for $W^{*}=1$ using the same functional $\mu(\cdot,\cdot)$ that was used to characterize the class of estimands we analyze.
-This property will be used when studying the mapping between weighted estimands and average effects for regular subpopulations of $W_{0}$.
-We conclude this subsection by showing that regular subpopulations are transitive, or that regular subpopulations of a regular subpopulation of $W_{0}$ are also regular subpopulations of $W_{0}$.
-Lemma 3.1 (Transitivity of regular subpopulations).
-Let $W^{*}$ be a regular subpopulation of $W^{\prime}$ and let $W^{\prime}$ be a regular subpopulation of $W_{0}$. Then, $W^{*}$ is a regular subpopulation of $W_{0}$.
+For example, this holds under unconfoundedness for the subpopulation of treated units, $W^{*}=D$.
+The second property allows us to write the average treatment effect for $W^{*}=1$ using the same functional $\mu(\cdot,\cdot)$ that was used to characterize the class of estimands we analyze.
+This property will be used when studying the mapping between weighted estimands and average treatment effects for regular subpopulations of $W_{0}$.
 3.3   Existence of a Causal Representation for Weighted Estimands
@@
 This is the set of regular subpopulations of $W_{0}$ such that the estimand $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for all $\tau_{0}$ functions in the set $\mathcal{T}$.
-If it is empty, then estimand $\mu(a,\tau_{0})$ cannot be written as an average treatment effect over a regular subpopulation of $W_{0}$ uniformly in $\tau_{0}\in\mathcal{T}$.
+If the set $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty, then estimand $\mu(a,\tau_{0})$ cannot be written as an average treatment effect over a regular subpopulation of $W_{0}$ uniformly in $\tau_{0}\in\mathcal{T}$.
 We use this set to formally define a notion of uniform causal representation.
@@
 Recall that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu(\underline{w}^{*},\tau_{0})$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$, so $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})$ if $\mu(a,\tau_{0})=\mu(\underline{w}^{*},\tau_{0})$ for all $\tau_{0}\in\mathcal{T}$.
-We examine further two main cases for the set $\mathcal{T}$.
+We further examine several cases for the set $\mathcal{T}$.
 3.3.1   Existence Uniformly in $\tau_{0}$
@@
 We denote this class by
-$\mathcal{T}_{\text{all}}$$\equiv\{\tau_{0}:\mathbb{E}[\tau_{0}(X)^{2}]<\infty\}.$
-In this function class, the existence of a causal representation is equivalent to the weights being nonnegative.
+$\mathcal{T}_{\text{all}}$$\coloneqq\{\tau_{0}:\mathbb{E}[\tau_{0}(X)^{2}]<\infty\}.$
+In this function class, we show the existence of a causal representation is equivalent to the estimand’s weights being nonnegative.
+In what follows, let $a_{\max}\coloneqq\sup(\text{supp}(a(X)\mid W_{0}=1))$ be the essential supremum of $a(X)$ given $W_{0}=1$.
 Theorem 3.1.
-Let Assumption 3.1 hold and let $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$.
-Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$ if and only if $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$.
-A uniform (in $\mathcal{T}_{\text{all}}$) causal representation exists if and only if $a(X)$ is nonnegative on the support of $X\mid W_{0}=1$.
-To give some intuition on why the sign of $a(X)$ must be nonnegative with probability 1, partition $\text{supp}(X\mid W_{0}=1)$ in $\mathcal{X}^{+}\equiv\{x\in\text{supp}(X\mid W_{0}=1):a(x)\geq 0\}$ and $\mathcal{X}^{-}\equiv\{x\in\text{supp}(X\mid W_{0}=1):a(x)<0\}$.
-If $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$, then $\mathbb{P}(X\in\mathcal{X}^{+}\mid W_{0}=1)>0$ and $\mathbb{P}(X\in\mathcal{X}^{-}\mid W_{0}=1)>0$ under Assumption 3.1.
-Then, setting $\tau^{-}(X)=\mathbbm{1}(X\in\mathcal{X}^{-})$ and $\tau^{+}(X)=\mathbbm{1}(X\in\mathcal{X}^{+})$ yields estimands of different signs, i.e.
-$\mu(a,\tau^{-})<0<\mu(a,\tau^{+})$, despite $\tau^{-}$ and $\tau^{+}$ being of the same sign.
-In other words, as is well known, nonnegative CATEs can be combined into a negative estimand when weights are negative.
-Hence, such an estimand is not weakly causal (Blandhol, Bonney, Mogstad, and Torgovitsky, 2022).
-Conversely, if $a(X)\geq 0$, our proof constructively defines a regular subpopulation $W^{*}$ for which the average effect is equal to the weighted estimand $\mu(a,\tau_{0})$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$.
+Let $\mu(a,\tau_{0})$ be an estimand satisfying equation (1.1).
+Suppose Assumption 3.1 holds and that $a_{\max}<\infty$.
+Then, $\mu(a,\tau_{0})$ has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$ if and only if
+$\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1.$
+A uniform (in $\mathcal{T}_{\text{all}}$) causal representation exists if and only if $a(X)$ is nonnegative when $W_{0}=1$.
+To give some intuition on why the sign of $a(X)$ must be nonnegative with probability 1, we present a contradiction that occurs when $a(X)$ can be negative.
+Suppose that $\mathbb{P}(a(X)<0\mid W_{0}=1)>0$ and consider the “adversarial” CATE function $\tau^{-}(X)=\mathbbm{1}(a(X)<0)$.
+This CATE function is nonnegative for all $X$, and implies a positive average effect only for units with negative weights.
+However, it yields a strictly negative weighted estimand, $\mu(a,\tau^{-})=\mathbb{E}[a(X)\mathbbm{1}(a(X)<0)\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]<0$.
+Clearly, $\mu(a,\tau^{-})$ cannot be the average treatment effect for any subpopulation of $W_{0}$, because averaging a nonnegative CATE function over any subpopulation cannot yield a negative average.
+Conversely, if $a(X)\geq 0$, our proof constructively defines a regular subpopulation $W^{*}$ for which the average treatment effect is equal to the weighted estimand $\mu(a,\tau_{0})$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$.
 Let
-$W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0},$
+$W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{a_{\max}}\right)\cdot W_{0},$
 where $U\sim\text{Unif}(0,1)\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$.
 This is a regular subpopulation of $W_{0}$ for which the probability of inclusion, conditional on $X$ and $W_{0}=1$, is proportional to $a(X)$.
-The condition $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$ restricts our attention to subpopulations with positive mass.
-From this construction, we can see that $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ is proportional to $a(X)$, and therefore $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu(\underline{w}^{*},\tau_{0})=\mu(a,\tau_{0})$ uniformly in $\tau_{0}$.
+We can also interpret $\mu(a,\tau_{0})$ as the average effect of an intervention in which units with covariate value $X$ are treated with probability $a(X)/a_{\max}$ given $W_{0}=1$.
+From this construction, we can see that $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ is proportional to $a(X)$, and therefore $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu(\underline{w}^{*},\tau_{0})=\mu(a,\tau_{0})$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$.
+The condition $a_{\max}<\infty$ restricts our attention to subpopulations with positive mass.
+We note that $a_{\max}$ is bounded above in each of our theoretical examples in Sections 2 and 5, implying that this condition trivially holds in these cases.
+As mentioned earlier, the condition that weights are nonnegative is well established.
+Blandhol, Bonney, Mogstad, and Torgovitsky (2022) show that it is equivalent to an estimand being “weakly causal,” which means that it will match the sign of $\tau_{0}$ whenever that sign is the same across all units.
+Thus, in the class of weighted estimands we consider, estimands have a causal representation uniformly in $\mathcal{T}_{\text{all}}$ if and only if they are weakly causal.
+This connection is formally established in Appendix B.
 3.3.2   Existence for a Given $\tau_{0}$
 We now provide an existence result that requires the causal representation to exist only for the given $\tau_{0}$, rather than uniformly for $\tau_{0}$ in the larger set $\mathcal{T}_{\text{all}}$.
-This result’s assumptions depend on the nature of the CATE function $\tau_{0}$ in the population, whereas Theorem 3.1’s condition depended only on the weight function $a(X)$, and the nature of the covariates’ support.
+The following result depends on the CATE function $\tau_{0}$ in the population, whereas Theorem 3.1’s condition depended only on the weight function $a(X)$.
 Thus, the distribution of the potential outcomes will have an impact on the existence of a causal representation given $\tau_{0}$.
@@
 Theorem 3.2.
-Let Assumption 3.1 hold. Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ if and only if
-$\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W_{0})\equiv\{t\in\mathbb{R}:\mathbb{P}(\tau_{0}(X)\leq t\mid W_{0}=1)>0\text{ and }\mathbb{P}(\tau_{0}(X)\geq t\mid W_{0}=1)>0\}.$ (3.5)
-The existence condition in Theorem 3.2 is weaker than the one in Theorem 3.1 since we no longer require this representation to be valid for all $\tau_{0}\in\mathcal{T}_{\text{all}}$, but rather for just one its element.
-The necessary and sufficient condition in this theorem is rather weak, since it only requires that the estimand is in the convex hull of the support of the CATEs.
+Let $\mu(a,\tau_{0})$ be an estimand satisfying equation (1.1). Suppose Assumption 3.1 holds. Then, $\mu(a,\tau_{0})$ has a causal representation for $\tau_{0}$ if and only if
+$\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W_{0})\coloneqq\{t\in\mathbb{R}:\;$ $\mathbb{P}(\tau_{0}(X)\leq t\mid W_{0}=1)>0$
+[…]
+The existence condition in Theorem 3.2 is weaker than the one in Theorem 3.1 since we no longer require this representation to be valid for any CATE function, but rather just for the one that is identified from the population.
+The necessary and sufficient condition in this theorem only requires that the estimand is in the convex hull of the support of the CATEs.
 This means $\mu(a,\tau_{0})$ has a causal representation even with negative weights, as long as there are CATEs smaller and greater than $\mu(a,\tau_{0})$.
-We can see this support condition holds for all $\tau_{0}\in\mathcal{T}_{\text{all}}$ if and only if $\mu(a,\tau_{0})$ is in the support of $\tau_{0}(X)$ for all $\tau_{0}$.
-This is precisely the case when the weights $a(X)$ are nonnegative since it guarantees $\inf(\text{supp}(\tau_{0}(X)\mid W_{0}=1))\leq\mu(a,\tau_{0})\leq\sup(\text{supp}(\tau_{0}(X)\mid W_{0}=1))$ for any function $\tau_{0}$.
-3.3.3 Intermediate Cases
-Analyzing the representativeness of an estimand under no restrictions on $\tau_{0}$ could be viewed as unnecessarily conservative in some settings.
+We can see this support condition holds for all $\tau_{0}\in\mathcal{T}_{\text{all}}$ if and only if $\mu(a,\tau_{0})$ is in the support of $\tau_{0}(X)$ for any $\tau_{0}\in\mathcal{T}_{\text{all}}$.
+This is precisely the case when the weights $a(X)$ are nonnegative since it guarantees $\inf(\text{supp}(\tau_{0}(X)\mid W_{0}=1))\leq\mu(a,\tau_{0})\leq\sup(\text{supp}(\tau_{0}(X)\mid W_{0}=1))$ for any $\tau_{0}$.
+3.3.3   Existence in Intermediate Cases
+Analyzing the causal representation of an estimand under no restrictions on $\tau_{0}$ could be viewed as unnecessarily conservative in some settings.
 At the other extreme, assuming knowledge of $\tau_{0}$ may be unrealistic, especially in scenarios where $X$ has many components which makes the estimation of $\tau_{0}$ more challenging.
@@
 In some economic applications one may posit that $\tau_{0}$ is monotonic or convex in some components of $X$, or positive/negative over a subset of $\text{supp}(X\mid W_{0}=1)$.
-In these cases, the existence of a causal representation may occur under weaker conditions than Theorem 3.1, but stronger than Theorem 3.2.
+In these cases, the existence of a causal representation may occur under weaker conditions than those in Theorem 3.1, but stronger than those in Theorem 3.2.
 In particular, one may be able to relax the requirement that $a(X)\geq 0$ without requiring that $\tau_{0}$ be completely known to the researcher.
@@
 Proposition 3.2.
-Let Assumption 3.1 hold and define
-$\mathcal{T}_{\text{lin}}$ =$=\{\tau_{0}(X)=c+d^{\prime}X:(c,d)\in\mathbb{R}^{1+d_{X}}\}.$
-Then, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$ if and only if
+Let $\mu(a,\tau_{0})$ be an estimand satisfying equation (1.1). Suppose $X$ has finite support. Suppose Assumption 3.1 holds and define
+$\mathcal{T}_{\text{lin}}$ =$=\{\tau_{0}\in\mathcal{T}_{\text{all}}:\tau_{0}(X)=c+d^{\prime}X:(c,d)\in\mathbb{R}^{1+d_{X}}\}.$
+Then, $\mu(a,\tau_{0})$ has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{lin}}$ if and only if
 $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}\in\text{conv}(\text{supp}(X\mid W_{0}=1)),$
 where $\text{conv}(\cdot)$ denotes the convex hull.
-The above proposition shows that once the class of CATE functions is restricted, the requirement that $a(X)$ be nonnegative is no longer needed for the existence of a uniform causal representation for an estimand.
+The above proposition shows that placing restrictions on $\mathcal{T}$ may remove the requirement that $a(X)\geq 0$ for the existence of a uniform causal representation for an estimand.
 In particular, the requirement here is that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$ lies in the convex hull of the support of $X$ given $W_{0}=1$.
 When $X$ is scalar, this consists of an interval.
-This condition does not require $a(X)$ be nonnegative: for example, if $\text{supp}(X)=\{0,1,2\}$, $\mathbb{P}(X=x)=1/3$ for $x\in\{0,1,2\}$, and $W_{0}=1$ almost surely, then any combination of $(a(0),a(1),a(2))$ values such that $\frac{a(1)+2a(2)}{a(0)+a(1)+a(2)}\in[0,2]$ implies a causal representation.
-For example, $(a(0),a(1),a(2))=(1,-1,1)$ contains a negative element, and thus $\mathbb{P}(a(X)\geq 0)=2/3<1$, but $\frac{a(1)+2a(2)}{a(0)+a(1)+a(2)}=1\in[0,2]$, implying that the corresponding weighted estimand has a causal representation.
+This condition does not require $a(X)$ be nonnegative.
+For example, if $\text{supp}(X)=\{0,1,2\}$ and $W_{0}=1$ almost surely, then any combination of values of $(a(0),a(1),a(2))$ such that $\mathbb{E}[a(X)X]/\mathbb{E}[a(X)]\in[0,2]=\text{conv}(\text{supp}(X))$ implies a causal representation.
+Let $\mathbb{P}(X=x)=1/3$ for $x\in\{0,1,2\}$ and $(a(0),a(1),a(2))=(1,-1,1)$.
+Here units with $X=1$ have a negative weight, but $\mathbb{E}[a(X)X]/\mathbb{E}[a(X)]=1\in[0,2]$, implying that the corresponding weighted estimand has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{lin}}$.
+The result is stated for discrete $X$, but an estimand with negative weights can have a causal representation even when $X$ has continuous components.
 We consider another class of CATE functions that restricts their heterogeneity. For $K\geq 0$, let
-$\mathcal{T}_{\text{BV}}(K)=\left\{\tau\in\mathcal{T}_{\text{all}}:\sup_{x,x^{\prime}\in\text{supp}(X\mid W_{0}=1)}|\tau(x)-\tau(x^{\prime})|\leq K\right\}.$
-This function class bounds the variation of the CATE function.
+$\mathcal{T}_{\text{BD}}(K)=\left\{\tau_{0}\in\mathcal{T}_{\text{all}}:\sup_{x,x^{\prime}\in\text{supp}(X\mid W_{0}=1)}|\tau_{0}(x)-\tau_{0}(x^{\prime})|\leq K\right\}.$
+This function class uniformly bounds differences of the CATE function.
 When $K=0$, the CATE function is constant, and thus equal to $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
 When $K>0$, CATEs may differ in value, but the maximum discrepancy between two CATEs is bounded above by $K$.
-We show that restricting the CATEs to satisfy this bounded variation assumption does not remove the requirement that $a(X)$ be nonnegative, unless $K=0$, in which case all $a$ functions yield a causal representation uniformly in $\mathcal{T}_{\text{BV}}(0)$.
+We show that restricting the CATEs to satisfy this bounded difference assumption does not remove the requirement that $a(X)$ be nonnegative, unless $K=0$, in which case all $a(\cdot)$ functions yield a causal representation uniformly in $\mathcal{T}_{\text{BD}}(0)$.
 We formalize this in the next proposition.
 Proposition 3.3.
-Let Assumption 3.1 hold.
-If $K>0$, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ if and only if $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$.
-If $K=0$, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(0))$ for all $a$.
-The last two propositions show that the impact of restrictions on $\tau_{0}$ on the requirement that $a(X)$ be nonnegative depends on the nature of these restrictions.
-Generalizations to additional or empirically motivated function classes is left for future work.
+Let $\mu(a,\tau_{0})$ be an estimand satisfying equation (1.1).
+Suppose Assumption 3.1 holds.
+Then, $\mu(a,\tau_{0})$ has a causal representation uniformly in $\mathcal{T}_{\text{BD}}(K)$ when $K>0$ if and only if
+$\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1.$
+The estimand $\mu(a,\tau_{0})$ has a causal representation uniformly in $\mathcal{T}_{\text{BD}}(0)$ for any $a(\cdot)$.
+To understand this proposition, consider the adversarial CATE function $\tau^{-}(X)=K\cdot\mathbbm{1}(a(X)<0)$, a member of $\mathcal{T}_{\text{BD}}(K)$, and assume $\mathbb{P}(a(X)\geq 0)<1$.
+Then we obtain the same contradiction we discussed after Theorem 3.1, where the CATE is nonnegative for all covariate values but the estimand is negative.
+These last two propositions show that the impact of restrictions on $\tau_{0}$ on the requirement that $a(X)$ be nonnegative critically depends on the nature of these restrictions.
+Generalizations to additional or empirically motivated function classes are left for future work.
 4   Quantifying the Internal Validity of Weighted Estimands
 Many estimands will admit causal representations, but their associated subpopulations $\{W^{*}=1\}$ will generally differ.
-Also, a weighted estimand may not always correspond to the target estimand a researcher is interested in, such as $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ if the average effect in $\{W_{0}=1\}$ is the parameter of interest.
-However, the set of $W^{*}$ corresponding to a weighted estimand can be used to understand how representative the weighted estimand is of the target.
+Also, a weighted estimand may not always correspond to the target estimand a researcher is interested in.
+For example, a researcher may be interested in setting $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the average effect in population $\{W_{0}=1\}$, as the target parameter.
+In general, this parameter differs from $\mu(a,\tau_{0})$.
+However, the set of subpopulations corresponding to a weighted estimand can be used to understand how representative the weighted estimand is of the target.
 For example, we may seek estimands for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ attains values closest to 1, since they have a higher degree of internal validity with respect to the target $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
-At one extreme, an estimand for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)=1$ is perfectly representative of $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ and would be deemed to have the highest degree of internal validity for this target parameter.
-Formally, we ask what is the sharp upper bound on $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ for any regular subpopulation $W^{*}$ of $W_{0}$ such that the weighted estimand $\mu(a,\tau_{0})$ has a causal representation as the average treatment effect over subpopulation $W^{*}$.
-We denote this object using the following notation:
-$\overline{P}(a,W_{0};\mathcal{T})=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})}\mathbb{P}(W^{*}=1\mid W_{0}=1).$ (4.1)
-We also set $\overline{P}(a,W_{0};\mathcal{T})=0$ when $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty.
+At one extreme, an estimand for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)=1$ would be deemed to have the highest degree of internal validity for this target parameter since it would equal $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
+We convert this interpretation in a formal measure of internal validity that we define here.
+Definition 4.1 (Internal validity).
+Let
+$\overline{P}(a,W_{0};\mathcal{T})=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})}\mathbb{P}(W^{*}=1\mid W_{0}=1)$
+denote the measure of internal validity of weighted estimand $\mu(a,\tau_{0})$ over function class $\mathcal{T}$.
+Formally, $\overline{P}(a,W_{0};\mathcal{T})$ is the sharp upper bound on $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ for any regular subpopulation $W^{*}$ of $W_{0}$ such that the weighted estimand $\mu(a,\tau_{0})$ has a causal representation as the average treatment effect over subpopulation $W^{*}$.
+Note that we set $\overline{P}(a,W_{0};\mathcal{T})=0$ when $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty.
 This object depends on the chosen function class $\mathcal{T}$, as did Theorems 3.1 and 3.2 in the previous section.
-Given the above terminology, we call $\overline{P}(a,W_{0};\mathcal{T})$ a measure of the internal validity of estimand $\mu(a,\tau_{0})$, assuming that $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target.
-We can also compute the maximum value of $\mathbb{P}(W^{*}=1)$ across $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})$, which equals $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W_{0}=1)\cdot\mathbb{P}(W_{0}=1)$ since $W^{*}$ is a subpopulation of $W_{0}$.
-This maximum value of $\mathbb{P}(W^{*}=1)$ gives the internal validity of the weighted estimand with respect to target estimand $\mathbb{E}[Y(1)-Y(0)]$, the average treatment effect in the population from which the sample is drawn.
-It can also be understood as a measure of representativeness of the weighted estimand with respect to that population.
-As earlier, we break down our results in two cases, the first being when $\tau_{0}$ is unrestricted.
+Given the above terminology and assuming that $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target, we call $\overline{P}(a,W_{0};\mathcal{T})$ a measure of the internal validity of estimand $\mu(a,\tau_{0})$, and we use this definition in the remainder of the paper.
+We can also compute the maximum value of $\mathbb{P}(W^{*}=1)$ across $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})$, which measures the largest share of the entire population for which the weighted estimand has a causal representation.
+We refer to this measure as a measure of representativeness.
+The measures of internal validity and representativeness are the same when $W_{0}=1$ almost surely.
+Definition 4.2 (Representativeness).
+Let
+$\overline{P}(a,W_{0};\mathcal{T})\cdot\mathbb{P}(W_{0}=1)=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T})}\mathbb{P}(W^{*}=1)$
+denote the measure of representativeness of weighted estimand $\mu(a,\tau_{0})$ over function class $\mathcal{T}$.
+Note that $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W_{0}=1)\cdot\mathbb{P}(W_{0}=1)$ since $W^{*}$ is a subpopulation of $W_{0}$.
+The maximum value of $\mathbb{P}(W^{*}=1)$ gives the internal validity of the weighted estimand with respect to target estimand $\mathbb{E}[Y(1)-Y(0)]$, the average treatment effect in the population from which the sample is drawn.
+Our measures of internal validity and representativeness are closely linked and a subpopulation will maximize $\mathbb{P}(W^{*}=1)$ if and only if it maximizes $\mathbb{P}(W^{*}=1\mid W_{0}=1)$.
+We will also show how to use these measures to obtain simple bounds on target estimands.
+We now derive explicit expressions for $\overline{P}(a,W_{0};\mathcal{T})$. We focus on two cases, the first being when $\tau_{0}$ is unrestricted.
 4.1   Quantifying Internal Validity Uniformly in $\tau_{0}$
@@
 Theorem 4.1.
-Let Assumption 3.1 hold and let $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. Then
-$\sup_{W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})}\mathbb{P}(W^{*}=1\mid W_{0}=1)=\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\begin{cases}\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\sup(\text{supp}(a(X)\mid W_{0}=1))}&\text{ if }\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1\\ 0&\text{ if }\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1.\end{cases}$ (4.2)
-Here we see that the maximum size of a subpopulation characterizing the estimand $\mu(a,\tau_{0})$ depends on $a(X)$ through two terms: the denominator of the estimand, $\mathbb{E}[a(X)\mid W_{0}=1]$, and the maximum value of $a(x)$ over the support of $X\mid W_{0}=1$.
-Note that $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ is invariant to the scale of $a(X)$, as is the estimand $\mu(a,\tau_{0})$.
+Let $\mu(a,\tau_{0})$ be an estimand satisfying equation (1.1). Suppose Assumption 3.1 holds and that $a_{\max}<\infty$. If $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$, then
+$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{a_{\max}}.$
+If $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$, then $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=0$.
+Here we see that the maximum size of a subpopulation characterizing the estimand $\mu(a,\tau_{0})$ depends on $a(X)$ through two terms: its conditional mean in the numerator, and its supremum $a_{\max}$ in the denominator.
+This bound can be computed at what Imbens and Rubin (2015) call the “design stage” of the study, that is, without any knowledge of the conditional distribution of the outcome.
+{adjustwidth}
+-1.25in-1.25in Refer to caption Note: $X$ is a single continuously distributed covariate with density $f_{X}$.
+Figure 1: Characterizing a Representative Subpopulation Uniformly in $\mathcal{T}_{\text{all}}$
 To understand the supremum’s role in this expression, let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$ and note that $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ is equivalent to writing
 $\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}=\mu(\underline{w}^{*},\tau_{0})$ (4.1)
-for all $\tau_{0}$.
-Equation (4.3) holding for all $\tau_{0}$ requires $\underline{w}^{*}(X)$ be exactly proportional to $a(X)$.
-While $a(X)$’s range is unconstrained, $\underline{w}^{*}(X)$ must lie in $[0,1]$ to be a valid conditional probability.
+for all $\tau_{0}\in\mathcal{T}_{\text{all}}$.
+Equation (4.1) holding for all $\tau_{0}$ requires $\underline{w}^{*}(X)$ to be exactly proportional to $a(X)$.
+While the range of $a(X)$ is unconstrained, $\underline{w}^{*}(X)$ must lie in $[0,1]$ to be a valid conditional probability.
 Since we seek to maximize $\mathbb{P}(W^{*}=1\mid W_{0}=1)=\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$, we let $\underline{w}^{*}(X)$ be the largest multiple of $a(X)$ that lies in $[0,1]$ with probability 1, which is defined below:
-$W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0}\qquad\text{ and }\qquad\underline{w}^{*}(X)=\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}.$
+$W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{a_{\max}}\right)\cdot W_{0}\qquad\text{ and }\qquad\underline{w}^{*}(X)=\frac{a(X)}{a_{\max}}.$
 Here, $U\sim\text{Unif}(0,1)$ and $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$.
@@
 Thus, units with larger values of $a(X)$ are more likely to be included in $W^{*}$.
-All units in $\{W_{0}=1\}$ with $X$ such that $a(X)=\sup(\text{supp}(a(X)\mid W_{0}=1))$ are included in $W^{*}$, whereas no units where $a(X)=0$ are included.
-Given a data-generating process, this bound can be computed without any knowledge of the outcome’s conditional distribution and depends solely on weight function $a(\cdot)$ and the distribution of $X\mid W_{0}=1$.
-Remark 4.1 (Defining the target estimand from the weighted estimand).
-Suppose we consider $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ to be the target estimand, where $W^{*}=\mathbbm{1}\left(U\leq\frac{a(X)}{\sup(\text{supp}(a(X)\mid W_{0}=1))}\right)\cdot W_{0}$ is the subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ uniformly in $\tau_{0}$.
-For example, under Example 2.1, this consists of a subpopulation where the probability of inclusion is proportional to $\operatorname{var}(D\mid X)$, the conditional variance of treatment.
-If this subpopulation is the target, it would be reasonable to infer that the measure of internal validity of the estimand $\mu(a,\tau_{0})$ is the maximum value of 1.
-Indeed, this is the case because the estimand can be written as
-$\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
-and $\{W^{*}=1\}$ is trivially the largest regular subpopulation of $\{W^{*}=1\}$.
-This illustrates how the internal validity of an estimand is entirely dependent on the target estimand.
-On the other hand, the representativeness of the weighted estimand, as measured by the largest value of $\mathbb{P}(W^{*}=1)$, is less than one unless $W^{*}=W_{0}$ almost surely, or that the weighted estimand actually equals the ATE.
-Theorem 4.4 below can also be applied to obtain this intuition.
+All units in $\{W_{0}=1\}$ with $X$ such that $a(X)=a_{\max}$ are included in $W^{*}$, whereas no units where $a(X)=0$ are included.
+[…]
+Several further comments about Theorem 4.1 are in order.
+Remark 4.1 (Estimands and their corresponding interventions).
+We note that estimand $\mu(a,\tau_{0})$ is invariant to the scale of $a(\cdot)$ and thus can be written as
+$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}\left[\frac{a(X)}{a_{\max}}\tau_{0}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{a(X)}{a_{\max}}\mid W_{0}=1\right]}=\mu(a/a_{\max},\tau_{0}),$
+where $a(X)/a_{\max}\in[0,1]$ almost surely when weights are nonnegative.
+From this representation, we can link estimand $\mu(a,\tau_{0})$ with an intervention where fraction $a(X)/a_{\max}$ of units with covariate value $X$ and $W_{0}=1$ are treated.
+For example, if $W_{0}=1$ and $a(X)=a_{\max}$ almost surely, then $\mu(a,\tau_{0})$ is the average treatment effect, $\mathbb{E}[Y(1)-Y(0)]$, and it measures the average effect of treatment among all units.
+Under unconfoundedness, we also note that the average treatment effect on the treated (ATT) can be written as $\mathbb{E}[Y(1)-Y(0)\mid D=1]=\mu(a,\tau_{0})$, a weighted estimand with weights $a(X)=\mathbb{P}(D=1\mid X)$.
+This means that it can be interpreted either as the effect of an intervention where fraction $\mathbb{P}(D=1\mid X)$ of units with covariate $X$ are treated or as the effect of an intervention where all treated units are treated.
+In our setting, we can interpret any weighted estimand with nonnegative weights as the effect of treatment for a feasible intervention defined only in terms of $X$, $W_{0}$, and independent noise $U\sim\text{Unif}(0,1)$ via $W^{*}=\mathbbm{1}(U\leq a(X)/a_{\max})\cdot W_{0}$.
+Remark 4.2 (Uniqueness of representative subpopulations).
+It is also worth noting that the subpopulation maximizing the level of internal validity is generally not unique.
+The population $W^{*}=\mathbbm{1}\left(U\leq a(X)/a_{\max}\right)\cdot W_{0}$ will generally change if $U$ is replaced by another draw from a uniform distribution.
+The probability (conditional on $X$) of any unit being part of $W^{*}$ does not change with the draw of $U$, but whether any given unit is included in subpopulation $\{W^{*}=1\}$ cannot be determined.
+Remark 4.3 (Distributional characteristics of representative subpopulations).
+We can generally identify distributional characteristics of units within the population $W^{*}$.
+For example, when $W_{0}$ is set to 1 almost surely, we can write the average values of $g(X)$ among subpopulation $\{W^{*}=1\}$ as
+$\mathbb{E}[g(X)\mid W^{*}=1]=\frac{\mathbb{E}[g(X)\underline{w}^{*}(X)]}{\mathbb{E}[\underline{w}^{*}(X)]}=\frac{\mathbb{E}[g(X)a(X)]}{\mathbb{E}[a(X)]},$
+a simple function of weights $a(\cdot)$ and the marginal distribution of $X$.
+We can recover the average covariate values in $\{W^{*}=1\}$ by setting $g(X)=X$, or the entire distribution by considering $g(X)=\mathbbm{1}(X\leq x)$ for all $x\in\mathbb{R}^{d_{X}}$.
+Reporting the average covariate values of units within and outside of $W^{*}$ can be of interest when assessing the representativeness of $\mu(a,\tau_{0})$.
 We now consider a simple example to give further intuition for Theorem 4.1.
-4.1.1 Illustrative Example: A Single Binary Covariate
-Consider an estimand $\mu(a,\tau_{0})$ where $W_{0}=1$ almost surely, $a(X)\geq 0$, and where $X$ is binary with support $\text{supp}(X)=\{0,1\}$.
-Let $p_{x}=\mathbb{P}(X=x)>0$ for $x\in\{0,1\}$.
-As in Section 3.1, the weighted estimand can be written as a linear combination of the two CATEs:
-$\mu(a,\tau_{0})$ $=\frac{a(0)p_{0}}{\mathbb{E}[a(X)]}\tau_{0}(0)+\frac{a(1)p_{1}}{\mathbb{E}[a(X)]}\tau_{0}(1)\equiv\omega_{0}\tau_{0}(0)+\omega_{1}\tau_{0}(1).$
-Let $\mathbb{E}[Y(1)-Y(0)]=\text{ATE}$ be the target estimand, which can be written as
-ATE $=p_{0}\tau_{0}(0)+p_{1}\tau_{0}(1).$
-If $a(0)=a(1)$, the relative weights placed on $\{X=0\}$ and $\{X=1\}$ by the estimand are equal to $p_{0}/p_{1}$, the ratio of the weights placed by the ATE.
+Illustrative Example: A Single Binary Covariate
+Consider an estimand $\mu(a,\tau_{0})$ where $W_{0}=1$ almost surely, $a(X)\geq 0$, and where $X$ is binary with support $\text{supp}(X)=\{1,2\}$.
+Let $p_{x}=\mathbb{P}(X=x)$ for $x\in\{1,2\}$.
+As in equation (3.1), $\mu(a,\tau_{0})$ can be written as a linear combination of the two CATEs:
+$\mu(a,\tau_{0})$ $=\frac{a(1)p_{1}}{\mathbb{E}[a(X)]}\tau_{0}(1)+\frac{a(2)p_{2}}{\mathbb{E}[a(X)]}\tau_{0}(2)\coloneqq\omega_{1}\tau_{0}(1)+\omega_{2}\tau_{0}(2).$
+Let $\text{ATE}=\mathbb{E}[Y(1)-Y(0)]$ be the target estimand, which can be written as
+ATE $=p_{1}\tau_{0}(1)+p_{2}\tau_{0}(2).$
+If $a(1)=a(2)$, the relative weights placed on $\{X=1\}$ and $\{X=2\}$ by the estimand are equal to $p_{1}/p_{2}$, the ratio of the weights placed by the ATE.
 Therefore, the estimand equals the ATE and thus clearly has the maximum degree of internal validity with respect to the ATE.
-Applying Theorem 4.1, we can directly see that
-$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)]}{\sup_{x\in\{0,1\}}a(x)}=\frac{a(0)p_{0}+a(1)p_{1}}{\max\{a(0),a(1)\}}=\frac{a(1)(p_{0}+p_{1})}{a(1)}=1$
-when $a(0)=a(1)$.
-However, when $a(0)\neq a(1)$, the estimand’s weights differ from $(p_{0},p_{1})$, the population weights for the two covariate cells.
-For concreteness, let $a(1)>a(0)$.
-In this case, the estimand overrepresents the population with $X=1$ and underrepresents the population with $X=0$, relative to the ATE.
-The largest subpopulation $\{W^{*}=1\}$ that places weights $(a(0),a(1))$ can be constructed by combining subsets of the subpopulations defined by $\{X=0\}$ and $\{X=1\}$.
+Applying Theorem 4.1, we can directly see that, when $a(1)=a(2)$, $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\mathbb{E}[a(X)]/\sup_{x\in\{1,2\}}a(x)=a(2)/a(2)=1$.
+However, when $a(1)\neq a(2)$, the estimand’s weights differ from $(p_{1},p_{2})$, the population weights for the two covariate cells.
+For concreteness, let $(p_{1},p_{2})=(0.2,0.8)$ and $(a(1),a(2))=(0.24,0.09)$, where the latter correspond, for example, to the OLS weights of Proposition 2.1 when the propensity score is $(p(1),p(2))=(0.4,0.1)$.
+In this case, $(\omega_{1},\omega_{2})=(0.4,0.6)$ and thus
+$\mu(a,\tau_{0})=0.4\cdot\tau_{0}(1)+0.6\cdot\tau_{0}(2)\qquad\text{ and }\qquad\text{ATE}$ $=0.2\cdot\tau_{0}(1)+0.8\cdot\tau_{0}(2).$
+Relative to the ATE, the weighted estimand overrepresents the population with $X=1$ and underrepresents the population with $X=2$.
+The largest subpopulation $\{W^{*}=1\}$ that causally represents the estimand can be constructed by combining subsets of the subpopulations defined by $\{X=1\}$ and $\{X=2\}$.
 Specifically, let
-$W^{*}=\mathbbm{1}(X=1)+\mathbbm{1}\left(U\leq\frac{a(0)}{a(1)},X=0\right),$
-where $U\sim\text{Unif}(0,1)$ is independent from $(Y(1),Y(0),X)$.
-This is a subpopulation that contains all units with $X=1$ and fraction $a(0)/a(1)<1$ of units with $X=0$, selected uniformly at random.
-Its size is
-$\mathbb{P}(W^{*}=1)=\mathbb{P}(X=1)+\mathbb{P}\left(U\leq\frac{a(0)}{a(1)},X=0\right)=p_{1}+\frac{a(0)}{a(1)}p_{0}<1,$
-which is the same as $\frac{\mathbb{E}[a(X)]}{\sup_{x\in\{0,1\}}a(x)}=\frac{a(1)p_{1}+a(0)p_{0}}{a(1)}$ obtainable from Theorem 4.1. The average treatment effect in this subpopulation is given by
-$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mathbb{P}(X=0\mid W^{*}=1)\tau_{0}(0)+\mathbb{P}(X=1\mid W^{*}=1)\tau_{0}(1)$
-$=\frac{\frac{a(0)}{a(1)}p_{0}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}\tau_{0}(0)+\frac{p_{1}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}\tau_{0}(1)$
-$=\frac{a(0)p_{0}}{a(0)p_{0}+a(1)p_{1}}\tau_{0}(0)+\frac{a(1)p_{1}}{a(0)p_{0}+a(1)p_{1}}\tau_{0}(1)$
-$=\mu(a,\tau_{0}).$
-The relative weights placed on $\{X=0\}$ and $\{X=1\}$ in subpopulation $\{W^{*}=1\}$ are given by
-$\frac{\mathbb{P}(X=0\mid W^{*}=1)}{\mathbb{P}(X=1\mid W^{*}=1)}$ $=\frac{\frac{a(0)}{a(1)}p_{0}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}/\frac{p_{1}}{\frac{a(0)}{a(1)}p_{0}+p_{1}}=\frac{\omega_{0}}{\omega_{1}},$
-matching the ratio of the weights on $\{X=0\}$ and $\{X=1\}$ assigned by the estimand.
-The subpopulation $\{W^{*}=1\}$ cannot expand while preserving this ratio since it already encompasses all units with $X=1$.
-Therefore, $W^{*}$ is the largest subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$.
-4.1.2 Partial Identification of Average Effects
-This subpopulation size can be used to bound target estimand $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ using only the weighted estimand and bounds on $Y$.
-Given knowledge of the weighted estimand $\mu(a,\tau_{0})$ and its representativeness, given by $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$, we can decompose the target estimand as
-$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\cdot\mathbb{P}(W^{*}=1\mid W_{0}=1)$
-$\qquad+\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]\cdot(1-\mathbb{P}(W^{*}=1\mid W_{0}=1))$
-$=\mu(a,\tau_{0})\cdot\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]\cdot(1-\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$
-for $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$.
-Therefore, $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is point-identified given knowledge of $(\mu(a,\tau_{0}),\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$ if $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=1$, and partially identified if $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})<1$.
-If we have knowledge of bounds for the treatment effect $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0,W_{0}=1]$, e.g.
-from the support of the potential outcomes, we can obtain bounds on $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
-For example, if $\text{supp}(Y(1)-Y(0))=[B_{\ell},B_{u}]$, bounds for the target estimand are given by
-$\left[(\mu(a,\tau_{0})-B_{\ell})\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+B_{\ell},(\mu(a,\tau_{0})-B_{u})\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})+B_{u}\right].$
-The width of these bounds is $(B_{u}-B_{\ell})(1-\overline{P}(a,W_{0};\mathcal{T}_{\text{all}}))$.
-Hence for fixed $(B_{\ell},B_{u})$, this width decreases linearly with $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ and values of $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ close to 1, or high degrees of internal validity, lead to narrow bounds.
+$W^{*}=\mathbbm{1}(X=1)+\mathbbm{1}\left(U\leq\frac{a(2)}{a(1)},X=2\right)=\mathbbm{1}(X=1)+\mathbbm{1}\left(U\leq\frac{3}{8},X=2\right),$
+where $U\sim\text{Unif}(0,1)$ is independent of $(Y(1),Y(0),X)$.
+This is a regular subpopulation that contains all units with $X=1$ and three eighths of units with $X=2$, selected uniformly at random.
+Therefore
+$\mathbb{P}(W^{*}=1\mid X=1)=\underline{w}^{*}(1)=1\qquad\text{ and }\qquad\mathbb{P}(W^{*}=1\mid X=2)=\underline{w}^{*}(2)=3/8,$
+which yields $\mathbb{P}(W^{*}=1)=p_{1}\underline{w}^{*}(1)+p_{2}\underline{w}^{*}(2)=0.5$.
+The same quantity can be obtained from Theorem 4.1, which implies that $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\mathbb{E}[a(X)]/\left(\sup_{x\in\{1,2\}}a(x)\right)=\left(a(1)p_{1}+a(2)p_{2}\right)/a(1)=0.5$.
+The average effect in this subpopulation is given by
+$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)]/\mathbb{P}(W^{*}=1)$
+$=2\left(1\cdot\tau_{0}(1)\cdot 0.2+3/8\cdot\tau_{0}(2)\cdot 0.8\right)$
+$=0.4\cdot\tau_{0}(1)+0.6\cdot\tau_{0}(2),$
+which equals $\mu(a,\tau_{0})$ for any choice of $\tau_{0}$.
+Note that the relative weights placed on $\{X=1\}$ and $\{X=2\}$ in subpopulation $\{W^{*}=1\}$ are given by $\frac{\mathbb{P}(X=1\mid W^{*}=1)}{\mathbb{P}(X=2\mid W^{*}=1)}=0.4/0.6=\omega_{1}/\omega_{2}$, matching the ratio of the weights on $\{X=1\}$ and $\{X=2\}$ assigned by the estimand.
+The subpopulation $\{W^{*}=1\}$ cannot expand while preserving this ratio since it already includes all units with $X=1$.
+Therefore, $W^{*}$ is the largest subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any $\tau_{0}$.
+4.2   Using Internal Validity to Bound Average Effects
+The subpopulation size in Theorem 4.1 can be used to bound the target estimand $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
+Consider a scenario where only the weighted estimand $\mu(a,\tau_{0})$, which we assume has a causal representation uniformly in $\mathcal{T}_{\text{all}}$, and its internal validity are known.
+For example, this could be the case if a researcher uses a weighted estimand (e.g., OLS) and reports the measure we propose in Definition 4.1 to quantify its degree of internal validity for the ATE.
+To simplify notation, assume that $W_{0}=1$ almost surely and denote $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ by $\overline{P}_{\text{all}}(a)$.
+Abstracting from sample uncertainty, we only assume knowledge of the weighted estimand and its internal validity.
+We can decompose the target estimand, here $\mathbb{E}[Y(1)-Y(0)]$, as
+$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\cdot\mathbb{P}(W^{*}=1)\;+\;\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0]\cdot\mathbb{P}(W^{*}=0)$
+for a $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$.
+If we have knowledge of bounds for the treatment effect $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=0]$, e.g., from the support of the potential outcomes, we can obtain bounds on $\mathbb{E}[Y(1)-Y(0)]$.
+For example, if $\text{supp}(Y(1)-Y(0))\subseteq[B_{\ell},B_{u}]$, bounds for the target estimand are given by
+$\mathbb{E}[Y(1)-Y(0)]\;$ $\in\;\left[\mu(a,\tau_{0})\cdot\mathbb{P}(W^{*}=1)+B_{\ell}\cdot(1-\mathbb{P}(W^{*}=1)),\right.$
+$\qquad\qquad\left.\;\mu(a,\tau_{0})\cdot\mathbb{P}(W^{*}=1)+B_{u}\cdot(1-\mathbb{P}(W^{*}=1))\right].$ (4.2)
+The width of these bounds is minimized when $\mathbb{P}(W^{*}=1)$ is maximized, or when it equals the measure of internal validity for $\mu(a,\tau_{0})$, given by $\overline{P}_{\text{all}}(a)$.
+The resulting bounds for $\mathbb{E}[Y(1)-Y(0)]$ can be written as
+$\left[\mu(a,\tau_{0})\cdot\overline{P}_{\text{all}}(a)+B_{\ell}\cdot(1-\overline{P}_{\text{all}}(a)),\;\mu(a,\tau_{0})\cdot\overline{P}_{\text{all}}(a)+B_{u}\cdot(1-\overline{P}_{\text{all}}(a))\right]$ (4.3)
+$=\left[\mu(a,\tau_{0})\cdot\frac{\mathbb{E}[a(X)]}{a_{\max}}+B_{\ell}\cdot\left(1-\frac{\mathbb{E}[a(X)]}{a_{\max}}\right),\;\mu(a,\tau_{0})\cdot\frac{\mathbb{E}[a(X)]}{a_{\max}}+B_{u}\cdot\left(1-\frac{\mathbb{E}[a(X)]}{a_{\max}}\right)\right].$ (4.4)
+If $\overline{P}_{\text{all}}(a)=1$, it is easy to see that the estimand equals the ATE and that the bounds in (4.3) collapse to a point.
+However, the ATE is not uniquely determined from $(\mu(a,\tau_{0}),\overline{P}_{\text{all}}(a))$ when $\overline{P}_{\text{all}}(a)<1$.
+The width of these bounds is $(B_{u}-B_{\ell})\cdot(1-\overline{P}_{\text{all}}(a))$.
+Hence for fixed $(B_{\ell},B_{u})$, this width decreases linearly with $\overline{P}_{\text{all}}(a)$.
+Moreover, values of $\overline{P}_{\text{all}}(a)$ close to 1, or high degrees of internal validity, lead to narrow bounds.
+It is easy to obtain a sample analog of these bounds by combining estimators for $\mu(a,\tau_{0})$ and our proposed estimator for $\overline{P}_{\text{all}}(a)$ from Section 6 below.
+We note that bounds on $\mathbb{E}[Y(1)-Y(0)]$ may be tightened by assuming knowledge of other aspects of the joint distribution of $(Y(1),Y(0),D,X)$.
+For example, if knowledge of $a(\cdot)$ is assumed, additional constraints on $\mathbb{E}[Y(1)-Y(0)]$ can help narrow the bounds given in (4.2).
+We focus here on the case where we add a single piece of additional information to $\mu(a,\tau_{0})$, namely its internal validity, and how simple bounds can be obtained from the estimand and our proposed measure.
+We leave refinements of such bounds under different information sets to future work.
+Bounding Average Effects with Negative Weights
+Now consider a case where the weighted estimand $\mu(a,\tau_{0})$ has weights that can be negative, i.e., $\mathbb{P}(a(X)<0)>0$.
+We continue to assume that $W_{0}=1$ almost surely.
+In this case, we know that $\mu(a,\tau_{0})$ does not have a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$ and thus we cannot write $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ uniformly in $\tau_{0}$.
+However, simple algebra reveals that an estimand with negative weights can be written as a weighted difference of two nonnegatively weighted estimands:
+$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[(a(X)\mathbbm{1}(a(X)\geq 0)+a(X)\mathbbm{1}(a(X)\leq 0))\tau_{0}(X)]}{\mathbb{E}[a(X)]}$
+$=\omega^{+}\cdot\mu(\max\{a,0\},\tau_{0})-\omega^{-}\cdot\mu(-\min\{a,0\},\tau_{0}),$
+where $\omega^{+}\coloneqq\mathbb{E}[\max\{a(X),0\}]/\mathbb{E}[a(X)]$ and $\omega^{-}\coloneqq\mathbb{E}[-\min\{a(X),0\}]/\mathbb{E}[a(X)]$ are both nonnegative, and $\omega^{+}-\omega^{-}=1$.
+We note that $(\omega^{+},\omega^{-})=(1,0)$ when the estimand’s weights are nonnegative, so this decomposition can be obtained regardless of the sign of $a(\cdot)$.
+Thus, by Theorem 3.1, we can write
+$\mu(a,\tau_{0})\;$ $=\;\omega^{+}\cdot\mathbb{E}[Y(1)-Y(0)\mid W^{+}=1]\;-\;\omega^{-}\cdot\mathbb{E}[Y(1)-Y(0)\mid W^{-}=1],$ (4.5)
+where $W^{+}$ and $W^{-}$ characterize two disjoint, regular subpopulations.
+As above, suppose we want to bound $\mathbb{E}[Y(1)-Y(0)]$, the ATE.
+Using the law of iterated expectations, we can write $\mathbb{E}[Y(1)-Y(0)]$ as
+$\mathbb{E}[Y(1)-Y(0)\mid W^{+}=1]\cdot\mathbb{P}(W^{+}=1)\;+\;\mathbb{E}[Y(1)-Y(0)\mid W^{-}=1]\cdot\mathbb{P}(W^{-}=1)$
+$+\;\mathbb{E}[Y(1)-Y(0)\mid W^{+}+W^{-}=0]\cdot\left[1-\mathbb{P}(W^{+}=1)-\mathbb{P}(W^{-}=1)\right].$ (4.6)
+Substituting equation (4.5) in (4.6) and assuming that $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=1]$ and $\mathbb{E}[Y(1)-Y(0)\mid W^{+}+W^{-}=0]$ lie in $[B_{\ell},B_{u}]$ yields
+$\mathbb{E}[Y(1)-Y(0)]\;$ $\leq\;\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\cdot\mu(a,\tau_{0})\;+\;\left(1-\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\right)\cdot B_{u}$
+as an upper bound for the ATE. A lower bound is obtained by replacing $B_{u}$ with $B_{\ell}$. Thus, the ATE lies in the interval
+$\mathbb{E}[Y(1)-Y(0)]\;$ $\in\;\left[\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\cdot\mu(a,\tau_{0})+\left(1-\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\right)\cdot B_{\ell},\right.$
+$\left.\qquad\;\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\cdot\mu(a,\tau_{0})+\left(1-\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\right)\cdot B_{u}\right].$ (4.7)
+This interval is similar to the interval in (4.2), but the latter is only valid when weights are nonnegative.
+These intervals are identical when weights are nonnegative because $\omega^{+}=1$ and $\mathbb{P}(W^{+}=1)=\mathbb{P}(W^{*}=1)$ in that case.
+In order to compute the interval in (4.7) and minimize its length, one needs to maximize the value of $\mathbb{P}(W^{+}=1)$, which corresponds to the level of internal validity of the estimand $\mu(\max\{a,0\},\tau_{0})$ where $\max\{a,0\}\geq 0$, and compute the value of $\omega^{+}$.
+The resulting interval depends only on the ratio of the two quantities, which can be written as
+$\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\;$ $\leq\;\frac{\overline{P}_{\text{all}}(\max\{a,0\})}{\omega^{+}}\;=\;\frac{\mathbb{E}[\max\{a(X),0\}]/\sup(\text{supp}(\max\{a(X),0\}))}{\mathbb{E}[\max\{a(X),0\}]/\mathbb{E}[a(X)]}$
+$=\;\mathbb{E}[a(X)]/a_{\max}.$
+This last expression equals the level of internal validity of the original estimand $\mu(a,\tau_{0})$ when it is assumed (perhaps incorrectly) to have nonnegative weights.
+It follows that the bounds in (4.4) are valid regardless of whether weights are nonnegative, because minimizing the length of the interval in (4.7) yields precisely the bounds in (4.4).
+As mentioned earlier, these bounds do not make use of the entire distribution of $(Y,D,X)$, but simply of the original estimand $\mu(a,\tau_{0})$ and of $\mathbb{E}[a(X)]/a_{\max}$, the expression for the level of internal validity under nonnegative weights.
 4.3   Quantifying Internal Validity Given $\tau_{0}$
-As in the previous subsection, we can also ask how internally valid a weighted estimand can be, given knowledge of the CATE function. In this case, the object of interest is
-$\overline{P}(a,W_{0};\{\tau_{0}\})=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})}\mathbb{P}(W^{*}=1\mid W_{0}=1).$ (4.4)
-where $\tau_{0}$ is a given CATE function. Since $\tau_{0}$ is known, the condition $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ can be written as
-$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
-or, equivalently,
-$\mathbb{E}[(\tau_{0}(X)-\mu(a,\tau_{0}))\underline{w}^{*}(X)\mid W_{0}=1]=0,$ (4.5)
-where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$.
-Equation (4.5) is a linear constraint on the conditional probability of being in subpopulation $W^{*}$.
+We can also ask how internally valid a weighted estimand can be, given knowledge of the CATE function. In this case, the object of interest is
+$\overline{P}(a,W_{0};\{\tau_{0}\})=\sup_{W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})}\mathbb{P}(W^{*}=1\mid W_{0}=1),$ (4.8)
+where $\tau_{0}$ is a given CATE function.
+Since $\tau_{0}$ is known, the condition $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ can be written as $\mu_{0}=\mu(\underline{w}^{*},\tau_{0})$, where we let $\mu_{0}\coloneqq\mu(a,\tau_{0})$ to simplify the notation.
+This condition is equivalent to $\mathbb{E}[(\tau_{0}(X)-\mu_{0})\underline{w}^{*}(X)\mid W_{0}=1]=0$, a linear constraint on the conditional probability of being in subpopulation $W^{*}$.
 Additionally, the objective function $\mathbb{P}(W^{*}=1\mid W_{0}=1)=\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ is linear in $\underline{w}^{*}$.
 Thus, the optimization in (4.8) can be cast as a linear program.
-To see this, consider as an example the case where $W_{0}=1$ almost surely and where $X$ is discrete with finite support, i.e.
-$\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$.
-Let $f_{k}=\mathbb{P}(W^{*}=1,X=x_{k})$ denote the probability of being in subpopulation $W^{*}$ and having covariate value $x_{k}$.
-Note that $f_{k}\in[0,p_{k}]$ where $p_{k}\equiv\mathbb{P}(X=x_{k})$.
+To see this, consider as an example the case where $W_{0}=1$ almost surely and where $X$ is discrete with finite support, i.e., $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$.
+Let $f_{k}\coloneqq\mathbb{P}(W^{*}=1,X=x_{k})$ and note that $f_{k}\in[0,p_{k}]$ where $p_{k}=\mathbb{P}(X=x_{k})$.
 We can write the above optimization problem as
-$\max_{(f_{1},\ldots,f_{K})\geq\textbf{0}}\sum_{k=1}^{K}f_{k}\qquad\text{ such that }\qquad$ $f_{k}\leq p_{k}\text{ for }k\in\{1,\ldots,K\},$
-$\sum_{k=1}^{K}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))f_{k}=0,$
+$\max_{(f_{1},\ldots,f_{K})\geq\textbf{0}}\sum_{k=1}^{K}f_{k}\;\text{ s.t. }f_{k}\leq p_{k}\text{ for }k\in\{1,\ldots,K\}\;\text{ and }\;\sum_{k=1}^{K}(\tau_{0}(x_{k})-\mu_{0})f_{k}=0,$
 a finite-dimensional linear program.
-This program has a feasible solution if $\tau_{0}(x_{k})-\mu(a,\tau_{0})$ is not strictly positive or strictly negative for all $k$, which is precisely stated in the condition for Theorem 3.2.
+This program has a feasible solution if $\tau_{0}(x_{k})-\mu_{0}$ is not strictly positive or strictly negative for all $k$, meaning that the weighted estimand lies in the convex hull of CATE values, which is precisely stated in the condition for Theorem 3.2.
 While there exist many methods for solving linear programs, the value function can be obtained through an algorithm that is simple to describe analytically.
-Let $\textbf{f}=(f_{1},\ldots,f_{K})$, $\textbf{p}=(p_{1},\ldots,p_{K})$, and $\textbf{t}_{\mu}=(\tau_{0}(x_{1})-\mu(a,\tau_{0}),\ldots,\tau_{0}(x_{K})-\mu(a,\tau_{0}))$.
-Without loss of generality, assume that $\tau_{0}(x_{1})-\mu(a,\tau_{0})\leq\tau_{0}(x_{2})-\mu(a,\tau_{0})\leq\cdots\leq\tau_{0}(x_{K})-\mu(a,\tau_{0})$.
+Algorithm 4.1 (Internal validity for fixed $\tau_{0}$).
+Without loss of generality, let $\tau_{0}(x_{1})\leq\ldots\leq\tau_{0}(x_{K})$.
   1. 1.
-[…]
+    Set $(f_{1},\ldots,f_{K})=(p_{1},\ldots,p_{K})$.
   2. 2.
-    If $\textbf{t}_{\mu}^{\prime}\textbf{f}=0$, end the algorithm and report $\sum_{k=1}^{K}f_{k}$.
+    If $\sum_{k=1}^{K}(\tau_{0}(x_{k})-\mu_{0})f_{k}=0$, end the algorithm and report $\sum_{k=1}^{K}f_{k}$.
   3. 3.
-    If $\textbf{t}_{\mu}^{\prime}\textbf{f}\neq 0$:
+    If $\sum_{k=1}^{K}(\tau_{0}(x_{k})-\mu_{0})f_{k}\neq 0$:
     1. (a)
-      If $\textbf{t}_{\mu}^{\prime}\textbf{f}>0$, let $k^{*}=\max\{k\in\{1,\ldots,K\}:f_{k}=p_{k}\}$ and set $f_{k^{*}}=\max\left\{0,\frac{-\sum_{k=1}^{k^{*}-1}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))p_{k}}{\tau_{0}(x_{k^{*}})-\mu(a,\tau_{0})}\right\}$.
+      If $\sum_{k=1}^{K}(\tau_{0}(x_{k})-\mu_{0})f_{k}>0$, let $k^{*}=\max\{k\in\{1,\ldots,K\}:f_{k}=p_{k}\}$ and set $f_{k^{*}}=\max\left\{0,-\sum_{k=1}^{k^{*}-1}(\tau_{0}(x_{k})-\mu_{0})p_{k}/(\tau_{0}(x_{k^{*}})-\mu_{0})\right\}$.
     2. (b)
-      If $\textbf{t}_{\mu}^{\prime}\textbf{f}<0$, let $k^{*}=\min\{k\in\{1,\ldots,K\}:f_{k}=p_{k}\}$ and set $f_{k^{*}}=\max\left\{0,\frac{-\sum_{k=k^{*}+1}^{K}(\tau_{0}(x_{k})-\mu(a,\tau_{0}))p_{k}}{\tau_{0}(x_{k^{*}})-\mu(a,\tau_{0})}\right\}$.
+      If $\sum_{k=1}^{K}(\tau_{0}(x_{k})-\mu_{0})f_{k}<0$, let $k^{*}=\min\{k\in\{1,\ldots,K\}:f_{k}=p_{k}\}$ and set $f_{k^{*}}=\max\left\{0,-\sum_{k=k^{*}+1}^{K}(\tau_{0}(x_{k})-\mu_{0})p_{k}/(\tau_{0}(x_{k^{*}})-\mu_{0})\right\}$.
   4. 4.
     Go to step 2.
-When $\mu(a,\tau_{0})$ exceeds $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, this algorithm reduces the weights associated with largest CATEs until $\mu(a,\tau_{0})$ equals $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for some subpopulation.
-When $\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the same procedure is instead applied to the smallest CATEs.
+When $\mu_{0}$ exceeds $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, this algorithm reduces the weights associated with smallest CATEs until $\mu_{0}$ equals $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for some subpopulation.
+When $\mu_{0}<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the same procedure is instead applied to the largest CATEs.
 The support assumption of Theorem 3.2 guarantees that this algorithm ends.
-When $X$ is not discretely supported, the problem can still be cast as a linear program, but its dimension may be infinite, which generates difficulties in implementation.
-However, we show this program has an analytical solution for the size of the subpopulation of interest, $\mathbb{P}(W^{*}=1\mid W_{0}=1)$.
-The following theorem gives its expression and is valid when vector $X$ has an arbitrary kind of distribution, with continuous, discrete, and mixed components, as is often the case in empirical applications.
+When $X$ is not discretely supported, the problem can still be cast as a linear program, but its dimension is infinite, which generates difficulties in implementation.
+However, we show this program has an analytical solution even when $X$’s components are allowed to be discrete, continuous, and mixed, as is often the case in practice.
 Theorem 4.2.
-Let Assumption 3.1 hold. Let $T_{\mu}=\tau_{0}(X)-\mu(a,\tau_{0})$. If $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W_{0})$,
-$\overline{P}(a,W_{0};\{\tau_{0}\})$
-=$=\begin{cases}\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}\\ \qquad\text{ where }\alpha^{+}=\inf\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]\geq 0\}\text{ if }\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]\\ \mathbb{P}(T_{\mu}\geq\alpha^{-}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}\\ \qquad\text{ where }\alpha^{-}=\sup\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]\leq 0\}\text{ if }\mu(a,\tau_{0})>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]\\ 1\qquad\text{ if }\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1].\end{cases}$ (4.6)
-If $\mu(a,\tau_{0})\notin\mathcal{S}(\tau_{0};W_{0})$, then $\overline{P}(a,W_{0};\{\tau_{0}\})=0$.
+Let $\mu(a,\tau_{0})$ be an estimand satisfying equation (1.1).
+Suppose Assumption 3.1 holds.
+If $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$, then $\overline{P}(a,W_{0};\{\tau_{0}\})=0$.
+If $\mu_{0}\in\mathcal{S}(\tau_{0};W_{0})$, then
+$\overline{P}(a,W_{0};\{\tau_{0}\})=$$\begin{cases}\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}&\text{ if }\mu_{0}<E_{0}\\ \mathbb{P}(T_{\mu}\geq\alpha^{-}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}&\text{ if }\mu_{0}>E_{0}\\ 1&\text{ if }\mu_{0}=E_{0},\end{cases}$ (4.9)
+where $T_{\mu}\coloneqq\tau_{0}(X)-\mu_{0}$, $E_{0}\coloneqq\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, $\alpha^{+}\coloneqq\inf\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]\geq 0\}$, and $\alpha^{-}\coloneqq\sup\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]\leq 0\}$.
 The computation of these bounds can be done using a linear programming algorithm when $X$ is discrete, or through plug-in estimators of the terms in equation (4.9) regardless of the nature of the support of $X$.
+To illustrate this theorem, let $\tau_{0}(X)$ be continuously distributed with support $[\underline{\tau},\overline{\tau}]$, and suppose $\mu_{0}\in[\underline{\tau},\overline{\tau}]$.
+Without loss of generality, assume $E_{0}\geq\mu_{0}$.
+If $E_{0}=\mu_{0}$, then the estimand is perfectly representative of the population since it equals the average treatment effect over it.
+In the case where $E_{0}>\mu_{0}$, the estimand is not representative of the entire population.
+We are searching for the largest subpopulation $\{W^{*}=1\}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu_{0}$.
+Initializing $W^{*}$ at $W_{0}$, removing the subpopulation with the largest values of $\tau_{0}(x)$ yields the steepest decrease in $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$.
+Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})$ is obtained by removing a subpopulation of the kind $W^{-}(\alpha)=\mathbbm{1}(\tau_{0}(X)>\alpha)\cdot W_{0}$ for a given threshold $\alpha$.
+This threshold is determined by the constraint
+$\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq\alpha]=\mu_{0}.$ (4.10)
+Thus, the remaining subpopulation $W^{*}$ corresponds to $W^{*}=\mathbbm{1}(\tau_{0}(X)\leq\alpha^{*})\cdot W_{0}$ where $\alpha^{*}$ is the unique solution to (4.10).
 In this setting, the value $\overline{P}(a,W_{0};\{\tau_{0}\})$ is larger when the truncated subpopulations are smaller.
 In particular, this is the case when there are a few units with extreme values of $\tau_{0}$ whose removal has a large impact on the estimand, but a small impact on the share of the population.
-4.2.1 Illustrative Example: A Single Continuous Covariate
-To illustrate the previous theorem, let $X$ be a continuously distributed covariate with support $[x_{L},x_{U}]$.
-Let $W_{0}=1$ almost surely, and let $\mu\equiv\mu(a,\tau_{0})$ denote the estimand.
-Also, for simplicity assume that $\tau_{0}(x)$ is increasing in $x$ and that $\tau_{0}(x_{L})<\mu<\tau_{0}(x_{U})$.
-Without loss of generality, let $\mathbb{E}[Y(1)-Y(0)]\geq\mu$.
-If $\mathbb{E}[Y(1)-Y(0)]=\mu$, then the estimand is perfectly representative of the population since it equals the average treatment effect over it.
-Now consider the case where $\mathbb{E}[Y(1)-Y(0)]>\mu$.
-In this case, the estimand is larger than the ATE so it is not representative of the entire population.
-We are seeking the largest subpopulation $\{W^{*}=1\}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu$.
-Equivalently, we can seek the smallest subpopulation $\{W^{-}=1\}$ such that, when it is removed from the original population, the average treatment effect equals $\mu$.
-The subpopulation $\{W^{-}=1\}$ is related to $\{W^{*}=1\}$ via $W^{-}+W^{*}=1$.
-We will search for $W^{-}$ such that $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=0]=\mu$ and $\mathbb{P}(W^{-}=1)$ is minimized.
-Since the ATE exceeds $\mu$, we must have that $\mathbb{E}[\tau_{0}(X)\mid W^{-}=1]>\mu$.
-For a given subpopulation of size $\mathbb{P}(W^{-}=1)$, removing the subpopulation with the largest values of $\tau_{0}(x)$ yields the largest decrease in $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=0]$.
-Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})$ is obtained by removing a subpopulation of the kind $W^{-}(\delta)=\mathbbm{1}(\tau_{0}(X)\geq\delta)$ for a given threshold $\delta$.
-This subpopulation consists of units whose CATEs exceed a threshold $\delta$.
-This threshold is determined by the constraint $\mathbb{E}[\tau_{0}(X)\mid W^{-}(\delta)=0]=\mu$, or
-$\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq\delta]=\mu.$ (4.7)
-This constraint states that once $\{W^{-}(\delta)=1\}$ is removed, the average treatment effect for the remaining units equals the weighted estimand.
-Since we assumed $\tau_{0}$ is increasing, $W^{*}$ corresponds to all units with covariate values below threshold $\delta=\tau_{0}^{-1}(\delta)$.
-Therefore,
-$W^{*}=\mathbbm{1}(X\leq\tau_{0}^{-1}(\delta)),$
-where $\delta$ is the unique solution to (4.7).
-How large $\mathbb{P}(W^{*}=1)$ is depends on the size of this fraction of removed units.
-A smaller subpopulation needs to be removed when $\tau_{0}(X)$ has a long right tail, and a larger subpopulation needs to be removed when the distribution of $\tau_{0}(X)$ in this right tail is shorter.
-4.3 Generalizing to Subpopulations
-The results from the previous two sections can be generalized to cases where we hold the average effect for a subset of $W_{0}$ as the object of interest.
-Specifically, let $W^{\prime}$ be a regular subpopulation of $W_{0}$ and $\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]$ be the target parameter.
-In analogy to the previous subsections, we ask (i) when does there exist a regular subpopulation $W^{*}$ of $W^{\prime}$ such that $\mu(a,\tau_{0})$ can be written as a causal estimand and (ii) how representative of $W^{\prime}$ is this estimand.
-For concreteness, in the IV model of Section 2.2, we may be interested in the average treatment effect for compliers/defiers with covariates in given set $\mathcal{X}_{0}$: $\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X\in\mathcal{X}_{0}]$.
-In this case $W_{0}=\mathbbm{1}(D(1)>D(0))$ and $W^{\prime}=\mathbbm{1}(D(1)>D(0),X\in\mathcal{X}_{0})$, a regular subpopulation of $W_{0}$.
-We first present a result showing conditions for the existence of such population $W^{*}$ in the case where $\tau_{0}$ is unrestricted, and when $\tau_{0}$ is fixed.
-Theorem 4.3.
-Let Assumption 3.1 hold and let $W^{\prime}$ be a regular subpopulation of $W_{0}$.
-  1. 1.
-    (Uniformly in $\tau_{0}$) Let $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$.
-[^5] We use the conventions $0/0=1$ and $a/0=\infty$ when $a>0$.
-Then, there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ if and only if $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$.
-  2. 2.
-    (Given $\tau_{0}$) There exists $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$ if and only if $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W^{\prime})$.
-Note that letting $W^{\prime}=W_{0}$ yields Theorems 3.1 and 3.2 as special cases since $W_{0}$ is a regular subpopulation of $W_{0}$ by definition.
-The condition $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ rules out uniform causal representations if there are covariate values for which $a(x)$ is strictly positive but are not represented in $W^{\prime}$, i.e., $\mathbb{P}(W^{\prime}=1\mid X=x,W_{0}=1)=0$ when $a(x)>0$.
-For example, if $W^{\prime}$ is a subpopulation of $W_{0}$ defined by a subset of covariate values such that $W^{\prime}=\mathbbm{1}(X\in\mathcal{X}_{0})\cdot W_{0}$, then $W^{\prime}$ is a regular subpopulation but will fail the above inequality if $a(X)>0$ for $X\in\text{supp}(X\mid W_{0}=1)\setminus\mathcal{X}_{0}$.
-In the case where $\tau_{0}$ is known, the existence of a causal subpopulation of $W^{\prime}$ is equivalent to the estimand being in the convex hull of $\text{supp}(\tau_{0}(X)\mid W^{\prime}=1)$.
-Regarding the representativeness of a weighted estimand, we generalize Theorems 4.1 and 4.2 and ask how large $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ can be given that $W^{*}$ is a regular subpopulation of $W^{\prime}$, which is itself a regular subpopulation of $W_{0}$.
-In what follows, we let $\underline{w}^{\prime}(X)=\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$.
-Theorem 4.4.
-Let Assumption 3.1 hold and let $W^{\prime}$ be a regular subpopulation of $W_{0}$.
-  1. 1.
-    (Uniformly in $\tau_{0}$) Let $\sup\left(\text{supp}\left(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1\right)\right)<\infty$. If $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$, then
-    $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})=\mathbb{E}[a(X)\mid W_{0}=1]\cdot\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}\cdot\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right).$ (4.8)
-    If $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)<1$, $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})=0$.
-  2. 2.
-    (Given $\tau_{0}$) If $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W^{\prime})$, then
-    $\overline{P}(a,W^{\prime};\{\tau_{0}\})$
-    =$=\begin{cases}\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W^{\prime}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W^{\prime}=1]}{\alpha^{+}}\\ \qquad\text{ where }\alpha^{+}=\inf\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W^{\prime}=1]\geq 0\}\text{ if }\mu(a,\tau_{0})<\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]\\ \mathbb{P}(T_{\mu}\geq\alpha^{-}\mid W^{\prime}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W^{\prime}=1]}{\alpha^{-}}\\ \qquad\text{ where }\alpha^{-}=\sup\{\alpha\in\mathbb{R}:\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]\leq 0\}\text{ if }\mu(a,\tau_{0})>\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1]\\ 1\qquad\text{if }\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1].\end{cases}$ (4.9)
-    If $\mu(a,\tau_{0})\notin\mathcal{S}(\tau_{0};W^{\prime})$, then $\overline{P}(a,W^{\prime};\{\tau_{0}\})=0$.
-We can also use this result to obtain bounds on $\mathbb{P}(W^{*}=1)$, or $\mathbb{P}(W^{*}=1\mid W_{0}=1)$.
-This can be done by noting that $W^{*}$ is a subpopulation of both $W^{\prime}$ and $W_{0}$, so
-$\mathbb{P}(W^{*}=1)$ $=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\cdot\mathbb{P}(W^{\prime}=1)\quad\text{ and }\quad\mathbb{P}(W^{*}=1\mid W_{0}=1)=\frac{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\cdot\mathbb{P}(W^{\prime}=1)}{\mathbb{P}(W_{0}=1)}.$
-Thus bounds on these probabilities are trivially obtained from the bound on $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ from Theorem 4.4 and knowledge of $(\mathbb{P}(W^{\prime}=1),\mathbb{P}(W_{0}=1))$.
-4.4 Examples Revisited
-For simplicity, we assume that $\sup(\text{supp}(a(X)\mid W_{0}=1))=\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)$ for the remainder of the paper.
-This condition is satisfied for example when $a(\cdot)$ is continuous.
-4.4.1 Unconfoundedness
-Recall the expression for the coefficient on $D$ in a population regression of $Y$ on $(1,D,X)$:
+5   Applications to Common Estimands
+Here we consider three identification strategies where commonly used estimands follow the structure of equation (1.1).
+We show how the results in Sections 3 and 4 apply in each of these cases.
+For simplicity, we assume that $a_{\max}=\sup(\text{supp}(a(X)\mid W_{0}=1))=\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)$ in this section.
+This condition is satisfied when $a(\cdot)$ is continuous or when $X$ has finite support, among other cases.
+We also note that our assumption $a_{\max}<\infty$ holds trivially in every case considered below.
+5.1   Unconfoundedness
+In Section 2, we provided the expression for the coefficient on $D$ in a population regression of $Y$ on $(1,D,X)$:
 $\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[p(X)(1-p(X))\tau_{0}(X)]}{\mathbb{E}[p(X)(1-p(X))]}.$
-Suppose the target estimand is the average treatment effect, i.e. $W_{0}=1$ almost surely.
-By Theorem 3.1, there exists a regular subpopulation $W^{*}$ such that $\beta_{\text{OLS}}$ equals the average treatment effect over $W^{*}$ since the weight function $a(X)\equiv p(X)(1-p(X))$ is non-negative.
-By Theorem 4.1, the upper bound on the size of subpopulation $W^{*}$ is given by
+Suppose the target estimand is the average treatment effect, i.e., $W_{0}=1$ almost surely.
+By Theorem 3.1, there exists a regular subpopulation $W^{*}$ such that $\beta_{\text{OLS}}$ equals the average treatment effect over $W^{*}$ since the weight function $a(X)=p(X)(1-p(X))$ is nonnegative.
+By Theorem 4.1, the upper bound on the size of subpopulation $W^{*}$ is
 $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[p(X)(1-p(X))]}{\sup_{x\in\text{supp}(X)}p(x)(1-p(x))}.$
-The corresponding subpopulation $W^{*}$ satisfies
-$\mathbb{P}(W^{*}=1\mid X)$ $=\frac{p(X)(1-p(X))}{\sup_{x\in\text{supp}(X)}p(x)(1-p(x))}.$
-This is a subpopulation that places greater weights on units which have a larger variation in treatment given their covariate values.
+A corresponding subpopulation $W^{*}$ can be written as
+$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{p(X)(1-p(X))}{\sup_{x\in\text{supp}(X)}p(x)(1-p(x))}\right),$
+where $U\perp\!\!\!\perp(Y(1),Y(0),X)$ and $U\sim\text{Unif}(0,1)$.
+This is a subpopulation where units with a larger variation in treatment given their covariate values are more likely to be included.
 The size of this subpopulation is largest when $\operatorname{var}(D\mid X)=p(X)(1-p(X))$ is constant, in which case $\mathbb{P}(W^{*}=1\mid X)=1$.
-This is the case if and only if $p(X)$ has support equal to $\{b,1-b\}$ for some $b\in[0,1]$.
-This is implied by $D\perp\!\!\!\perp X$, or random assignment.
-It can also be achieved if there exists a partition of $\text{supp}(X)$ where $\mathbb{P}(D=1\mid X)=b$ on one element and $\mathbb{P}(D=1\mid X)=1-b$ on its complement.
+This is the case if and only if $p(X)$ has support contained in $\{b,1-b\}$ for some $b\in(0,1)$.
 Whenever $\operatorname{var}(p(X)(1-p(X)))>0$, $\{W^{*}=1\}$ will be a strict subpopulation.
@@
 Its numerator is bounded above by $\operatorname{var}(D)=\mathbb{P}(D=1)\cdot\mathbb{P}(D=0)$, which is particularly simple to estimate.
-As for the denominator, it is a non-smooth functional of $p(\cdot)$ which will lead to nonstandard inference.
-However, if $X$ is continuously distributed, it may be likely that $p(X)$ is continuously distributed as well.
-In this case, it is likely that $1/2\in\text{supp}(p(X))$.
+As for the denominator, it is a nonsmooth functional of $p(\cdot)$.
+However, if $X$ is continuously distributed, it may be likely that $p(X)$ is continuously distributed and thus that $1/2\in\text{supp}(p(X))$.
 If this is the case, $\sup_{x\in\text{supp}(X)}p(x)(1-p(x))=1/4$.
-Combining these approximations yields that
-$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $\leq 4\cdot\mathbb{P}(D=1)\cdot\mathbb{P}(D=0),$
+Combining these two approximations yields
+$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})\;$  $\leq\;4\cdot\mathbb{P}(D=1)\cdot\mathbb{P}(D=0),$
 when the support of $p(X)$ includes 1/2.
-This bound is trivial when $\mathbb{P}(D=1)=1/2$, but is informative when the treatment probability is close to 0 or 1.
+This bound is trivial when $\mathbb{P}(D=1)=1/2$, but is informative when the unconditional treatment probability is close to 0 or 1.
 For example, if $\mathbb{P}(D=1)=0.1$, the OLS estimand cannot causally represent more than 36% of the population.
-When $1/2\in\text{supp}(p(X))$, we can also compute bounds on the ATE derived from the OLS estimand, bounds on the support of $(Y(1),Y(0))$, and our measure of representativeness $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$.
-Following Remark 4.1.2, bounds on the ATE are given by
-$\left[(\beta_{\text{OLS}}-B_{\ell})4\mathbb{E}[\operatorname{var}(D\mid X)]+B_{\ell},(\beta_{\text{OLS}}-B_{u})4\mathbb{E}[\operatorname{var}(D\mid X)]+B_{u}\right].$
-Estimating these bounds requires the estimation of one additional quantity beyond the OLS estimator, which is the expectation of $\operatorname{var}(D\mid X)$.
-Of course, these bounds’ width depends crucially on $B_{u}-B_{\ell}$, or the width of the support for unit-level treatment effects.
-Alternatively, we can assess the representativeness of $\beta_{\text{OLS}}$ with respect to an alternative estimand such as $\mathbb{E}[Y(1)-Y(0)\mid D=1]$, the average treatment effect on the treated.
-In this case, we can write
-$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))w_{0}(X)\tau_{0}(X)]}{\mathbb{E}[(1-p(X))w_{0}(X)]},$
-where $w_{0}(X)=\mathbb{P}(D=1\mid X)=p(X)$, the propensity score, and $a(X)=\mathbb{P}(D=0\mid X)=1-p(X)$. Applying Theorem 4.1 (or 4.4) yields that
-$\overline{P}(a,D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[1-p(X)\mid D=1]}{\sup_{x\in\text{supp}(X\mid D=1)}(1-p(x))}=\frac{\mathbb{E}[p(X)(1-p(X))]}{\mathbb{P}(D=1)\cdot\sup_{x\in\text{supp}(X\mid D=1)}(1-p(x))}$
-is the largest value that $\mathbb{P}(W^{*}=1\mid D=1)$ can take.
-Once again, this bound depends only on the propensity score and the distribution of $X$.
-The same bound is obtained by using Theorem 4.4 instead and setting $W_{0}=1$ and $W^{\prime}=D$.
-This subpopulation satisfies
-$\mathbb{P}(W^{*}=1\mid X,D=1)=\frac{1-p(X)}{1-\inf_{x\in\text{supp}(X\mid D=1)}p(X)}$
+This is consistent with the result in Słoczyński (2022) that the OLS estimand is more similar to the ATE when $\mathbb{P}(D=1)$ is close to 1/2.
+When $1/2\in\text{supp}(p(X))$, we can also compute bounds on the ATE derived from $\beta_{\text{OLS}}$, bounds on the support of $(Y(1),Y(0))$, and our measure of internal validity $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$.
+Following the expression in (4.4), bounds on the ATE are given by
+$\left[(\beta_{\text{OLS}}-B_{\ell})\cdot 4\mathbb{E}[\operatorname{var}(D\mid X)]+B_{\ell},\;(\beta_{\text{OLS}}-B_{u})\cdot 4\mathbb{E}[\operatorname{var}(D\mid X)]+B_{u}\right].$
+Estimating these bounds requires the estimation of one additional quantity beyond the OLS estimand, which is the expectation of $\operatorname{var}(D\mid X)$.
+The width of these bounds depends crucially on $B_{u}-B_{\ell}$, the width of the support for unit-level treatment effects.
+Alternatively, we can assess the internal validity of $\beta_{\text{OLS}}$ with respect to an alternative estimand such as $\mathbb{E}[Y(1)-Y(0)\mid D=1]$, the average treatment effect on the treated.
+In this case, we consider an alternative representation of the estimand:
+$\beta_{\text{OLS}}$ $=\frac{\mathbb{E}[(1-p(X))\tau_{0}(X)\mid D=1]}{\mathbb{E}[1-p(X)\mid D=1]}.$
+Applying Theorem 4.1 yields that
+$\overline{P}(1-p,D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[1-p(X)\mid D=1]}{\sup_{x\in\text{supp}(X\mid D=1)}(1-p(x))}=\frac{\mathbb{E}[p(X)(1-p(X))]}{\mathbb{P}(D=1)\cdot\sup_{x\in\text{supp}(X\mid D=1)}(1-p(x))}$
+is the largest value that $\mathbb{P}(W^{*}=1\mid D=1)$ can take. Once again, this bound depends only on the propensity score and the distribution of $X$. This subpopulation satisfies
+$\mathbb{P}(W^{*}=1\mid X,D=1)=\frac{1-p(X)}{1-\inf_{x\in\text{supp}(X\mid D=1)}p(X)},$
 so units with smaller propensity scores are more likely to be included in $W^{*}$, given that they are treated.
-This probability is maximized at 1 when $p(X)$ is constant, or if $D\perp\!\!\!\perp X$.
+$\overline{P}(1-p,D;\mathcal{T}_{\text{all}})$ is maximized at 1 when $p(X)$ is constant, or if $D\perp\!\!\!\perp X$.
 In this case, $\mathbb{P}(W^{*}=1\mid D=1)=1$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(D=1)$.
-If $p(X)$ takes values close to 0, this bound equals
-$\overline{P}(a,D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[p(X)(1-p(X))]}{\mathbb{P}(D=1)}\leq\frac{\mathbb{P}(D=1)\cdot\mathbb{P}(D=0)}{\mathbb{P}(D=1)}=\mathbb{P}(D=0).$
+If $p(X)$ takes values close to 0, the bound satisfies
+$\overline{P}(1-p,D;\mathcal{T}_{\text{all}})$ $\approxeq\frac{\mathbb{E}[p(X)(1-p(X))]}{\mathbb{P}(D=1)}\leq\frac{\mathbb{P}(D=1)\cdot\mathbb{P}(D=0)}{\mathbb{P}(D=1)}=\mathbb{P}(D=0).$
 This suggests that the OLS estimand is more representative of the ATT when the fraction of untreated units is larger.
-We can also assess the representativeness of $\beta_{\text{OLS}}$ given $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$.
-For simplicity, assume that $X$ has a continuous distribution and, without loss of generality, assume that $\beta_{\text{OLS}}>\text{ATE}$.
+This again echoes the results in Słoczyński (2022) on the relationship between $\mathbb{P}(D=1)$ and the interpretation of the OLS estimand.
+We can also assess the internal validity of $\beta_{\text{OLS}}$ given $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$.
+For simplicity, assume that $\tau_{0}(X)$ has a continuous distribution and, without loss of generality, assume that $\text{ATE}>\beta_{\text{OLS}}$.
 Then, using Theorem 4.2, we obtain
-$\overline{P}(a,W_{0};\{\tau_{0}\})$ $=\mathbb{P}(\tau_{0}(X)\leq b^{*}),$
-where $b^{*}$ satisfies $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq b^{*}]=\beta_{\text{OLS}}$.
+$\overline{P}(a,W_{0};\{\tau_{0}\})$ $=\mathbb{P}(\tau_{0}(X)\leq\alpha^{*}),$
+where $\alpha^{*}$ satisfies $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq\alpha^{*}]=\beta_{\text{OLS}}$.
 The quantity $\overline{P}(a,W_{0};\{\tau_{0}\})$ is largest when the least amount of trimming needs to be applied.
-This is the case when the trimmed values are largest, or when $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\geq b]$ is large when $b$ is near the maximum of $\text{supp}(\tau_{0}(X))$.
-More concretely, if some covariate values in $\text{supp}(X)$ have large CATEs, then considering a subpopulation that removes only a small subset of $\text{supp}(X)$ — those corresponding to large CATE values — can allow $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq b^{*}]$ and $\beta_{\text{OLS}}$ to be equal.
-4.4.2 Instrumental Variables
-Consider again the setting of Section 2.2.
-The estimand $\beta_{\text{2SLS}}$ can be characterized as $\mu(a_{\text{2SLS}},\tau_{0})$ for $a_{\text{2SLS}}(X)=\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)\neq D(0)\mid X)$, $W_{0}=\mathbbm{1}(D(1)\neq D(0))$, $w_{0}(X)=\mathbb{P}(D(1)\neq D(0)\mid X)$, and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$.
-Since $a_{\text{2SLS}}(X)\geq 0$, there exists a subpopulation of $\{W_{0}=1\}$ such that $\beta_{\text{2SLS}}$ is an average treatment effect over that subpopulation.
+This is the case when the trimmed values are largest, or when $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\geq\alpha]$ is large for large $\alpha$.
+5.2   Instrumental Variables
+Now consider a binary treatment $D\in\{0,1\}$ and a binary instrument $Z\in\{0,1\}$.
+Potential treatments, denoted by $(D(1),D(0))$, are linked to the realized treatment through $Z$, that is, $D=D(Z)$.
+Potential outcomes, $Y(d,z)$ for $d,z\in\{0,1\}$, may depend on both $D$ and $Z$ in the absence of an exclusion restriction.
+Let $Y=Y(D,Z)$ be the realized outcome.
+As before, let $X$ denote covariates.
+We make the following assumptions.
+Assumption 5.1 (Instrument validity).
+Almost surely, the following hold:
+  1. 1.
+    Exogeneity: $(Y(0,0),Y(1,0),Y(0,1),Y(1,1),D(1),D(0))\perp\!\!\!\perp Z\mid X$;
+  2. 2.
+    Exclusion: $\mathbb{P}(Y(d,0)=Y(d,1)\mid X)=1$ for $d\in\{0,1\}$;
+  3. 3.
+    First stage: $\mathbb{P}(Z=1\mid X)\in(0,1)$ and $\mathbb{P}(D(1)=1\mid X)\neq\mathbb{P}(D(0)=1\mid X)$;
+  4. 4.
+    Strong monotonicity: $\mathbb{P}(D(1)\geq D(0)\mid X)=1$.
+The first instrumental variables estimand we consider was originally studied by Angrist and Imbens (1995).
+In addition to Assumption 5.1, suppose that the model for $X$ is saturated, with $K$ possible combinations of covariate values, i.e., let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$.
+Let $X_{S}=\left(1,\mathbbm{1}(X=x_{1}),\ldots,\mathbbm{1}(X=x_{K-1})\right)$ and $Z_{S}=\left(Z,Z\cdot\mathbbm{1}(X=x_{1}),\ldots,Z\cdot\mathbbm{1}(X=x_{K-1})\right)=ZX_{S}$, where $Z_{S}$ is the constructed instrument vector.
+The estimand in Angrist and Imbens (1995) is given by
+$\beta_{\text{2SLS}}\coloneqq\left[\left(\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}W_{S}\right]\right)^{-1}\mathbb{E}\left[W_{S}^{\prime}Q_{S}\right]\left(\mathbb{E}\left[Q_{S}^{\prime}Q_{S}\right]\right)^{-1}\mathbb{E}\left[Q_{S}^{\prime}Y\right]\right]_{1},$
+where $W_{S}=\left(D,X_{S}\right)$, $Q_{S}=\left(Z_{S},X_{S}\right)$, and $\left[\cdot\right]_{k}$ denotes the $k$th element of the corresponding vector.
+This estimand has been studied by Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022), and the representation in Proposition 5.1 follows from Słoczyński (2020).
+Proposition 5.1.
+Suppose Assumption 5.1 holds. Suppose $X$ is discrete with finite support. Then
+$\beta_{\text{2SLS}}$ $=\frac{\mathbb{E}\left[|\operatorname{cov}(D,Z\mid X)|\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]\mid D(1)>D(0)\right]}{\mathbb{E}[|\operatorname{cov}(D,Z\mid X)|\mid D(1)>D(0)]}.$
+Thus, $\beta_{\text{2SLS}}$ satisfies the representation in (1.1) with $a_{\text{2SLS}}(X)=|\operatorname{cov}(D,Z\mid X)|$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]$, and $W_{0}=\mathbbm{1}(D(1)>D(0))$.
+Note that $\beta_{\text{2SLS}}=\text{LATE}\coloneqq\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0)]$ if and only if $a_{\text{2SLS}}(X)$ is uncorrelated with $\tau_{0}(X)$ given $D(1)>D(0)$.
+The practical limitation of focusing on $\beta_{\text{2SLS}}$ is that applied researchers rarely create multiple interacted instruments (cf.
+Blandhol, Bonney, Mogstad, and Torgovitsky, 2022), which is how $Z_{S}$ is defined and used to obtain $\beta_{\text{2SLS}}$ above.
+A more practically relevant estimand is the “noninteracted” IV estimand,
+$\beta_{\text{IV}}\coloneqq\left[\left(\mathbb{E}\left[Q^{\prime}W\right]\right)^{-1}\mathbb{E}\left[Q^{\prime}Y\right]\right]_{1},$
+where $Q=\left(Z,X\right)$ and $W=\left(D,X\right)$.
+We also make the following “rich covariates” assumption on the instrument propensity score, which is implied by the saturated specification in Proposition 5.1.
+Assumption 5.2 (Rich covariates).
+$\mathbb{P}(Z=1\mid X)$ is linear in $X$.
+Under the instrument validity assumption and the rich covariates assumption, Słoczyński (2020) obtains the following representation of the “noninteracted” IV estimand.
+Proposition 5.2.
+Suppose Assumptions 5.1 and 5.2 hold. Then
+$\beta_{\text{IV}}$ $=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\cdot\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]\mid D(1)>D(0)]}{\mathbb{E}[\operatorname{var}(Z\mid X)\mid D(1)>D(0)]}.$
+It follows that $\beta_{\text{IV}}$ is a weighted estimand satisfying (1.1) with weights $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$, CATE function $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0),X]$, and where the average is again taken over the complier subpopulation, i.e., $W_{0}=\mathbbm{1}(D(1)>D(0))$.
+Causal Representation and Internal Validity of 2SLS
+First consider the estimand $\beta_{\text{2SLS}}$, which can be characterized as $\mu(a_{\text{2SLS}},\tau_{0})$.
+Since $a_{\text{2SLS}}(X)\geq 0$, there exists a subpopulation of $\{D(1)>D(0)\}$ such that $\beta_{\text{2SLS}}$ is an average treatment effect over that subpopulation.
 The maximum size of that subpopulation is given by
-$\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{2SLS}}(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}$
-$=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)\neq D(0)\mid X)^{2}]}{\mathbb{P}(D(1)\neq D(0))\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\operatorname{var}(Z\mid X=x)\cdot\mathbb{P}(D(1)\neq D(0)\mid X=x)}.$
-Under Assumptions 2.2 and 2.4, $\mathbb{P}(D(1)\neq D(0)\mid X)$ is identified from the distribution of $(D,X,Z)$ as $w_{0}(X)\equiv|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|$, which is the fraction of compliers or defiers, i.e.
-units for which $D(1)\neq D(0)$, among those with covariates $X$.
-The maximum value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is obtained when $\operatorname{var}(Z\mid X)\cdot|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|=|\operatorname{cov}(Z,D\mid X)|$ does not depend on $X$.
-This occurs, for example, when the instrument and the fraction of units for which $D(1)\neq D(0)$ are independent of $X$.
-In this case, we have that $\beta_{\text{2SLS}}=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, the average effect of treatment in the complier/defier subpopulation.
-The IV estimand has the same $W_{0}$, $w_{0}(X)$, but has $a_{\text{IV}}(X)=\text{sign}(\mathbb{P}(D(1)\geq D(0)\mid X)-\mathbb{P}(D(1)\leq D(0)\mid X))\cdot\operatorname{var}(Z\mid X)\equiv c(X)\operatorname{var}(Z\mid X)$ instead.
-First, we notice that if $\mathbb{P}(c(X)=-1)>0$, then weights are negative with positive probability and there does not exist a causal representation for the estimand $\beta_{\text{IV}}$ that is uniform in $\tau_{0}\in\mathcal{T}_{\text{all}}$.
-However, there will exist a causal representation given $\tau_{0}$ if the support condition of Theorem 3.2 holds, in that $\beta_{\text{IV}}$ lies in the support of $\tau_{0}(X)$.
-If we assume strong monotonicity (Assumption 2.3), then $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)\geq 0$ and
-$\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{IV}}(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}$
-$=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\cdot\mathbb{P}(D(1)>D(0)\mid X)]}{\mathbb{P}(D(1)>D(0))\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\operatorname{var}(Z\mid X=x)}.$
-Here the representativeness of the IV estimand is maximized when $\operatorname{var}(Z\mid X)$ is constant, which occurs when $Z$ is independent of $X$.
+$\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{2SLS}}(X)\mid W_{0}=1]}{\sup_{x\in\text{supp}(X\mid W_{0}=1)}a_{\text{2SLS}}(x)}=\frac{\mathbb{E}[|\operatorname{cov}(D,Z\mid X)|\mid D(1)>D(0)]}{\sup_{x\in\text{supp}(X\mid D(1)>D(0))}|\operatorname{cov}(D,Z\mid X=x)|}.$
+The maximum value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is obtained when $|\operatorname{cov}(D,Z\mid X)|$ does not depend on $X$.
+This occurs, for example, when the instrument and the fraction of units for which $D(1)>D(0)$ are independent of $X$.
+In this case, we have that $\beta_{\text{2SLS}}=\mathbb{E}[Y(1)-Y(0)\mid D(1)>D(0)]$, the average treatment effect for compliers.
+Under the representation in Proposition 5.2, the IV estimand has the same $W_{0}$, but has $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ instead. Here, $a_{\text{IV}}(X)\geq 0$ and
+$\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{\text{IV}}(X)\mid W_{0}=1]}{\sup_{x\in\text{supp}(X\mid W_{0}=1)}a_{\text{IV}}(x)}=\frac{\mathbb{E}[\operatorname{var}(Z\mid X)\mid D(1)>D(0)]}{\sup_{x\in\text{supp}(X\mid D(1)>D(0))}\operatorname{var}(Z\mid X=x)}.$
+The internal validity of the IV estimand is maximized when $\operatorname{var}(Z\mid X)$ is constant, which occurs when $Z$ is independent of $X$.
 In this case, $\beta_{\text{IV}}$ equals LATE.
 The quantities $\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$ and $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})$ are not ranked uniformly in the distributions of $(D(1),D(0),X,Z)$ as there are data-generating processes that make each of these two quantities larger than the other.
-For example, if $\operatorname{var}(Z\mid X)$ is constant but $\mathbb{P}(D(1)\neq D(0)\mid X)$ is not, then $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})<\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$.
+For example, if $\operatorname{var}(Z\mid X)$ is constant but $\mathbb{P}(D(1)>D(0)\mid X)$ is not, then $\overline{P}(a_{\text{2SLS}},W_{0};\mathcal{T}_{\text{all}})<\overline{P}(a_{\text{IV}},W_{0};\mathcal{T}_{\text{all}})$.
 This scenario is plausible if $Z$ is randomly assigned and $X$ is a vector of pre-assignment characteristics.
-This inequality is reversed if $a_{\text{2SLS}}(X)=|\operatorname{cov}(Z,D\mid X)|$ is constant but $\mathbb{P}(D(1)\neq D(0)\mid X)$ is not.
-They are equally representative when $\mathbb{P}(D(1)\neq D(0)\mid X)$ is constant.
+This inequality is reversed if $a_{\text{2SLS}}(X)=|\operatorname{cov}(D,Z\mid X)|$ is constant but $\mathbb{P}(D(1)>D(0)\mid X)$ is not.
+They are equally representative when $\mathbb{P}(D(1)>D(0)\mid X)$ is constant.
 In this case, the estimands are equal, so this is not unexpected.
-4.4.3 Differences-in-Differences
+5.3   Difference-in-Differences
+Now suppose units are observed for $T$ periods and, for $t\in\{1,\ldots,T\}$, denote binary treatment by $D_{t}\in\{0,1\}$, potential outcomes $(Y_{t}(1),Y_{t}(0))$, and realized outcome $Y_{t}=Y_{t}(D_{t})$.
+We assume units are untreated prior to period $G\in\{2,3,\ldots,T\}\cup\{+\infty\}$, receive the treatment in period $G$, and remain treated thereafter.
+We assume no units are treated in the first time period.
+This may include a group that remains untreated throughout, for which $G=+\infty$.
+Thus, $D_{t}=\mathbbm{1}(G\leq t)$.
+The panel is balanced, that is, no group appears or disappears over time.
+The two-way fixed effects estimand is often used in this setting, and consists of regressing the outcome on the treatment indicator, group indicators, and period indicators.
+By partitioned regression results, the coefficient on treatment indicator is
+$\beta_{\text{TWFE}}\coloneqq\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}Y_{t}\right]\big{/}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\left[\ddot{D}_{t}^{2}\right],$
+where $\ddot{D}_{t}=D_{t}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]$.
+We assume a version of parallel trends most similar to the one in de Chaisemartin and D’Haultfœuille (2020).
+Assumption 5.3 (Difference-in-differences).
+We have
+  1. 1.
+    $\text{supp}(G)=\{2,3,\ldots,T\}\cup\{+\infty\}$;
+  2. 2.
+    For all $t\in\{2,\ldots,T\}$ and $g,g^{\prime}\in\text{supp}(G)$, we have that $\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g]=\mathbb{E}[Y_{t}(0)-Y_{t-1}(0)\mid G=g^{\prime}]$.
+We use a proposition that is essentially a special case of Theorem 1 in de Chaisemartin and D’Haultfœuille (2020) to obtain a representation of $\beta_{\text{TWFE}}$ as a weighted average.
+Proposition 5.3.
+Suppose Assumption 5.3 holds. Then
+$\beta_{\text{TWFE}}=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}a_{\text{TWFE}}(G,t)\cdot\mathbb{P}(D_{t}=1\mid G)\cdot\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid G,D_{t}=1]\Big{]}}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\Big{[}a_{\text{TWFE}}(G,t)\cdot\mathbb{P}(D_{t}=1\mid G)\Big{]}},$
+where $a_{\text{TWFE}}(g,t)=1-\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}\mid G=g]-\mathbb{E}[D_{t}]+\frac{1}{T}\sum_{s=1}^{T}\mathbb{E}[D_{s}]$.
+We show the above representation satisfies equation (1.1) by introducing an auxiliary variable $P$ that is uniformly distributed on $\{1,\ldots,T\}$ independently from $\{(Y_{t}(0),Y_{t}(1),G)\}_{t=1}^{T}$.
+This period variable denotes the time period and we use it to define $(Y(1),Y(0),Y,D)\coloneqq(Y_{P}(1),Y_{P}(0),Y_{P},D_{P})$, which are potential outcomes, the realized outcome, and treatment at random period $P$, respectively.
+Letting $X=(G,P)$, this means we can write $\beta_{\text{TWFE}}$ as
+$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a_{\text{TWFE}}(X)\cdot\tau_{0}(X)\mid D=1]}{\mathbb{E}[a_{\text{TWFE}}(X)\mid D=1]},$
+[…]
+Proposition 5.4.
+Suppose Assumption 5.3 holds and that $\mathbb{E}[Y_{t}(1)-Y_{t}(0)\mid D=1,G]=\mathbb{E}[Y_{s}(1)-Y_{s}(0)\mid D=1,G]$ for any $s,t\in\{1,\ldots,T\}$. Then
+$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[a_{\text{TWFE,H}}(G)\cdot\mathbb{E}[Y(1)-Y(0)\mid D=1,G]\mid D=1]}{\mathbb{E}[a_{\text{TWFE,H}}(G)\mid D=1]},$
+where $a_{\text{TWFE,H}}(g)=\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\geq 0$ for $g\in\{2,\ldots,T\}$.
+As is the case of the representation in Proposition 5.3, the two-way fixed effects estimand in Proposition 5.4 satisfies the representation in (1.1), with $X=G$, $W_{0}=D$, $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid D=1,G]$, and the weight function $a_{\text{TWFE,H}}(G)\geq 0$.
+This weight function is derived in Appendix E, with additional comparisons with the weights in Goodman-Bacon (2021) in Appendix H.
+Causal Representation and Internal Validity of TWFE
 We now consider the weights obtained in Proposition 5.4 under its assumptions.
 These weights are nonnegative and therefore Theorem 3.1 guarantees the existence of a causal representation for $\beta_{\text{TWFE}}$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$.
-Using Theorem 4.1, the level of internal validity of $\beta_{\text{TWFE}}$ relative to $\mathbb{E}[Y(1)-Y(0)\mid D=1]$ is given by
-$\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ $=\frac{\mathbb{E}[a_{B}(G)w_{0}(G)]}{\mathbb{E}[w_{0}(G)]\cdot\sup_{g\in\text{supp}(G\mid D=1)}a_{B}(g)}$
+Using Theorem 4.1, the internal validity of $\beta_{\text{TWFE}}$ relative to target parameter $\mathbb{E}[Y(1)-Y(0)\mid D=1]$ is given by
+$\overline{P}(a_{\text{TWFE,H}},D;\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a_{\text{TWFE,H}}(G)\mid D=1]}{\sup_{g\in\text{supp}(G\mid D=1)}a_{\text{TWFE,H}}(g)}$
 $=\frac{\sum_{g=2}^{T}\operatorname{var}(D\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))\cdot\mathbb{P}(G=g)}{\mathbb{P}(D=1)\cdot\sup_{g\in\{2,\ldots,T\}}\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=0\mid P\geq g)+\mathbb{P}(D=1\mid P<g))}.$
-Due to the absorbing nature of treatment in our setting, all expressions involving the distribution of $D$ given $P$ or $G$ can be derived as a function of the marginal distribution of $G$.
-Therefore, $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ depends only on $\{\mathbb{P}(G=g)\}_{g\in\{2,\ldots,T\}}$.
-To give some intuition, consider the case where $T=3$ and therefore $G\in\{2,3,+\infty\}$.
-In this case, $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ will simply be a function of $(\mathbb{P}(G=2),\mathbb{P}(G=3))$.
-Calculations yield that
-$\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ =$=\begin{cases}\frac{2\mathbb{P}(G=2)+\omega\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)}&\text{ if }\omega<1\\ \frac{\omega^{-1}2\mathbb{P}(G=2)+\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)}&\text{ if }\omega>1\\ 1&\text{ if }\omega=1,\end{cases}$
-where
-$\omega$ $=\frac{4-2\mathbb{P}(G=2)-4\mathbb{P}(G=3)}{2-2\mathbb{P}(G=2)-\mathbb{P}(G=3)}=\frac{a_{B}(2)}{a_{B}(3)}.$
-Therefore, the TWFE estimand is perfectly representative of the ATT if and only if $\omega=1$.
-With some algebra, we can see that $\omega<1$ if and only if $\mathbb{P}(G=3)>2/3$, while $\omega>1$ if and only if $\mathbb{P}(G=3)<2/3$.
-We can also see that $\overline{P}(a_{B},D;\mathcal{T}_{\text{all}})$ is equal to 1 when $\mathbb{P}(G=3)=2/3$, and that the representativeness of the TWFE estimand declines as $|\mathbb{P}(G=3)-2/3|$ increases.
-This is due to weight function $a_{B}(g)$ being constant in $g$ if and only if $\mathbb{P}(G=3)$ equals 2/3.
-Whenever the weight function is constant, we have that $\mu(a,\tau_{0})=\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]=\mathbb{E}[\tau_{0}(X)\mid W_{0}=1]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, and therefore the weighted estimand equals the average effect over population $\{W_{0}=1\}$.
-5 Discussion of Estimation and Inference
-We now briefly discuss procedures to estimate $\overline{P}(a,W_{0};\mathcal{T})$.
-First consider the case when $\mathcal{T}=\mathcal{T}_{\text{all}}$. Then, we seek to estimate
-$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}=\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}a(x)}.$ (5.1)
-Suppose we observe a random sample of size $n$, $\{(W_{i},X_{i})\}_{i=1}^{n}$, where $W_{i}$ are a set of variables that allow the estimation of $a(\cdot)$ and $w_{0}(\cdot)$.
-For example, under unconfoundedness we can let $W_{i}=(Y_{i},D_{i})$ or $W_{i}=D_{i}$ since only the distribution of $(D,X)$ is needed to identify $a(\cdot)$.
-In our instrumental variables examples, one can let $W_{i}=(Z_{i},D_{i})$.
-Assuming the existence of estimators for $a(\cdot)$ and $w_{0}(\cdot)$, we can propose the following analog estimator:
-$\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})\cdot\sup_{x\in\text{supp}(X\mid W_{0}=1)}\widehat{a}(x)}.$
-This estimator requires knowledge of the support of $X$ given $W_{0}=1$, but can be implemented by taking the maximum over the empirical distribution of $X\mid W_{0}=1$.
-Consistency of $\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$ for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ is achieved when
+Due to the absorbing nature of the treatment in our setting, all expressions involving the distribution of $D$ given $P$ or $G$ can be derived as a function of the marginal distribution of $G$.
+Therefore, $\overline{P}(a_{\text{TWFE,H}},D;\mathcal{T}_{\text{all}})$ depends only on $\{\mathbb{P}(G=g)\}_{g\in\{2,\ldots,T\}}$.
+To give some intuition, consider the case where $T=3$ and therefore $G\in\{2,3,+\infty\}$. In this case, calculations yield that
+$\overline{P}(a_{\text{TWFE,H}},D;\mathcal{T}_{\text{all}})$ $=\min\left\{\frac{2\mathbb{P}(G=2)+\omega\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)},\frac{\omega^{-1}2\mathbb{P}(G=2)+\mathbb{P}(G=3)}{2\mathbb{P}(G=2)+\mathbb{P}(G=3)}\right\},$
+where $\omega=\frac{4-2\mathbb{P}(G=2)-4\mathbb{P}(G=3)}{2-2\mathbb{P}(G=2)-\mathbb{P}(G=3)}=\frac{a_{\text{TWFE,H}}(2)}{a_{\text{TWFE,H}}(3)}$.
+Therefore, $\beta_{\text{TWFE}}$ is perfectly representative of the ATT if and only if $\omega=1$, which occurs if and only if $\mathbb{P}(G=3)=2/3$.
+Thus, $\overline{P}(a_{\text{TWFE,H}},D;\mathcal{T}_{\text{all}})=1$ when $\mathbb{P}(G=3)=2/3$ and the internal validity of $\beta_{\text{TWFE}}$ declines as $|\mathbb{P}(G=3)-2/3|$ increases.
+This is due to the weight function $a_{\text{TWFE,H}}(g)$ being constant in $g$ if and only if $\mathbb{P}(G=3)=2/3$.
+As before, constant weights imply that the weighted estimand equals the average treatment effect over $\{W_{0}=1\}$.
+6   Estimation and Inference
+We now consider estimation and inference for our measures of internal validity and representativeness.
+We focus our attention on the case when $\mathcal{T}=\mathcal{T}_{\text{all}}$ and briefly discuss the case where $\mathcal{T}=\{\tau_{0}\}$ in Appendix C.
+To measure internal validity, we seek to estimate
+$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{a_{\max}}=\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot a_{\max}}.$
+Suppose we observe a random sample of size $n$, $\{(W_{i},X_{i})\}_{i=1}^{n}$, where $W_{i}$ is a set of variables needed to estimate $a(\cdot)$ and $w_{0}(\cdot)$.
+For example, under unconfoundedness we can let $W_{i}=D_{i}$ since the distribution of $(D,X)$ is sufficient to identify $a(\cdot)$; the outcome’s distribution does not affect $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$.
+In our instrumental variables examples, we let $W_{i}=(D_{i},Z_{i})$.
+Assuming the existence of estimators for $a(\cdot)$ and $w_{0}(\cdot)$, we consider the following analog estimator of $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$:
+$\widehat{\overline{P}}$ $\coloneqq\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})\cdot\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}.$
+Here $c_{n}$ is a tuning parameter that converges to 0 as $n$ diverges.
+We start by noting that we can estimate $\mathbb{E}[a(X)\mid W_{0}=1]$ via $\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, which will be consistent under standard conditions on $\widehat{a}(\cdot)$ and $\widehat{w}_{0}(\cdot)$.
+Estimating $a_{\max}=\sup(\text{supp}(a(X)\mid W_{0}=1))$ is more delicate.
+In some of our examples, this supremum is known or can be bounded above without using data.
+For example, the OLS estimand under unconfoundedness has weights $a(X)=\operatorname{var}(D\mid X)$ which are bounded above by $1/4$.
+If $X$ is continuously distributed, then 1/2 may lie in the support of $p(X)$, and thus we may avoid the estimation of $a_{\max}$.
+The IV estimand of Section 5 has weights $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ which are similarly bounded above by $1/4$.
+Similarly, $a_{\text{2SLS}}(X)=|\operatorname{cov}(D,Z\mid X)|\leq 1/4$ by the Cauchy–Schwarz inequality.
+If knowledge of $a_{\max}$ is not assumed, but $\text{supp}(X\mid W_{0}=1)$ is known and $a(x)$ is continuous, [^4] Note that $a(x)$ is trivially continuous on finite support.
+then $\sup_{x\in\text{supp}(X\mid W_{0}=1)}\widehat{a}(x)$ will be consistent for $a_{\max}$ when $\widehat{a}(x)$ is consistent for $a(x)$ uniformly in $x\in\text{supp}(X\mid W_{0}=1)$.
+Many parametric and nonparametric estimators for $a(\cdot)$ satisfy this requirement.
+In Appendix C, we prove the consistency of $\widehat{\overline{P}}$ and derive its limiting distribution.
+We also provide a step-by-step bootstrap algorithm that can be employed to conduct inference and prove its validity.
+This bootstrap approach is based on Fang and Santos (2019) and is nonstandard, but yields valid inferences even when $a_{\max}$ is estimated, as opposed to standard bootstrap approaches, such as the empirical bootstrap.
+7   Empirical Application
+In this section, we implement the proposed tools in an application to the effects of unilateral divorce laws in the U.S.
+on female suicide, as in Stevenson and Wolfers (2006).
+Between 1969 and 1985, 37 states (including D.C.) reformed their law by enabling each spouse to seek divorce without the other spouse’s consent.
+Stevenson and Wolfers (2006) argue that these “unilateral” or “no-fault” divorce laws reduced female suicide, domestic violence, and spousal homicide.
+The results on female suicide are also replicated by Goodman-Bacon (2021), whose analysis we follow here.
+Our sample consists of 41 states observed over the 1964–1996 period.
+The outcome of interest is the state- and year-specific female suicide rate, as computed by the National Center for Health Statistics.
+The treatment is whether the state allowed unilateral divorce in a given year.
+Following Goodman-Bacon (2021), our sample omits Alaska and Hawaii.
+We also omit eight further states which had unilateral divorce laws preceding 1964 and are therefore always treated within our timeframe.
+Panel A of Table 1 reports our baseline estimates of the average effects of unilateral divorce laws on female suicide.
+After we drop the eight always-treated states, the TWFE estimate, –0.604, becomes much smaller in absolute value than the corresponding estimate in Goodman-Bacon (2021), –3.080.
+Unlike that estimate, ours is also statistically insignificant, with $p$-value = 0.819.
+The conclusion changes, however, when we explicitly target the average treatment effect on the treated (ATT), that is, the average effect for the largest subpopulation for which such an effect is identified under standard assumptions.
+Using the approach of Wooldridge (2021), we obtain an estimate of –5.530 with a $p$-value of 0.138.
+The approach of Callaway and Sant’Anna (2021) produces an estimate of –10.220 and a $p$-value of 0.001.
+These estimates are more strongly suggestive of a causal effect of unilateral divorce laws than the TWFE estimate.
+{adjustwidth}
+-1.25in-1.25in A.
+Estimates of the effects of unilateral divorce laws TWFE ATT Callaway and Sant’Anna Wooldridge –0.604 –10.220 –5.530 (2.622) (3.086) (3.650) B.
+Internal validity of the TWFE estimand based on Proposition 5.3 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216 C.
+Internal validity of the TWFE estimand based on Proposition 5.4 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
+Table 1: Internal Validity of TWFE Estimand for Effects of Unilateral Divorce Laws
+  • Notes: The data are a panel of the 1964–1996 U.S.
+The outcome is the state- and year-specific female suicide rate (per million women).
+The treatment is whether the state allowed unilateral divorce in a given year.
+The sample includes D.C.
+but excludes the states excluded by Goodman-Bacon (2021) as well as eight additional always-treated states.
+The measures of internal validity “given $\tau_{0}$” require an estimate of the CATE function, which we obtain using the approach of Wooldridge (2021).
+While the TWFE estimate and the two estimates of the ATT are quite different, this paper focuses on another implication of the nonuniformity of the TWFE weight function.
+We ask: How representative of the underlying population is the TWFE estimand?
+What is the internal validity of this estimand if we are interested in the treated subpopulation?
+Panel B of Table 1 reports our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$, based on the representation of the TWFE estimand in de Chaisemartin and D’Haultfœuille (2020), revisited in our Proposition 5.3.
+First, because the weights on some group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, ^$\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$.
+Second, when we estimate the CATE function and use these estimates in constructing the bounds, we conclude that the TWFE estimand corresponds to the average treatment effect for at most 62.16% of the treated units or 38.73% of the entire population.
+Panel C of Table 1 revisits these questions on the basis of the representation of the TWFE estimand in Proposition 5.4.
+Here, we assume that group-time average treatment effects are constant over time, which eliminates the problem of negative weights.
+Indeed, we now conclude that the TWFE estimand has a causal interpretation uniformly in $\tau_{0}$, even if it is still not particularly representative of the underlying population or the treated subpopulation.
+Our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$ are equal to 14.00% and 22.46%, respectively.
+When we use the estimated CATE function in constructing the bounds, these estimates increase to 48.02% and 77.07%.
+This is obviously much more than our initial estimate of 0, but still substantially less than 1, guaranteed in the case of $\mathbb{P}(W^{*}=1\mid D=1)$ when using the estimation methods in Callaway and Sant’Anna (2021), Wooldridge (2021), and other recent papers, each of which explicitly targets the ATT.
+8   Conclusion
+In this paper, we studied the representativeness and internal validity of a class of weighted estimands, which includes the popular OLS, 2SLS, and TWFE estimands in additive linear models.
+We examined the conditions under which such estimands can be written as the average treatment effect over a subpopulation.
+When a given estimand can be shown to correspond to the average treatment effect for a large subset of the population of interest, we say its internal validity is high.
+In our main results, we derived the sharp upper bound on the size of that subpopulation under different assumptions on treatment effect heterogeneity, which offers a practical tool to quantify the internal validity of weighted estimands.
+Appendix A Proofs of Key Results
+This appendix contains the proofs for our key results in Sections 3 and 4. Proofs for other results, including Theorem 4.2, can be found in the Supplemental Appendix.
+Proof of Proposition 3.1.
+To show the first claim, note that the equation $\mathbb{E}[W^{*}(Y(1)-Y(0))\mid W_{0}=1,X]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\cdot\underline{w}^{*}(X)$ holds since $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$ by Definition 3.1.
+Since $\underline{w}^{*}(X)>0$, we can obtain
+$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ $=\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]/\underline{w}^{*}(X)$ (A.1)
+$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1,X]$
+$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X],$
+where the third equality holds from $W^{*}$ being a subpopulation of $W_{0}$.
+The proposition’s second claim is established below:
+$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1]$
+$=\frac{\mathbb{E}[\mathbb{E}[W^{*}(Y(1)-Y(0))\mid X,W_{0}=1]\mid W_{0}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\mid W_{0}=1]}$
+$=\frac{\mathbb{E}[\underline{w}^{*}(X)\cdot\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
+$=\mu(\underline{w}^{*},\tau_{0}).$
+The first equality follows from $W^{*}$ being a subpopulation of $W_{0}$, the second from the law of iterated expectations and $W^{*}\leq W_{0}$, and the third from equation (A.1). ∎
+Proof of Theorem 3.1.
+()($\Longrightarrow$)( ⟹ ) First, suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$.
+By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{all}}$.
+Let $\tau^{*}(X)=\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}$.
+We have $\mathbb{E}[a(X)^{2}]<\infty$ by Assumption 3.1 and $\mathbb{E}[\underline{w}^{*}(X)^{2}]\leq 1$, which holds by construction.
+Hence, $\mathbb{E}[\tau^{*}(X)^{2}]<\infty$, and thus $\tau^{*}\in\mathcal{T}_{\text{all}}$.
+Thus, we must have $\mu(a,\tau^{*})-\mu(\underline{w}^{*},\tau^{*})=0$.
+Expanding this equality yields $0=\mu(a,\tau^{*})-\mu(\underline{w}^{*},\tau^{*})=\mathbb{E}[\tau^{*}(X)^{2}\mid W_{0}=1]$, which implies that
+11 11 $=\mathbb{P}(\tau^{*}(X)=0\mid W_{0}=1)=\mathbb{P}\left(a(X)=\frac{\underline{w}^{*}(X)\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}\mid W_{0}=1\right).$ (A.2)
+Note that $\underline{w}^{*}(X)\geq 0$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)>0$ by construction.
+Also, $\mathbb{E}[a(X)\mid W_{0}=1]>0$ by Assumption 3.1.
+Therefore, $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$.
 […]
-and $\widehat{a}(x)$ is uniformly consistent for $a(x)$ over $x\in\text{supp}(X\mid W_{0}=1)$.
-These conditions can be achieved under many standard estimators of $a(\cdot)$ and $w_{0}(\cdot)$.
-For example, when $a(X)$ and $w_{0}(X)$ consist of continuous functions of conditional means, as is the case when $a(X)=\operatorname{var}(D\mid X)$ or, for the IV estimator, $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ and $w_{0}(X)=|\mathbb{P}(D=1\mid X,Z=1)-\mathbb{P}(D=1\mid X,Z=0)|$, consistent estimators of these conditional means will result in the consistency of $\widehat{\overline{P}}(a,W_{0};\mathcal{T}_{\text{all}})$.
-Cell-by-cell estimators can be used when $X_{i}$ is discrete, or many nonparametric estimators when $X_{i}$ contains continuous components.
-See Hansen (2022) for a recent textbook treatment of such methods.
-Inference will generally be nonstandard.
-This is because the asymptotic behavior of the denominator will usually be non-Gaussian due to the lack of Hadamard differentiability of the supremum operator.
-This lack of differentiability will invalidate standard bootstrap inference: see Fang and Santos (2019).
-However, the supremum is a Hadamard directionally differentiable functional and nonstandard bootstrap approaches, such as those proposed in Hong and Li (2018) and Fang and Santos (2019), can be used.
-We can also view these inferential problems through the lens of intersection or union bounds. For example, we can write
-$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\inf_{x\in\text{supp}(X\mid W_{0}=1)}\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot a(x)}\equiv\inf_{x\in\text{supp}(X\mid W_{0}=1)}\overline{P}(x).$
+$\mathbb{P}(W^{*}=1)$ $=\mathbb{E}\left[\mathbb{P}\left(U\leq a(X)/a_{\max}\mid W_{0}=1,X\right)\mid W_{0}=1\right]\cdot\mathbb{P}(W_{0}=1)$
+$=\mathbb{E}\left[a(X)/a_{\max}\mid W_{0}=1\right]\cdot\mathbb{P}(W_{0}=1)>0.$
+The first equality follows from $\mathbb{P}(W^{*}=1\mid W_{0}=0)=0$ and $U\perp\!\!\!\perp(X,W_{0})$, the second equality from $a(X)/a_{\max}\in[0,1]$ almost surely given $W_{0}=1$, and the inequality from $\mathbb{P}(W_{0}=1)$ and $\mathbb{E}[a(X)\mid W_{0}=1]$ being positive by assumption, and by $a_{\max}<\infty$.
+That $W^{*}$ satisfies the two properties of Definition 3.1 holds immediately.
+Therefore, $W^{*}$ is a regular subpopulation of $W_{0}$.
+Finally, let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)=a(X)/a_{\max}$. Using Proposition 3.1, we can see that for a given $\tau_{0}\in\mathcal{T}_{\text{all}}$
+$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mu(\underline{w}^{*},\tau_{0})=\mu(a/a_{\max},\tau_{0})=\mu(a,\tau_{0}),$
+where the last equality follows from the scale invariance of the estimands.
+Since $\tau_{0}\in\mathcal{T}_{\text{all}}$ was arbitrary, we have that $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$.
+∎
+Proof of Theorem 3.2.
+To simplify the notation in the proof, we let $\mu_{0}\coloneqq\mu(a,\tau_{0})$.
+()($\Longrightarrow$)( ⟹ ) First, suppose $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$.
+Since $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$, we can without loss of generality suppose that $\mathbb{P}(\tau_{0}(X)>\mu_{0}\mid W_{0}=1)=1$.
+Since $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$, we can write by Proposition 3.1
+$\mu_{0}$ $=\mu(\underline{w}^{*},\tau_{0})\geq\mu(\underline{w}^{*},\mu_{0})=\mu_{0}.$ (A.3)
+The inequality is strict unless $\mathbb{E}[(\underbrace{\tau_{0}(X)-\mu_{0}}_{>0\text{ w.p.1}})\underbrace{\underline{w}^{*}(X)}_{\in[0,1]}\mid W_{0}=1]=0$ holds.
+This holds if $\mathbb{P}((\tau_{0}(X)-\mu_{0})\underline{w}^{*}(X)=0\mid W_{0}=1)=1$, which in turns occurs if and only if $\mathbb{P}(\underline{w}^{*}(X)=0\mid W_{0}=1)=1$.
+This implies $\mathbb{P}(W^{*}=1\mid W_{0}=1)=0$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W_{0}=1)\cdot\mathbb{P}(W_{0}=1)=0$, a contradiction of $W^{*}\in\mathcal{W}(a,W_{0},\{\tau_{0}\})$.
+Therefore, the inequality in (A.3) is strict and yields $\mu_{0}>\mu_{0}$, a contradiction.
+Thus, $\mathcal{W}(a;W_{0},\{\tau_{0}\})=\emptyset$ when $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$.
+[…]
+Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$. For $u\in[0,1]$, let
+$W^{*}(u)$ $=(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+}))\cdot W_{0}.$
+We show $W^{*}(u)$ is a regular subpopulation of $W_{0}$ for all $u\in[0,1]$.
+We can see that $W^{*}(u)\in\{0,1\}$, that $W^{*}(u)\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, and that $\mathbb{P}(W_{0}=1\mid W^{*}(u)=1)=1$.
+To show that $W^{*}(u)$ characterizes a regular subpopulation of $W_{0}$, we also show that it is nonzero with positive probability:
+$\mathbb{P}(W^{*}(u)=1\mid W_{0}=1)$ $=(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W_{0}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W_{0}=1)>0$
+for all $u\in[0,1]$, which implies $\mathbb{P}(W^{*}(u)=1)>0$ by $\mathbb{P}(W_{0}=1)>0$.
+Hence, $W^{*}(u)\in\text{SP}(W_{0})$ for all $u\in[0,1]$.
+For $u\in[0,1]$, we have that $\underline{w}^{*}(X;u)\coloneqq\mathbb{P}(W^{*}(u)=1\mid X,W_{0}=1)=(1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+})$.
+Therefore, using Proposition 3.1,
+$\mathbb{E}[Y(1)-$ $Y(0)\mid W^{*}(u)=1]=\mu(\underline{w}^{*}(\cdot;u),\tau_{0})$
+$=\frac{(1-u)\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W_{0}=1]+u\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W_{0}=1]}{(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W_{0}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W_{0}=1)}.$
+By construction, $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{-})\leq\mu_{0}\mathbbm{1}(X\in\mathcal{X}^{-})$ and $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{+})\geq\mu_{0}\mathbbm{1}(X\in\mathcal{X}^{+})$ almost surely.
+Therefore, $\mathbb{E}[Y(1)-Y(0)\mid W^{*}(0)=1]\leq\mu_{0}\leq\mathbb{E}[Y(1)-Y(0)\mid W^{*}(1)=1]$.
+By the continuity of $\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ in $u$ and the intermediate value theorem, there exists $u^{*}\in[0,1]$ such that $\mu_{0}=\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u^{*})=1]$ and $W^{*}(u^{*})\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$.
+∎
+Proof of Proposition 3.2.
+()($\Longrightarrow$)( ⟹ ) First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$.
+By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{lin}}$.
+Therefore,
+00 0 $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
+$=\frac{\mathbb{E}[a(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
+$=d^{\prime}\left(\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}\right)$
+for all $d\in\mathbb{R}^{d_{X}}$, which implies $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}=\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$.
+Let $u(x)=\underline{w}^{*}(x)\mathbb{P}(X=x\mid W_{0}=1)/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$.
+We have that $\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}=\sum_{x\in\text{supp}(X\mid W_{0}=1)}xu(x)$, a convex combination of values in $\text{supp}(X\mid W_{0}=1)$ because $u(\cdot)\geq 0$ and $\sum_{x\in\text{supp}(X\mid W_{0}=1)}u(x)=1$.
+Thus, $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}\in\text{conv}(\text{supp}(X\mid W_{0}=1))$.
+[…]
+$\mu(\underline{w}^{*},\tau_{0})$ $=\mu(u/\max(\text{supp}(u(X)\mid W_{0}=1)),\tau_{0})$
+$=\mu(u,\tau_{0})$
+$=\mathbb{E}\left[u(X)(c+d^{\prime}X)\mid W_{0}=1\right]/\mathbb{E}\left[u(X)\mid W_{0}=1\right]$
+$=c+d^{\prime}\mathbb{E}\left[a(X)X\mid W_{0}=1\right]/\mathbb{E}\left[a(X)\mid W_{0}=1\right]$
+$=\mathbb{E}\left[a(X)(c+d^{\prime}X)\mid W_{0}=1\right]/\mathbb{E}\left[a(X)\mid W_{0}=1\right]$
+$=\mu(a,\tau_{0}).$
+Therefore, by Proposition 3.1 we have that $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. ∎
+Proof of Proposition 3.3.
+We consider the $K>0$ case first and the $K=0$ case second.
+Case 1: $K>0$.
+()($\Longrightarrow$)( ⟹ ) First, let $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. We will show that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))=\emptyset$ by way of contradiction.
+Suppose there is a $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))$ and let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\in[0,1]$.
+Let $\tau^{*}(x)=K\cdot\mathbbm{1}(a(x)<0)$, which implies $\tau^{*}\in\mathcal{T}_{\text{BD}}(K)$.
+We have that $\mu(a,\tau^{*})=K\mathbb{E}[a(X)\mathbbm{1}(a(X)<0)\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]$.
+$\mathbb{E}[a(X)\mathbbm{1}(a(X)<0)\mid W_{0}=1]$ is strictly negative because it is weakly negative and because $\mathbb{E}[a(X)\mathbbm{1}(a(X)<0)\mid W_{0}=1]=0$ implies $\mathbb{P}(a(X)=0\mid a(X)<0,W_{0}=1)=1$, a contradiction of $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$.
+Since $K>0$ and $\mathbb{E}[a(X)\mid W_{0}=1]>0$, we conclude that $\mu(a,\tau^{*})<0$.
+However, $\mu(\underline{w}^{*},\tau^{*})\geq 0$ since $\tau^{*}\geq 0$.
+By Proposition 3.1, $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))$ implies $0>\mu(a,\tau^{*})=\mu(\underline{w}^{*},\tau^{*})\geq 0$, a contradiction.
+Therefore $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))=\emptyset$.
+[…]
+Case 2: $K=0$.
+When $K=0$, the function class $\mathcal{T}_{\text{BD}}(K)$ is the set of all constant functions.
+In this case, $\tau_{0}(X)=t_{0}$, where $t_{0}\in\mathbb{R}$ denotes a constant.
+Thus $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(0))\neq\emptyset$ for all weight functions $a(\cdot)$ since $W_{0}\in\text{SP}(W_{0})$ and because $\mu(a,\tau_{0})=\mathbb{E}[a(X)t_{0}\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]=t_{0}=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ for any $a(\cdot)$.
+∎
+Proof of Theorem 4.1.
+First, suppose $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$.
+From Theorem 3.1, there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$.
+From derivations in the proof of Theorem 3.1 (see equation (A.2)) we have that $\mathbb{P}(C\cdot a(X)=\underline{w}^{*}(X)\mid W_{0}=1)=1$ for some positive constant $C>0$.
+Since $\underline{w}^{*}(X)\leq 1$, we must have $C\cdot a(X)\leq 1$ almost surely given $W_{0}=1$.
+Thus $C$ is bounded above by $\inf(\text{supp}(1/a(X)\mid W_{0}=1))=1/a_{\max}$, which is strictly positive by assumption.
+Therefore,
+$\mathbb{P}(W^{*}=1\mid W_{0}=1)$ $=\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]=\mathbb{E}[C\cdot a(X)\mid W_{0}=1]\leq\frac{\mathbb{E}\left[a(X)\mid W_{0}=1\right]}{a_{\max}}.$
+This upper bound is sharp because it is attained by setting $W^{*}=\mathbbm{1}\left(U\leq a(X)/a_{\max}\right)\cdot W_{0}$ and noting that $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$ from the proof of Theorem 3.1.
+Second, suppose $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. By Theorem 3.1, $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})=\emptyset$ and therefore $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})=0$. ∎
+Supplemental Appendix
+This document contains appendices that supplement the main text.
+Appendix B formalizes the connection between uniform causal representations and weakly causal estimands, defined in Blandhol, Bonney, Mogstad, and Torgovitsky (2022).
+Appendix C supplements the estimation and inference section (Section 6) and provides an algorithm for constructing confidence intervals.
+Appendices DG provide proofs for results in the main text and for results in Appendices B and C.
+Appendix H contains additional calculations related to the weights for the TWFE estimand.
+Appendix B Weakly Causal Estimands and Uniform Causal Representations in $\mathcal{T}_{\text{all}}$
+We now establish equivalence between weakly causal estimands as defined in Blandhol, Bonney, Mogstad, and Torgovitsky (2022) (henceforth, BBMT) and estimands that have uniform causal representations as in Theorem 3.1.
+As in BBMT, consider the case where $X$ has finite support and, as in this paper, assume the treatment is binary.
+We also abstract from choice groups denoted by $G$ in BBMT.
+Since $X$ has finite support, let $\text{supp}(X\mid W_{0}=1)=\{x_{1},\ldots,x_{K}\}$ and let $\tau\coloneqq(\tau(x_{1}),\ldots,\tau(x_{K}))\in\mathbb{R}^{K}$ be the collection of CATEs.
+For $d\in\{0,1\}$ let $\nu_{d}(x)\coloneqq\mathbb{E}[Y(d)\mid W_{0}=1,X=x]$ denote the average structural function (ASF) which also conditions on $W_{0}=1$, let $\nu_{d}\coloneqq(\nu_{d}(x_{1}),\ldots,\nu_{d}(x_{K}))\in\mathbb{R}^{K}$, and let $\mathcal{M}\subseteq\mathbb{R}^{2K}$ be a set of possible ASFs such that $(\nu_{0},\nu_{1})\in\mathcal{M}$.
+We now state the definition of weakly causal estimands from BBMT (i.e., their Definition WC) in our setting which features binary treatments.
+Definition B.1.
+The estimand $\beta$ is weakly causal if the following statements are true for all $(\nu_{0},\nu_{1})\in\mathcal{M}$:
+  1. 1.
+    If $\nu_{1}-\nu_{0}\geq\textbf{0}_{K}$, [^5] Vector inequalities hold if they hold component-wise. then $\beta\geq 0$.
+  2. 2.
+    If $\nu_{1}-\nu_{0}\leq\textbf{0}_{K}$, then $\beta\leq 0$.
+Thus, an estimand is weakly causal if all CATEs having the same sign implies the estimand also has that sign.
+Whether an estimand satisfies this condition also depends on $\mathcal{M}$, the set of allowed ASFs.
+To compare weak causality to our result on uniform causal representations, we consider $\mathcal{M}_{\text{all}}\coloneqq\mathbb{R}^{2K}$, the unrestricted set of ASFs.
+The corresponding unrestricted set of CATE functions, which we denoted by $\mathcal{T}_{\text{all}}$, allows $\tau$ to be any vector in $\mathbb{R}^{K}$.
+We consider estimands characterized by our equation (1.1).
+We note that these estimands rule out “level dependence,” i.e., that the estimand changes if potential outcomes $(Y(0),Y(1))$ are translated to $(Y(0)+c,Y(1)+c)$ for some constant $c\in\mathbb{R}$.
+For example, the IV estimand is generally level dependent when the propensity score $\mathbb{P}(Z=1\mid X)$ is nonlinear in $X$ [^6] See p.
+17 in BBMT.
+With these choices, we can show the two definitions are equivalent.
+Proposition B.1.
+Let $\mu(a,\tau_{0})$ be an estimand satisfying equation (1.1).
+Suppose Assumption 3.1 holds and that $a_{\max}<\infty$.
+Then $\mu(a,\tau_{0})$ is weakly causal with $\mathcal{M}=\mathcal{M}_{\text{all}}$ if and only if it has a causal representation uniformly in $\mathcal{T}_{\text{all}}$.
+The proof of this proposition hinges on the equivalence, under level independence, of weakly causal estimands and estimands with nonnegative weights, as in Proposition 4 of BBMT.
+Also, as shown in Theorem 3.1, estimands with nonnegative weights have a uniform causal representation in $\mathcal{T}_{\text{all}}$.
+Therefore, a weighted estimand has nonnegative weights if and only if it is weakly causal and if and only if it has a causal representation uniformly in $\mathcal{T}_{\text{all}}$.
+Thus, a weakly causal estimand admits a regular subpopulation $W^{*}$ such that the estimand measures the average effect of treatment over that subpopulation.
+Appendix C Details on Estimation and Inference
+This appendix complements Section 6 in the main text.
+In it, we compute the limiting distribution of our estimated measure of internal validity and prove the validity of a nonstandard bootstrap algorithm for constructing confidence intervals around it.
+This is done for the case where $\mathcal{T}=\mathcal{T}_{\text{all}}$.
+Estimation and inference for $\overline{P}(a,W_{0};\{\tau_{0}\})$ is related to the question of estimation and inference in linear programs with estimated constraints.
+See Andrews, Roth, and Pakes (2023), Cox and Shi (2023), Fang, Santos, Shaikh, and Torgovitsky (2023), and Cho and Russell (2024) for recent advances on this topic.
+As in Section 6, we consider the case where $X$ is discrete.
+We assume the existence of estimators for $a(\cdot)$ and $w_{0}(\cdot)$, but we do not assume knowledge of $\text{supp}(X\mid W_{0}=1)$.
+In this case, $a(x)$ and $w_{0}(x)$ are usually estimated “cell-by-cell” and their estimators are $\sqrt{n}$-consistent with a limiting Gaussian distribution.
+We will see that inference on $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ is generally nonstandard and, as a result, most common bootstrap procedures fail.
+We consider the following simple plug-in estimator:
+$\widehat{\overline{P}}$ $=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})\cdot\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})},$
+where $c_{n}$ is a tuning parameter that converges to 0 as $n\rightarrow\infty$.
+Note that this tuning parameter is absent when $w_{0}$ is known, for example when $W_{0}=1$ almost surely.
+This estimator does not assume knowledge of the support of $X$ given $W_{0}=1$, but it can also be implemented by taking the maximum over $\text{supp}(X\mid W_{0}=1)$ when it is known.
+Let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, denote by $p_{j}=\mathbb{P}(X=x_{j})$ the frequency of cell $j$, and let $\widehat{p}_{j}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})$ denote its sample frequency.
+Let $\widehat{\theta}=(\widehat{\textbf{a}},\widehat{\textbf{w}}_{0},\widehat{\textbf{p}})$ where $\widehat{\textbf{a}}=(\widehat{a}(x_{1}),\ldots,\widehat{a}(x_{K}))$, $\widehat{\textbf{w}}_{0}=(\widehat{w}_{0}(x_{1}),\ldots,\widehat{w}_{0}(x_{K}))$, and $\widehat{\textbf{p}}=(\widehat{p}_{1},\ldots,\widehat{p}_{K})$.
+Let $\theta=(\textbf{a},\textbf{w}_{0},\textbf{p})$ denote their population counterparts.
+We make the following assumption on the behavior of the first-step estimators.
+Assumption C.1 (Preliminary estimators).
+Let
+[…]
+where $\mathbb{Z}\coloneqq(\mathbb{Z}_{\textbf{a}},\mathbb{Z}_{\textbf{w}_{0}},\mathbb{Z}_{X})\in\mathbb{R}^{3K}$ has a Gaussian distribution.
+The above assumption is often satisfied when $X$ has finite support since estimators for $a(x_{j})$ and $w_{0}(x_{j})$ can be obtained using only the observations for which $X_{i}=x_{j}$.
+Note that the limiting distribution of $\mathbb{Z}_{\textbf{w}_{0}}$ may be degenerate.
+For example, if $W_{0}=1$ almost surely, then $\widehat{w}_{0}(x)=w_{0}(x)=1$ and thus $\mathbb{Z}_{\textbf{w}_{0}}=\textbf{0}_{K}$, a $K$-vector of zeros.
+The next theorem shows the consistency of $\widehat{\overline{P}}$ and establishes the limiting distribution of this estimator.
+To simplify the exposition, we use $\overline{P}$ to denote $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ in what follows.
+Theorem C.1 (Consistency and asymptotic distribution).
+Suppose Assumption C.1 holds.
+Suppose $c_{n}=o(1)$ and $c_{n}\sqrt{n}\rightarrow\infty$ as $n\rightarrow\infty$.
+Suppose $\overline{P}\neq 0$.
+Then, $\widehat{\overline{P}}$ is consistent for $\overline{P}$ and
+[…]
+where $\psi$ is a continuous mapping defined by
+$\psi(\mathbb{Z})\;$ $=\;\sum_{j=1}^{K}\frac{w_{0}(x_{j})p_{j}}{\mathbb{P}(W_{0}=1)a_{\max}}\mathbb{Z}_{\textbf{a}}(j)\;-\;\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{a_{\max}^{2}}\max_{j\in\Psi_{\mathcal{X}^{+}}}\mathbb{Z}_{\textbf{a}}(j)$
+$+\;\sum_{j=1}^{K}\frac{(a(x_{j})-\mathbb{E}[a(X)\mid W_{0}=1])p_{j}}{\mathbb{P}(W_{0}=1)a_{\max}}\mathbb{Z}_{\textbf{w}_{0}}(j)$
+$+\;\sum_{j=1}^{K}\frac{(a(x_{j})-\mathbb{E}[a(X)\mid W_{0}=1])w_{0}(x_{j})}{\mathbb{P}(W_{0}=1)a_{\max}}\mathbb{Z}_{X}(j),$ (A.1)
+where $\Psi_{\mathcal{X}^{+}}=\{j\in\{1,\ldots,K\}:a(x_{j})=a_{\max}\}$.
+The mapping $\psi$ is linear if and only if $a(x)$ is maximized at a unique value $x\in\text{supp}(X\mid W_{0}=1)$, and nonlinear if multiple values maximize $a(x)$.
+The linearity of this mapping crucially affects the choice of the inference procedure.
+When $\psi$ is linear, the limiting distribution of $\widehat{\overline{P}}$ is Gaussian and common bootstrap procedures, such as the empirical bootstrap, are valid whenever they are valid for $\widehat{\theta}$.
+However, when $a(x)$ is maximized at more than one value, the limiting distribution of $\widehat{\overline{P}}$ is nonlinear in $\mathbb{Z}$ and thus non-Gaussian.
+In this case, it can be shown (see Theorem 3.1 in Fang and Santos (2019)) that standard bootstrap approaches are invalid.
+However, the fact that the estimand $\overline{P}$ can be written as a Hadamard directionally differentiable mapping of $\theta$ implies that alternative bootstrap procedures, such as those proposed by Hong and Li (2018) and Fang and Santos (2019), can be applied to obtain valid inferences on $\overline{P}$.
+We propose a bootstrap procedure that can be applied regardless of the linearity of $\psi$.
+In order to show its validity, we assume that the limiting distribution $\mathbb{Z}$ can be approximated via a bootstrap procedure.
+Assumption C.2 (Bootstrap for first-step estimators).
+[…]
+This assumption is easily satisfied when $\widehat{\textbf{p}}$ are sample frequencies, $(\widehat{\textbf{a}},\widehat{\textbf{w}}_{0})$ are cell-by-cell estimators that are asymptotically linear and Gaussian, and when $\mathbb{Z}^{*}$ is the distribution of these estimators under a standard bootstrap approach.
+For example, for the empirical bootstrap we can let $\mathbb{Z}_{X}^{*}(j)=\sqrt{n}(\widehat{p}_{j}^{*}-\widehat{p}_{j})$ where $\widehat{p}_{j}^{*}=\frac{1}{n}\sum_{i=1}^{N}\mathbbm{1}(X_{i}^{*}=x_{j})$, where $(X_{1}^{*},\ldots,X_{n}^{*})$ are drawn from the empirical distribution of $(X_{1},\ldots,X_{n})$.
+Theorem C.2 (Bootstrap validity).
+Suppose the assumptions of Theorem C.1 hold and that Assumption C.2 holds. Then,
+[…]
+as $n\rightarrow\infty$, where $\widehat{\psi}$ is defined by
+$\widehat{\psi}(\mathbb{Z}^{*})\;$ $=\;\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\widehat{a}_{\max}}\mathbb{Z}^{*}_{\textbf{a}}(j)\;-\;\frac{\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]}{\widehat{a}_{\max}^{2}}\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}\mathbb{Z}^{*}_{\textbf{a}}(j)$
+$+\;\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\widehat{a}_{\max}}\mathbb{Z}^{*}_{\textbf{w}_{0}}(j)$
+$+\;\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{w}_{0}(x_{j})}{\widehat{\mathbb{P}}(W_{0}=1)\widehat{a}_{\max}}\mathbb{Z}^{*}_{X}(j),$ (A.2)
+where $\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, ^$\widehat{\mathbb{P}}(W_{0}=1)=\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})$, $\widehat{a}_{\max}=\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$, and
+$\widehat{\Psi}_{\mathcal{X}^{+}}=\left\{k\in\{1,\ldots,K\}:\widehat{a}(x_{k})\geq\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\xi_{n}\right\},$
+where $\xi_{n}$ is a positive sequence satisfying $\xi_{n}=o(1)$ and $\xi_{n}\sqrt{n}\rightarrow\infty$ as $n\rightarrow\infty$.
+The proof of Theorem C.2 shows that $\widehat{\Psi}_{\mathcal{X}^{+}}$ consistently estimates $\Psi_{\mathcal{X}^{+}}$ and thus satisfies the conditions of Theorem 3.2 in Fang and Santos (2019).
+The bootstrap procedure is valid whether the limiting distribution is Gaussian or not.
+If we assume $a(x)$ is maximized at a single value, standard bootstrap procedures can also be used to approximate the limiting distribution of $\widehat{\overline{P}}$.
+We propose the following bootstrap procedure to compute a one-sided $(1-\alpha)$ confidence interval for $\overline{P}$.
+Algorithm C.1 (One-sided confidence interval for $\overline{P}$).
+We compute the confidence interval in three steps:
+  1. 1.
+    Compute $\widehat{\theta}$ and $\widehat{\overline{P}}$ using the random sample $\{(W_{i},X_{i})\}_{i=1}^{n}$;
+  2. 2.
+    For bootstrap samples $b=1,\ldots,B$, compute $\widehat{\theta}^{\ast,b}=(\widehat{\textbf{a}}^{\ast,b},\widehat{\textbf{w}}_{0}^{\ast,b},\widehat{\textbf{p}}^{\ast,b})$ and $\mathbb{Z}^{\ast,b}=\sqrt{n}(\widehat{\theta}^{\ast,b}-\widehat{\theta})$;
+  3. 3.
+    Compute $\widehat{q}_{\alpha}$, the $\alpha$ quantile of $\widehat{\psi}(\mathbb{Z}^{\ast,b})$, and report the interval $\left[0,\widehat{\overline{P}}-\widehat{q}_{\alpha}/\sqrt{n}\right]$.
+We could also view these inferential problems through the lens of intersection or union bounds. For example, we can write
+$\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\inf_{x\in\text{supp}(X\mid W_{0}=1)}\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot a(x)}=:\inf_{x\in\text{supp}(X\mid W_{0}=1)}\overline{P}(x).$
 Computing a one-sided confidence interval for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ of the kind $[0,\widehat{\overline{P}}^{+}]$ can be cast as doing inference on intersection bounds.
@@
 We leave all details for future work.
-Estimation and inference of $\overline{P}(a,W_{0};\{\tau_{0}\})$ is related to the question of estimation and inference in linear programs with estimated constraints.
-See Fang, Santos, Shaikh, and Torgovitsky (2023) and Cho and Russell (2024) for recent advances on this topic.
-6 Empirical Application
-In this section, we implement the proposed tools in an application to the effects of unilateral divorce laws in the U.S.
-on female suicide, as in Stevenson and Wolfers (2006).
-Between 1969 and 1985, 37 states (including the District of Columbia) reformed their law by enabling each spouse to seek divorce without the other spouse’s consent.
-Stevenson and Wolfers (2006) argue that these “unilateral” or “no-fault” divorce laws reduced female suicide, domestic violence, and spousal homicide.
-The results on female suicide are also replicated by Goodman-Bacon (2021), whose analysis we follow here.
-Our sample consists of 41 states observed over the 1964–1996 period.
-The outcome of interest is the state- and year-specific female suicide rate (per million women), as computed by the National Center for Health Statistics (NCHS).
-The treatment is whether the state allowed unilateral divorce in a given year.
-Following Goodman-Bacon (2021), our sample omits Alaska and Hawaii.
-Additionally, we omit Louisiana, Maryland, North Carolina, Oklahoma, Utah, Vermont, Virginia, and West Virginia.
-These eight states (and Alaska) had unilateral divorce laws preceding 1964 and are therefore always treated within our timeframe.
-{adjustwidth}
--1in-1in A.
-Estimates of the effects of unilateral divorce laws TWFE ATT Callaway and Sant’Anna Wooldridge –0.604 –10.220 –5.530 (2.622) (3.086) (3.650) B.
-Representativeness of the TWFE estimand based on Proposition 2.4 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216 C.
-Representativeness of the TWFE estimand based on Proposition 2.5 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
-Table 1: Representativeness of the TWFE estimand of the effects of unilateral divorce laws
-Panel A of Table 1 reports our baseline estimates of the average effects of unilateral divorce laws on female suicide.
-After we drop the eight always-treated states, the TWFE estimate, –0.604, becomes much smaller in absolute value than the corresponding estimate in Goodman-Bacon (2021), –3.080.
-Unlike that estimate, ours is also statistically insignificant, with $p$-value = 0.819.
-The conclusion changes, however, when we explicitly target the average treatment effect on the treated (ATT), that is, the average effect for the largest subpopulation for which such an effect is identified.
-Using the approach of Wooldridge (2021), we obtain an estimate of –5.530 with a $p$-value of 0.138.
-The approach of Callaway and Sant’Anna (2021) produces an estimate of –10.220 and a $p$-value of 0.001.
-As is clear, the choice of an estimation approach has fundamental implications for our conclusions in this study.
-While the TWFE estimate and both estimates of the ATT are clearly different, this paper focuses on another implication of the nonuniformity of the TWFE weight function.
-We ask: How representative of the underlying population is the TWFE estimand?
-What is the internal validity of this estimand if we are interested in the treated subpopulation?
-Panel B of Table 1 reports our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)=\mathbb{P}(W^{*}=1\mid W_{0}=1)$, based on the representation of the TWFE estimand in de Chaisemartin and D’Haultfœuille (2020), restated in our Proposition 2.4.
-First, because the weights on some of the group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, ^$\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$.
-Second, when we estimate the CATE function and use these estimates in constructing the bounds, we conclude that the TWFE estimand corresponds to the average treatment effect for at most 62.16% of the treated units or 38.73% of the entire population.
-Panel C of Table 1 revisits these questions on the basis of the representation of the TWFE estimand in Proposition 2.5, based on Goodman-Bacon (2021).
-Here, we assume that group-level average treatment effects are constant over time, which eliminates the problem of negative weights.
-Indeed, we now conclude that the TWFE estimand has a causal interpretation uniformly in $\tau_{0}$, even if it is still not particularly representative of the underlying population or the treated subpopulation.
-Our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$ are equal to 14.00% and 22.46%, respectively.
-When we use the estimated CATE function in constructing the bounds, these estimates increase to 48.02% and 77.07%.
-This is obviously much more than our initial estimate of 0, but still substantially less than 1, guaranteed in the case of $\mathbb{P}(W^{*}=1\mid D=1)$ when using the estimation approaches in Callaway and Sant’Anna (2021), Wooldridge (2021), and other recent papers.
-7 Conclusion
-In this paper, we studied the representativeness and internal validity of a class of weighted estimands, which includes the popular OLS, 2SLS, and TWFE estimands in additive linear models.
-We examined the conditions under which such estimands can be written as the average treatment effect for some (possibly latent) subpopulation.
-In our main result, we derived the sharp upper bound on the size of that subpopulation.
-We consider this bound to be a valuable diagnostic for empirical research.
-When a given estimand can be shown to correspond to the average treatment effect for a large subset of the population of interest, we say its internal validity is high.
-In an application to the effects of unilateral divorce laws in the U.S.
-on female suicide, as in Stevenson and Wolfers (2006) and Goodman-Bacon (2021), we showed that the TWFE estimand has a low degree of internal validity (assuming that the treated subpopulation is of interest), even when we assume away the existence of negative weights.
-Because this result is then necessarily driven by the nonuniformity of the TWFE weight function, it corroborates the negative view of both negative and nonuniform weights in Callaway, Goodman-Bacon, and Sant’Anna (2024).
-Appendix
-This appendix is organized as follows: Appendix A contains proofs for Section 2, Appendix B contains proofs for Section 3, Appendix C contains proofs for Section 4, and Appendix D contains additional results and derivations regarding difference-in-differences and associated weighted estimands.
-Appendix A Proofs for Section 2
+Appendix D Proof of Theorem 4.2
+We begin with a technical lemma that we use in the proof of Theorem 4.2.
+Lemma D.1.
+Suppose Assumption 3.1 holds. Then,
+  1. 1.
+    The functions $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$ and $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]$ are left-continuous.
+  2. 2.
+    The functions $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}>\alpha)\mid W_{0}=1]$ and $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]$ are right-continuous.
+Proof of Lemma D.1.
+The function $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$ is left-continuous if for any strictly increasing sequence $\alpha_{n}\nearrow\alpha$ we have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})\mid W_{0}=1]\rightarrow\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$.
+To see this holds, note that $f_{n}(t)\coloneqq t\mathbbm{1}(t<\alpha_{n})\rightarrow t\mathbbm{1}(t<\alpha)$ since $t\mathbbm{1}(t<\alpha_{n})=0$ for all $t\geq\alpha$, and $t\mathbbm{1}(t<\alpha_{n})=t$ whenever $t<\alpha$ for sufficiently large $n$.
+The random variable $|T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})|$ is dominated by $|T_{\mu}|$ and $\mathbb{E}[|T_{\mu}|\mid W_{0}=1]<\infty$ by Assumption 3.1 and by $\mathbb{P}(W_{0}=1)>0$.
+Therefore, by dominated convergence, $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})\mid W_{0}=1]\rightarrow\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$ hence $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$ is left-continuous.
+The function $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]$ is also left-continuous because $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]=\mathbb{E}[T_{\mu}\mid W_{0}=1]-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$.
+The lemma’s second claim can be similarly shown by considering a sequence $\alpha_{n}\searrow\alpha$.
+∎
+Proof of Theorem 4.2.
+We break down this proof into four cases.
+Case 1: $\mu_{0}\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu_{0}<E_{0}$
+We want to maximize $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ over the subpopulations $W^{*}$ in $\mathcal{W}(a;W_{0},\{\tau_{0}\})$.
+Recall that $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ if $\mu_{0}=\mu(\underline{w}^{*},\tau_{0})$ and $W^{*}\in\text{SP}(W_{0})$ hold, where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$.
+Therefore,
+$\overline{P}(a,W_{0};\{\tau_{0}\})$ $=\max_{W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})}\mathbb{P}(W^{*}=1\mid W_{0}=1)$
+$=\max_{W^{*}\in\{0,1\}:\mu_{0}=\mu(\underline{w}^{*},\tau_{0}),W^{*}\in\text{SP}(W_{0})}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$
+$\leq\max_{W^{*}\in\{0,1\}:\mu_{0}=\mu(\underline{w}^{*},\tau_{0})}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$
+$=\max_{\underline{w}^{*}:\mu_{0}=\mu(\underline{w}^{*},\tau_{0}),\underline{w}^{*}(X)\in[0,1]}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1].$
+We will first show a closed-form expression for an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$.
+Then, we will show that this upper bound can be attained by a corresponding $W^{+}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$, and therefore it equals $\overline{P}(a,W_{0};\{\tau_{0}\})$.
+Before proceeding, let $\alpha^{+}\coloneqq\inf\{\alpha\in\mathbb{R}:R(\alpha)\geq 0\}$ where $R(\alpha)\coloneqq\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]$.
+By construction, $\alpha^{+}\geq 0$.
+By $\mu_{0}<E_{0}$ we also have that $\alpha^{+}<\infty$.
+By Lemma D.1, $R$ is a right-continuous function, and therefore $R(\alpha^{+})=\lim_{\alpha\searrow\alpha^{+}}R(\alpha)\geq 0$.
+We now claim that $\alpha^{+}>0$.
+To show this claim, assume $\alpha^{+}=0$.
+Then, $0\leq R(\alpha^{+})=R(0)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq 0)\mid W_{0}=1]\leq 0$, which implies $\mathbb{P}(\tau_{0}(X)=\mu_{0}\mid W_{0}=1)=1$.
+This is ruled out by the assumption that $\mu_{0}<E_{0}=\mathbb{E}[\tau_{0}(X)\mid W_{0}=1]=\mu_{0}$.
+Therefore, $\alpha^{+}>0$.
+Second, we show an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$.
+For all $\underline{w}^{*}$ such that $\mu_{0}=\mu(\underline{w}^{*},\tau_{0})$ and $\underline{w}^{*}(X)\in[0,1]$, we have that
+$\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$
+$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[\underline{w}^{*}(X)T_{\mu}\mid W_{0}=1]}{\alpha^{+}}$
+$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mid W_{0}=1]}{\alpha^{+}}$
+$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}>\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
+$\leq\frac{\mathbb{E}[1\cdot(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[0\cdot(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}>\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
+$=\mathbb{E}[\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
+$=:P^{+}.$
+The second equality follows from $\mu_{0}=\mu(\underline{w}^{*},\tau_{0})$.
+The inequality follows from $\{0,1\}$ being lower/upper bounds for $\underline{w}^{*}(X)$.
+Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})\leq P^{+}$.
+Third, and finally, we show this inequality is binding by finding $W^{+}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ such that $\mathbb{P}(W^{+}=1\mid W_{0}=1)=P^{+}$.
+We start by defining
+$\underline{w}^{+}(X)$ $=\mathbbm{1}(T_{\mu}<\alpha^{+})+\left(1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\mathbbm{1}(T_{\mu}=\alpha^{+}).$
+This function is bounded above by 1 because $R(\alpha^{+})\geq 0$ and $\alpha^{+}>0$.
+To show $\underline{w}^{+}$ is bounded below by 0, consider cases where $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$ or $R(\alpha^{+})$ equal and differ from 0.
+If $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0$ or $R(\alpha^{+})=0$, then $\underline{w}^{+}(X)\in\{0,1\}$ and it is therefore bounded below by 0.
+If $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)>0$ and $R(\alpha^{+})>0$, then
+$1-\frac{R(\alpha^{+})}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$ $=\frac{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$
+$=\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}=\alpha^{+})\mid W_{0}=1]-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$
+$=\frac{-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}.$
+By the definition of $\alpha^{+}$ as an infimum, we must have that $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, implying that $R(\alpha)$ is discontinuous at $\alpha^{+}$.
+By Lemma D.1, $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]$ is left-continuous and satisfies $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W_{0}=1]\leq R(\alpha)$.
+Therefore, since $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, we must have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+}-\varepsilon)\mid W_{0}=1]<0$ for all $\varepsilon>0$.
+Letting $\varepsilon\searrow 0$ yields that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]\leq 0$.
+Therefore $-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]/(\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1))\geq 0$ and $\underline{w}^{+}(X)\geq 0$.
+Next, we compute
+$\mathbb{E}[\underline{w}^{+}(X)\mid W_{0}=1]$
+$=\mathbb{P}(T_{\mu}<\alpha^{+}\mid W_{0}=1)+\left(1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$
+$=\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)$
+$=\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
+$=P^{+}.$
+The indicator function disappears in the third equality because $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0$ implies $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]=0$ as shown above.
+We next verify $\mu(\underline{w}^{*},\tau_{0})=\mu_{0}$. This condition is equivalent to $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]=0$, which we verify here:
+$\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ $=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]$
+$\quad-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$
+$=R(\alpha^{+})-R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)$
+$=R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0).$
+Therefore, $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ equals 0 when $R(\alpha^{+})=0$.
+When $R(\alpha^{+})>0$, we have that $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)>0$ as shown earlier.
+So $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ is also equal to 0 in this case.
+We conclude by showing that $\underline{w}^{+}(X)$ corresponds to $\mathbb{P}(W^{+}=1\mid X,W_{0}=1)$ for some $W^{+}\in\text{SP}(W_{0})$.
+Let $U\sim\text{Unif}(0,1)$ satisfy $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$ and define
+$W^{+}$ $=\left(\mathbbm{1}(T_{\mu}<\alpha^{+})+\mathbbm{1}\left(T_{\mu}=\alpha^{+},U\leq 1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\right)\cdot W_{0}.$
+[…]
+Case 2: $\mu_{0}\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu_{0}>E_{0}$
+As in Case 1, $\overline{P}(a,W_{0};\{\tau_{0}\})\leq\max_{\underline{w}^{*}:\mu(\underline{w}^{*},\tau_{0})=\mu_{0},\underline{w}^{*}(X)\in[0,1]}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$.
+Let $\alpha^{-}=\sup\{\alpha\in\mathbb{R}:L(\alpha)\leq 0\}$ where $L(\alpha)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]$.
+By construction, $\alpha^{-}\leq 0$ and by $\mu_{0}>E_{0}$ we have that $\alpha^{-}>-\infty$.
+By Lemma D.1, $L$ is a left-continuous function, and therefore $L(\alpha^{-})=\lim_{\alpha\nearrow\alpha^{-}}L(\alpha)\leq 0$.
+Similarly to Case 1, we can show that $\alpha^{-}<0$.
+We now show an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$. For all $\underline{w}^{*}$ such that $\mu(\underline{w}^{*},\tau_{0})=\mu_{0}$ and $\underline{w}^{*}(X)\in[0,1]$, we have that
+$\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$
+$\leq\frac{\mathbb{E}[1\cdot(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[0\cdot(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}<\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}$
+$=\mathbb{E}[\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}$
+$=:P^{-},$
+which follows a similar argument as above.
+This implies $\overline{P}(a,W_{0};\{\tau_{0}\})\leq P^{-}$.
+We now show that this inequality is an equality by finding $W^{-}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ such that $\mathbb{P}(W^{-}=1\mid W_{0}=1)=P^{-}$.
+Let
+$\underline{w}^{-}(X)$ $=\mathbbm{1}(T_{\mu}>\alpha^{-})+\left(1-\frac{L(\alpha^{-})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{-}\mid W_{0}=1)\neq 0)}{\alpha^{-}\mathbb{P}(T_{\mu}=\alpha^{-}\mid W_{0}=1)}\right)\mathbbm{1}(T_{\mu}=\alpha^{-}).$
+The rest of the proof symmetrically follows the one for the previous case.
+Case 3: $\mu_{0}=E_{0}\in\mathcal{S}(\tau_{0};W_{0})$
+Since $W^{*}=W_{0}\in\text{SP}(W_{0})$, we have that $\mu_{0}=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ and thus $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is maximized at 1.
+Case 4: $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$
+By Theorem 3.2, there does not exist a regular subpopulation $W^{*}$ satisfying $\mu_{0}=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ and therefore the supremum equals 0 by its definition. ∎
+Appendix E Proofs for Section 5
 Proof of Proposition 5.3.
 We begin by noting that
-$\beta_{\text{TWFE}}$ $=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}]}$
-$=\frac{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}\mid P=t]\mathbb{P}(P=t)}{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}\mid P=t]\mathbb{P}(P=t)}$
-$=\frac{\mathbb{E}[\ddot{D}Y]}{\mathbb{E}[\ddot{D}^{2}]},$
+$\beta_{\text{TWFE}}$ $=\frac{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}]}{\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}]}=\frac{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}Y_{t}\mid P=t]\mathbb{P}(P=t)}{\sum_{t=1}^{T}\mathbb{E}[\ddot{D}_{t}^{2}\mid P=t]\mathbb{P}(P=t)}=\frac{\mathbb{E}[\ddot{D}Y]}{\mathbb{E}[\ddot{D}^{2}]},$
 where the second equality follows from the uniform distribution of $P$ which is independent from $(\ddot{D}_{t},Y_{t})$ for all $t\in\{1,\ldots,T\}$.
-The third equality follows from defining $\ddot{D}\equiv\ddot{D}_{P}$.
+The third equality follows from defining $\ddot{D}\coloneqq\ddot{D}_{P}$.
 We also note that
-$\ddot{D}_{P}$ $=D_{P}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\sum_{t=1}^{T}\mathbb{E}[D_{t}]\mathbbm{1}(P=t)+\sum_{s=1}^{T}\mathbb{E}[D_{s}]\mathbb{E}[\mathbbm{1}(P=s)]$
+$\ddot{D}$ $=D_{P}-\frac{1}{T}\sum_{s=1}^{T}D_{s}-\sum_{t=1}^{T}\mathbb{E}[D_{t}]\mathbbm{1}(P=t)+\sum_{s=1}^{T}\mathbb{E}[D_{s}]\mathbb{E}[\mathbbm{1}(P=s)]$
 $=D-\frac{1}{T}\sum_{s=1}^{T}\mathbbm{1}(G\leq s)-\mathbb{E}[D\mid P]+\mathbb{E}[D]$
@@
 We have that
-$\mathbb{E}[\ddot{D}Y]=\mathbb{E}[\ddot{D}(Y(0)+D(Y(1)-Y(0)))]=\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]+\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid P,G]]$
+$\mathbb{E}[\ddot{D}Y]$ $=\mathbb{E}[\ddot{D}(Y(0)+D(Y(1)-Y(0)))]$
+$=\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]+\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]].$
 The first equality follows from $Y=Y(0)+D(Y(1)-Y(0))$ and the second from iterated expectations and $\mathbb{E}[D\mid G,P]=D$.
@@
 We have that
-$\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]$ $=\mathbb{E}[\ddot{D}\theta(G,P)]=\mathbb{E}[\ddot{D}\ddot{\theta}(G,P)]$
+$\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]$ $=\mathbb{E}[\ddot{D}\theta(G,P)]=\mathbb{E}[\ddot{D}\ddot{\theta}(G,P)],$
 where $\theta(G,P)=\mathbb{E}[Y(0)\mid G,P]$. The second equality follows by properties of projections and from defining $\ddot{\theta}(G,P)$ as follows:
-$\ddot{\theta}(G,P)$ $\equiv\theta(G,P)-\mathbb{E}[\theta(G,P)\mid G]-\mathbb{E}[\theta(G,P)\mid P]+\mathbb{E}[\theta(G,P)]$
+$\ddot{\theta}(G,P)$ $\coloneqq\theta(G,P)-\mathbb{E}[\theta(G,P)\mid G]-\mathbb{E}[\theta(G,P)\mid P]+\mathbb{E}[\theta(G,P)]$
 $=\mathbb{E}[Y(0)\mid G,P]-\mathbb{E}[Y(0)\mid G]-\mathbb{E}[Y(0)\mid P]+\mathbb{E}[Y(0)].$
 Then, we note that
-$\theta(g^{\prime},t^{\prime})$ $=\mathbb{E}[Y(0)\mid G=g^{\prime},P=t^{\prime}]-\mathbb{E}[Y(0)\mid G=g^{\prime}]-\mathbb{E}[Y(0)\mid P=t^{\prime}]+\mathbb{E}[Y(0)]$
-$=\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g^{\prime}]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
-$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g).$
+$\ddot{\theta}(g^{\prime},t^{\prime})$ $=\mathbb{E}[Y(0)\mid G=g^{\prime},P=t^{\prime}]-\mathbb{E}[Y(0)\mid G=g^{\prime}]-\mathbb{E}[Y(0)\mid P=t^{\prime}]+\mathbb{E}[Y(0)]$
+$=\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g^{\prime}]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g^{\prime}]$
+$-\sum_{g\in\mathcal{G}}\left(\mathbb{E}[Y_{t^{\prime}}(0)\mid G=g]-\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
+$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G]\right].$
 Assumption 5.3.2 implies that for any pair $t,t^{\prime}\in\{1,\ldots,T\}$ and any $g^{\prime}\in\mathcal{G}$
@@
 Therefore,
-$\theta(g^{\prime},t^{\prime})$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g]\right)\mathbb{P}(G=g)$
-$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]-\sum_{g\in\mathcal{G}}\left(\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]\right)\mathbb{P}(G=g)$
+$\theta(g^{\prime},t^{\prime})$ $=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G=g^{\prime}]-\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)\mid G]\right]$
+$=\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]-\mathbb{E}\left[\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}[Y_{t^{\prime}}(0)-Y_{t}(0)]\right]$
 $=0$
@@
 We can write
-$\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$ $=\mathbb{E}[(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
-$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
-$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[D(Y(1)-Y(0))\mid G,P]]$
+$\ddot{D}D$ $=(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D$
+$=(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)$ (A.1)
+by $D^{2}=D=\mathbb{E}[D\mid G,P]$. Thus,
+$\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]]$
 $=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\mathbb{P}(D=1\mid G,P)].$
-The first equality is by definition, the second by $D^{2}=D$, the third by $\mathbb{E}[D\mid G,P]=D$, and the fourth by law of total probability.
 Step 4: Denominator
 In this step, we show that
-$\mathbb{E}[\ddot{D}^{2}]$ $=\mathbb{E}[(D-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]$
-$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]$
-$=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)].$
-The first line is obtained from properties of linear projections, the second from $D^{2}=D$, and the third from $D=\mathbb{P}(D=1\mid G,P)$.
-We conclude the proof by noting the equivalence of integrating over the distribution of $P$ and averages over time periods, which shows the equivalence between
-$\beta_{\text{TWFE}}=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\mathbb{P}(D=1\mid G,P)]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)]}$
-and the expression in Proposition 2.4. ∎
+$\mathbb{E}[\ddot{D}^{2}]$ $=\mathbb{E}[\ddot{D}D]=\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\mathbb{P}(D=1\mid G,P)].$
+The first equality is obtained from properties of linear projections and the second follows from equation (A.1).
+We conclude the proof by noting the equivalence of integrating over the distribution of $P$ and averages over time periods, which shows the equivalence between $\beta_{\text{TWFE}}$, the expression in Proposition 5.3, and
+$\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]\cdot\mathbb{P}(D=1\mid G,P)]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot\mathbb{P}(D=1\mid G,P)]}.$
 Proof of Proposition 5.4.
-Proposition 2.4 and $\mathbb{P}(D=1\mid G,P)=D$ yielded
-$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D]}.$
+Proposition 5.3 and $\mathbb{P}(D=1\mid G,P)=D$ yields
+$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D\cdot\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]]}{\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D]}.$
 Since $\mathbb{E}[Y(1)-Y(0)\mid G,P,D=1]=\mathbb{E}[Y(1)-Y(0)\mid G,D=1]$ by assumption, we can use the law of iterated expectations to obtain
-$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G]\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G]]}.$
-We now calculate the conditional expectation $\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G=g]$ for $g\in\mathcal{G}$.
-If $g=+\infty$, then this conditional expectation is 0 by construction, so we let $g\in\{2,\ldots,T\}$ in what follows:
+$\beta_{\text{TWFE}}$ $=\frac{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D\mid G]\cdot\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D\mid G]]}.$
+We now calculate the conditional expectation $\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])\cdot D\mid G=g]$ for $g\in\mathcal{G}$.
+If $g=+\infty$, then this conditional expectation is 0 by construction, so we focus on the case where $g\in\{2,\ldots,T\}$.
+For these derivations, we let $F_{G}(p)\coloneqq\mathbb{P}(G\leq p)$ denote the cdf of $G$ at $p$.
+We have that:
 $\mathbb{E}[(1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D])D\mid G=g]$
 $=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[D\mid G=g]-\frac{\mathbb{E}[\mathbb{E}[D\mid P]D\mid G=g]}{\mathbb{E}[D\mid G=g]}+\mathbb{E}[D]\right)$
-$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(G\leq P)\mid G=g]}{\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]}+\mathbb{E}[\mathbb{E}[\mathbbm{1}(G\leq P)\mid P]]\right)$
+$=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(G\leq P)\mid G=g]}{\mathbb{E}[\mathbbm{1}(G\leq P)\mid G=g]}\right.$
+$\quad\left.+\;\mathbb{E}[\mathbb{E}[\mathbbm{1}(G\leq P)\mid P]]\right)$
 $=\mathbb{E}[D\mid G=g]\left(1-\mathbb{E}[\mathbbm{1}(g\leq P)]-\frac{\mathbb{E}[F_{G}(P)\mathbbm{1}(g\leq P)]}{\mathbb{E}[\mathbbm{1}(g\leq P)]}+\mathbb{E}[F_{G}(P)]\right)$
@@
 $=\mathbb{E}[D\mid G=g](1-\mathbb{E}[D\mid G=g])(1+\mathbb{E}[D\mid g>P]-\mathbb{E}[D\mid g\leq P])$
-$=\mathbb{P}(D=1\mid G=g)\mathbb{P}(D=0\mid G=g)(\mathbb{P}(D=1\mid P<g)+\mathbb{P}(D=0\mid P\geq g)).$
-The first equality follows from $\mathbb{E}[D\mid G=g]>0$ for $g\in\{2,\ldots,T\}$.
-The second follows from $D=\mathbbm{1}(G\leq P)$ and the law of iterated expectations, the third from $G\perp\!\!\!\perp P$, the fourth from definitions of conditional expectations and the law of iterated expectations, the fifth from combining terms, the seventh from the law of iterated expectations again, and the last line is obtained by the fact that $D\in\{0,1\}$.
+$=\mathbb{P}(D=1\mid G=g)\cdot\mathbb{P}(D=0\mid G=g)\cdot(\mathbb{P}(D=1\mid P<g)+\mathbb{P}(D=0\mid P\geq g)).$
+The first equality follows from $\mathbb{E}[D\mid G=g]>0$ for $g\in\{2,\ldots,T\}$, the second from $D=\mathbbm{1}(G\leq P)$ and the law of iterated expectations, the third from $G\perp\!\!\!\perp P$, the fourth from the law of iterated expectations, the fifth from combining terms, the sixth from the law of iterated expectations again, and the last line is obtained by the fact that $D\in\{0,1\}$.
 The representation in Proposition 5.4 follows.
 ∎
-Appendix B Proofs for Section 3
-Proof of Proposition 3.1.
-We begin by showing the first claim of the proposition. The equation
-$\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]$ $=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ (B.1)
-holds since $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, which holds by Definition 3.1.
-Since $\mathbb{P}(W^{*}=1\mid W_{0}=1,X)$ is assumed positive, we can divide both sides of equation (B.1) by it and obtain
-$\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]$ $=\frac{\mathbb{E}[(Y(1)-Y(0))W^{*}\mid W_{0}=1,X]}{\mathbb{P}(W^{*}=1\mid W_{0}=1,X)}$
-$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1,X]$
-$=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,X],$
-where the second equality holds by definition, and the third holds from $W^{*}$ being a subpopulation of $W_{0}$.
-We now show the second claim of the proposition. Equation (3.4) can be obtained as follows:
-$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1,W_{0}=1]$
-$=\frac{\mathbb{E}[W^{*}(Y(1)-Y(0))\mid W_{0}=1]}{\mathbb{E}[W^{*}\mid W_{0}=1]}$
-$=\frac{\mathbb{E}[\mathbb{E}[W^{*}(Y(1)-Y(0))\mid X,W_{0}=1]\mid W_{0}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\mid W_{0}=1]}$
-$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid W_{0}=1,X)\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W_{0}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid W_{0}=1,X)\mid W_{0}=1]}$
-$=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
-$=\mu(\underline{w}^{*},\tau_{0}).$
-The first equality follows from $W^{*}$ being a subpopulation of $W_{0}$, the second from the definition of conditional expectation and $\mathbb{P}(W^{*}=1)>0$, the third from the law of iterated expectations and $W^{*}=1$ implying $W_{0}=1$, the fourth from $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$, and the fifth and sixth follow immediately.
+Appendix F Proofs for Appendix B
+Proof of Proposition B.1.
+By Theorem 3.1, $\mu(a,\tau_{0})$ has a uniform causal representation in $\mathcal{T}_{\text{all}}$ if and only if $a(x_{k})\geq 0$ for $k\in\{1,\ldots,K\}$.
+Therefore, it is sufficient to show the equivalence between weakly causal estimands and estimands with nonnegative weights.
+A similar result was shown in Proposition 4 of BBMT, but we nevertheless provide a proof here to account for the slight differences in notation.
+Suppose $\mu(a,\tau_{0})$ is weakly causal.
+Let $\nu_{1}=(\mathbbm{1}(a(x_{1})<0),\ldots,\mathbbm{1}(a(x_{K})<0))$ and $\nu_{0}=\textbf{0}_{K}$.
+Trivially, $(\nu_{1},\nu_{0})\in\mathcal{M}_{\text{all}}$ and $\tau^{-}\coloneqq\nu_{1}-\nu_{0}\in\mathcal{T}_{\text{all}}$, where $\tau^{-}\geq\textbf{0}_{K}$.
+Since $\mu(a,\tau_{0})$ is weakly causal, $\mu(a,\tau^{-})\geq 0$ where
+$\mu(a,\tau^{-})$ $=\frac{\mathbb{E}[a(X)\tau^{-}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
+$=\frac{1}{\mathbb{E}[a(X)\mid W_{0}=1]}\sum_{k=1}^{K}a(x_{k})\mathbbm{1}(a(x_{k})<0)\mathbb{P}(X=x_{k}\mid W_{0}=1)\geq 0.$
+This implies $a(x_{k})\geq 0$ for all $k\in\{1,\ldots,K\}$. Thus, $\mu(a,\tau_{0})$ has a uniform causal representation in $\mathcal{T}_{\text{all}}$.
+Now suppose $\mu(a,\tau_{0})$ has a uniform causal representation in $\mathcal{T}_{\text{all}}$, or that $a(x_{k})\geq 0$ for $k\in\{1,\ldots,K\}$.
+Then, for any $(\nu_{0},\nu_{1})\in\mathcal{M}_{\text{all}}$ such that $\tau\coloneqq\nu_{1}-\nu_{0}\geq\textbf{0}_{K}$,
+$\mu(a,\tau)$ $=\frac{1}{\mathbb{E}[a(X)\mid W_{0}=1]}\sum_{k=1}^{K}a(x_{k})\tau(x_{k})\mathbb{P}(X=x_{k}\mid W_{0}=1)\geq 0.$
+The inequality holds because $a(x_{k})$ and $\tau(x_{k})$ are nonnegative for all $k\in\{1,\ldots,K\}$.
+This last inequality is reversed if we instead assume that $\tau\leq\textbf{0}_{K}$.
+Thus, $\mu(a,\tau_{0})$ is weakly causal.
 ∎
-Proof of Lemma 3.1.
-We verify that $W^{*}$ satisfies the two conditions in Definition 3.1.
-Condition 1 holds since $\mathbb{P}(W^{*}=1\mid W_{0}=0)=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid W_{0}=1)\leq\mathbb{P}(W^{\prime}=1\mid W_{0}=1)=0$, The first equality follows from $W^{*}\leq W^{\prime}$ almost surely, and the last one from $W^{\prime}\in\text{SP}(W_{0})$.
-To verify condition 2, note that
-$\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W_{0}=1)$ $=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
-$=\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W^{\prime}=1,W_{0}=1)\mathbb{P}(W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
-$=\mathbb{P}(W^{*}=1\mid Y(1),Y(0),X,W^{\prime}=1)\mathbb{P}(W^{\prime}=1\mid Y(1),Y(0),X,W_{0}=1)$
-$=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$
-$=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1,W_{0}=1)\mathbb{P}(W^{\prime}=1\mid X,W_{0}=1)$
-$=\mathbb{P}(W^{*}=1,W^{\prime}=1\mid X,W_{0}=1)$
-$=\mathbb{P}(W^{*}=1\mid X,W_{0}=1).$
-The first and seventh line follow from $W^{*}\leq W^{\prime}$ almost surely.
-The second and sixth line follow from factoring conditional probabilities.
-The third and fifth line follow from $W^{\prime}\leq W_{0}$ almost surely.
-The fourth line follows from $W^{\prime}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$ and $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W^{\prime}=1$.
+Appendix G Proofs for Appendix C
+We use the following lemma in the proof of Theorem C.1.
+Lemma G.1.
+Let $\theta=(\theta(1),\ldots,\theta(K))\in\mathbb{R}^{K}$ and define the mapping $\phi:\mathbb{R}^{K}\to\mathbb{R}$ by $\phi(\theta)=\max_{j\in\{1,\ldots,K\}}\theta(j)$.
+Then, $\phi$ is Hadamard directionally differentiable for all $\theta\in\mathbb{R}^{K}$ tangentially to $\mathbb{R}^{K}$ with directional derivative at $\theta$ in direction $h\in\mathbb{R}^{K}$
+$\phi_{\theta}^{\prime}(h)$ $=\max_{j\in\operatorname*{arg\,max}_{k\in\{1,\ldots,K\}}\theta(k)}h(j).$
+Proof of Lemma G.1.
+Let $h_{n}\rightarrow h\in\mathbb{R}^{K}$ and $t_{n}\searrow 0$ as $n\rightarrow\infty$. Then,
+$t_{n}^{-1}(\phi(\theta+t_{n}h_{n})-\phi(\theta))$ $=t_{n}^{-1}\left(\max_{k\in\{1,\ldots,K\}}(\theta(k)+t_{n}h_{n}(k))-\max_{k\in\{1,\ldots,K\}}\theta(k)\right).$
+Let $\Theta_{\max}=\{j\in\{1,\ldots,K\}:\theta(j)=\max_{k\in\{1,\ldots,K\}}\theta(k)\}$ and let $j_{\max}$ be an element of $\Theta_{\max}$.
+Then, $\max_{k\in\{1,\ldots,K\}}\theta(k)=\theta(j_{\max})$ and thus
+$\frac{\phi(\theta+t_{n}h_{n})-\phi(\theta)}{t_{n}}$ $=\max\left\{\frac{\theta(1)-\theta(j_{\max})}{t_{n}}+h_{n}(1),\ldots,\frac{\theta(K)-\theta(j_{\max})}{t_{n}}+h_{n}(K)\right\}.$
+For each $j\in\Theta_{\max}$, $(\theta(j)-\theta(j_{\max}))/t_{n}+h_{n}(j)=h_{n}(j)\rightarrow h(j)$.
+For each $j\notin\Theta_{\max}$, $(\theta(j)-\theta(j_{\max}))/t_{n}\rightarrow-\infty$ since $\theta(j)-\theta(j_{\max})<0$ and $t_{n}\searrow 0$.
+Therefore, by continuity of the maximum operator in its arguments, $t_{n}^{-1}(\phi(\theta+t_{n}h_{n})-\phi(\theta))\rightarrow\max_{j\in\Theta_{\max}}h(j)$.
 ∎
-Proof of Theorem 3.1.
-This theorem follows as a special case of the first part of Theorem 4.3 when $W^{\prime}=W_{0}$. This is because $W_{0}$ trivially regular subpopulation of $W_{0}$, and because the condition
-$\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$
-is equivalent to $\sup(\text{supp}(a(X)\mid W_{0}=1))<\infty$. They are equivalent since $\underline{w}^{\prime}(X)=\mathbb{P}(W_{0}=1\mid X,W_{0}=1)=1$ almost surely. ∎
-Proof of Theorem 3.2.
-This is a special case of the second part of Theorem 4.3 where we set $W^{\prime}=W_{0}$. ∎
-Proof of Proposition 3.2.
-First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$.
-Therefore, by Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{lin}}$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$.
-Therefore,
-00 0 $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
-$=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
-$=\frac{\mathbb{E}[a(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
-$=d^{\prime}\left(\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}\right)$
-for all $d\in\mathbb{R}^{d_{X}}$, which implies that $\frac{\mathbb{E}[a(X)X\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}=\frac{\mathbb{E}[\underline{w}^{*}(X)X\mid W_{0}=1]}{\mathbb{E}[w^{*}(X)\mid W_{0}=1]}$.
-Letting $u(x)=\underline{w}^{*}(x)/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$, we have that
+Proof of Theorem C.1.
+We begin by showing the consistency of $\widehat{\overline{P}}$ for $\overline{P}$.
+Part 1: Consistency
+The estimator $\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})$ is consistent for $\mathbb{E}[a(X)w_{0}(X)]$ since its components are assumed consistent by Assumption C.1, and by the continuous mapping theorem.
+The consistency of $\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})$ for $\mathbb{E}[w_{0}(X)]$ is similarly established.
+We now consider the maximum term in the denominator. We can write
+$\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$ $=\max_{x:\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x)>0,\widehat{w}_{0}(x)>c_{n}}\widehat{a}(x).$
+Let $\mathcal{X}^{+}\coloneqq\text{supp}(X\mid W_{0}=1)=\{x\in\text{supp}(X):w_{0}(x)>0\}$ and let $\widehat{\mathcal{X}}^{+}=\{x:\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x)>0,\widehat{w}_{0}(x)>c_{n}\}$.
+We first show that $\mathbb{P}(\widehat{\mathcal{X}}^{+}=\mathcal{X}^{+})\rightarrow 1$ as $n\rightarrow\infty$.
+To see this, first consider $x_{j}\in\mathcal{X}^{+}$.
+Then,
+$\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})$ =($=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right\}\cap\left\{\widehat{w}_{0}(x_{j})>c_{n}\right\}\right)$
+$\geq\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right)+\mathbb{P}\left(\widehat{w}_{0}(x_{j})>c_{n}\right)-1,$
+following an application of Bonferroni’s inequality.
 […]
+Now let $x_{j}\notin\mathcal{X}^{+}$. Then
+$\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})$ =($=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})=0\right\}\cup\left\{\widehat{w}_{0}(x_{j})\leq c_{n}\right\}\right)$
+$\geq\mathbb{P}(\widehat{w}_{0}(x_{j})\leq c_{n})=\mathbb{P}(\sqrt{n}\widehat{w}_{0}(x_{j})\leq\sqrt{n}c_{n}).$
 […]
+$\mathbb{P}(\widehat{\mathcal{X}}^{+}=\mathcal{X}^{+})$ $=\mathbb{P}\left(\bigcap_{x_{j}\in\mathcal{X}^{+}}\{x_{j}\in\widehat{\mathcal{X}}^{+}\}\cap\bigcap_{x_{j}\notin\mathcal{X}^{+}}\{x_{j}\notin\widehat{\mathcal{X}}^{+}\}\right)$
+$=1-\mathbb{P}\left(\bigcup_{x_{j}\in\mathcal{X}^{+}}\{x_{j}\notin\widehat{\mathcal{X}}^{+}\}\cup\bigcup_{x_{j}\notin\mathcal{X}^{+}}\{x_{j}\in\widehat{\mathcal{X}}^{+}\}\right)$
+$\geq 1-\left(\sum_{j:x_{j}\in\mathcal{X}^{+}}\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})+\sum_{j:x_{j}\notin\mathcal{X}^{+}}\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})\right)$
+$\rightarrow 1-0=1.$
+Thus, we obtain $\mathbb{P}\left(\max_{x\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x)=\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\right)\geq\mathbb{P}(\widehat{\mathcal{X}}^{+}\in\mathcal{X}^{+})\rightarrow 1$, which yields
+$\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$ $=\max_{x\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x)=\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)+o_{p}(1).$
 […]
-$W^{*}=\mathbbm{1}\left(U\leq\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\right)\cdot W_{0}$
-where $U\perp\!\!\!\perp(X,Y(1),Y(0),W_{0})$.
-Then $W^{*}$ is a regular subpopulation of $W_{0}$ and $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)=\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}$ since $\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\in[0,1]$ with probability 1 given $W_{0}=1$.
-Therefore, for all $\tau_{0}(x)=c+d^{\prime}x\in\mathcal{T}_{\text{lin}}$ we have that
-$\mu(\underline{w}^{*},\tau_{0})$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
-$=\frac{\mathbb{E}\left[\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\tau_{0}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{u(X)}{\sup(\text{supp}(u(X)\mid W_{0}=1))}\mid W_{0}=1\right]}$
-$=\frac{\mathbb{E}\left[u(X)\tau_{0}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
-$=\frac{\mathbb{E}\left[u(X)(c+d^{\prime}X)\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
-$=c+d^{\prime}\frac{\mathbb{E}\left[u(X)X\mid W_{0}=1\right]}{\mathbb{E}\left[u(X)\mid W_{0}=1\right]}$
-$=c+d^{\prime}\frac{\mathbb{E}\left[a(X)X\mid W_{0}=1\right]}{\mathbb{E}\left[a(X)\mid W_{0}=1\right]}$
-$=\frac{\mathbb{E}\left[a(X)(c+d^{\prime}X)\mid W_{0}=1\right]}{\mathbb{E}\left[a(X)\mid W_{0}=1\right]}$
-$=\mu(a,\tau_{0}).$
-Therefore, by Proposition 3.1 we have that $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$. ∎
-Proof of Proposition 3.3.
-We consider the $K>0$ case first and the $K=0$ case second.
-Case 1: $K>0$.
-First, let $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. We will show that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))=\emptyset$ by way of contradiction.
-Suppose there is a $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ and let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\in[0,1]$.
-Let $\tau^{*}(x)=K\cdot\mathbbm{1}(a(x)<0)$.
-This definition implies $\tau^{*}\in\mathcal{T}_{\text{BV}}(K)$.
-Since $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$ we must have that $\mu(a,\tau_{0})=\mu(\underline{w}^{*},\tau_{0})$ for all $\tau_{0}\in\mathcal{T}_{\text{BV}}(K)$.
-Since $\tau^{*}\in\mathcal{T}_{\text{BV}}(K)$,
-00 0 $=\mu(a,\tau^{*})-\mu(\underline{w}^{*},\tau^{*})$
-$=\mathbb{E}\left[\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\tau^{*}(X)\mid W_{0}=1\right]$
-$=K\cdot\mathbb{E}\left[\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\mathbbm{1}(a(X)<0)\mid W_{0}=1\right].$ (B.2)
-The right-hand side of (B.2) is $K$ times the expectation of a nonpositive function.
-This follows from $a(X)\mathbbm{1}(a(X)<0)\leq 0$, $\underline{w}^{*}(X)\mathbbm{1}(a(X)<0)\geq 0$ by $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\geq 0$, $\mathbb{E}[a(X)\mid W_{0}=1]>0$ by Assumption 3.1, and $\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]=\mathbb{P}(W^{*}=1\mid W_{0}=1)>0$ by $W^{*}\in\text{SP}(W_{0})$.
-Since $K>0$, equation (B.2) implies the nonpositive function must equal 0 with probability 1 given $W_{0}=1$:
-11 11 $=\mathbb{P}\left(\left(\frac{a(X)}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{E}[\underline{w}^{*}(X)]}\right)\mathbbm{1}(a(X)<0)=0\mid W_{0}=1\right)$
-$=\mathbb{P}\left(a(X)=\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)]}\cdot\underline{w}^{*}(X)=0\mid a(X)<0,W_{0}=1\right),$
-where the second equality follows from $\mathbb{P}(a(X)<0\mid W_{0}=1)>0$.
-This implies $a(X)$ equals a positive multiple of $\underline{w}^{*}(X)$, a nonnegative quantity, with probability 1 given $\{a(X)<0,W_{0}=1\}$, an event with positive probability that implies $a(X)$ is strictly negative.
-This is a contradiction and therefore $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))=\emptyset$.
-Second, suppose $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)=1$.
-By Theorem 3.1, $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})\neq\emptyset$.
-Since $\mathcal{T}_{\text{BV}}(K)\subseteq\mathcal{T}_{\text{all}}$, we have that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})\subseteq\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))$.
-Therefore, $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(K))\neq\emptyset$ thus $\mu(a,\tau_{0})$ has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{BV}}(K)$.
-Case 2: $K=0$.
-When $K=0$, the function class $\mathcal{T}_{\text{BV}}(K)$ is the set of all constant functions.
-In this case, $\tau_{0}(X)=t_{0}$, where $t_{0}\in\mathbb{R}$ denotes a constant.
-Thus $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BV}}(0))\neq\emptyset$ for all weight functions $a(\cdot)$ since $W_{0}\in\text{SP}(W_{0})$ and because $\mu(a,\tau_{0})=\mathbb{E}[a(X)t_{0}\mid W_{0}=1]/\mathbb{E}[a(X)\mid W_{0}=1]=t_{0}=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ for all $a(\cdot)$.
-∎
-Appendix C Proofs for Section 4
-Proof of Theorem 4.1.
-This is a special case of the first part of Theorem 4.4 where we set $W^{\prime}=W_{0}$. ∎
-We begin with a technical lemma that we use to prove the subsequent theorems.
-Lemma C.1.
-Let Assumption 3.1 hold. Let $T_{\mu}=\tau_{0}(X)-\mu$. Then, for any $W^{\prime}\in\text{SP}(W_{0})$, we have that
-  1. 1.
-    The functions $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ and $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]$ are left-continuous.
-  2. 2.
-    The functions $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}>\alpha)\mid W^{\prime}=1]$ and $\alpha\mapsto\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W^{\prime}=1]$ are right-continuous.
-Proof of Lemma C.1.
-The function $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ is left-continuous if for any strictly increasing sequence $\alpha_{n}\uparrow\alpha$ we have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})\mid W^{\prime}=1]\rightarrow\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$.
-To see this holds, note that $f_{n}(t)\equiv t\mathbbm{1}(t<\alpha_{n})\rightarrow t\mathbbm{1}(t<\alpha)$ since $t\mathbbm{1}(t<\alpha_{n})=0$ for all $t\geq\alpha$, and $t\mathbbm{1}(t<\alpha_{n})=t$ whenever $t<\alpha$ for sufficiently large $n$.
-The random variable $|T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})|$ is dominated by $|T_{\mu}|$ and $\mathbb{E}[|T_{\mu}|\mid W^{\prime}=1]<\infty$ by Assumption 3.1 and by $\mathbb{P}(W^{\prime}=1)>0$.
-Therefore, by dominated convergence, $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha_{n})\mid W^{\prime}=1]\rightarrow\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ hence $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$ is left-continuous.
-The function $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]$ is also left-continuous because $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W^{\prime}=1]=\mathbb{E}[T_{\mu}\mid W^{\prime}=1]-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)\mid W^{\prime}=1]$.
-The lemma’s second claim can be similarly shown by considering a sequence $\alpha_{n}\downarrow\alpha$. ∎
-Proof of Theorem 4.2.
-To simplify the notation in the proof, let $\mu\equiv\mu(a,\tau_{0})$. We break down this proof into four cases.
-Case 1: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
-We want to maximize $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ over the subpopulations $W^{*}$ in $\mathcal{W}(a;W_{0},\{\tau_{0}\})$.
-Recall that $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ if $\mu=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$ and $W^{*}\in\text{SP}(W_{0})$ hold, where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$.
-Therefore,
-$\overline{P}(a,W_{0};\{\tau_{0}\})$ $=\max_{W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})}\mathbb{P}(W^{*}=1\mid W_{0}=1)$
-$=\max_{W^{*}\in\{0,1\}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],W^{*}\in\text{SP}(W_{0})}\mathbb{P}(W^{*}=1\mid W_{0}=1)$
-$\leq\max_{W^{*}\in\{0,1\}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\mid W_{0}=1]$
-$=\max_{\underline{w}^{*}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],\underline{w}^{*}(X)\in[0,1]}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1].$
-We will first show a closed-form expression for an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$.
-Then, we will show that this upper bound can be attained by a corresponding $W^{+}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$, and therefore it equals $\overline{P}(a,W_{0};\{\tau_{0}\})$.
-Before proceeding, let $\alpha^{+}=\inf\{\alpha\in\mathbb{R}:R(\alpha)\geq 0\}$ where $R(\alpha)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha)\mid W_{0}=1]$.
-By construction, $\alpha^{+}\geq 0$.
-By $\mu<\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ we also have that $\alpha^{+}<+\infty$.
-By Lemma C.1, $R$ is a right-continuous function, and therefore $R(\alpha^{+})=\lim_{\alpha\downarrow\alpha^{+}}R(\alpha)\geq 0$.
-We now claim that $\alpha^{+}>0$.
-To show this claim, assume $\alpha^{+}=0$.
-Then, $0\leq R(\alpha^{+})=R(0)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq 0)\mid W_{0}=1]\leq 0$, which implies $\mathbb{P}(\tau_{0}(X)=\mu\mid W_{0}=1)=1$.
-This is ruled out by the assumption that $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]=\mathbb{E}[\tau_{0}(X)\mid W_{0}=1]$.
-Therefore, $\alpha^{+}>0$.
-We now show an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$.
-For all $\underline{w}^{*}$ such that $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ and $\underline{w}^{*}(X)\in[0,1]$, we have that
-$\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\alpha^{+}\mid W_{0}=1]}{\alpha^{+}}$
-$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[\underline{w}^{*}(X)T_{\mu}\mid W_{0}=1]}{\alpha^{+}}$
-$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mid W_{0}=1]}{\alpha^{+}}$
-$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}>\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
-$\leq\frac{\mathbb{E}[1\cdot(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}+\frac{\mathbb{E}[0\cdot(\alpha^{+}-T_{\mu})\mathbbm{1}(T_{\mu}>\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
-$=\mathbb{E}[\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
-$\equiv P^{+}.$
-The third equality follows from $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$.
-The inequality follows from $\{0,1\}$ being lower/upper bounds for $\underline{w}^{*}(X)$.
-Therefore, $\overline{P}(a,W_{0};\{\tau_{0}\})\leq P^{+}$.
-We now show that this inequality is binding by finding $W^{+}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ such that $\mathbb{P}(W^{+}=1\mid W_{0}=1)=P^{+}$.
-Let
-$\underline{w}^{+}(X)$ =$=\begin{cases}1&\text{ if }T_{\mu}<\alpha^{+}\\ 1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}&\text{ if }T_{\mu}=\alpha^{+}\\ 0&\text{ if }T_{\mu}>\alpha^{+}.\end{cases}$
-This function is bounded above by 1 because $R(\alpha^{+})\geq 0$ and $\alpha^{+}>0$.
-To show it is bounded below by 0, consider cases where $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$ or $R(\alpha^{+})$ equal and differ from 0.
-If $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0$ or $R(\alpha^{+})=0$, then $\underline{w}^{+}(X)\in\{0,1\}\subseteq[0,1]$ and it is therefore bounded below by 0.
-If $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)>0$ and $R(\alpha^{+})>0$, then
-$1-\frac{R(\alpha^{+})}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$ $=\frac{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$
-$=\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}=\alpha^{+})\mid W_{0}=1]-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}$
-$=\frac{-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}.$
-By the definition of $\alpha^{+}$ as an infimum, we must have that $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, implying that $R(\alpha)$ is discontinuous at $\alpha^{+}$.
-By Lemma C.1, $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)]$ is left-continuous and satisfies $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha)]\leq R(\alpha)$.
-Therefore, since $R(\alpha^{+}-\varepsilon)<0$ for all $\varepsilon>0$, we must have that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+}-\varepsilon)]<0$ for all $\varepsilon>0$.
-Letting $\varepsilon\downarrow 0$ yields that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})]\leq 0$.
-Therefore $-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]/(\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1))\geq 0$ and $\underline{w}^{*}(X)\geq 0$ in that case as well.
-Next, we compute
-$\mathbb{E}[\underline{w}^{+}(X)\mid W_{0}=1]$ $=\mathbb{E}[\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]$
-$\qquad+\left(1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\mathbb{E}[\mathbbm{1}(T_{\mu}=\alpha^{+})\mid W_{0}=1]$
-$=\mathbb{E}[\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]+\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
-$=\mathbb{P}(T_{\mu}\leq\alpha^{+}\mid W_{0}=1)-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]}{\alpha^{+}}$
-$=P^{+}.$
-Finally we verify that $\frac{\mathbb{E}[\underline{w}^{+}(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{+}(X)\mid W_{0}=1]}=\mu$.
-This condition is equivalent to $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]=0$, which we verify here:
-$\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ $=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]$
-$\qquad+\left(1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\mathbb{E}[T_{\mu}1(T_{\mu}=\alpha^{+})\mid W_{0}=1]$
-$=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\leq\alpha^{+})\mid W_{0}=1]-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)$
-$=R(\alpha^{+})-R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)$
-$=R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)=0).$
-Therefore, $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ equals 0 when $R(\alpha^{+})=0$.
-When $R(\alpha^{+})>0$, we have that $\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)>0$ as shown earlier.
-So $\mathbb{E}[\underline{w}^{+}(X)T_{\mu}\mid W_{0}=1]$ is also equal to 0 in this case.
-We conclude the proof by showing that $\underline{w}^{+}(X)$ corresponds to $\mathbb{P}(W^{+}=1\mid X,W_{0}=1)$ for some $W^{+}\in\text{SP}(W_{0})$.
-Let $U\sim\text{Unif}(0,1)$ satisfy $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0})$ and define
-$W^{+}$ $=\left(\mathbbm{1}(T_{\mu}<\alpha^{+})+\mathbbm{1}\left(T_{\mu}=\alpha^{+},U\leq 1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\right)\cdot W_{0}.$
-By construction, $W^{+}\in\{0,1\}$, $\mathbb{P}(W^{+}=1\mid X,W_{0}=1)=\underline{w}^{+}(X)$, $\mathbb{P}(W^{+}=1\mid W_{0}=0)=0$, and
-$W^{+}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$.
-Also, since $\mu\in\mathcal{S}(\tau_{0};W_{0})$, $\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu\mid W_{0}=1)>0$.
-Since $\alpha^{+}>0$ we have that $\mathbb{P}(W^{+}=1\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}<\alpha^{+}\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)>0$.
-Therefore $W^{+}$ is a regular subpopulation of $W^{+}$ for which $\mathbb{P}(W^{*}=1\mid W_{0}=1)=P^{+}$, hence $P^{+}$ is the maximum.
-Case 2: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
-As in case 1,
-$\overline{P}(a,W_{0};\{\tau_{0}\})$ $\leq\max_{\underline{w}^{*}:\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1],\underline{w}^{*}(X)\in[0,1]}\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1].$
-Before proceeding, let $\alpha^{-}=\sup\{\alpha\in\mathbb{R}:L(\alpha)\leq 0\}$ where $L(\alpha)=\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha)\mid W_{0}=1]$.
-By construction, $\alpha^{-}\leq 0$ and by $\mu>\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ we have that $\alpha^{-}>-\infty$.
-By Lemma C.1, $L$ is a left-continuous function, and therefore $L(\alpha^{-})=\lim_{\alpha\uparrow\alpha^{-}}L(\alpha)\leq 0$.
-Similarly to case 1, we can show that $\alpha^{-}<0$.
-We now show an upper bound for $\overline{P}(a,W_{0};\{\tau_{0}\})$.
-For all $\underline{w}^{*}$ such that $\mu=\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W_{0}=1]/\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ and $\underline{w}^{*}(X)\in[0,1]$, we have that
-$\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\alpha^{-}\mid W_{0}=1]}{\alpha^{-}}$
-$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[\underline{w}^{*}(X)T_{\mu}\mid W_{0}=1]}{\alpha^{-}}$
-$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mid W_{0}=1]}{\alpha^{-}}$
-$=\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[\underline{w}^{*}(X)(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}<\alpha^{-})\mid W_{0}=1]}{\alpha^{+}}$
-$\leq\frac{\mathbb{E}[1\cdot(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}\geq\alpha^{+})\mid W_{0}=1]}{\alpha^{-}}+\frac{\mathbb{E}[0\cdot(\alpha^{-}-T_{\mu})\mathbbm{1}(T_{\mu}<\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}$
-$=\mathbb{E}[\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]-\frac{\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}\geq\alpha^{-})\mid W_{0}=1]}{\alpha^{-}}$
-$\equiv P^{-},$
-which follows a similar argument as above.
-This implies $\overline{P}(a,W_{0};\{\tau_{0}\})\leq P^{-}$.
-We now show that this inequality is an equality by finding $W^{-}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$ such that $\mathbb{P}(W^{-}=1\mid W_{0}=1)=P^{-}$.
-Let
-$\underline{w}^{-}(X)$ =$=\begin{cases}1&\text{ if }T_{\mu}>\alpha^{+}\\ 1-\frac{L(\alpha^{-})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{-}\mid W_{0}=1)\neq 0)}{\alpha^{-}\mathbb{P}(T_{\mu}=\alpha^{-}\mid W_{0}=1)}&\text{ if }T_{\mu}=\alpha^{-}\\ 0&\text{ if }T_{\mu}<\alpha^{+}.\end{cases}$
-The rest of the proof symmetrically follows the one for the previous case.
-Case 3: $\mu\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$
-Note that $W^{*}=W_{0}$ is the largest regular subpopulation of $W_{0}$.
-Since $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$, we have that $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ is trivially maximized at $1$.
-Case 4: $\mu\notin\mathcal{S}(\tau_{0};W_{0})$
-By Theorem 3.2, there does not exist a regular subpopulation $W^{*}$ satisfying $\mu=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ and therefore the supremum equals 0 by its definition. ∎
-Proof of Theorem 4.3.
-Part 1: $\mathcal{T}=\mathcal{T}_{\text{all}}$
-()($\Longrightarrow$)( ⟹ ) First, suppose there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$.
-Using the law of iterated expectations and letting $\underline{w}^{\prime}(X)=\mathbb{P}(W^{\prime}=1\mid W_{0}=1,X)$, we can write
-$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
-$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\underline{w}^{\prime}(X)\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\underline{w}^{\prime}(X)\mid W_{0}=1\right]}$
-$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)W^{\prime}\mid W_{0}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\mid W_{0}=1\right]}$
-$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1,W_{0}=1\right]\mathbb{P}(W^{\prime}=1\mid W_{0}=1)}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1,W_{0}=1\right]\mathbb{P}(W^{\prime}=1\mid W_{0}=1)}$
-$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
-$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
-$\equiv\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right).$
-The second equality is valid due to $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$.
-The third and fourth follow from the law of iterated expectations, and the fifth from $W^{\prime}$ being a subpopulation of $W_{0}$.
-The second to last line follows from Proposition 3.1 and from $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ implying $\mathbb{P}(\underline{w}^{\prime}(X)>0\mid W_{0}=1)>0$.
-Similarly, we can write $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\mu^{\prime}(\underline{w}^{*},\tau_{0})$ where $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid W^{\prime}=1,X)$.
-Therefore, by Proposition 3.1, we have that $\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right)-\mu^{\prime}(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{all}}$.
-Let $\tau^{*}(X)=\frac{a(X)/\underline{w}^{\prime}(X)}{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}$.
-We have $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ and $\mathbb{E}[\underline{w}^{*}(X)^{2}]\leq 1$ by construction.
-Hence, $\mathbb{E}[\tau^{*}(X)^{2}]<\infty$ and thus $\tau^{*}\in\mathcal{T}_{\text{all}}$.
-Thus, we must have $\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau^{*}\right)-\mu^{\prime}(\underline{w}^{*},\tau^{*})=0$. Expanding this equality yields
-00 0 $=\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau^{*}\right)-\mu^{\prime}(\underline{w}^{*},\tau^{*})$
-$=\mathbb{E}\left[\tau^{*}(X)\left(\frac{a(X)/\underline{w}^{\prime}(X)}{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}-\frac{\underline{w}^{*}(X)}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}\right)\mid W^{\prime}=1\right]$
-$=\mathbb{E}[\tau^{*}(X)^{2}\mid W^{\prime}=1].$
-This implies that $\mathbb{P}(\tau^{*}(X)=0\mid W^{\prime}=1)=1$. In turn, this implies that
-$\mathbb{P}\left(a(X)=\frac{\underline{w}^{\prime}(X)\underline{w}^{*}(X)\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]}{\mathbb{P}(W^{*}=1\mid W^{\prime}=1)}\mid W^{\prime}=1\right)=1.$
-We have that $\underline{w}^{*}(X)\geq 0$ and $\underline{w}^{\prime}(X)\geq 0$ almost surely, and $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)>0$ by the assumption that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$.
-Also, $\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]=\mathbb{E}[a(X)\mid W_{0}=1]>0$ by Assumption 3.1.
-Therefore, $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$.
+Part 2: Asymptotic Distribution
+We first establish the joint limiting distribution of terms (i) $\sqrt{n}(\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})-\mathbb{E}[a(X)w_{0}(X)])$, (ii) $\sqrt{n}(\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})-\mathbb{E}[w_{0}(X_{i})])$, and (iii) $\sqrt{n}(\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\max_{x\in\mathcal{X}^{+}}a(x))$.
+The terms (i) and (ii) can be expanded as follows:
+$\sqrt{n}$ $\left(\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})-\mathbb{E}[a(X)w_{0}(X)]\right)$
+$=\sqrt{n}\left(\sum_{j=1}^{K}\widehat{a}(x_{j})\widehat{w}_{0}(x_{j})\widehat{p}_{j}-\sum_{j=1}^{K}a(x_{j})w_{0}(x_{j})p_{j}\right)$
+$=\sum_{j=1}^{K}w_{0}(x_{j})p_{j}\sqrt{n}(\widehat{a}(x_{j})-a(x_{j}))+\sum_{j=1}^{K}a(x_{j})p_{j}\sqrt{n}(\widehat{w}_{0}(x_{j})-w_{0}(x_{j}))$
+$+\sum_{j=1}^{K}a(x_{j})w_{0}(x_{j})\sqrt{n}(\widehat{p}_{j}-p_{j}))+o_{p}(1)$ (A.1)
+and
+$\sqrt{n}$ $\left(\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})-\mathbb{E}[w_{0}(X)]\right)=\sqrt{n}\left(\sum_{j=1}^{K}\widehat{w}_{0}(x_{j})\widehat{p}_{j}-\sum_{j=1}^{K}w_{0}(x_{j})p_{j}\right)$
+$=\sum_{j=1}^{K}\left(p_{j}\sqrt{n}(\widehat{w}_{0}(x_{j})-w_{0}(x_{j}))+w_{0}(x_{j})\sqrt{n}(\widehat{p}_{j}-p_{j})\right)+o_{p}(1).$ (A.2)
+For term (iii), we use the expansion
+$\sqrt{n}\left(\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\max_{x\in\mathcal{X}^{+}}a(x)\right)$ $=\sqrt{n}\left(\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\right)$ (A.3)
+$+\sqrt{n}\left(\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)-\max_{x\in\mathcal{X}^{+}}a(x)\right).$ (A.4)
+The term in (A.3) is of order $o_{p}(1)$ because
+$\mathbb{P}\left(\sqrt{n}\left(\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\right)=0\right)$ $=\mathbb{P}\left(\max_{x\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x)=\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\right)$
+$\geq\mathbb{P}(\widehat{\mathcal{X}}^{+}\in\mathcal{X}^{+})\rightarrow 1,$
+as shown earlier.
+The term in (A.4) can be analyzed using Theorem 2.1 in Fang and Santos (2019), which generalizes the delta method to the class of Hadamard directionally differentiable functions.
+Using Lemma G.1, we have that
+$\sqrt{n}\left(\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)-\max_{x\in\mathcal{X}^{+}}a(x)\right)$ $=\max_{x_{j}\in\operatorname*{arg\,max}_{x\in\mathcal{X}^{+}}a(x)}\sqrt{n}\left(\widehat{a}(x_{j})-a(x_{j})\right)+o_{p}(1)$
+$=:\max_{j\in\Psi_{\mathcal{X}^{+}}}\sqrt{n}\left(\widehat{a}(x_{j})-a(x_{j})\right)+o_{p}(1).$ (A.5)
+Combining the expressions in (A.1), (A.2), and (A.5) with the delta method yields
+$\sqrt{n}(\widehat{\overline{P}}-\overline{P})$ $=\frac{1}{\mathbb{P}(W_{0}=1)\max_{x\in\mathcal{X}^{+}}a(x)}\sum_{j=1}^{K}\left(w_{0}(x_{j})p_{j}\sqrt{n}(\widehat{a}(x_{j})-a(x_{j}))\right.$
+$\left.+a(x_{j})p_{j}\sqrt{n}(\widehat{w}_{0}(x_{j})-w_{0}(x_{j}))+a(x_{j})w_{0}(x_{j})\sqrt{n}(\widehat{p}_{j}-p_{j})\right)$
+$-\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{P}(W_{0}=1)\max_{x\in\mathcal{X}^{+}}a(x)}\sum_{j=1}^{K}\left(p_{j}\sqrt{n}(\widehat{w}_{0}(x_{j})-w_{0}(x_{j}))+w_{0}(x_{j})\sqrt{n}(\widehat{p}_{j}-p_{j})\right)$
+$-\frac{\mathbb{E}[a(X)\mid W_{0}=1]}{\max_{x\in\mathcal{X}^{+}}a(x)^{2}}\max_{j\in\Psi_{\mathcal{X}^{+}}}\sqrt{n}\left(\widehat{a}(x_{j})-a(x_{j})\right)+o_{p}(1)$
+$=\psi(\sqrt{n}(\widehat{\textbf{a}}-\textbf{a}),\sqrt{n}(\widehat{\textbf{w}}_{0}-\textbf{w}_{0}),\sqrt{n}(\widehat{\textbf{p}}-\textbf{p}))+o_{p}(1)$
 […]
-$\mu(a,\tau_{0})$ $=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mathbb{E}[Y(1)-Y(0)\mid W^{\prime}=1,X]\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
-$\equiv\mu^{\prime}\left(\frac{a}{\underline{w}^{\prime}},\tau_{0}\right).$
-Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W_{0},W^{\prime})$, and define
-$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}.$
-We verify that $W^{*}$ is a regular subpopulation of $W^{\prime}$. First, we see that $\mathbb{P}(W^{*}=1)>0$ because
-$\mathbb{P}(W^{*}=1)$ $=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)+\mathbb{P}(W^{*}=1\mid W^{\prime}=0)\mathbb{P}(W^{\prime}=0)$
-$=\mathbb{P}\left(\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}=1\mid W^{\prime}=1\right)\mathbb{P}(W^{\prime}=1)$
-$=\mathbb{E}\left[\mathbb{P}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid X,W^{\prime}=1\right)\mid W^{\prime}=1\right]\mathbb{P}(W^{\prime}=1)$
-$=\mathbb{E}\left[\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid W^{\prime}=1\right]\mathbb{P}(W^{\prime}=1)$
-$=\frac{\mathbb{E}[a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1]\mathbb{P}(W^{\prime}=1)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}$
-$=\frac{\mathbb{E}[a(X)\mid W_{0}=1]\mathbb{P}(W^{\prime}=1)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}$
-$>0.$
-The second equality follows from $\mathbb{P}(W^{*}=1\mid W^{\prime}=0)=0$, which holds by the construction of $W^{*}$.
-The fourth equality holds from $U\perp\!\!\!\perp(X,W^{\prime})$ and from $\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\in[0,1]$ almost surely given $W^{\prime}=1$.
-To establish the final inequality, recall that $\mathbb{P}(W^{\prime}=1)$ and $\mathbb{E}[a(X)\mid W_{0}=1]$ are positive by assumption.
-Also $\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))\leq\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1))<\infty$ since $\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1)$ is a subset of $\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W_{0}=1)$.
-That $W^{*}$ satisfies the two properties of Definition 3.1 holds immediately.
-Therefore, $W^{*}$ is a regular subpopulation of $W^{\prime}$.
-Finally, let
-$\underline{w}^{*}(X)$ $=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)$
-$=\mathbb{P}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid X,W^{\prime}=1\right)$
-$=\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}.$
-Letting $\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]=\tau_{0}(X)\in\mathcal{T}_{\text{all}}$ and using Proposition 3.1, we can see that
-$\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X)\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[\underline{w}^{\prime}(X)\mid W^{\prime}=1]}$
-$=\frac{\mathbb{E}\left[\frac{\frac{a(X)}{\underline{w}^{\prime}(X)}}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{\frac{a(X)}{\underline{w}^{\prime}(X)}}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\mid W^{\prime}=1\right]}$
-$=\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$
-$=\frac{\mathbb{E}[a(X)\tau_{0}(X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}$
-$=\mu(a,\tau_{0}).$
-Since $\tau_{0}\in\mathcal{T}_{\text{all}}$ was arbitrary, we have that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ which concludes the proof.
-Part 2: $\mathcal{T}=\{\tau_{0}\}$
-To simplify the notation in the proof, we let $\mu\equiv\mu(a,\tau_{0})$.
-()($\Longrightarrow$)( ⟹ ) First, let $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$.
-Since $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$, we can without loss of generality suppose that $\mathbb{P}(\tau_{0}(X)\leq\mu\mid W^{\prime}=1)=0$, which implies $\mathbb{P}(\tau_{0}(X)>\mu\mid W^{\prime}=1)=1$.
-Since $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$, we can write by Proposition 3.1
-$\mu$ $=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
-$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}$
-$=\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}$
-$\geq\frac{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mu\mid W^{\prime}=1]}{\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]}=\mu.$ (C.1)
-The inequality is strict unless $\mathbb{E}[(\underbrace{\tau_{0}(X)-\mu}_{>0\text{ w.p.1}})\underbrace{\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)}_{\in[0,1]}\mid W^{\prime}=1]=0$ holds.
-This holds if $\mathbb{P}((\tau_{0}(X)-\mu)\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)=0\mid W^{\prime}=1)=1$, which in turns occurs if and only if $\mathbb{P}(\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)=0\mid W^{\prime}=1)=1$.
-This implies $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)=0$ and $\mathbb{P}(W^{*}=1)=\mathbb{P}(W^{*}=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)=0$, a contradiction of $W^{*}\in\mathcal{W}(a,W^{\prime},\{\tau_{0}\})$.
-Therefore, the inequality in (C.1) is strict and yields $\mu>\mu$, a contradiction.
-Therefore, $\mathcal{W}(a;W^{\prime},\{\tau_{0}\})=\emptyset$ when $\mu\notin\mathcal{S}(\tau_{0};W^{\prime})$.
+by the continuity of $\psi$ and Assumption C.1. ∎
+Proof of Theorem C.2.
 […]
-$\mathcal{X}^{-}$ =$=\{x\in\text{supp}(X):\tau_{0}(x)\leq\mu\}\qquad\text{ and }\qquad\mathcal{X}^{+}=\{x\in\text{supp}(X):\tau_{0}(x)\geq\mu\}.$
-By $\mu\in\mathcal{S}(\tau_{0};W^{\prime})$, $\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu\mid W^{\prime}=1)>0$.
-Similarly, $\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)>0$.
-Let $U\sim\text{Unif}(0,1)$ where $U\perp\!\!\!\perp(Y(1),Y(0),X,W^{\prime},W_{0})$. For $u\in[0,1]$ let
-$W^{*}(u)$ $=(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+}))\cdot W^{\prime}.$
-We can see that $W^{*}(u)\in\{0,1\}$, that $W^{*}(u)\perp\!\!\!\perp(Y(1),Y(0))\mid X,W^{\prime}=1$, and that $\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=0)=0$.
-To show that $W^{*}(u)$ characterizes a regular subpopulation of $W^{\prime}$, we also show that it is nonzero with positive probability:
-$\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=1)$ $=\mathbb{P}(\mathbbm{1}(U>u,X\in\mathcal{X}^{-})+\mathbbm{1}(U\leq u,X\in\mathcal{X}^{+})=1\mid W^{\prime}=1)$
-$=\mathbb{P}(U>u,X\in\mathcal{X^{-}}\mid W^{\prime}=1)+\mathbb{P}(U\leq u,X\in\mathcal{X^{-}}\mid W^{\prime}=1)$
-$=(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)$
-$>0$
-for all $u\in[0,1]$.
-Therefore, $\mathbb{P}(W^{*}(u)=1)=\mathbb{P}(W^{*}(u)=1\mid W^{\prime}=1)\mathbb{P}(W^{\prime}=1)>0$ by $\mathbb{P}(W^{\prime}=1)>0$.
-Hence, $W^{*}(u)\in\text{SP}(W^{\prime})$ for all $u\in[0,1]$.
-For a given $u$, we have that $\underline{w}^{*}(X;u)\equiv\mathbb{P}(W^{*}(u)=1\mid X,W^{\prime}=1)=(1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+})$.
-Therefore, using Proposition 3.1,
-$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ $=\frac{\mathbb{E}[\underline{w}^{*}(X;u)\mathbb{E}[Y(1)-Y(0)\mid X,W^{\prime}=1]\mid W^{\prime}=1]}{\mathbb{E}[\underline{w}^{*}(X;u)\mid W^{\prime}=1]}$
-$=\frac{\mathbb{E}[((1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+}))\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{E}[((1-u)\mathbbm{1}(X\in\mathcal{X}^{-})+u\mathbbm{1}(X\in\mathcal{X}^{+}))\mid W^{\prime}=1]}$
-$=\frac{(1-u)\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W^{\prime}=1]+u\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W^{\prime}=1]}{(1-u)\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)+u\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}.$
-By construction, $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{-})\leq\mu\mathbbm{1}(X\in\mathcal{X}^{-})$ and $\tau_{0}(X)\mathbbm{1}(X\in\mathcal{X}^{+})\geq\mu\mathbbm{1}(X\in\mathcal{X}^{+})$ almost surely.
-Therefore,
-$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(0)=1]$ $=\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)}\leq\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{-})\mu\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{-}\mid W^{\prime}=1)}=\mu$
-and
-$\mathbb{E}[Y(1)-Y(0)\mid W^{*}(1)=1]$ $=\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\tau_{0}(X)\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}\geq\frac{\mathbb{E}[\mathbbm{1}(X\in\mathcal{X}^{+})\mu\mid W^{\prime}=1]}{\mathbb{P}(X\in\mathcal{X}^{+}\mid W^{\prime}=1)}=\mu.$
-By the continuity of $\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u)=1]$ in $u$ and the intermediate value theorem, there exists $u^{*}\in[0,1]$ such that $\mu=\mathbb{E}[Y(1)-Y(0)\mid W^{*}(u^{*})=1]$ and $W^{*}(u^{*})\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$.
-∎
-Proof of Theorem 4.4.
-Part 1: $\mathcal{T}=\mathcal{T}_{\text{all}}$
-First suppose $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)=1$. From Theorem 4.3, there exists $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$. Written differently, we have that
-$\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]}$ $=\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]=\frac{\mathbb{E}\left[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\tau_{0}(X)\mid W^{\prime}=1\right]}{\mathbb{E}\left[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1\right]}$
-for all $\tau_{0}\in\mathcal{T}_{\text{all}}$.
-From derivations in the proof of Theorem 4.3, we have that $\mathbb{P}(C\cdot\frac{a(X)}{\underline{w}^{\prime}(X)}=\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1)=1$ for some positive constant $C>0$.
-Since $\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\leq 1$ almost surely given $W^{\prime}=1$, we must have $C\cdot a(X)/\underline{w}^{\prime}(X)\leq 1$ almost surely given $W^{\prime}=1$.
-This means $C$ is bounded above by $\inf(\text{supp}(\underline{w}^{\prime}(X)/a(X)\mid W^{\prime}=1))$, which is strictly positive by assumption.
-Therefore,
-$\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ $=\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]$
-$\leq\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}\mid W^{\prime}=1\right]$
-$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\frac{\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\right]}{\mathbb{P}(W^{\prime}=1)}$
-$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)$
-$\qquad\cdot\mathbb{E}\left[\mathbb{E}\left[\frac{a(X)}{\underline{w}^{\prime}(X)}W^{\prime}\mid X,W_{0}=1\right]\mid W_{0}=1\right]\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}$
-$=\inf\left(\text{supp}\left(\frac{\underline{w}^{\prime}(X)}{a(X)}\mid W^{\prime}=1\right)\right)\cdot\mathbb{E}[a(X)\mid W_{0}=1]\frac{\mathbb{P}(W_{0}=1)}{\mathbb{P}(W^{\prime}=1)}.$
-The fifth line follows from $W^{\prime}$ being a subpopulation of $W_{0}$, and the last line follows from the law of iterated expectations.
-This upper bound is sharp because it is attained by setting
-$W^{*}$ $=\mathbbm{1}\left(U\leq\frac{a(X)/\underline{w}^{\prime}(X)}{\sup(\text{supp}(a(X)/\underline{w}^{\prime}(X)\mid W^{\prime}=1))}\right)\cdot W^{\prime}$
-and noting that $W^{*}\in\mathcal{W}(a;W^{\prime},\mathcal{T}_{\text{all}})$ from Theorem 4.3.
-Now suppose $\mathbb{P}(a(X)\geq 0\mid W^{\prime}=1)<1$. By Theorem 4.3, $\mathcal{W}(a;W^{\prime},\mathcal{T})=\emptyset$ and therefore $\overline{P}(a,W^{\prime};\mathcal{T}_{\text{all}})$ is zero.
-Part 2: $\mathcal{T}=\{\tau_{0}\}$
-In this case, we seek to maximize $\mathbb{P}(W^{*}=1\mid W^{\prime}=1)$ subject to $W^{*}\in\mathcal{W}(a;W^{\prime},\{\tau_{0}\})$.
-Using Proposition 3.1, we can write $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ as $\mathbb{E}[\tau_{0}(X)\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]/\mathbb{E}[\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)\mid W^{\prime}=1]$.
-The result then follows from a direct application of Theorem 4.2 that replaces $W_{0}$ by $W^{\prime}$ in its statement and $\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$ by $\mathbb{P}(W^{*}=1\mid X,W^{\prime}=1)$ in the proofs.
-∎
-Appendix D Difference-in-Differences
-Goodman-Bacon (2021) provides the following representation of the two-way fixed effects estimand under the assumption that group-level average treatment effects are constant over time:
+Let $h=(h_{1},h_{2},h_{3})$ and $h^{\prime}=(h_{1}^{\prime},h_{2}^{\prime},h_{3}^{\prime})$.
+$|\widehat{\psi}(h^{\prime})-\widehat{\psi}(h)|$ $\leq\left|\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}(h_{1}^{\prime}(j)-h_{1}(j))\right|$
+$+\left|\frac{\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]}{\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\left(\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}^{\prime}(j)-\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}(j)\right)\right|$
+$+\left|\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}(h_{2}^{\prime}(j)-h_{2}(j))\right|$
+$+\left|\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{w}_{0}(x_{j})}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}(h_{3}^{\prime}(j)-h_{3}(j))\right|$
+$\leq\left(\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})^{2}\widehat{p}_{j}^{2}}{\widehat{\mathbb{P}}(W_{0}=1)^{2}\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\right)^{1/2}\|h_{1}^{\prime}-h_{1}\|$ (A.6)
+$+\frac{|\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]|}{\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\left|\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}^{\prime}(j)-\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}(j)\right|$ (A.7)
+$+\left(\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])^{2}\widehat{p}_{j}^{2}}{\widehat{\mathbb{P}}(W_{0}=1)^{2}\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\right)^{1/2}\|h_{2}^{\prime}-h_{2}\|$ (A.8)
+$+\left(\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])^{2}\widehat{w}_{0}(x_{j})^{2}}{\widehat{\mathbb{P}}(W_{0}=1)^{2}\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}}\right)^{1/2}\|h_{3}^{\prime}-h_{3}\|,$ (A.9)
+where we applied the Cauchy–Schwarz inequality several times. Note that the maximum function is Lipschitz with Lipschitz constant one and therefore
+$\left|\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}^{\prime}(j)-\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}(j)\right|$ $\leq\sum_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}|h_{1}^{\prime}(j)-h_{1}(j)|$
+$\leq\sum_{j=1}^{K}|h_{1}^{\prime}(j)-h_{1}(j)|$
+$\leq\sqrt{K}\|h_{1}^{\prime}-h_{1}\|.$
+[…]
+Thus we now consider the consistency of the different components of $\widehat{\psi}(h)$. Applying Theorem C.1 and the continuous mapping theorem, we can show that terms
+$\left(\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}h_{1}(j),\frac{\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]}{\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})^{2}},\right.$
+$\left.\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}h_{2}(j),\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{w}_{0}(x_{j})}{\widehat{\mathbb{P}}(W_{0}=1)\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}h_{3}(j)\right),$
+are all consistent for their counterparts in $\psi(h)$.
+It remains to show that $\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}h_{1}(j)=\max_{j\in\Psi_{\mathcal{X}^{+}}}h_{1}(j)+o_{p}(1)$.
+This holds if the set $\widehat{\Psi}_{\mathcal{X}^{+}}$ is consistent for $\Psi_{\mathcal{X}^{+}}$, which we establish here.
+Let $k\in\Psi_{\mathcal{X}^{+}}$.
+Then,
+$\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})$ $=\mathbb{P}\left(\widehat{a}(x_{k})\geq\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j})-\xi_{n}\right)$
+$=\mathbb{P}\left(\sqrt{n}(\widehat{a}(x_{k})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))\geq-\sqrt{n}\xi_{n}\right)$
+$=\mathbb{P}\left(\sqrt{n}(\max_{j\in\mathcal{X}^{+}}\widehat{a}(x_{j})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))\geq-\sqrt{n}\xi_{n}\right).$
+The third equality follows from $k\in\Psi_{\mathcal{X}^{+}}$.
+By the proof of Theorem C.1, $\sqrt{n}(\max_{j\in\mathcal{X}^{+}}\widehat{a}(x_{j})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))=o_{p}(1)$.
+Since $-\sqrt{n}\xi_{n}\rightarrow-\infty$, $\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})\to\mathbb{P}(0\geq-\infty)=1$ when $k\in\Psi_{\mathcal{X}^{+}}$.
+Now suppose that $k\notin\Psi_{\mathcal{X}^{+}}$. Then,
+$\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})$ $=\mathbb{P}(\widehat{a}(x_{k})\geq\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j})-\xi_{n})\rightarrow\mathbb{P}(a(x_{k})\geq\max_{j\in\mathcal{X}^{+}}a(x_{j})-0)=0,$
+[…]
+Appendix H Additional Derivations for Difference-in-Differences
+Goodman-Bacon (2021) provides the following representation of the two-way fixed effects estimand under the assumption that group-time average treatment effects are constant over time:
 $\beta_{\text{TWFE}}$ $=\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\Bigg{[}\sum_{j=1}^{k-1}\sigma_{jk}^{k}+\sum_{j=k+1}^{K}\sigma_{kj}^{k}\Bigg{]}\cdot\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1],$
@@
 […]
-$\beta_{\text{TWFE}}$ $=$ $\frac{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\mathbb{E}[Y(1)-Y(0)\mid G,D=1]]}{\mathbb{E}[a_{B}(G)\cdot\mathbb{P}(D=1\mid G)]}$
-$=$ $\frac{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)\cdot\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1]}{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)},$
-it becomes clear that, for each group $k$ other than the always treated and the never treated,
-$a_{B}(k)\cdot\mathbb{P}(D=1\mid G=k)$ $=$ $\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\mathbb{P}(D=1\mid G=k)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
+$\beta_{\text{TWFE}}$ $=$ $\frac{\mathbb{E}[a_{\text{TWFE,H}}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\tau_{0}(G)]}{\mathbb{E}[a_{\text{TWFE,H}}(G)\cdot\mathbb{P}(D=1\mid G)]}$
+$=$ $\frac{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{\text{TWFE,H}}(k)\cdot\mathbb{P}(D=1\mid G=k)\cdot\tau_{0}(k)}{\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\mathbb{P}(G=k)\cdot a_{\text{TWFE,H}}(k)\cdot\mathbb{P}(D=1\mid G=k)},$
+where $\tau_{0}(k)=\mathbb{E}[Y(1)-Y(0)\mid G=k,D=1]$, it becomes clear that, for each group $k$ other than the always treated and the never treated,
+$a_{\text{TWFE,H}}(k)\cdot\mathbb{P}(D=1\mid G=k)$
+$=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\mathbb{P}(D=1\mid G=k)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
 $+\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}1-\mathbb{P}(D=1\mid G=k)\big{]}\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]},$
 and this, in turn, implies that
-$a_{B}(k)$ $=$ $\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$ (D.1)
-$+$ $\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}.$
-D.1 Equivalence of Weight Functions
-We now show that the weights obtained in equation (D.1) are equivalent to those in Proposition 2.5. First, we rewrite the weights in (D.1) as follows:
-$a_{B}(k)$ $=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
+$a_{\text{TWFE,H}}(k)$
+$=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
+$+\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}.$ (A.1)
+Equivalence of Weight Functions
+We now show that the weights obtained in equation (A.1) are equivalent to those in Proposition 5.4. First, we rewrite the weights in (A.1) as follows:
+$a_{\text{TWFE,H}}(k)$
+$=\sum_{j=1}^{k-1}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=j)-\mathbb{P}(D=1\mid G=k)\big{]}$
 $+\;\sum_{j=k+1}^{K}\mathbb{P}(G=j)\cdot\big{[}\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big{]}\cdot\frac{1-\mathbb{P}(D=1\mid G=k)}{\mathbb{P}(D=1\mid G=k)}$
@@
 $=\mathbb{P}(D=1,G\neq k)-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
-$=(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]\mathbb{P}(G=k))-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
+$=(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]\mathbb{P}(G=k))-\mathbb{E}[D\mid G=k]\mathbb{P}(G\neq k)$
+$\quad+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)$
 $=\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right).$
 For $k\in\{2,\ldots,T\}$, the weights in Proposition 5.4 are equal to
-$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D].$ (D.2)
+$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D],$ (A.2)
 because they are the average of the weights in Proposition 5.3 conditional on $G=k$. The proof of Proposition 5.4 explicitly shows that
-$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$ $=\mathbb{P}(D=0\mid G=k)\cdot(\mathbb{P}(D=0\mid P\geq k)+\mathbb{P}(D=1\mid P<g)).$
-Let us look at the difference between the weights in (D.1) and (D.2). Fix $k\in\{2,\ldots,T\}$ and write
-$\left(1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D]\right)-\left(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)\right)$
+$\mathbb{E}[1-\mathbb{E}[D\mid G]-\mathbb{E}[D\mid P]+\mathbb{E}[D]\mid G=k]$
+$=\mathbb{P}(D=0\mid G=k)\cdot(\mathbb{P}(D=0\mid P\geq k)+\mathbb{P}(D=1\mid P<g)).$
+Let us look at the difference between the weights in (A.1) and (A.2). Fix $k\in\{2,\ldots,T\}$ and write:
+$\left(1-\mathbb{E}[D\mid G=k]-\mathbb{E}[D\mid P\geq k]+\mathbb{E}[D]\right)$
+$\quad-\left(\mathbb{E}[D]-\mathbb{E}[D\mid G=k]+\mathbb{P}(G>k)\left(1-\frac{\mathbb{E}[D\mid G>k]}{\mathbb{E}[D\mid G=k]}\right)\right)$
 $=1-\mathbb{E}[D\mid P\geq k]-\mathbb{P}(G>k)+\frac{\mathbb{E}[D\mathbbm{1}(G>k)]}{\mathbb{E}[D\mid G=k]}$
@@
 $=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[(F_{G}(P)-F_{G}(k))\mathbbm{1}(P\geq k)]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
-$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]-F_{G}(k)\mathbb{E}[\mathbbm{1}(P\geq k)]-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
+$=\frac{1}{\mathbb{E}[\mathbbm{1}(k\leq P)]}\left(F_{G}(k)\mathbb{E}[\mathbbm{1}(k\leq P)]+\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]-F_{G}(k)\mathbb{E}[\mathbbm{1}(P\geq k)]\right.$
+$\left.\quad-\mathbb{E}[F_{G}(P)\mathbbm{1}(P\geq k)]\right)$
 $=0.$
-Therefore, the weights in Proposition (2.5), (D.1), and (D.2) are all equal to one another.
+Therefore, the weights in Proposition 5.4 and equations (A.1) and (A.2) are all equal to one another.

### v4

@@
-Supp_Appendix_v4
+
 Abstract
@@
 Our paper develops practical tools to quantify the internal validity of weighted estimands.
+We also apply these tools to revisit a prominent study of the effects of unilateral divorce laws on female suicide.
 Keywords: internal validity, ordinary least squares, representativeness, treatment effects, two-stage least squares, two-way fixed effects, weakly causal estimands, weighted estimands
@@
 If $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is the target parameter, we interpret a large value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ as evidence of a high degree of internal validity of $\mu(a,\tau_{0})$ with respect to the target.
+[^2] We use the term “internal validity” because it is often associated with the question of whether the probability limit of an estimator is equal to the parameter of interest.
+If the researcher reports estimates of $\mu(a,\tau_{0})$ when $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$ is their ultimate target parameter, then the estimator is internally valid for the target when $\mathbb{P}(W^{*}=1\mid W_{0}=1)=1$.
+If $\mathbb{P}(W^{*}=1\mid W_{0}=1)<1$, then $\mu(a,\tau_{0})$ will be more informative about the target as the value of $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ increases.
 If $\mathbb{P}(W^{*}=1)$, the corresponding marginal probability, is large, we say that $\mu(a,\tau_{0})$ is highly representative of the underlying population.
 The answer to our questions about subpopulation existence and size depends on the information we have about the CATE function, $\tau_{0}$.
-Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any knowledge of this function.
+Specifically, in one case, we may want to know whether $\mu(a,\tau_{0})$ can be written as $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for any choice of $\tau_{0}$, or without any restrictions on this function.
 If this is the case, then we know that the interpretation of $\mu(a,\tau_{0})$ as a causal parameter is robust to heterogeneous treatment effects of any form, including the most adversarial CATE functions.
@@
 This expression can be used to derive plug-in estimators.
-Besides theoretical interest, we argue that $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ are a practically relevant measure that may be appealing to applied researchers.
-First, as stated above, our initial results demonstrate that two commonly invoked criteria for weighted estimands—that they lack negative weights and that they lie in the convex hull of CATE values—are necessary and sufficient (under different assumptions) for the existence of their causal representation, that is, for $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be strictly positive.
+Implications for Empirical Practice
+Besides theoretical interest, we argue that $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ represent a valuable diagnostic for empirical research.
+First, as stated above, our initial results demonstrate that two popular criteria for weighted estimands—that they lack negative weights and that they lie in the convex hull of CATE values—are necessary and sufficient (under different assumptions) for the existence of their causal representation, that is, for $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$ to be strictly positive.
 This suggests that researchers invoking these criteria are (indirectly) interested in whether their estimands can be represented as an average treatment effect for some subpopulation.
@@
 Indeed, large values of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)$, our primary measures of interest, guarantee that the weighted estimand is not “too different” from $\mathbb{E}[Y(1)-Y(0)]$ and $\mathbb{E}[Y(1)-Y(0)\mid W_{0}=1]$.
+This makes our measures a practical diagnostic tool to evaluate the robustness of weighted estimands to heterogeneous treatment effects of any form.
 Literature Review
 This paper is related to a large literature studying weighted average representations of common estimands, including ordinary least squares (OLS), 2SLS, and TWFE in additive linear models.
-Some of the contributions to this literature include Angrist (1998), Humphreys (2009), Aronow and Samii (2016), Słoczyński (2022), Chen (2024), and Goldsmith-Pinkham, Hull, and Kolesár (2024) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
+Some of the contributions to this literature include Angrist (1998), Aronow and Samii (2016), Słoczyński (2022), Chen (2024), Goldsmith-Pinkham, Hull, and Kolesár (2024), and Humphreys (2025) for OLS; Imbens and Angrist (1994), Angrist and Imbens (1995), Kolesár (2013), Słoczyński (2020), and Blandhol, Bonney, Mogstad, and Torgovitsky (2022) for 2SLS; and de Chaisemartin and D’Haultfœuille (2020), Goodman-Bacon (2021), Sun and Abraham (2021), Athey and Imbens (2022), Caetano and Callaway (2023), Borusyak, Jaravel, and Spiess (2024), and Callaway, Goodman-Bacon, and Sant’Anna (2024) for TWFE.
 A common view in much of this literature, attributable to Imbens and Angrist (1994), is that causal interpretability of weighted estimands requires all weights to be nonnegative.
@@
     Overlap: $p(X)\coloneqq\mathbb{P}(D=1\mid X)\in(0,1)$ almost surely.
-Following Angrist (1998), we can establish that $\beta_{\text{OLS}}$, the coefficient on $D$ in the linear projection of $Y$ on $(1,D,X)$, satisfies the representation in (1.1).
+Following Angrist (1998), we can establish that $\beta_{\text{OLS}}$, the coefficient on $D$ in the linear projection of $Y$ on $(1,D,X)$, satisfies the representation in (1.1) under a restriction on the propensity score.
 The following proposition summarizes Angrist’s (1998) result.
@@
 The overlap assumption can also be weakened since it is not required for $\beta_{\text{OLS}}$ to be defined.
-Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as a weighted estimand satisfying the representation in (1.1) where $a(X)=p(X)(1-p(X))$ and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$.
+Proposition 2.1 implies that we can write $\beta_{\text{OLS}}$ as a weighted estimand satisfying the representation in (1.1) where a(X)=p(X)(1p(X))a(X)=p(X)(1-p(X)) and $\tau_{0}(X)=\mathbb{E}[Y(1)-Y(0)\mid X]$.
 Here we implicitly set $W_{0}=1$ with probability 1.
-Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are $p(X)(1-p(X))$.
+Thus, the regression coefficient $\beta_{\text{OLS}}$ is a weighted average of CATEs whose weights are p(X)(1p(X))p(X)(1-p(X)).
 Note that $\beta_{\text{OLS}}=\text{ATE}\coloneqq\mathbb{E}[Y(1)-Y(0)]$ if and only if $a(X)$ and $\tau_{0}(X)$ are uncorrelated, which is the case, for example, when $p(X)$ or $\tau_{0}(X)$ is constant.
@@
 In this case, $\tau_{0}$ denotes the conditional local average treatment effect.
-[…]
+Note that $\tau_{0}$ is defined for all values of $X$ such that $w_{0}(X)=\mathbb{P}(W_{0}=1\mid X)>0$.
+[^3] While $\tau_{0}(X)$ is only defined when $w_{0}(X)>0$, we set $\tau_{0}(X)w_{0}(X)=0$ when $w_{0}(X)=0$.
+Throughout this paper, we assume that $\mathbb{P}(W_{0}=1)>0$, so that this subpopulation has a positive mass, which avoids technical issues associated with conditioning on zero-probability events.
 Also recall the weighted estimands of equation (1.1):
@@
     (Conditional independence) $W^{*}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$.
-For convenience, we will abbreviate this as “$W^{*}$ is a regular subpopulation of $W_{0}$”. We denote the set of regular subpopulations of $W_{0}$ as
-$\text{SP}(W_{0})$ =$=\{W^{*}\in\{0,1\}:W^{*}\text{ is a regular subpopulation of }W_{0}\}.$
+For convenience, we will abbreviate this as $W^{*}$ is a regular subpopulation of $W_{0}$”. We denote the set of regular subpopulations of $W_{0}$ as
+[…]
 These subpopulations have positive masses and are subsets of $\{W_{0}=1\}$.
@@
 However, if $\tau_{0}$ is nonconstant, the existence of causal representations will depend on the weight function $a(X)$.
-Among other cases, we will consider the case where no restrictions are placed on function $\tau_{0}$, and in this case, the existence of a causal representation of $\mu(a,\tau_{0})$ will require the sign of $a(X)$ to be constant.
+Among other cases, we will consider the case where no restrictions are placed on function $\tau_{0}$.
+In this case, the existence of a causal representation of $\mu(a,\tau_{0})$ will require the sign of $a(X)$ to be constant.
 To formalize this, let $\mathcal{T}$ denote a class of functions such that $\tau_{0}\in\mathcal{T}$ and define
-$\mathcal{W}(a;W_{0},\mathcal{T})$ =$=\{W^{*}\in\text{SP}(W_{0}):\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\text{ for all }\tau_{0}\in\mathcal{T}\}.$
-This is the set of regular subpopulations of $W_{0}$ such that the estimand $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for all $\tau_{0}$ functions in the set $\mathcal{T}$.
-If the set $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty, then estimand $\mu(a,\tau_{0})$ cannot be written as an average treatment effect over a regular subpopulation of $W_{0}$ uniformly in $\tau_{0}\in\mathcal{T}$.
+$\mathcal{W}(a;W_{0},\mathcal{T})$ $=\{W^{*}\in\text{SP}(W_{0}):\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]\text{ for all }\tau_{0}\in\mathcal{T}\}.$
+This is the set of regular subpopulations of $W_{0}$ such that the estimand $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ for all functions $\tau_{0}$ in the set $\mathcal{T}$.
+If the set $\mathcal{W}(a;W_{0},\mathcal{T})$ is empty, then the estimand $\mu(a,\tau_{0})$ cannot be written as an average treatment effect over a regular subpopulation of $W_{0}$ uniformly in $\tau_{0}\in\mathcal{T}$.
 We use this set to formally define a notion of uniform causal representation.
@@
 $\mu(a,\tau_{0})\in\mathcal{S}(\tau_{0};W_{0})\coloneqq\{t\in\mathbb{R}:\;$ $\mathbb{P}(\tau_{0}(X)\leq t\mid W_{0}=1)>0$
-[…]
+$\text{ and }\mathbb{P}(\tau_{0}(X)\geq t\mid W_{0}=1)>0\}.$
 The existence condition in Theorem 3.2 is weaker than the one in Theorem 3.1 since we no longer require this representation to be valid for any CATE function, but rather just for the one that is identified from the population.
@@
 Let $\mu(a,\tau_{0})$ be an estimand satisfying equation (1.1). Suppose $X$ has finite support. Suppose Assumption 3.1 holds and define
-$\mathcal{T}_{\text{lin}}$ =$=\{\tau_{0}\in\mathcal{T}_{\text{all}}:\tau_{0}(X)=c+d^{\prime}X:(c,d)\in\mathbb{R}^{1+d_{X}}\}.$
+$\mathcal{T}_{\text{lin}}$ $=\{\tau_{0}\in\mathcal{T}_{\text{all}}:\tau_{0}(X)=c+d^{\prime}X:(c,d)\in\mathbb{R}^{1+d_{X}}\}.$
 Then, $\mu(a,\tau_{0})$ has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{lin}}$ if and only if
@@
 For example, if $\text{supp}(X)=\{0,1,2\}$ and $W_{0}=1$ almost surely, then any combination of values of $(a(0),a(1),a(2))$ such that $\mathbb{E}[a(X)X]/\mathbb{E}[a(X)]\in[0,2]=\text{conv}(\text{supp}(X))$ implies a causal representation.
-Let $\mathbb{P}(X=x)=1/3$ for $x\in\{0,1,2\}$ and $(a(0),a(1),a(2))=(1,-1,1)$.
+Let $\mathbb{P}(X=x)=1/3$ for $x\in\{0,1,2\}$ and (a(0),a(1),a(2))=(1,1,1)(a(0),a(1),a(2))=(1,-1,1).
 Here units with $X=1$ have a negative weight, but $\mathbb{E}[a(X)X]/\mathbb{E}[a(X)]=1\in[0,2]$, implying that the corresponding weighted estimand has a causal representation uniformly in $\tau_{0}\in\mathcal{T}_{\text{lin}}$.
@@
 This bound can be computed at what Imbens and Rubin (2015) call the “design stage” of the study, that is, without any knowledge of the conditional distribution of the outcome.
-{adjustwidth}
--1.25in-1.25in Refer to caption Note: $X$ is a single continuously distributed covariate with density $f_{X}$.
-Figure 1: Characterizing a Representative Subpopulation Uniformly in $\mathcal{T}_{\text{all}}$
 To understand the supremum’s role in this expression, let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)$ and note that $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ is equivalent to writing
@@
 All units in $\{W_{0}=1\}$ with $X$ such that $a(X)=a_{\max}$ are included in $W^{*}$, whereas no units where $a(X)=0$ are included.
-[…]
+The construction of this subpopulation is illustrated in Figure 1 for the case where $x$ is continuous and where we omit the conditioning on $W_{0}=1$ for simplicity.
+We seek to maximize $\mathbb{P}(W^{*}=1)=\int\underline{w}^{*}(x)f_{X}(x)\,dx$ with the requirement that $\underline{w}^{*}(x)\leq 1$ (or, equivalently, $\underline{w}^{*}(x)f_{X}(x)\leq f_{X}(x)$) and that $\underline{w}^{*}(x)$ is a multiple of $a(x)$.
+In the figure, we see that $a_{\max}>1$ and thus the largest multiple of $a(x)$ that is weakly smaller than 1 is illustrated by the gray curve.
+The area under this curve is precisely $\mathbb{P}(W^{*}=1)$.
+Note that the area under $f_{X}(x)$ is one, so closer alignment of the gray curve and the density $f_{X}(x)$ corresponds to more representative estimands.
+Figure 1: Characterizing a Representative Subpopulation Uniformly in $\mathcal{T}_{\text{all}}$
+Note: $X$ is a single continuously distributed covariate with density $f_{X}$.
 Several further comments about Theorem 4.1 are in order.
@@
 Reporting the average covariate values of units within and outside of $W^{*}$ can be of interest when assessing the representativeness of $\mu(a,\tau_{0})$.
+Remark 4.4 (Meaning of internal validity and representativeness).
+Suppose we consider $\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ to be the target parameter, where $W^{*}=\mathbbm{1}\left(U\leq a(X)/a_{\max}\right)\cdot W_{0}$ is the subpopulation for which $\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$ uniformly in $\tau_{0}$.
+For example, in the case of the OLS estimand in Section 2, this consists of a subpopulation where the probability of inclusion is proportional to the conditional variance of treatment.
+If this subpopulation is the target, it would be reasonable to infer that the measure of internal validity of the estimand $\mu(a,\tau_{0})$ is the maximum value of 1.
+Indeed, this is the case because the estimand can be written as
+$\mu(a,\tau_{0})=\mathbb{E}[Y(1)-Y(0)\mid W^{*}=1]$
+and $\{W^{*}=1\}$ is trivially the largest regular subpopulation of $\{W^{*}=1\}$.
+This example illustrates how the internal validity of $\mu(a,\tau_{0})$ is entirely dependent on the target estimand.
+This is appropriate because we associate the term “internal validity” with the question of whether the probability limit of an estimator is equal to the parameter of interest.
+Thus, the degree of internal validity should naturally depend on the parameter choice.
+On the other hand, the representativeness of the weighted estimand, as measured by the largest value of $\mathbb{P}(W^{*}=1)$, is independent of the target parameter.
+In fact, it will be less than one unless $W^{*}=1$ almost surely, or that the weighted estimand actually equals the ATE.
+If, in addition, we wish to operationalize the concept of external validity in our framework, we need to consider whether the weighted estimand can be written as the average treatment effect within a subpopulation of another, possibly arbitrary population.
+This extension is left for future work.
 We now consider a simple example to give further intuition for Theorem 4.1.
@@
 $+\;\mathbb{E}[Y(1)-Y(0)\mid W^{+}+W^{-}=0]\cdot\left[1-\mathbb{P}(W^{+}=1)-\mathbb{P}(W^{-}=1)\right].$ (4.6)
-Substituting equation (4.5) in (4.6) and assuming that $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=1]$ and $\mathbb{E}[Y(1)-Y(0)\mid W^{+}+W^{-}=0]$ lie in $[B_{\ell},B_{u}]$ yields
+Substituting equation (4.5) in (4.6) and assuming that $\mathbb{E}[Y(1)-Y(0)\mid W^{-}=1]$ and $\mathbb{E}[Y(1)-Y(0)\mid W^{+}+W^{-}=0]$ lie in $B_{\ell},B_{u}]$ yields
 $\mathbb{E}[Y(1)-Y(0)]\;$ $\leq\;\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\cdot\mu(a,\tau_{0})\;+\;\left(1-\frac{\mathbb{P}(W^{+}=1)}{\omega^{+}}\right)\cdot B_{u}$
@@
 The computation of these bounds can be done using a linear programming algorithm when $X$ is discrete, or through plug-in estimators of the terms in equation (4.9) regardless of the nature of the support of $X$.
-To illustrate this theorem, let $\tau_{0}(X)$ be continuously distributed with support $[\underline{\tau},\overline{\tau}]$, and suppose $\mu_{0}\in[\underline{\tau},\overline{\tau}]$.
+To illustrate this theorem, let $\tau_{0}(X)$ be continuously distributed with support $\underline{\tau},\overline{\tau}]$, and suppose $\mu_{0}\in[\underline{\tau},\overline{\tau}]$.
 Without loss of generality, assume $E_{0}\geq\mu_{0}$.
@@
 In particular, this is the case when there are a few units with extreme values of $\tau_{0}$ whose removal has a large impact on the estimand, but a small impact on the share of the population.
+Figure 2: Characterizing a Representative Subpopulation When $\tau_{0}$ Is Known
+Note: The figure assumes that $\tau_{0}(X)$ is continuously distributed, that $W_{0}=1$ almost surely, and that $\mu<\mathbb{E}[\tau_{0}(X)]=\text{ATE}$.
+Theorem 4.2 can also be illustrated visually.
+In Figure 2, the probability density function of $\tau_{0}(X)$ is drawn.
+In this figure, it is assumed that $\tau_{0}(X)$ is continuously distributed, that $W_{0}=1$ almost surely, and that $\mu<\mathbb{E}[\tau_{0}(X)]=\text{ATE}$.
+The representative subpopulation is obtained by trimming away covariate values that correspond to $\tau_{0}(X)\geq\alpha^{+}$, where $\alpha^{+}$ is determined by the equation $\mathbb{E}[\tau_{0}(X)\mid\tau_{0}(X)\leq\alpha^{+}]=\mu$.
+The size of the shaded area is the measure of internal validity.
 5  Applications to Common Estimands
@@
 Suppose the target estimand is the average treatment effect, i.e., $W_{0}=1$ almost surely.
-By Theorem 3.1, there exists a regular subpopulation $W^{*}$ such that $\beta_{\text{OLS}}$ equals the average treatment effect over $W^{*}$ since the weight function $a(X)=p(X)(1-p(X))$ is nonnegative.
+By Theorem 3.1, there exists a regular subpopulation $W^{*}$ such that $\beta_{\text{OLS}}$ equals the average treatment effect over $W^{*}$ since the weight function a(X)=p(X)(1p(X))a(X)=p(X)(1-p(X)) is nonnegative.
 By Theorem 4.1, the upper bound on the size of subpopulation $W^{*}$ is
@@
 These weights are nonnegative and therefore Theorem 3.1 guarantees the existence of a causal representation for $\beta_{\text{TWFE}}$ uniformly in $\tau_{0}\in\mathcal{T}_{\text{all}}$.
+[^5] In the context of Proposition 5.4, $\tau_{0}$ is a function of $G$ only, thus $\mathcal{T}_{\text{all}}$ denotes the set of all functions of $G$ with finite second moments.
+Note that this is a strict subset of all “time-heterogeneous” conditional average treatment effects, $\mathbb{E}[Y(1)-Y(0)\mid D=1,G,P]$.
 Using Theorem 4.1, the internal validity of $\beta_{\text{TWFE}}$ relative to target parameter $\mathbb{E}[Y(1)-Y(0)\mid D=1]$ is given by
@@
 $\widehat{\overline{P}}$ $\coloneqq\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})\cdot\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})}.$
-Here $c_{n}$ is a tuning parameter that converges to 0 as $n$ diverges.
-We start by noting that we can estimate $\mathbb{E}[a(X)\mid W_{0}=1]$ via $\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, which will be consistent under standard conditions on $\widehat{a}(\cdot)$ and $\widehat{w}_{0}(\cdot)$.
-Estimating $a_{\max}=\sup(\text{supp}(a(X)\mid W_{0}=1))$ is more delicate.
-In some of our examples, this supremum is known or can be bounded above without using data.
-For example, the OLS estimand under unconfoundedness has weights $a(X)=\operatorname{var}(D\mid X)$ which are bounded above by $1/4$.
-If $X$ is continuously distributed, then 1/2 may lie in the support of $p(X)$, and thus we may avoid the estimation of $a_{\max}$.
-The IV estimand of Section 5 has weights $a_{\text{IV}}(X)=\operatorname{var}(Z\mid X)$ which are similarly bounded above by $1/4$.
-Similarly, $a_{\text{2SLS}}(X)=|\operatorname{cov}(D,Z\mid X)|\leq 1/4$ by the Cauchy–Schwarz inequality.
-If knowledge of $a_{\max}$ is not assumed, but $\text{supp}(X\mid W_{0}=1)$ is known and $a(x)$ is continuous, [^4] Note that $a(x)$ is trivially continuous on finite support.
-then $\sup_{x\in\text{supp}(X\mid W_{0}=1)}\widehat{a}(x)$ will be consistent for $a_{\max}$ when $\widehat{a}(x)$ is consistent for $a(x)$ uniformly in $x\in\text{supp}(X\mid W_{0}=1)$.
-Many parametric and nonparametric estimators for $a(\cdot)$ satisfy this requirement.
+[…]
 In Appendix C, we prove the consistency of $\widehat{\overline{P}}$ and derive its limiting distribution.
@@
 The conclusion changes, however, when we explicitly target the average treatment effect on the treated (ATT), that is, the average effect for the largest subpopulation for which such an effect is identified under standard assumptions.
-Using the approach of Wooldridge (2021), we obtain an estimate of –5.530 with a $p$-value of 0.138.
-The approach of Callaway and Sant’Anna (2021) produces an estimate of –10.220 and a $p$-value of 0.001.
+Using the approach of Callaway and Sant’Anna (2021), we obtain an estimate of –10.220 with a $p$-value of 0.001.
+The approach of Wooldridge (2025) produces an estimate of –5.530 and a $p$-value of 0.138.
 These estimates are more strongly suggestive of a causal effect of unilateral divorce laws than the TWFE estimate.
-{adjustwidth}
--1.25in-1.25in A.
-Estimates of the effects of unilateral divorce laws TWFE ATT Callaway and Sant’Anna Wooldridge –0.604 –10.220 –5.530 (2.622) (3.086) (3.650) B.
-Internal validity of the TWFE estimand based on Proposition 5.3 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216 C.
-Internal validity of the TWFE estimand based on Proposition 5.4 uniformly in $\tau_{0}$ given $\tau_{0}$$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802 ^$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
 Table 1: Internal Validity of TWFE Estimand for Effects of Unilateral Divorce Laws
+A. Estimates of the effects of unilateral divorce laws
+TWFE ATT
+Callaway and Sant’Anna Wooldridge
+–0.604 –10.220 –5.530
+(2.622) (3.086) (3.650)
+B. Internal validity of the TWFE estimand based on Proposition 5.3
+uniformly in $\tau_{0}$ given $\tau_{0}$
+$\widehat{\mathbb{P}}(W^{*}=1)$ 0 0.3873
+$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0 0.6216
+C. Internal validity of the TWFE estimand based on Proposition 5.4
+uniformly in $\tau_{0}$ given $\tau_{0}$
+$\widehat{\mathbb{P}}(W^{*}=1)$ 0.1400 0.4802
+$\widehat{\mathbb{P}}(W^{*}=1\mid D=1)$ 0.2246 0.7707
   • Notes: The data are a panel of the 1964–1996 U.S.
@@
 Panel B of Table 1 reports our estimates of $\mathbb{P}(W^{*}=1)$ and $\mathbb{P}(W^{*}=1\mid D=1)$, based on the representation of the TWFE estimand in de Chaisemartin and D’Haultfœuille (2020), revisited in our Proposition 5.3.
-First, because the weights on some group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, ^$\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$.
+First, because the weights on some group-time average treatment effects are negative, the TWFE estimand does not have a causal interpretation uniformly in $\tau_{0}$, $\widehat{\mathbb{P}}(W^{*}=1)=\widehat{\mathbb{P}}(W^{*}=1\mid D=1)=0$.
 Second, when we estimate the CATE function and use these estimates in constructing the bounds, we conclude that the TWFE estimand corresponds to the average treatment effect for at most 62.16% of the treated units or 38.73% of the entire population.
@@
 Proof of Theorem 3.1.
-()($\Longrightarrow$)( ⟹ ) First, suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$.
+$(\Longrightarrow)$ First, suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{all}})$.
 By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{all}}$.
@@
 Expanding this equality yields $0=\mu(a,\tau^{*})-\mu(\underline{w}^{*},\tau^{*})=\mathbb{E}[\tau^{*}(X)^{2}\mid W_{0}=1]$, which implies that
-11 11 $=\mathbb{P}(\tau^{*}(X)=0\mid W_{0}=1)=\mathbb{P}\left(a(X)=\frac{\underline{w}^{*}(X)\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}\mid W_{0}=1\right).$ (A.2)
+$1$ $=\mathbb{P}(\tau^{*}(X)=0\mid W_{0}=1)=\mathbb{P}\left(a(X)=\frac{\underline{w}^{*}(X)\mathbb{E}[a(X)\mid W_{0}=1]}{\mathbb{P}(W^{*}=1\mid W_{0}=1)}\mid W_{0}=1\right).$ (A.2)
 Note that $\underline{w}^{*}(X)\geq 0$ and $\mathbb{P}(W^{*}=1\mid W_{0}=1)>0$ by construction.
@@
 To simplify the notation in the proof, we let $\mu_{0}\coloneqq\mu(a,\tau_{0})$.
-()($\Longrightarrow$)( ⟹ ) First, suppose $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$.
+$(\Longrightarrow)$ First, suppose $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$ and suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\{\tau_{0}\})$.
 Since $\mu_{0}\notin\mathcal{S}(\tau_{0};W_{0})$, we can without loss of generality suppose that $\mathbb{P}(\tau_{0}(X)>\mu_{0}\mid W_{0}=1)=1$.
@@
 Proof of Proposition 3.2.
-()($\Longrightarrow$)( ⟹ ) First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$.
+$(\Longrightarrow)$ First, we suppose there exists $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{lin}})$.
 By Proposition 3.1, we have that $\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})=0$ for all $\tau_{0}\in\mathcal{T}_{\text{lin}}$.
 Therefore,
-00 0 $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
+$0$ $=\mu(a,\tau_{0})-\mu(\underline{w}^{*},\tau_{0})$
 $=\frac{\mathbb{E}[a(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[a(X)\mid W_{0}=1]}-\frac{\mathbb{E}[\underline{w}^{*}(X)(c+d^{\prime}X)\mid W_{0}=1]}{\mathbb{E}[\underline{w}^{*}(X)\mid W_{0}=1]}$
@@
 Case 1: $K>0$.
-()($\Longrightarrow$)( ⟹ ) First, let $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. We will show that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))=\emptyset$ by way of contradiction.
+$(\Longrightarrow)$ First, let $\mathbb{P}(a(X)\geq 0\mid W_{0}=1)<1$. We will show that $\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))=\emptyset$ by way of contradiction.
 Suppose there is a $W^{*}\in\mathcal{W}(a;W_{0},\mathcal{T}_{\text{BD}}(K))$ and let $\underline{w}^{*}(X)=\mathbb{P}(W^{*}=1\mid X,W_{0}=1)\in[0,1]$.
@@
 We note that these estimands rule out “level dependence,” i.e., that the estimand changes if potential outcomes $(Y(0),Y(1))$ are translated to $(Y(0)+c,Y(1)+c)$ for some constant $c\in\mathbb{R}$.
-For example, the IV estimand is generally level dependent when the propensity score $\mathbb{P}(Z=1\mid X)$ is nonlinear in $X$ [^6] See p.
+For example, the IV estimand is generally level dependent when the propensity score $\mathbb{P}(Z=1\mid X)$ is nonlinear in XX.
+[^8] See p.
 17 in BBMT.
@@
 This estimator does not assume knowledge of the support of $X$ given $W_{0}=1$, but it can also be implemented by taking the maximum over $\text{supp}(X\mid W_{0}=1)$ when it is known.
-Let $\text{supp}(X)=\{x_{1},\ldots,x_{K}\}$, denote by $p_{j}=\mathbb{P}(X=x_{j})$ the frequency of cell $j$, and let $\widehat{p}_{j}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})$ denote its sample frequency.
-Let $\widehat{\theta}=(\widehat{\textbf{a}},\widehat{\textbf{w}}_{0},\widehat{\textbf{p}})$ where $\widehat{\textbf{a}}=(\widehat{a}(x_{1}),\ldots,\widehat{a}(x_{K}))$, $\widehat{\textbf{w}}_{0}=(\widehat{w}_{0}(x_{1}),\ldots,\widehat{w}_{0}(x_{K}))$, and $\widehat{\textbf{p}}=(\widehat{p}_{1},\ldots,\widehat{p}_{K})$.
-Let $\theta=(\textbf{a},\textbf{w}_{0},\textbf{p})$ denote their population counterparts.
+[…]
 We make the following assumption on the behavior of the first-step estimators.
@@
 Let
-[…]
+$\sqrt{n}(\widehat{\theta}-\theta)\xrightarrow{d}\mathbb{Z}$
 where $\mathbb{Z}\coloneqq(\mathbb{Z}_{\textbf{a}},\mathbb{Z}_{\textbf{w}_{0}},\mathbb{Z}_{X})\in\mathbb{R}^{3K}$ has a Gaussian distribution.
@@
 Then, $\widehat{\overline{P}}$ is consistent for $\overline{P}$ and
-[…]
+$\sqrt{n}(\widehat{\overline{P}}-\overline{P})$ $\xrightarrow{d}\psi(\mathbb{Z}),$
 where $\psi$ is a continuous mapping defined by
@@
 Assumption C.2 (Bootstrap for first-step estimators).
-[…]
+Let $\mathbb{Z}^{*}\coloneqq(\mathbb{Z}^{*}_{\textbf{a}},\mathbb{Z}^{*}_{\textbf{w}_{0}},\mathbb{Z}^{*}_{X})\in\mathbb{R}^{3K}$ be a random vector such that $\mathbb{Z}^{*}\overset{p}{\rightsquigarrow}\mathbb{Z}$, where $\overset{p}{\rightsquigarrow}$ denotes convergence in probability conditioning on the data used to compute $\widehat{\theta}$.
 This assumption is easily satisfied when $\widehat{\textbf{p}}$ are sample frequencies, $(\widehat{\textbf{a}},\widehat{\textbf{w}}_{0})$ are cell-by-cell estimators that are asymptotically linear and Gaussian, and when $\mathbb{Z}^{*}$ is the distribution of these estimators under a standard bootstrap approach.
@@
 Suppose the assumptions of Theorem C.1 hold and that Assumption C.2 holds. Then,
+$\widehat{\psi}(\mathbb{Z}^{*})\overset{p}{\rightsquigarrow}\psi(\mathbb{Z})$
+as $n\rightarrow\infty$, where $\widehat{\psi}$ is defined by
 […]
-as $n\rightarrow\infty$, where $\widehat{\psi}$ is defined by
-$\widehat{\psi}(\mathbb{Z}^{*})\;$ $=\;\sum_{j=1}^{K}\frac{\widehat{w}_{0}(x_{j})\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\widehat{a}_{\max}}\mathbb{Z}^{*}_{\textbf{a}}(j)\;-\;\frac{\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]}{\widehat{a}_{\max}^{2}}\max_{j\in\widehat{\Psi}_{\mathcal{X}^{+}}}\mathbb{Z}^{*}_{\textbf{a}}(j)$
 $+\;\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{p}_{j}}{\widehat{\mathbb{P}}(W_{0}=1)\widehat{a}_{\max}}\mathbb{Z}^{*}_{\textbf{w}_{0}}(j)$
 $+\;\sum_{j=1}^{K}\frac{(\widehat{a}(x_{j})-\widehat{\mathbb{E}}[a(X)\mid W_{0}=1])\widehat{w}_{0}(x_{j})}{\widehat{\mathbb{P}}(W_{0}=1)\widehat{a}_{\max}}\mathbb{Z}^{*}_{X}(j),$ (A.2)
-where $\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, ^$\widehat{\mathbb{P}}(W_{0}=1)=\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})$, $\widehat{a}_{\max}=\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$, and
+where $\widehat{\mathbb{E}}[a(X)\mid W_{0}=1]=\frac{\frac{1}{n}\sum_{i=1}^{n}\widehat{a}(X_{i})\widehat{w}_{0}(X_{i})}{\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})}$, $\widehat{\mathbb{P}}(W_{0}=1)=\frac{1}{n}\sum_{i=1}^{n}\widehat{w}_{0}(X_{i})$, $\widehat{a}_{\max}=\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$, and
 $\widehat{\Psi}_{\mathcal{X}^{+}}=\left\{k\in\{1,\ldots,K\}:\widehat{a}(x_{k})\geq\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})-\xi_{n}\right\},$
@@
   2. 2.
-    For bootstrap samples $b=1,\ldots,B$, compute $\widehat{\theta}^{\ast,b}=(\widehat{\textbf{a}}^{\ast,b},\widehat{\textbf{w}}_{0}^{\ast,b},\widehat{\textbf{p}}^{\ast,b})$ and $\mathbb{Z}^{\ast,b}=\sqrt{n}(\widehat{\theta}^{\ast,b}-\widehat{\theta})$;
+[…]
   3. 3.
@@
 $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ $=\inf_{x\in\text{supp}(X\mid W_{0}=1)}\frac{\mathbb{E}[a(X)w_{0}(X)]}{\mathbb{E}[w_{0}(X)]\cdot a(x)}=:\inf_{x\in\text{supp}(X\mid W_{0}=1)}\overline{P}(x).$
-Computing a one-sided confidence interval for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ of the kind $[0,\widehat{\overline{P}}^{+}]$ can be cast as doing inference on intersection bounds.
+Computing a one-sided confidence interval for $\overline{P}(a,W_{0};\mathcal{T}_{\text{all}})$ of the kind $0,\widehat{\overline{P}}^{+}]$ can be cast as doing inference on intersection bounds.
 Chernozhukov, Lee, and Rosen (2013) offer methods for such problems.
-Equivalently, the computation of a one-sided confidence interval $[\widehat{\overline{P}}^{-},1]$ is related to inferential questions in union bounds: see Bei (2024).
+Equivalently, the computation of a one-sided confidence interval $\widehat{\overline{P}}^{-},1]$ is related to inferential questions in union bounds: see Bei (2024).
 We leave all details for future work.
@@
 Letting $\varepsilon\searrow 0$ yields that $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]\leq 0$.
-Therefore $-\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]/(\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1))\geq 0$ and $\underline{w}^{+}(X)\geq 0$.
+Therefore $\mathbb{E}[T_{\mu}\mathbbm{1}(T_{\mu}<\alpha^{+})\mid W_{0}=1]/(\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1))\geq 0$ and $\underline{w}^{+}(X)\geq 0$.
 Next, we compute
@@
 $W^{+}$ $=\left(\mathbbm{1}(T_{\mu}<\alpha^{+})+\mathbbm{1}\left(T_{\mu}=\alpha^{+},U\leq 1-\frac{R(\alpha^{+})\mathbbm{1}(\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)\neq 0)}{\alpha^{+}\mathbb{P}(T_{\mu}=\alpha^{+}\mid W_{0}=1)}\right)\right)\cdot W_{0}.$
-[…]
+By construction, $W^{+}\in\{0,1\}$, $\mathbb{P}(W^{+}=\;1\mid X,W_{0}\;=\;1)\;=\;\underline{w}^{+}(X)$, $\mathbb{P}(W_{0}\;=\;1\mid W^{+}=\;1)\;=\;1$, and $W^{+}\perp\!\!\!\perp(Y(1),Y(0))\mid X,W_{0}=1$.
+Also, since $\mu_{0}\in\mathcal{S}(\tau_{0};W_{0})$, $\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)=\mathbb{P}(\tau_{0}(X)\leq\mu_{0}\mid W_{0}=1)>0$.
+Since $\alpha^{+}>0$ we have that $\mathbb{P}(W^{+}=1\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}<\alpha^{+}\mid W_{0}=1)\geq\mathbb{P}(T_{\mu}\leq 0\mid W_{0}=1)>0$.
+Therefore $W^{+}$ is a regular subpopulation of $W_{0}$ for which $\mathbb{P}(W^{+}=1\mid W_{0}=1)=P^{+}$, hence $P^{+}$ is the maximum.
 Case 2: $\mu_{0}\in\mathcal{S}(\tau_{0};W_{0})$ and $\mu_{0}>E_{0}$
@@
 $=\mathbb{E}[\ddot{D}\mathbb{E}[Y(0)\mid G,P]]+\mathbb{E}[\ddot{D}D\mathbb{E}[Y(1)-Y(0)\mid G,P]].$
-The first equality follows from $Y=Y(0)+D(Y(1)-Y(0))$ and the second from iterated expectations and $\mathbb{E}[D\mid G,P]=D$.
+The first equality follows from Y=Y(0)+D(Y(1)Y(0))Y=Y(0)+D(Y(1)-Y(0)) and the second from iterated expectations and $\mathbb{E}[D\mid G,P]=D$.
 Step 2: First Numerator Term
@@
 Then,
-$\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})$ =($=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right\}\cap\left\{\widehat{w}_{0}(x_{j})>c_{n}\right\}\right)$
+$\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})$ $=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right\}\cap\left\{\widehat{w}_{0}(x_{j})>c_{n}\right\}\right)$
 $\geq\mathbb{P}\left(\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0\right)+\mathbb{P}\left(\widehat{w}_{0}(x_{j})>c_{n}\right)-1,$
 following an application of Bonferroni’s inequality.
-[…]
+We have that $\mathbb{P}(\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0)\rightarrow 1$ since $\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})\xrightarrow{p}p_{j}>0$.
+We also have that $\mathbb{P}(\widehat{w}_{0}(x_{j})>c_{n})=\mathbb{P}(\widehat{w}_{0}(x_{j})-c_{n}>0)\rightarrow 1$ because $\widehat{w}_{0}(x_{j})-c_{n}\xrightarrow{p}w_{0}(x_{j})>0$ by $c_{n}=o(1)$ and $w_{0}(x_{j})>0$, which follows from $x_{j}\in\mathcal{X}^{+}$.
+Therefore, $\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})=1-\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})\leq 1-(\mathbb{P}(\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})>0)+\mathbb{P}(\widehat{w}_{0}(x_{j})>c_{n})-1)\rightarrow 0$ as $n\rightarrow\infty$.
 Now let $x_{j}\notin\mathcal{X}^{+}$. Then
-$\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})$ =($=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})=0\right\}\cup\left\{\widehat{w}_{0}(x_{j})\leq c_{n}\right\}\right)$
+$\mathbb{P}(x_{j}\notin\widehat{\mathcal{X}}^{+})$ $=\mathbb{P}\left(\left\{\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}(X_{i}=x_{j})=0\right\}\cup\left\{\widehat{w}_{0}(x_{j})\leq c_{n}\right\}\right)$
 $\geq\mathbb{P}(\widehat{w}_{0}(x_{j})\leq c_{n})=\mathbb{P}(\sqrt{n}\widehat{w}_{0}(x_{j})\leq\sqrt{n}c_{n}).$
-[…]
+By Assumption C.1, $\sqrt{n}\widehat{w}_{0}(x_{j})=\sqrt{n}(\widehat{w}_{0}(x_{j})-w_{0}(x_{j}))\xrightarrow{d}\mathbb{Z}_{\textbf{w}_{0}}(j)=O_{p}(1)$, since $w_{0}(x_{j})=0$ for $x_{j}\notin\mathcal{X}^{+}$.
+Also $\sqrt{n}c_{n}\rightarrow+\infty$ by the theorem’s assumption.
+Therefore, $\mathbb{P}(\sqrt{n}\widehat{w}_{0}(x_{j})\leq\sqrt{n}c_{n})\rightarrow 1$ and $\mathbb{P}(x_{j}\in\widehat{\mathcal{X}}^{+})\rightarrow 0$ as $n\rightarrow\infty$.
+Because of this,
 $\mathbb{P}(\widehat{\mathcal{X}}^{+}=\mathcal{X}^{+})$ $=\mathbb{P}\left(\bigcap_{x_{j}\in\mathcal{X}^{+}}\{x_{j}\in\widehat{\mathcal{X}}^{+}\}\cap\bigcap_{x_{j}\notin\mathcal{X}^{+}}\{x_{j}\notin\widehat{\mathcal{X}}^{+}\}\right)$
@@
 $\max_{i:\widehat{w}_{0}(X_{i})>c_{n}}\widehat{a}(X_{i})$ $=\max_{x\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x)=\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)+o_{p}(1).$
-[…]
+By the consistency of $\widehat{\textbf{a}}$ for a, the continuity of the maximum operator, and the continuous mapping theorem, $\max_{x\in\mathcal{X}^{+}}\widehat{a}(x)\xrightarrow{p}\max_{x\in\mathcal{X}^{+}}a(x)$.
+Because $\mathcal{X}^{+}=\text{supp}(X\mid W_{0}=1)$ is a finite set, we also have that $\max_{x\in\mathcal{X}^{+}}a(x)=\sup(\text{supp}(a(X)\mid W_{0}=1))$.
+Another application of the continuous mapping theorem suffices to show that $\widehat{\overline{P}}$ is consistent for $\overline{P}$.
 Part 2: Asymptotic Distribution
@@
 $=\psi(\sqrt{n}(\widehat{\textbf{a}}-\textbf{a}),\sqrt{n}(\widehat{\textbf{w}}_{0}-\textbf{w}_{0}),\sqrt{n}(\widehat{\textbf{p}}-\textbf{p}))+o_{p}(1)$
-[…]
+$\xrightarrow{d}\psi(\mathbb{Z})$
 by the continuity of $\psi$ and Assumption C.1. ∎
 Proof of Theorem C.2.
-[…]
+We verify the validity of the bootstrap by appealing to Theorem 3.2 in Fang and Santos (2019).
+We show that their Assumption 4 holds by showing the mapping $\widehat{\psi}$ satisfies $|\widehat{\psi}(h^{\prime})-\widehat{\psi}(h)|\leq C_{n}\|h^{\prime}-h\|$ for any $h^{\prime},h\in\mathbb{R}^{3K}$ and for $C_{n}=O_{p}(1)$, and by showing that $\widehat{\psi}(h)\xrightarrow{p}\psi(h)$ for all $h\in\mathbb{R}^{3K}$.
 Let $h=(h_{1},h_{2},h_{3})$ and $h^{\prime}=(h_{1}^{\prime},h_{2}^{\prime},h_{3}^{\prime})$.
@@
 By the proof of Theorem C.1, $\sqrt{n}(\max_{j\in\mathcal{X}^{+}}\widehat{a}(x_{j})-\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j}))=o_{p}(1)$.
-Since $-\sqrt{n}\xi_{n}\rightarrow-\infty$, $\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})\to\mathbb{P}(0\geq-\infty)=1$ when $k\in\Psi_{\mathcal{X}^{+}}$.
+Since $\sqrt{n}\xi_{n}\rightarrow-\infty$, $\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})\to\mathbb{P}(0\geq-\infty)=1$ when $k\in\Psi_{\mathcal{X}^{+}}$.
 Now suppose that $k\notin\Psi_{\mathcal{X}^{+}}$. Then,
 $\mathbb{P}(k\in\widehat{\Psi}_{\mathcal{X}^{+}})$ $=\mathbb{P}(\widehat{a}(x_{k})\geq\max_{j\in\widehat{\mathcal{X}}^{+}}\widehat{a}(x_{j})-\xi_{n})\rightarrow\mathbb{P}(a(x_{k})\geq\max_{j\in\mathcal{X}^{+}}a(x_{j})-0)=0,$
-[…]
+where the last equality holds from $k\notin\Psi_{\mathcal{X}^{+}}$.
+Therefore, $\mathbb{P}(\widehat{\Psi}_{\mathcal{X}^{+}}=\Psi_{\mathcal{X}^{+}})\rightarrow 1$ as $n\rightarrow\infty$.
+This implies $\widehat{\psi}(h)\xrightarrow{p}\psi(h)$, which concludes the proof.
+∎
 Appendix H Additional Derivations for Difference-in-Differences
@@
 $\sigma_{kj}^{k}=\frac{\mathbb{P}(G=j)\cdot\mathbb{P}(G=k)\cdot\big[1-\mathbb{P}(D=1\mid G=k)\big]\big[\mathbb{P}(D=1\mid G=k)-\mathbb{P}(D=1\mid G=j)\big]}{\operatorname{var}(D^{\perp\left(G_{t_{1}},\ldots,G_{t_{K-1}},P_{1},\ldots,P_{T}\right)})}.$
-[…]
+Here $A^{\perp B}$ is used to denote the residual in the linear projection of $A$ on $(1,B)$.
+It is also the case that $\sum_{k:\;\operatorname{var}(D\mid G=k)>0}\sum_{l>k}\big(\sigma_{kl}^{k}+\sigma_{kl}^{l}\big)=1$.
+[^9] The result in Goodman-Bacon (2021) technically also includes a weight $\sigma_{kU}$ attached to the contrast between group $k$ and the never-treated group.
+We subsume this weight under $\sigma_{kj}^{k}$, and likewise subsume the weight on the contrast with the always-treated group under $\sigma_{jk}^{k}$.
+When we compare this representation with Proposition 5.4, that is,
 $\beta_{\text{TWFE}}$ $=$ $\frac{\mathbb{E}[a_{\text{TWFE,H}}(G)\cdot\mathbb{P}(D=1\mid G)\cdot\tau_{0}(G)]}{\mathbb{E}[a_{\text{TWFE,H}}(G)\cdot\mathbb{P}(D=1\mid G)]}$